### v4

# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA Current address: ]Thomas Jefferson National Accelerator Facility, Newport News, Virginia 23606, USA

Abstract
The spin structure functions of the proton and the deuteron were measured during the EG4 experiment at Jefferson Lab in 2006. Data were collected for longitudinally polarized electron scattering off longitudinally polarized NH3 and ND3 targets, for values as small as 0.012 and 0.02 GeV2, respectively, using the CEBAF Large Acceptance Spectrometer (CLAS). This is the archival paper of the EG4 experiment that summaries the previously reported results of the polarized structure functions , , and their moments , , and , for both the proton and the deuteron. In addition, we report on new results on the neutron extracted by combining proton and deuteron data and correcting for Fermi smearing, and on the neutron moments , , and formed directly from those of the proton and the deuteron. Our data are in good agreement with the Gerasimov-Drell-Hearn sum rule for the proton, deuteron, and neutron. Furthermore, the isovector combination was formed for and the Bjorken integral , and compared to available theoretical predictions. All of our results provide for the first time extensive tests of spin observable predictions from chiral effective field theory (EFT) in a range commensurate with the pion mass. They motivate further improvement in EFT calculations from other approaches such as the lattice gauge method.
I Introduction
The study of nucleon structure has been an active field of research ever since the discovery that nucleons were composite [Frisch:1933, Esterman:1933, Hofstadter:1953zz]. It is now well known that the nucleon, like all hadrons, is made of partons (quarks, antiquarks, and gluons) and that its structure is dominantly ruled by the strong nuclear interaction described by quantum chromodynamics (QCD) [book:QCDhadron, book:Roberts]. However, the behavior of the QCD coupling constant [Prosperi:2006hx, Deur:2016tte, Deur:2023dzc] determines that our understanding of nucleon structure at long distance is not as developed as that at short distance. At high energy-momentum (short distance) typically well above the GeV scale, , which allows us to treat QCD perturbatively. Presently, the approximations of observables are typically known to order , and some to . On the other hand, at low energy-momentum, , and a perturbative expansion in becomes inapplicable. The problem is further complicated by the rise of complex phenomena, e.g. the confinements of partons, the emergence of hadronic degrees of freedom or the spontaneous breaking of QCD’s near chiral symmetry. In this domain, due to the lack of a mature analytic technique based on the QCD Lagrangian, effective theories or models are often used. In particular, chiral effective field theory (EFT) [Bernard:1995dp], based on the observed spontaneous breaking of the chiral symmetry in hadronic states, has successfully predicted hadronic observables at low energies.
The principle of EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon- mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared () where EFT applies, typically below a few tenths of GeV2. Especially important in this context are low nucleon spin structure data [Anselmino:1994gn, Bass:2004xa, Bass:2004xa, Chen:2005tda, Burkardt:2008jw, Kuhn:2008sy, Chen:2010qc, Aidala:2012mv, Deur:2018roz, Ji:2020ena] which are challenging to obtain experimentally but are crucial for the development of EFT [JeffersonLabE94010:2004ekh]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03-006 and E05-111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [CLAS:2003umf]. EG4 was designed to test EFT predictions by precisely measuring the longitudinal spin structure function and its moments for the proton and the deuteuron down to GeV2. These low data, reported in two letters [CLAS:2017ozc, CLAS:2021apd], provided benchmarks for EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [CLAS:2016qxj], but these were focused on the semi-inclusive channel and are not included here.
The article is arranged as follows: In Section II we present the formalism of unpolarized and polarized electron scattering. We introduce the nucleon structure functions, their moments, and polarized sum rules, which are relevant not only for testing EFT but also to form observables that have no known method of direct measurement. In Section III we describe EG4 and the special features that distinguish it from other CLAS experiments. In Section IV we describe the analysis procedure of extracting the polarized yield differences from the data. In Section V we describe the Monte-Carlo simulation, the normalization of the data using the elastic scattering yield, and the extraction of the structure function . In Section VI.1 we present results on , of the proton and the deuteron, followed by the procedure to extract the neutron results in Section VI.2. We present results for and discussions of all the moments in Section-VII. We then summarize, conclude, and provide perspectives in Section VIII.
II Formalism
All formalism presented in this section is based on Ref. [Anselmino:1994gn] or [Kuhn:2008sy], and the algebraic manipulations therein.
II.1 Inclusive Electron Scattering
Within the approximation of one-photon exchange between the lepton and the target (Fig. 1), the scattering process is described by two kinematic variables:
the virtuality of the exchanged photon and the invariant mass of the photon-target system . The virtual photon can be viewed as a probe of the substructure of the target nucleus (or nucleon) and describes the (inverse) space-time resolution of the probe. For the fixed-target case, one has
where is the scattering angle of the electron, i.e., the angle formed by and , and
Here is usually the mass of the nuclear target in the case of nuclear scattering. However, when we study inelastic scattering from the nucleons, the nucleon mass is used instead:
Unless indicated otherwise, Eq. (3) will be used hereafter for the definition of the invariant mass. Alternatively to , one may characterize the scattering process using the Bjorken scaling variable [Bjorken:1968dy] , or for fixed targets:
II.2 Cross Sections and Structure Functions
For inclusive unpolarized inelastic scattering off a spin-1/2 target, the differential cross section for detecting the final-state lepton in the solid angle and in the final energy range (, ) in the laboratory frame can be written as:
where and are the two unpolarized structure functions characterizing the target structure for unpolarized inclusive lepton scattering. The Mott cross section (representing scattering off a point-like, spinless, infinitely-heavy target) is
with the quantum electrodynamics (QED) coupling.
In the polarized case where the electrons and the spin-1/2 target are polarized along the beam direction, the helicity-dependent cross section difference is:
where the and represent the electron and the target spin directions being parallel (anti-parallel) to the beam direction, respectively. The spin-dependent properties of the target are characterized for inclusive lepton scattering by and , the longitudinal and transverse polarized structure functions, respectively. The second line on the right-hand side (RHS) of Eq. (7) indicates that one can extract from this cross section difference assuming a model for the product , where is the virtual photon asymmetry to be defined in the next section.
II.3 Virtual Photon Cross Sections
The formalism provided in the previous two sections focused on the interaction cross section between the incident electron and the target. It is often useful to also study cross sections for the virtual photon being absorbed by the target, and their dependence on the photon polarization. We follow the formalism of Ref. [book:Roberts] and Refs. [Drechsel:2000ct, Gorchtein:2004jd].
The photon polarization 4-vector, , includes two transverse polarization modes that satisfy and a longitudinal polarization mode that satisfies . The total cross section of the photon absorption by the spin-1/2 target can be separated into two terms: the longitudinal and transverse virtual photon cross sections
that are associated with the longitudinally and transversely polarized virtual photons, respectively. The virtual photon equivalent energy (sometimes called virtual photon flux) in Eqs. (8-9) is not a direct observable and different definitions have been proposed: the Hand convention , the Gilman convention , or using the photon energy . For now, we will continue to use instead of choosing a convention. The ratio of the longitudinal to transverse cross sections is:
where
Using the “relative longitudinal polarization” 111It is called “photon transverse polarization” in Ref. [Drechsel:2000ct]. However since this is the weighting factor of , not , we prefer “relative longitudinal polarization”. of the virtual photon , the total photon absorption cross section can be written as
with
and the photon flux factor
For the polarized case, one can similarly divide the virtual photon absorption cross section by a spin-1/2 target into several components based on the polarization of the photon, and the total spin of the photon and the target projected along the direction of the photon momentum ( and for the photon spin antiparallel and parallel to the target spin, respectively), and a component resulting from the interference between transverse and longitudinal virtual photon-nucleon amplitudes [book:Roberts]. From these components one further defines
These spin-dependent virtual photon cross sections relate to the nucleon polarized structure functions as:
Finally, two virtual photon asymmetries are also commonly employed. The first is the longitudinal asymmetry :
from which we obtain:
The second is the longitudinal-transverse interference asymmetry:
The asymmetries , , and are related as
Note that one can use either or to describe the polarized cross section, Eq. (7).
II.4 Sum Rules and Moments of Structure Functions
In this section we present the Bjorken, GDH, generalized GDH, and generalized polarizability sum rules, which all involve moments of polarized structure functions or, equivalently, and . For reviews of nucleon spin structure and additional sum rules, see Refs. [Chen:2005tda, Kuhn:2008sy, Chen:2010qc, Deur:2018roz].
II.4.1 Bjorken Sum Rule
The Bjorken sum rule, derived for using current algebra and isospin symmetry [Bjorken:1966jh, Bjorken:1969mm], predates QCD and is not a QCD prediction. However, its generalization to include the -dependence appearing at finite [Kataev:1994gd, Kataev:2005hv, Baikov:2010je] stems from pQCD. Studying the sum rule in the pQCD domain ( GeV2) therefore tests whether QCD correctly describes the strong force when spin degrees-of-freedom are explicit. The Bjorken sum has been actively measured at CERN [SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv], DESY [HERMES:2002gmr], JLab [Deur:2004ti, Deur:2008ej, Deur:2014vea], and SLAC [E143:1995rkd, E143:1995clm, E154:1997xfa, E143:1998hbs, E154:1997ysl, E155:2000qdr], with measurement agreeing to better than 10% with the sum rule expectation [Deur:2018roz]. Measurements performed with the higher-energy accelerators, CERN and SLAC, provide data at larger values, offering extensive coverage of the low- part of the Bjorken sum. Measurements at the lower-energy facilities, DESY and JLab, covered the smaller range while overlapping with the SLAC data. Due to their lower energy, DESY and JLab have a limited low- reach, which is supplemented by extrapolating the low- data obtained from CERN and SLAC.
The original Bjorken sum rule () reads:
where is the axial coupling constant measured in neutron beta decay. Equation (24) can be generalized for finite by accounting for gluon radiation and higher-twist (HT) effects:
where [Kataev:1994gd]
and HT contains higher-twist corrections proportional to powers of .
Beside the pQCD domain, the Bjorken sum rule is also important at low due to its close connection to the GDH sum rule (see next section). The behavior of the Bjorken sum at GeV2 was precisely mapped by several experiments at JLab [Deur:2004ti, Deur:2008ej, ResonanceSpinStructure:2008ceg, Deur:2021klh]. Note that Eqs. (24-26) are not valid in the low domain and Bjorken sum rule predictions are provided by either models using non-perturbative approaches [Burkert:1992tg, Burkert:1993ya, Soffer:2004ip, Pasechnik:2010fg, Pasechnik:2009yc, Brodsky:2010ur] or EFT calculations [Bernard:2012hb, Alarcon:2020icz].
II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
The GDH sum rule [Gerasimov:1965et, Drell:1966jv, Helbing:2006zp] is based on dispersion relations, unitarity, relativity, and gauge invariance and is derived for real photoproduction (). Like the Bjorken sum rule, the GDH sum rule is derived in a more general context than that of QCD and also predates it. In particular, the GDH sum rule applies to any type of target. For targets whose internal structure is governed by the strong interaction, such as nucleons and nuclei, the sum rule provides a path to study QCD [Deur:2018ntc]. For real photons, the GDH sum rule gives
where and are the photo-absorption cross sections for target and photon spins anti-parallel (A) and parallel (P). (For spin-1/2 targets, these are the same as and in Eqs. (15-16).) Here, and are the mass and the spin of the target, and is the minimal photon energy for inelastic excitation of the target. In the context of (virtual) photon absorption on a nucleon, this threshold energy corresponds to the pion production threshold:
with the pion mass. The anomalous magnetic moment of the target is defined by the total magnetic moment of the particle with the charge of the target in units of the elementary charge, and can be related to the gyromagnetic ratio as . The RHS of Eq. (27) is b for the proton (, and, ), b for the neutron (, , and ), and b for the deuteron (, , and ). Whereas the overall consensus [Helbing:2006zp] is that the GDH sum rule is theoretically very solid, the question of its validity has been debated, mainly regarding whether decreases fast enough with for the integral to converge. Concurrently, its experimental verification has been the focus of several experiments at MAMI and ELSA [GDH:2001zzk, Ahrens:2006yx, GDH:2003xhc, GDH:2005noz, Dutz:2004zz], BNL [LSC:2008wiu], JLab [CLAS:2017ozc, CLAS:2017ozc, JeffersonLabE97-110:2019fsc], GRAAL (ESRF) [Renard:2000jc], and HIGS (TUNL) [Weller:2002aa] providing results that show, given reasonable assumptions for the large behavior of its integrand, the GDH sum rule is valid for the proton to within the 10% precision of the world data [Helbing:2006zp, CLAS:2017ozc, CLAS:2021apd, Strakovsky:2022tvu].
The GDH integral for the nucleon, when expressed with spin structure functions using Eq. (19), is an integral over [Anselmino:1988hn]. Like the Bjorken sum, the small correction from vanishes when , see Eq. (11). Therefore, the formal expression of the (isovector) GDH and Bjorken integrals are proportional, whereas the domains of applicability of the sum rules are disjoint at high and , respectively. Several methods have been proposed to bridge the two sum rules. Generalization to finite values for the left-hand side (LHS) of Eq. (27) is straightforward since or equivalently and exist also at non-zero . A commonly used definition is [Gorchtein:2004jd, Anselmino:1994gn]:
Note that hereafter in this article, a line above a symbol signifies that it does not include the contribution from elastic scattering. At finite , Eq. (29) can be expanded using Eq. (21):
where was used in the last step. The integral can thus be treated as a generalized form of the GDH integral and can be determined from the measured values of .
The limit for the integral of Eq. (29) can be obtained by multiplying the RHS of Eq. (27) by the extra factor . Since at , , it reads:
Therefore, measuring at very low also allows one to test the GDH sum rule.
Using , the Bjorken sum can be extrapolated to the real photon point as:
which should have a positive slope given .
The above procedure generalizes only the GDH integral but not the GDH sum rule, i.e., the RHS of Eq. (27). Other possible generalizations of the GDH integral as reviewed in Ref. [Drechsel:2000ct] share the same caveat. To generalize the sum rule itself, one needs to note that the integral of can be related to the first forward virtual Compton scattering amplitude [Ji:1999mr, Kuhn:2008sy, Deur:2018roz] using the low energy theorem [Low:1954kd]. Accordingly, a full generalization of the GDH sum rule is [Ji:1999mr]:
where the bars and indicate that the elastic contribution (quasi-elastic for nuclear targets) is excluded. In the quark-parton model, this integral gives the quark polarization within the nucleon weighted by their electric charge squared, and provides one of the ingredients for Ellis-Jaffe sum rule [Ellis:1973kp] and Bjorken sum rule. For the generalized form of the GDH sum rule, Eq. (33), the amplitude can be calculated in EFT if is sufficiently low [Ji:1999pd, Ji:1999sv, Bernard:2002bs, Bernard:2002pw, Kao:2002cp, Bernard:2012hb, Lensky:2014dda, Lensky:2016nui, Alarcon:2020icz]. The generalized GDH integrals have been measured over an extensive range and on proton, neutron, deuteron and 3He by CERN [EuropeanMuon:1987isl, SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, COMPASS:2005xxc, COMPASS:2006mhr, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv, COMPASS:2017hef], DESY [HERMES:2000apm, HERMES:1998pau, HERMES:2002gmr], JLab [Amarian:2002ar, JeffersonLabE94-010:2003dvv, CLAS:2002knl, CLAS:2003rjt, RSS:2006tbm, ResonanceSpinStructure:2008ceg, CLAS:2006ozz, CLAS:2008xos, JeffersonLabE01-012:2008kgk, E01-012:2013fky, CLAS:2014qtg, CLAS:2017qga, CLAS:2015otq, CLAS:2017ozc, CLAS:2021apd, JeffersonLabE97-110:2019fsc, E97-110:2021mxm], and SLAC [E142:1996thl, E143:1994vcg, E143:1996vck, E143:1995rkd, E143:1995clm, E154:1997xfa, E155:1999pwm, E143:1998hbs, E154:1997ysl, E155:2000qdr]. CERN and SLAC provide the higher coverage, with extensive contributions to the moments from the low- region. JLab data cover the lowest domain. All accelerators provide data points in the intermediate region of a few GeV or less, including the COMPASS experiment [COMPASS:2005xxc, COMPASS:2006mhr, COMPASS:2010wkz, COMPASS:2015mhb, COMPASS:2016jwv, COMPASS:2017hef] for CERN, E143 [E142:1996thl, E143:1994vcg, E143:1996vck, E143:1995rkd, E143:1995clm] for SLAC and HERMES for DESY [HERMES:2000apm, HERMES:1998pau, HERMES:2002gmr].
II.4.3 Forward Spin Polarizability
The interaction of a particle with an electromagnetic field, in our case the reaction of the target particle to the photons exchanged with the electron beam, is dictated at first order in the photon energy by the particle electric charge. If the beam and target are polarized, the target particle’s anomalous magnetic moment also comes into play. Both the electric charge and the magnetic moment reflect the pure elastic reaction expected from a pointlike or perfectly rigid object. At second order, the particle compositeness and deformability must be considered. The deformation of the particle, viz its internal rearrangement under the electromagnetic field, are described by its electromagnetic polarizabilities [Hagelstein:2015egb]. The forward spin polarizability arises for the case when the beam and target are polarized longitudinally to the beam direction. Polarizabilities were initially defined for real photons but, just like the GDH sum, they can be generalized to finite , becoming -dependent generalized polarizabilities.
There is no practical method known that allows us to directly measure generalized polarizabilities. Instead, sum rules are used to access them. For , the relevant sum rule involves a higher moment of and [Gell-Mann:1954ttj, Guichon:1995pu, Drechsel:2002ar]:
which can also be calculated in EFT [Ji:1999pd, Ji:1999sv, Bernard:2002bs, Bernard:2002pw, Kao:2002cp, Bernard:2012hb, Lensky:2014dda, Lensky:2016nui, Alarcon:2020icz]. Equivalently, we can write this integral as:
Measurement of has been carried out for the proton at the real photon point at MAMI [Dutz:2004zz]. Measurements at small and intermediate have been carried out at JLab on the proton [CLAS:2008xos, CLAS:2017qga, CLAS:2021apd] and neutron [JeffersonLabE94010:2004ekh, CLAS:2008xos, CLAS:2015otq, E97-110:2021mxm]. A short review of the experimental data and predictions from EFT, MAID [Drechsel:1998hk], and SAID [Workman:2012jf] models can be found in Ref. [Strakovsky:2022tvu].
II.5 Deuteron Nuclear Model
As discussed above, the GDH sum rule holds for any target with arbitrary spin, once we replace the expressions and with the cross sections for target and photon spins aligned vs. anti-aligned.
At the same time, experimental information on free neutrons is not realistically accessible, necessitating measurements on (neutrons bound in) nuclei to access the various structure functions and integrals defined above. For the EG4 experiment, we used the deuteron as a source of information on (bound) neutrons. In the most naïve model, the rather loosely bound deuteron can be thought of as a combination of one neutron and one proton, allowing us in principle to extract neutron structure functions after subtracting the contribution from the proton. In reality, this interpretation is complicated by various nuclear binding effects, including the Fermi motion of both bound nucleons, the depolarization of both nucleons through the D-state admixture in the deuteron ground state wave function, and potential off-shell corrections to the (virtual-) photon-nucleon interaction. In addition, the measured cross section (differences) also contain contributions not present for single free nucleons, including two-body currents, final state interactions and coherent production as well as the two-body breakup of the deuteron, D.
Ideally, a rigorous test of theoretical models like EFT would compare observables measured on the deuteron with calculations taking into account all of the above effects. In practice, such calculations do not (yet) exist, and experimentally, covering the entire kinematic range from two-body breakup (2.2 MeV for real photons) to the limit of large GeV is also extremely challenging. For example, the GDH sum rule on the deuteron predicts a very small value for the integral of Eq. (27) because the anomalous magnetic moment of the deuteron is much smaller than that of the proton or neutron (0.14 vs. 1.8-1.9). This is due to a very strong cancellation between the contribution from the two-body () breakup of the deuteron, dominated by low photon energy of order a few MeV, and the incoherent sum from inelastic channels on both nucleons (e.g., production with a threshold of over 140 MeV).
In lieu of a full theoretical description of the scattering process on the deuteron, we use an approximation to express structure functions of the deuteron as a sum of structure functions of the proton and the neutron, convoluted with their momentum and spin distribution inside the deuteron. Here we follow the prescription of Ref. [Kahn:2008nq]. Our approach corrects for the effects of Fermi motion and the effective nucleon polarization due to the deuteron D-state contribution. At the same time, experimentally, we avoid the kinematic region where coherent and two-body breakup effects are expected to be most significant, i.e., low energy transfer. For this reason, we extract neutron (and deuteron) structure function results only in the region GeV from our data, corresponding to virtual photon energies in excess of 240 MeV even at our lowest point. This also minimizes the contribution from quasi-elastic scattering and its radiative tail to our data, so that we should be dominated by the sum of contributions from the two individual nucleons. Since our lowest bin, with a mean value of GeV2, corresponds to a resolution of fm, we can also assume that these contributions largely add incoherently, given the root mean square (RMS) separation of the two nucleons in deuterium of nearly 4 fm. The region 1.073 GeV 1.15 GeV, when needed, can be augmented by a parametrization of single-pion production off protons and neutrons from the MAID partial wave analysis [Drechsel:1998hk], where multi-particle final states do not contribute and the MAID parametrization is very reliable.
For moments of spin structure functions, the most conspicuous nuclear effect, viz. smearing of the resonances by Fermi motion, is integrated out. It vanishes exactly for first moments and to a good approximation for higher moments. We studied this assertion using a model for both smeared and unsmeared spin structure functions and found no difference in the first moments, and only minute differences in the higher moments () (up to 4.8%, which is small compared to our statistical and systematic uncertainties). Thus, no corrections were applied to the data; we only account for the effective polarization of the nucleons in deuterium when relating integrals for protons, neutrons, and deuterons, while supplementing the region 1.073 GeV 1.15 GeV with the MAID parametrization:
with [Lacombe:1980dr, Machleidt:1987hj, Zuilhof:1980ae, Kotthoff:1976fg, Desplanques:1988mp] accounting for the nucleon depolarization caused by the deuteron D-state.
III The EG4 Experiment
The CLAS EG4 experiment [PR03006, PR05111] was carried out at JLab’s experimental Hall B in 2006. Data in the range GeV were collected using six beam energies and on longitudinally polarized NH3 and ND3 targets. Since the physics goal of EG4 was focused on the EFT domain, data were collected with high statistics at very low values, which enabled us to form moments down to GeV2 for the NH3 target. As a comparison, the previous CLAS spin structure experiment EG1b [CLAS:2017qga] measured down to GeV2. Therefore, the data below GeV2 were improved significantly from this experiment.
Figure 2 shows the EG4 kinematic coverage for polarized NH3 and ND3 targets. The beam energies used during the experiment were 1.054, 1.338, 1.989, 2.260, and 2.999 GeV. For the ND3 target, only two beam energies 1.34 and 1.99 GeV were used. The beam polarization was on average 85% throughout the experiment, as measured by the Hall B Möller polarimeter. The ranges in the target polarization were and for NH3 and ND3, respectively, measured by the NMR polarimetry.
III.1 The Polarized Electron Beam and Beam Monitoring
For each of the three experimental halls during the JLab 6 GeV era, a 499 MHz beam was generated with the desired beam current and polarization at the source. They were then intercalated to form a 1497 MHz beam and accelerated to about 45 MeV by superconducting radio-frequency (SRF) cavities. The beam was then injected into the two 0.4-km LINACs consisting of 20 SRF cavities each. The two LINACs are connected by recirculation arcs, forming a racetrack-shaped accelerator. Each revolution increased the beam energy by up to 1.2 GeV, determined by the settings of the SRF cavities. The beam could be circulated up to 5 times before being separated and dispatched to the three halls.
The EG4 experiment employed beams circulated between 1 and 3 times, with the beam current ranging from 1 to 3.5 nA. The Hall B beamline is instrumented to measure and monitor the beam properties, namely, position, transverse distribution (beam profile), current, polarization, and charge asymmetry [CLAS:2003umf]. Nine RF cavities grouped in three beam position monitor (BPM) sets, located 36.0, 24.6, and 8.2 m upstream from the CLAS center, measured both the beam position and current. The latter was also monitored by synchrotron light monitors (SLM) and measured absolutely by a Faraday cup positioned at the end of the beamline, 29.0 m downstream from the CLAS center. The Faraday cup consisted of a 15-cm diameter Pb cylinder of 75 radiation lengths thickness and measured the accumulated charge. Time-derivation of the accumulated charge provided the beam current. The beam profile was periodically verified by running thin wires (“harps”) through the beam and detecting scattered electrons using photomultipliers (PMTs) located 10 cm from the beamline. The harps also served to calibrate the BPMs.
The EG4 experiment used a longitudinally polarized beam. The beam helicity was flipped at 30 Hz, taking the quartet structure of either + - - + or - + + -, with the first helicity window selected from a pseudo-random sequence. The beam polarization was regularly measured by a Möller polarimeter set at the entrance of the Hall. It consisted of detectors, transport and polarizing magnets, and a 25 m thick Permendur foil (49% Fe, 49% Co, 2% Va) acting as a polarized electron target and oriented at with respect to the beamline. A pair of 120-G Helmholtz coils thermally polarized the foil longitudinally to 7.5%. For each beam polarization measurement, the foil was inserted in the beam for about 30 minutes. The scattered (recoil) electrons were transported to the detectors (two lead-glass blocks) by two quadrupoles. The Möller asymmetry was measured by detecting the two electrons in coincidence and compared to the well-known theoretical expectation to obtain the beam polarization. A 30-min measurement provided a 1% statistical precision. The beam polarization for EG4 was typically around 86% [CLAS:2016qxj], determined with a 2.5% systematic uncertainty. Two of the Möller measurements were cross-checked with a Mott polarimeter set in the CEBAF injector that measured the polarization of 5 MeV electrons [Price:1998xd]. These direct measurements were used to check the stability of the beam polarization, while for the data analysis the product of beam and target polarization was directly determined in elastic or quasi-elastic scattering, see Sections V.1 and V.4.
The beam charge asymmetry, viz the relative difference in the number of electrons per helicity bunch, was monitored by the BPMs, SLM, and Faraday cup and minimized by feedback to the electron polarized source. Most individual runs had beam charge asymmetry lower than , with a maximum of several for some runs.
The Hall B beamline was not equipped with a beam energy measurement device. Instead, the beam energy was calculated from the setting of the CEBAF magnets together with the beam position [Krafft:2000nm]. The relative accuracy reached on the beam energy measurement was a few . In addition to its instrumentation, the Hall B beamline had two pairs of magnetic coils to raster the beam over the target. This was to spread the heat uniformly over the target to avoid depolarization, see next section. The raster used a spiral pattern of 1.2 cm outer diameter.
III.2 The Polarized Targets
During EG4 the polarized target was placed 1.01 m upstream from the CLAS center to increase the acceptance at low by reducing the minimum angle for scattered electrons. The targets included frozen 15NH3 and 15ND3 dynamically polarized at 1 K with a 5-T field, as well as two carbon targets and an empty cell for reference. These were the same as the targets used for previous CLAS double-polarization measurements [Crabb:1995xi]. The beam was rastered to protect the target from overheating and to minimize depolarization. The target density was 0.917 and 1.056 g/cm3 for NH3 and ND3, respectively.
The structure of the target insert for EG4 is shown in Fig. 3. The target cells were either 1.0 cm or 0.5 cm in length, for systematic studies, and were referred to as the long and short targets, respectively. The insert was placed into the target “banjo”, an approximately 1-liter vessel of 1-K liquid helium at the center of a 5-T superconducting split coil magnet. A complete description of the polarized target can be found in Ref. [Keith:2003ca].
III.3 The CLAS Spectrometer and the Cherenkov Detector
The basic structure of CLAS is shown in Fig. 4. The design of the CLAS detector was based on a toroidal magnetic field that was generated by six superconducting coils arranged around the beamline. In the EG4 experiment, the torus field setting was to bend electrons away from the beamline (outbending configuration). The magnet coils naturally separated the detector into six “sectors”, each functioning as an independent magnetic spectrometer. Each sector included four sub-detectors: drift chambers (DC), Cherenkov counters (CC), time-of-flight scintillator counters (SC), and electromagnetic calorimeters (EC). The drift chambers were located before, within, and after the torus magnetic field and performed charged particle tracking, allowing for the determination of the particle momentum from the curvature of their trajectories. The other sub-detectors were located outside the magnetic field region. In order to cover the very low region with the high detector efficiency necessary for the absolute cross section measurements, a new CC was built by the INFN-Genova group and was installed in Sector 6. The main elements of this detector are shown in Fig. 5. Basic information on the new CC can be found in Ref. [CLAS:2016qxj].
Another special feature adopted by EG4 was a new tungsten cone for the Möller shield, shown in Fig. 4, which allowed direct line-of-sight from the target at smaller scattering angles down to 5∘.
III.4 Trigger and Data Acquisition System
The main electron trigger for EG4 was formed using a coincidence between the signals from the EC and the new CC; consequently only 1/6 of the full azimuthal acceptance of CLAS was used. For calibration of the new CC performance, data were also taken using EC-only triggers.
All photomultiplier-tube (PMT) time-to-digital-converter (TDC) and analog-to-digital-converter (ADC) signals (i.e., SC, EC, and CC signals) generated within 90 ns of the trigger were recorded, along with DC TDC signals [11]. A trigger supervisor (TS) directed all the signals to the data acquisition system. The offline physics event reconstruction code used geometric parameters and calibration constants to convert the TDC and ADC data into kinematic and particle identification data. The code cycled through particles in the event to search for a single trigger electron – a negatively charged particle that produced a shower in the EC. If more than one candidate was found, the one with the highest momentum was selected. This particle was traced along its geometric path back to its intersection in the target to determine the path length, which, with the assumption that its velocity , determined the event start time. From this start time, the TOF of other particles could then be determined from the SC TDC values. The TDC values from the EC were used when SC values were not available for a given particle.
IV Data Analysis
The analysis presented in this section lead to the extraction of both the proton and the deuteron spin structure functions , , and their moments, previously reported in Refs. [CLAS:2017ozc, CLAS:2021apd]. We will present the general method (Section IV.1), event selection criteria and kinematic corrections (Section IV.2), particle identification performance and efficiency analysis of the new CC (Section IV.3), and background (Section IV.4).
IV.1 Overview of the Analysis Methods
In this work, we first extracted the event yield difference between the positive and negative beam helicity states. The polarized cross section differences, Eq. (7), were extracted from the yield differences as follows:
where is the number of incident electrons with helicity , respectively, as recorded by the Faraday cup or beam charge monitors; is the number of scattered electrons with incident helicity , within the bin , that pass through all analysis cuts; is the number of polarized target nucleons per cm2; is the product of the beam and the target polarizations; and finally, is the product of the acceptance and efficiency of the detector and the trigger. In practice, we calculated the yield difference in each bin and compared to the simulated value that accounts for efficiency and acceptance, and with the same acceptance and fiducial cuts as data.
One important advantage of our method is that contributions from unpolarized material cancel in the yield differences. This is important for experiments utilizing polarized NH3 and ND3 targets because of the presence of significant amounts of unpolarized material, including target windows, support, insulation, liquid helium for cooling, and the nitrogen in the ammonia beads (the nitrogen can be slightly polarized and is treated as a background contribution). The remaining unknowns on the RHS of Eq. (39), , were obtained indirectly by normalizing the yield differences of elastic scattering of the data to simulation, which provided more accurate information than combining beam and target polarimetry data (, ) along with information on the polarized material thickness of the target.
After event selection and kinematic correction (see next two sections), the data were divided into the bins shown in Table 1, with bin limits determined logarithmically (the upper edge of each bin is of the lower edge ). In practice, proton data exist for all 25 bins, and deuteron data from bin 3 to 24. Our neutron results (as presented in Section VI.2) were extracted from combining proton with deuteron data, and thus also exist from bins 3 to 24 only.
| bin ID | 1 | 2 | 3 | 4 | 5 |
| min (GeV/)2 | 0.0110 | 0.0131 | 0.0156 | 0.0187 | 0.0223 |
| mean (GeV/)2 | 0.0120 | 0.0143 | 0.0171 | 0.0204 | 0.0244 |
| max (GeV/)2 | 0.0131 | 0.0156 | 0.0187 | 0.0223 | 0.0266 |
| bin ID | 6 | 7 | 8 | 9 | 10 |
| min (GeV/)2 | 0.0266 | 0.0317 | 0.0379 | 0.0452 | 0.0540 |
| mean (GeV/)2 | 0.0290 | 0.0347 | 0.0414 | 0.0494 | 0.0590 |
| max (GeV/)2 | 0.0317 | 0.0379 | 0.0452 | 0.0540 | 0.0645 |
| bin ID | 11 | 12 | 13 | 14 | 15 |
| min (GeV/)2 | 0.0645 | 0.0770 | 0.0919 | 0.110 | 0.131 |
| mean (GeV/)2 | 0.0705 | 0.0841 | 0.1005 | 0.1200 | 0.143 |
| max (GeV/)2 | 0.0770 | 0.0919 | 0.110 | 0.131 | 0.156 |
| bin ID | 16 | 17 | 18 | 19 | 20 |
| min (GeV/)2 | 0.156 | 0.187 | 0.223 | 0.266 | 0.317 |
| mean (GeV/)2 | 0.171 | 0.204 | 0.244 | 0.290 | 0.347 |
| max (GeV/)2 | 0.187 | 0.223 | 0.266 | 0.317 | 0.379 |
| bin ID | 21 | 22 | 23 | 24 | 25 |
| min (GeV/)2 | 0.379 | 0.452 | 0.540 | 0.645 | 0.770 |
| mean (GeV/)2 | 0.414 | 0.494 | 0.590 | 0.705 | 0.841 |
| max (GeV/)2 | 0.452 | 0.540 | 0.645 | 0.770 | 0.919 |
IV.2 Event Selection and Kinematic Corrections
The primary signal of this analysis was the scattered electrons in the inclusive mode. Multiple selection criteria were used to select good electron events, requiring the correct particle charge () and valid detector signals in DC, SC, EC, and CC. The quality of the helicity signal recorded in the data stream was examined closely and any quartet sequence with inconsistency was rejected. Kinematic corrections were then applied to each event that consisted of: incoming energy loss correction, raster correction, tracking correction, momentum correction, and outgoing ionization loss correction. These corrections were checked by the position of the peak of elastic scattering, which is required to coincide with the proton mass perfectly.
After kinematic corrections, more selection criteria were applied. First, events must be detected in sector 6 of CLAS. The particle momentum must be between 20% and 100% of the incoming beam energy, and must be higher than GeV, corresponding to the trigger threshold set for the experiment. The distribution of the position along the beamline of the electron vertex was then studied for different bins and fitted using a Gaussian function. Only events within three standard deviation () of the average position of that bin were accepted, except for those few bins where the -distribution was rather wide and a tighter cut was used; see Fig. 6. These vertex cuts ensured that data statistics were maximized while reducing backgrounds from the beamline, and that events scattered from the insulation layers of the target were rejected. A series of cuts were then applied on the event’s momentum vs. polar angle as well as vs. the azimuthal angle to avoid the fiducial regions where the detector efficiencies were low.
IV.3 Particle Identification and Cherenkov Efficiency
Particle identification cuts were applied on the EC and CC signals. For the EC, the ratio of the total energy deposited over the particle momentum within each bin was fit with a Gaussian peak and cuts were applied to select events within from the peak center. Furthermore, the energy deposited in the inner layer of the EC, , was required to be greater than 0.06 GeV. The same and cuts were also applied to the simulation such that a direct comparison with data could be made to account for the efficiency of these PID selection criteria.
Further PID cuts were applied to the number of photoelectrons in the CC signal, requiring for NH3 and for ND3 analysis. The CC signals must also pass conditions that match their timing and geometrical location with those of the EC and DC [Osipenko]. Unlike the EC, the CC efficiency due to these PID selection criteria could not be easily obtained from detector simulation, and had to be extracted from data and applied to the simulation. We discuss below the details of the CC efficiency study.
The event distribution in the hit position of the CC, in terms of , the polar angle at the interaction point, and , the azimuthal angle in the first layer of the DC, was studied to determine both yield and the high efficiency region of the new CC. Both the event distribution and the depend on the event hit position in the CC, and thus vary with the ratio of the CLAS torus current over particle momentum , where A for beam energies 2.0, 2.3, and 3.0 GeV and A for beam energies 1.1 and 1.3 GeV. Figure 7 shows the distribution for one representative bin, for the 2.3 GeV data collected on NH3.
For a specific bin, the CC efficiency for a given bin was determined by taking calibration data that utilized EC alone for electron identification. For each bin, we denote the number of electrons recorded in EC-only calibration data to be and among those events, electrons also pass the CC selection criteria. The efficiency, and its uncertainty for case, were calculated as
For the case of or if , we must consider the uncertainty due to being an integer (that is, we cannot detect half of an electron!). If , we assign:
If we assign:
Note that the uncertainty due to being integer also exists for the general case, but was completely negligible compared with Eq. (40) for as small as .
Once we extracted the CC efficiency and its uncertainty map for given range and bin, we matched it to the map for the same bin. Then we fit the values of the efficiency as a function of using the Poisson function form, see Fig. 8 top panel. We studied the efficiency fit for both individual bins as well as all bins combined. (When combining bins, we combined and first and then evaluated the efficiency and its uncertainty.) We found the fit to achieve good results for all cases, and the fit results for individual bins were consistent with those achieved by combining all bins, see Fig. 8 bottom panel. These results were used in the simulation to account for the CC efficiency, see Section V.
IV.4 Background Contamination
For inelastic scattering processes, charged pions can be misidentified as electrons. Similarly, neutral pion decay can produce photons and subsequently electrons via the pair production process. Contamination from these two background processes were studied using charged pion yields, and positron yields taken with the in-bending torus configuration. The pion contamination was found to be less than 1%, and pair production yield was on the order of for all kinematics. We did not correct these backgrounds, but instead incorporated them into the systematic uncertainties. An additional background is the electron scattering off the 15N in the target. The 15N is slightly polarized, and thus we discuss this background along with its correction in Section V.8 as part of the systematic uncertainties.
V Simulation and Extraction of Physics Results
As described in the previous section, this analysis was based on extraction of polarized yield differences from data. This was done for all and ranges covered by the experiment, including both elastic and inelastic scattering. By comparing (“normalizing”) the yield difference from simulation and data in the region of the elastic (proton) or quasi-elastic (deuteron) peak, factors such as beam and target polarizations, luminosity, average trigger efficiency and detector efficiencies of Eq. (39) were accounted for and did not need to be treated or corrected separately. Then, the same normalization was applied to the yield difference from simulation of inelastic scattering. By comparing these differences to the measured spectra, the structure functions and can be extracted. In this section we describe the simulation procedure and the extraction of the structure functions.
V.1 Simulation of (Quasi-)Elastic Polarized Yield Differences
A simulation for elastic and quasi-elastic scattering events was performed using GSIM [GSIM:2002a, GSIM:2002b], the standard CLAS simulation program, combined with an event generator for (quasi-)elastic scattering. Events were generated for ranges of and . The and ranges ensured the whole Sector 6 was covered, while keeping a reasonable generator efficiency.
Two functions were used to generate separately the proton elastic peak and its radiative tail with appropriate relative weights. The proton form factors were from Ref. [Arrington:2007ux], and the radiative tail included both internal and external radiation effects following the prescription of Refs. [Mo:1968cg][Tsai:1973py]. The depolarization effect of the Bremsstrahlung photons was calculated following Ref. [Olsen:1959zz].
For the quasi-elastic events, we used a convolution prescription [Kahn:2008nq] to calculate the inclusive structure function of the deuteron stemming only from scattering elastically off a moving proton or neutron. Radiative effects were included using the code RCSLACPOL [CLAS:2017qga], which was based on the treatments of Refs. [Mo:1968cg] and [Kukhto:1983pv]. For the proton contribution, we used the same form factors as for elastic scattering; for the neutron form factors, we used a continued fraction parameterization [Kubon:2001rj] for and the two-parameter Galster fit [E93-038:2003ixb] for .
The simulated data were then passed through the GSIM Post Processor (GPP) [GSIM:2002a] to account for smearing factors in the detectors, and RECSIS [RECSIS:1997], the standard CLAS simulation program for event reconstruction. Detailed inputs to the simulation, GPP, and corrections applied in the reconstruction are given below.
V.2 Energy Loss Correction
The beam energy measured in the experiment (Section III.1) was used in the simulations. The quality of the simulation was evaluated by comparing the reconstructed spectra of the elastic peak to data in all bins. Special attention was paid to the position and the width of the elastic peak because both can affect the normalization. Because kinematic corrections were applied to the data, the elastic peak position extracted from data was very close to the proton mass (within 1-2 MeV). However, the elastic peak position extracted from the reconstructed simulated events was shifted from the proton mass by several MeV due to energy loss of scattering electrons in the simulation that were not corrected completely in the reconstruction. To correct for this effect, we calculated the correction that should be added to in order to align the value of the reconstructed simulated elastic peak to the proton mass. A linear dependence on was observed when combining simulation of all five beam energies: (MeV) with in GeV, and the factor was to account for the possible -dependence due to most of material being oriented perpendicular to the beamline.
V.3 Smearing Factor for DC Resolution
By comparing the peaks of the elastic simulation with those obtained from data, it was found necessary to apply DC smearing in the GPP stage of the simulation. For the ND3 analysis, a single DC smearing factor for each beam energy was used. For the NH3 analysis, we found it necessary to adjust the smearing factor separately for each beam energy. This was done by adopting a wire-number dependence in the DC smearing factor of GPP. The functional form used for the DC smearing factor is
where is the DC wire number and are fit parameters, see Table 2. The smearing factor starts from a value greater than 1 for low wire numbers (small scattering angle) to approximately 1.0 for high wire numbers (large scattering angle). The large smearing factor for small scattering angle was possibly due to limits in the knowledge of the magnetic field of CLAS in that region.
| 1.1 | |||
| 1.3 | |||
| 2.0 | |||
| 2.23 | |||
| 3.0 |
V.4 Comparison of Elastic Simulation with Data
Figure 9 shows a representative bin for the comparison between data and simulation for elastic scattering. Data from low bins were not used for normalization because of the 15N elastic contamination (see Section V.8). Both the energy loss correction and the DC smearing were applied to the simulation.
Careful studies were done for the elastic simulation quality for individual bins. Gaussian fits were used to compare the peak location and width of simulation with data to ensure that the simulated elastic peak agrees almost perfectly with the data (and the peak position agrees with the proton mass). Bins where the simulated peak position or shape deviate from data were excluded from the normalization input. This happened only at very low bins due to 15N contamination (see Section V.8), and at very high bins due to sensitivity to the low-efficiency area of the DC as shown in Fig. 7. After rejecting the very low and very high bins, we normalized the simulated elastic peak to that from data, summed over all bins. The normalization for the deuteron data proceeded along the same lines, using the simulation of the quasi-elastic peak instead.
The uncertainty in the elastic peak normalization included the uncertainty in the proton form factor input [Arrington:2003qk], which was found to be less than at 1.1 GeV and increased slightly for higher beam energies. The uncertainty due to radiative corrections was found to be for all energies. In addition, we compared [Arrington:2007ux] with a recent proton form factor fit [Griffioen:2015hta] that accounted for the uncertainty in the proton radius, and found the effect to be less than 1%. Overall, we estimated the uncertainty in elastic normalization to be with the beam energy in GeV. For the deuteron normalization based on the quasi-elastic peak, we estimated a total uncertainty of 10% for each beam energy, accounting for form factor uncertainties, the observed variation of the measured yield to simulated yield ratio with , and uncertainties in the folding and radiative correction procedure. These uncertainties were taken into account in our estimate of the systematic uncertainties (see Section V.9).
V.5 Simulation of Inelastic Polarized Yield Differences
We used an event generator originally developed for Ref. [CLAS:2003iiq] and improved for EG4 to generate events distributed according to the polarized cross section model from RCSLACPOL [CLAS:2017qga]. The generator worked in two steps: first, it generated separate maps for radiated inclusive polarized cross section differences
for polarized electron scattering off a longitudinally polarized proton or deuteron target. The map covered the 3D space of for ranges of , and GeV. The and ranges were identical to the elastic scattering generator. Two such maps were generated for any given model of the double spin asymmetries ; one for positive values of and one for negative values. Next, events were thrown according to the two cross section maps. The events were given vertex coordinates that were uniformly distributed over the volume of a 1-cm-long cylinder with a radius of 0.01 cm around the beamline – with the center of this volume being at the EG4 target position of cm. Equal numbers of events were generated for each map of . The events generated were then fed into the GSIM/GPP/RECSIS chain to create the reconstructed spectra from the simulation. These spectra were normalized according to the total integrated cross section of each map. The positive and negative spectra were then combined to form the full spectra for inelastic scattering. Careful studies were carried out to ensure the radiative tail of elastic scattering was subtracted from the data before they were compared with the inelastic simulation.
V.6 Extraction of and
Extraction of and was done by varying the structure function model used in the inelastic simulation as follows: first, we constructed the standard simulation spectra using the best models for the unpolarized structure functions and asymmetries:
The choice for and was based on the latest fit by Bosted/Christy/Kalantarians (dated 2014), interpolating to the real photon point at very low . The most up-to-date references are [Bosted:2007xd] and [Christy:2007ve].
For asymmetries, the latest fit performed by the EG1b collaboration [CLAS:2017qga] was used. The parameterization itself is unpublished. The asymmetry model includes three options:
The asymmetries and in the resonance region are based on a combination of the deep inelastic scattering (DIS) fit with extensions into the resonance region. A parametrized fit from MAID is used to get the resonance behavior. All JLab data as of the early 2010’s, as well as some MIT BATES and NIKHEF data, were used in determining the parametrization. Sum rules are obeyed as much as possible (such as the Burkhardt-Cottingham sum rule [Burkhardt:1970ti] and the Soffer inequality [Soffer:1994ww]).
The asymmetry in the DIS region was based on a fit to world data on inclusive spin structure functions in the DIS region (COMPASS, SLAC, HERMES, and JLab early 2010’s) [SpinMuon:1993gcv, SpinMuonSMC:1997voo, SpinMuonSMC:1994met, SpinMuon:1995svc, SpinMuonSMC:1997ixm, SpinMuonSMC:1997mkb, HERMES:1997hjr, HERMES:2006jyl, E142:1996thl, E143:1994vcg, E143:1995rkd, E143:1995clm, E143:1998hbs, E154:1997xfa, E155:1999pwm, E155:2000qdr, COMPASS:2010wkz, COMPASS:2017hef, CLAS:2017qga, CLAS:2008xos, CLAS:2015otq, JeffersonLabHallA:2004tea]. The measured values were fit, with simple Regge (powers of ) behavior assumed as . The full uncertainty matrix of the fit was used to evaluate the systematic uncertainties.
The asymmetry in the DIS region was determined from [Wandzura:1977qf]. To evaluate systematic uncertainties, we added an extra twist-3 term obtained by fitting to the SLAC E155x [E155:2002iec] data.
After establishing the standard simulation as above, we varied by an arbitrary value, or , for all inelastic values, and generated the difference in the polarized cross section:
After applying the same normalization factor as the standard simulation, the resulting simulated event count was denoted and was added to the standard simulation to form the non-standard spectra , which corresponds to the model with changed by :
The polarized yield obtained from data, , would nominally agree with neither the standard nor the non-standard , since neither model would match our data perfectly.
By comparing the polarized yield from our data () with that from the standard simulation () and the non-standard simulation (), the value for was extracted as:
where and are the values used in the standard and the non-standard simulations, respectively. Similarly, the value for was extracted as:
The ratios of and on the RHS of Eqs. (47) and (48) describe how the yield varies with the input or of the model, and directly affect the extraction of these structure functions. These ratios were sensitive to the statistical fluctuations in , which can be large if small bins are used. In our analysis, these ratios were smoothed out by taking a 5-point average of adjacent bins.
We emphasize that our method of extracting and described above does not depend strongly on the specific models used for the Monte Carlo generator, since a larger deviation of the model from the data would simply lead to a larger “correction” in Eqs. (47-48), bringing the result to the same final value.
V.7 Systematic Uncertainty due to Models Used In the Extraction
In the extraction of or described above, there can be systematic uncertainties due to the models used for , , , and in particular, for , because only longitudinally polarized targets were used in this experiment. We evaluated the model dependence as follows: a simulation corresponding to a variation of the input structure function was performed, following the same method as the non-standard simulation (where is changed by ). This variation of simulation gave , with “var” for “variation in the input model”. The value for with this model variation was also calculated, denoted as . The systematic uncertainty in the extracted was calculated as:
and similarly for :
where and are the same values as those on the RHS of Eqs. (47-48), extracted from comparison between the standard vs. the non-standard simulations. A total of six variations was performed:
Variation 1: Both and were increased by 3%;
Variation 2: was changed by with the uncertainty of given by Refs. [Bosted:2007xd] and [Christy:2007ve];
Variation 3: The value of in the resonance region was changed to the fit before including the EG1b data ( unchanged);
Variation 4: The value of in the resonance was changed to the fit before 2009 ( unchanged);
Variation 5: The value of in the DIS region was changed by the uncertainty of the DIS fit used in the the standard version. The full uncertainty matrix of the fit is used.
Variation 6: The value of was changed by adding an extra twist-3 term, obtained by fitting to the SLAC E155x data [E155:2002iec].
Note that the systematic uncertainty evaluated this way can be either positive or negative. When integrating the measured or for the moments, the sign of the systematic uncertainty was retained and there can be cancellations. On the other hand, when combining these variations to form the total systematic uncertainty, the uncertainty from each variation was added in quadrature for , , and for the moments. Also note that model uncertainties affect both the extraction of and directly as well as indirectly through radiative corrections that are based on these models. Our method accounted properly for correlations between these two effects.
V.8 Contamination from Polarized Nitrogen
The 15N in the target was slightly polarized, and scattering from the 15N can contribute to the polarized count difference measured in the data. The 15N contribution had three effects: contribution of 15N nuclear elastic scattering to the proton elastic peak; contribution of radiative tail from 15N nuclear elastic scattering to proton inelastic scattering; and contribution from 15N inelastic scattering to proton inelastic scattering. The first affected normalization of the elastic simulation to the data, while the latter two affected the extraction of the structure functions from the inelastic polarized yield.
To account for each contribution from the 15N background, one first needs to obtain the polarization of 15N as follows: the ratio of 15N polarization over that of the proton is about [Crabb:1995xi], and there are three protons for every 15N nucleus in NH3. Therefore, the 15N contamination to either the inelastic or elastic polarized yield can be obtained by multiplying the calculated 15N unpolarized yield by the effective polarization ratio of 15N.
The contribution from 15N nuclear elastic scattering was calculated using the 15N form factors [Dally:1970lvj], and was found to be problematic only in the low bins. This is because the invariant mass for the 15N nuclear elastic scattering, calculated using Eq. (3), is separated from that of the proton elastic peak () by:
where is the 15N nuclear mass. Therefore in the spectra, the separation the between 15N nuclear elastic and proton elastic peaks was large enough for the two peaks to be distinguished from each other in all data except the very low bins. The contamination from 15N nuclear elastic scattering to the elastic normalization was kept below by excluding elastic data below GeV2 from the elastic normalization.
The contribution of radiative tails from 15N nuclear elastic scattering to the inelastic polarized yield was calculated and found to be no more than 2% of the proton elastic tail itself. Similarly, the relative contribution from 15N inelastic scattering to the proton inelastic polarized yield difference was found to be below 0.7%. We included both effects in the systematic uncertainty evaluation.
In the case of the deuteron, the possible corrections due to other components in the target were somewhat larger, since the deuteron polarization is typically a factor 2 smaller than that of protons. In addition to the contribution from 15N, there is also the possibility of contamination from 14N and ordinary hydrogen in the target. A detailed study of possible corrections from these contributions showed a maximum effect of 4.5% which we included in the systematic uncertainty budget for the deuteron results.
V.9 Summary of Systematic Uncertainties
In the following, we list all systematic uncertainties that enter our final results:
- 1.
2.
Model uncertainties, analyzed by using the model variations described in Section V.7.
3.
Systematic uncertainty due to the radiative tail from proton elastic scattering and from radiative corrections included in our Monte Carlo simulation. There are two potential contributions to this uncertainty: the models for the form factors and structure functions that enter these tails and radiative corrections, and the amount of target material traversed by the beam, which affects the external radiative tails. The former are “automatically” accounted for by our study of the overall systematic effect of using different structure function models, see Section V.7, and by trying different form factor parametrizations (see Section V.4). Note that we modified those model inputs simultaneously in the entire analysis chain, and retained the sign and magnitude of all variations to propagate them to the final results for the structure functions and moments, to properly account for correlations. The second uncertainty comes mostly from the uncertainty of the overall areal density of the target material, which in turn is determined by the packing fraction. The packing fraction (PF) is defined as the ratio of the actual quantity of ammonia target material contained in the target cell to the ideal quantity for a complete filling of the cell volume. While PF is not explicit in Eq. (39), it affects the total target thickness and thus the radiative effects in the simulation. To obtain the uncertainty due to PF, we re-ran the full simulation using a set of “alternate PF values” that differ from the nominal PF values. The data were re-analyzed using the alternate PF simulations and and were extracted. The differences in the structure function results and those obtained using nominal PF values were taken as the uncertainty due to PF values. The uncertainty in the PF values themselves is taken to be 0.1. The nominal PF values used in the analysis are shown in Table 3.
Beam Energy Target Cell PF Value 1.1 GeV long NH3 top 0.717 long NH3 bottom 0.625 1.3 GeV long NH3 bottom 0.657 short NH3 0.602 ND3 0.624 2.0 GeV long NH3 top 0.716 ND3 0.764 2.3 GeV long NH3 top 0.682 short NH3 0.720 3.0 GeV long NH3 top 0.782 Table 3: Nominal packing factor (PF) values used in the analysis for each combination of target type and beam energy. See Fig. 3 for the detailed naming convention of the target cells. - 4.
5.
Systematic uncertainty due to CC efficiency. This was obtained by re-running the analysis without applying the CC efficiency to the simulated spectra. The difference in the extracted and from the nominal analysis indicates the effect of the CC efficiency, and 20% of this difference was taken as the systematic uncertainty due to the CC efficiency. Details of CC efficiency analysis were described in Section IV.3.
6.
Systematic uncertainty due to background contributions. This includes effect from pion and pair production background (1% of the polarized yield, see Section IV.4); 15N elastic tail in the inelastic region (see Section V.8); and the 15N inelastic scattering background in the inelastic polarized yield, which contributes 0.7% of the polarized yield estimated by calculating the 15N to proton inelastic polarized cross section ratio and applying the polarization ratio of the two.
7.
Systematic uncertainty due to event reconstruction. Even with momentum corrections applied to both data and simulation, the simulated elastic peak can differ from data in its position by a small amount (typically 1-2 MeV or smaller). To estimate the uncertainty due to this effect, we shifted the simulated spectra by or MeV, and re-ran the analysis. The difference in the extracted and from the nominal analysis was then scaled by the actual shift in the elastic peak, obtained from figures such as Fig. 9. This uncertainty is typically very small except near the edge of the acceptance. When evaluating the integral of the structure functions, the data in the single bin on the edge are not used because of this reason. Another reason to exclude data from the edges is that they very likely do not cover the full range of the bin.
VI Results on the Structure Functions
We present in this section results on the proton and the deuteron structure functions and , which were reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd]. We also present new results on the neutron as extracted from the proton and deuteron data, as well as their isovector combination .
VI.1 Results on and for the Proton and Deuteron
In the previous section, we described the method of extracting and and the determination of systematic uncertainties on our results. The and results obtained from all target types and beam energies were then combined. When doing so, we first checked whether any two data sets were compatible with each other, using Student’s t-test to compare the distributions of both data sets in the region of common kinematics. The data were found to be consistent.
When combining the five beam energies, the spin structure function and its statistical uncertainty were determined as
where and are the value and its statistical uncertainty obtained from beam energy (). The systematic uncertainty was calculated as
where is the systematic uncertainty obtained for the data taken with beam energy . That is, both experimental systematic and model uncertainties were weighted by the relevant statistical uncertainties of each beam energy. The same method was applied when combining data from different target cell types.
Our final results for and of the proton and the deuteron are shown in Figs. 10 through 13. Our results are consistent with the previous data [CLAS:2015otq, CLAS:2017qga] where their coverage overlap, but extend the measured range to three times smaller values, below the pion mass squared (). This makes it possible to rigorously test EFT calculations for the nucleon spin structure functions. Comparing to our existing model, which was based on existing data prior to EG4, one can see that it tends to be more positive than the data points, especially at lower . Our new data can help improve this model.
VI.2 Results on neutron
From our results on of the proton and the deuteron, we further extracted values for the neutron spin structure function, , which were not available previously. This extraction is based on the simplified model of the deuteron structure functions described in Section II.5, where the structure functions of the deuteron are expressed as sum of the two nucleon structure functions, convoluted with their spin and momentum distributions in the deuteron [Kahn:2008nq]. To extract the deconvoluted structure functions for a neutron at rest, we followed a similar procedure as for the extraction of from the data on the proton and deuteron, described in detail below.
As a first step, we took our results and converted them to the contribution from the bound proton to the deuteron spin structure function using the folding technique of Ref. [Kahn:2008nq]. The bound were then subtracted from our results to obtain data for the bound neutron . The data were compared with expected values for the bound neutron, , produced using our model of the free neutron processed by the same folding technique, and any observed difference is attributed to how the input model of the free neutron differs from reality. We then produced a modification of the base model where was changed by () in the input, and the difference between the modified and the original model provides information on how much the model of the free neutron should be adjusted to match data. More specifically, for the free neutron is extracted as:
Note that the two versions () of the modification produce nearly identical results. The statistical and experimental systematic uncertainties of each of the input and were propagated to produce the experimental uncertainties in the extracted . The uncertainty in this extraction was studied using a similar method as in Section V.7, but now by changing the initial model for the neutron.
Our results for are shown in Fig. 14. Similar to the proton and the deuteron case, we see that the new results from EG4 are consistent with previous results in the region where they overlap, but the best model can be improved in the lowest bins, below about GeV2.
Finally, we show the isovector combination in Fig. 15. Comparing with Figs.10, 11, and 14, one clearly observes the suppression in the isovector component of the . This has important consequences for the Bjorken integral (i.e., the first moment of ), see Section VII.6, and its comparison to EFT [Burkert:2000qm]. Like the other EG4 results on and , Fig. 15 also makes clear that in the domain covered by EG4, the parameterization of the world data (continuous blue line) needs to be revised.
VII Results on Moments of Structure Functions
Similar to the previous section(s), we summarize the procedure to form the moments of the proton and the deuteron structure functions, and present the results that were reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd] but with updated PT calculations. We also present new results on the moments of the neutron extracted directly from the moments of the proton and the deuteron, as well as their isospin combinations and .
VII.1 General Procedure Used to Form the Integrals
To form moments of and , we integrated these quantities over the full range of ( corresponds to the electroproduction threshold), using our model for the low and high region beyond the data coverage. In addition, for a few high bins the coverage of our data has gaps at intermediate values because of disabled DC wires. The integrand over these gap regions was provided by the model as well. The full integral, e.g., was therefore evaluated as:
where and are the lower and upper limits of the data integration, is the estimate from our best model described in Section V.6 and is our data. The low- integration starts at and the step size was set to .
The range of the data integration in Eq. (56) is slightly narrower than the actual kinematic coverage of the experiment because the data uncertainties on the edge of the coverage are large: At large , near the pion threshold GeV, the polarized yield is dominated by the radiative tail from elastic scattering and the measured polarized yield is not sensitive to the value of or ; while for the low or large edge of the data, the spectrometer acceptance drops and the statistics of the data is low, see Section V.9. Thus, the value of was chosen to be at GeV and was set at a few tens of MeV below the maximum covered by the data. The upper -limits of the data integration ranges are shown in Figs. 10 and 11.
Uncertainties in the model enters the integral in two ways. First, the model enters the extraction of the data itself, as described in Section V.7. Second, the uncertainty of the model integration below and above was evaluated by variation of the model parameters, similar to the method described in Section V.7. When evaluating model uncertainties of the data integral, the sign of each model uncertainty of or is retained to allow partial cancellation when integrating. The total uncertainty on the full integral was formed by adding all experimental and model uncertainties in quadrature.
VII.2 Latest EFT Moment Predictions
In the following sections, we will compare the EG4 result on the moments to theoretical predictions, in particular from EFT. We will display only two recent predictions, from Ref. [Bernard:2012hb] and Ref. [Alarcon:2020icz], as they supersede earlier EFT calculations [Bernard:1992nz, Ji:1999pd, Bernard:2002bs, Kao:2002cp]. The crucial advance of the latest calculations is that they account rigorously for the excitation. At the same time, the approaches by Refs. [Bernard:2012hb] and [Alarcon:2020icz] have significant differences. The chief reason is the effective accounting of higher-order contributions of the EFT series by Ref. [Alarcon:2020icz], which uses a phenomenological form factor. It softens the steep -dependence otherwise seen in Ref. [Bernard:2012hb]. The softening form factor of Ref. [Alarcon:2020icz] accounts for their typically better description of the -dependence of moments. The methodological differences between the two calculations are detailed in Section VII.7.
VII.3 Moments of the Proton Spin Structure functions
They agree with the most recent EFT predictions [Bernard:2012hb, Alarcon:2020icz] up to GeV2, above which only that of Ref. [Alarcon:2020icz] agrees with the data. The new results generally agree with a previous experiment [CLAS:2017qga] in the overlapping region within their uncertainties. However, it is visible that the parameterization based on Ref. [CLAS:2017qga] (purple solid curve) can be improved by including the new results. The phenomenological models Burkert and Ioffe [Burkert:1992tg, Burkert:1993ya] and Pasechnik et al. [Pasechnik:2010fg] agree well with the new results for all values. They predict the behavior of in the non-perturbative domain of QCD and beyond: The Burkert and Ioffe model [Burkert:1992tg, Burkert:1993ya] extrapolates DIS data using vector meson dominance and a parameterization of resonance contributions; while the Pasechnik et al. model [Pasechnik:2010fg], improved upon the earlier Soffer-Teryaev model [Soffer:2004ip], employs the analytical perturbation theory (APT) approach that extrapolates DIS data to low , accounting for the mild -dependence of . Lastly, the MAID unitary isobar phenomenological model [Drechsel:1998hk] is a parametrization of experimental pion production data based on a partial wave analysis in the resonance region, and includes nonresonant background. MAID does not account for multi-particle final states and therefore provides a partial in the resonance region, without the low- (high-) contribution. Consequently, it can be roughly compared with the experimentally measured (open blue circles), although it does not necessarily have the same integration limits. With these caveats in mind, MAID and the EG4 measured sum agree well up to GeV2, after which the model is below the data.
Our results on (Eq. (30)) are shown in Fig. 17. Compared with the EFT predictions, similar conclusions as for are reached. The MAID model still agrees with the experimentally measured integral within uncertainties. Finally, the difference between the new results and the parameterization of previous data [CLAS:2017qga] is more visible and the parameterization can clearly be improved.
Extrapolating our results to yields [CLAS:2021apd]
This agrees well with the GDH sum rule prediction and the experimental photoproduction result (stat) (syst) [GDH:2001zzk, GDH:2003xhc, GDH:2005noz, Ahrens:2006yx]. We note that the uncertainties from Refs. [GDH:2001zzk, GDH:2003xhc, GDH:2005noz, Ahrens:2006yx] and Eq. (57) are comparable. For EG4, the uncertainty due to the extrapolation is compensated by the fact that inclusive electroproduction automatically sums over all reaction channels, removing uncertainties due to the detection of all final states needed in photoproduction. The EG4 data thus provide the first test of the GDH sum rule with a different technique than photoproduction.
One more moment that can be formed from the EG4 data is the generalized longitudinal spin polarizability (Eq. (35)). Compared to (Eq. (30)), the -weighting in Eq. (35) indicates that and have different contributions from both systematic and model uncertainties. Results for are shown in Fig. 18. They agree with EFT calculation from Bernard et al. [Bernard:2012hb] only at very low (below GeV2) and agree marginally with Alarcon et al. [Alarcon:2020icz], though both calculations have large uncertainties. On the other hand, when combined with the photoproduction data point [Gurevich:2017cpt], the EG4 results support the larger slope at very low predicted by Bernard et al. [Bernard:2012hb]. The agreement between the measured integrals and the MAID model is similar to both and . The recent calculation by Bigazzi et al. [Bigazzi:2023odl] (not shown in the figure) based on the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [Maldacena:1997re] applied to QCD [Brodsky:2014yha] recovers the sign and qualitative behavior of , albeit up to a normalization factor.
VII.4 Moments of the Deuteron Spin Structure Functions
Results for the deuteron moments , , and are shown in Figs. 19, 20, and 21, respectively. The bar now indicates that, in addition to the elastic reaction, the deuteron electro-disintegration contribution is also excluded from the integrals. Therefore, they approximately correspond to the sum of the proton and neutron moments rather than to the deuteron nucleus (see Section II.5).
Similar observations as for the proton measurements can be drawn: the EG4 data agree with the previous data from CLAS [CLAS:2017qga] (and from SLAC E143 [E143:1996vck], not shown in the figure). The measured and agree with the EFT results for the sum of proton and neutron (multiplied with the D-state depolarization, see Section II.5) of Alarcón et al. [Alarcon:2020icz], and of Bernard et al. [Bernard:2012hb] for the lowest points. The models of Pasechnik et al. [Pasechnik:2010fg] and Burkert-Ioffe [Burkert:1992tg, Burkert:1993ya] agree well with the data, while the parameterization [CLAS:2017qga] is systematically slightly larger. The comparison between the MAID model and the data is similar to that of the proton for all three integrals.
Extrapolating the data to yields [CLAS:2017ozc]
which agrees at the level with the GDH sum rule prediction , obtained using Eq. (37), where the uncertainty is from the deuteron D-state correction. This can be compared with the data using real photons of energy between GeV, (stat) (syst) [GDH:2001zzk]. Note that the systematic uncertainty of [GDH:2001zzk] does not include that from the unmeasured low and large contributions.
Regarding the generalized spin polarizability , the EG4 results are clearly outside the range predicted by Alarcón et al. [Alarcon:2020icz], but agree with the Bernard et al. calculation for the lowest points. The parameterization [CLAS:2017qga] does not describe the data well below GeV2, similarly to the case.
VII.5 Moments of the Neutron Spin Structure Functions
Moments of the neutron spin structure function can be extracted directly from those of the proton and neutron results, see Section II.5. Understanding the deuteron moments as “per nucleus” (rather than “per nucleon”), we rewrite Eqs. (36-38) as
with given previously under Eq. (38). When extracting the neutron moments as above, we added in quadrature the proton and deuteron statistical uncertainties, as well as the systematic uncertainties.
The resulting and are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [E143:1996vck], EG1a [CLAS:2002knl, CLAS:2003rjt] and EG1b [CLAS:2017qga, CLAS:2015otq] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94-010 [JeffersonLabE94-010:2003dvv] and E97-110 [JeffersonLabE97-110:2019fsc] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [Burkert:1992tg, Burkert:1993ya] and Pasechnik et al. [Pasechnik:2010fg] (when available) and with the EFT predictions from Bernard et al. [Bernard:2012hb] (up to GeV2) and, to a lesser extent, with Alarcón et al. [Alarcon:2020icz]. The MAID model prediction is systematically more positive than the data. Like in the case of , the calculation of [Bigazzi:2023odl] (not shown in the figure) agrees in sign and behavior with the , up to a similar normalization factor as for .
The real photon GDH sum for the neutron can be deduced from Eqs. (57) and (58):
where the uncertainty arises from the uncertainties of , and the D-state component added linearly. This agrees within from the sum rule expectation, .
Results on the generalized longitudinal spin polarizability on the neutron are shown in Fig. 24, along with earlier results from JLab EG1b [CLAS:2017qga, CLAS:2015otq]), E97-110 [JeffersonLabE97-110:2019fsc] and E94-010 [JeffersonLabE94-010:2003dvv]). The EG4 neutron results generally agree with the earlier data, though there is a tension between EG4 and E97-110 for GeV2. Adding the EG4 deuteron and proton systematic uncertainties linearly, rather than in quadrature, does not remove the tension. It contrasts with the agreement between EG4 and E97-110 for the neutron data on or , suggesting the discrepancy arises from the high- (near pion threshold) contribution. The EG4 results agree with EFT calculation from Ref. [Bernard:2012hb] up to GeV2 but not with that of Alarcón et al. [Alarcon:2020icz].
VII.6 Isospin Analysis on the Moments of Spin Structure Functions
Results on the proton and deuteron moments presented in the previous sections can be used to form various isospin combinations based on the description in Section II.5. The isoscalar () combination would be similar to the deuteron results, up to a factor from the deuteron D-state. For isovector () moments, the proton and deuteron results were combined as:
where is given under Eq. (38) and with the deuteron moments understood as “per nucleus”. Alternatively to using the EG4 proton and deuteron data, one can use the proton results from EG4 and the neutron results obtained from 3He data such as those from JLab E97-110 [JeffersonLabE97-110:2019fsc] to perform the isospin separation.
The isovector moments are expected to be simpler to compute than the isoscalar or nucleon moments. For lattice QCD simulations, this is because the disconnected diagrams are suppressed [Deur:2018roz]. For EFT calculations, e.g., Refs. [Burkert:2000qm, Alarcon:2020icz], the inclusion of the excitation is often difficult, which would cancel in the isovector combination. However, the consequence of the expectation that the EFT predictions are more robust has been upheld only with data on the Bjorken sum [Deur:2004ti, Deur:2008ej, Deur:2014vea], but not with [Deur:2008ej], nor with the longitudinal-transverse interference polarizability [JeffersonLabE94010:2004ekh, E97-110:2021mxm] for which the contribution is also expected to cancel. Our results provide improved tests of this expectation thanks to the lower coverage and higher precision of EG4 data.
VII.6.1 Results on the Bjorken Sum
The isovector combination is the Bjorken sum [Bjorken:1966jh, Bjorken:1969mm], see Eqs. (25)-(32). Our results on are shown in Fig. 25, compared with data from earlier experiments [E143:1994vcg, Deur:2004ti, Deur:2008ej, Deur:2014vea]. Besides the models already discussed in connection with Figs. 16-24, we also show predictions from holographic light-front QCD (HLFQCD) [Brodsky:2014yha]. In HLFQCD, the QCD coupling is derived in the scheme following the effective charge prescription [Grunberg:1980ja, Deur:2016tte, Deur:2023dzc], from which the Bjorken sum can be obtained [Deur:2005cf, Deur:2008rf, Deur:2022msf].
Note the comparison between the results obtained from EG4’s proton and deuteron data and those extracted from the EG4 proton and E97-110 neutron (3He) data. An agreement between the two different approaches was not assured since the nuclear structures of the deuteron and 3He are quite different: 3He is strongly bound compared to the deuteron, therefore more sensitive to potential nuclear modification of the proton. Our treatment of the deuteron is minimal and consists of accounting for its D-state [Lacombe:1980dr, Machleidt:1987hj, Zuilhof:1980ae, Kotthoff:1976fg, Desplanques:1988mp], while treatment of the 3He data accounted for the nucleon effective polarizations [CiofidegliAtti:1996cg]. As shown in Fig. 25, the two approaches agree well with each other. This suggests that, at least for , the nuclear corrections applied to the deuteron and 3He data to obtain the neutron information appear sufficient at these low values.
Both the EG4 and EG4/E97-110 results on are systematically above EFT and model predictions, except that of HLFQCD [Brodsky:2010ur] and [Gabdrakhmanov:2023rjt], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form to the entire world data set, based on Eq. (32) [Deur:2021klh], where is expected to be the GDH slope. Our results are shown in Table 4. The “” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that /n.d.f = 1 for the best fit. The “” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
We note that our value for deviates significantly from that expected from the GDH sum rule. This is partially due to the large cancellation for the GDH sum rule in the combination, which is only 1/4 of the value for the individual nucleons. Experimentally, because of the nearly perfect cancellation of the large contribution from the resonance in the difference, the isovector integral is much more strongly affected by small- contributions than the other moments, with relatively large uncertainties for . Of course, our results could also indicate a much stronger –dependence near the photon point than can be captured by a 2-parameter fit. In any case, they do not indicate a violation of the GDH sum rule since, as shown in the previous section, our results for the proton and neutron are individually consistent with the sum rule within (1 - 1.3). Further discussions on the possible discrepancy between the parameter and the GDH slope can be found in Ref. [Deur:2021klh].
| Data set | [GeV-2] | [GeV-4] | ||||
| World data |
| GDH sum rule [Gerasimov:1965et] | 0.0618 | - | ||||
| EFT Bernard et al. [Bernard:2012hb] | 0.07 | 0.3 | ||||
| EFT Alarcón et al. [Alarcon:2020icz] | 0.066(4) | 0.25(12) | ||||
| Burkert-Ioffe [Burkert:1992tg, Burkert:1993ya] | 0.09 | 0.3 | ||||
| Pasechnik et al. [Pasechnik:2010fg] | 0.09 | 0.4 | ||||
| HLFQCD [Brodsky:2010ur] | 0.177 | -0.067 |
VII.6.2 Isospin Study of
Our results on the isovector and isoscalar combination of from EG4 are shown in Fig. 26. We also provide results based on EG4 data and E97-110 (3He) data [E97-110:2021mxm]. The total experimental systematic uncertainties on the proton and deuteron/neutron(3He) are combined quadratically, as are the statistical uncertainties. The low- uncertainties are added linearly. The results shown in Fig. 26 agree reasonably well with those from the previous experiments EG1b and EG1b/E94-010 [Deur:2008ej]. However, unlike for , results from the EG4-only and EG4/E97-110 approaches agree only at the higher points. On the other hand, if we combine linearly the total systematic uncertainties of the two experiments, the tension at low vanishes.
For , both the EG4 and EG4/E97-110 results indicate that it stays positive in the measured range, conversely to the EFT and MAID predictions. In the case of , both the EG4 and EG4/E97-110 combinations agree with Bernard et al. in the lower range, but disagree with the Alarcón et al. prediction.
VII.7 Discussions
As discussed in previous sections, the success of EFT in describing our results on the moments of the spin structure function remains limited. One reason for this limited success may be coming from the difficulty to fully account for the degree of freedom in the calculation. In fact, early EFT calculations [Bernard:1992nz, Ji:1999pd] did not include explicitly the excitation, which slows the convergence of the EFT perturbation series, or they included it phenomenologically [Bernard:2002bs, Kao:2002cp]. This was thought to not be important for some observables for which the contribution was expected to be suppressed, such as the longitudinal-transverse spin polarizability and the isovector quantities [Burkert:2000qm] or . It therefore came as a surprise that much of the early nucleon spin structure function data [Amarian:2002ar] disagreed with calculations [Bernard:1992nz, Ji:1999pd, Bernard:2002bs, Kao:2002cp], noticeably and . This puzzle prompted refined EFT calculations [Bernard:2012hb, Alarcon:2020icz] and a new experimental program at JLab optimized to cover the EFT domain [CLAS:2017ozc, JeffersonLabE97-110:2019fsc, JeffersonLabHallAg2p:2022qap], including the EG4 experiment reported here.
The latest EFT calculations by Bernard et al. [Bernard:2012hb] and Alarcón et al. [Alarcon:2020icz] both included the but differ in their expansion method for the (pion)- corrections. In EFT, the general expansion parameter is , where GeV and GeV is the chiral symmetry breaking scale. To explicitly account for the , the nucleon- mass difference GeV must be included in the chiral expansion. Bernard et al. [Bernard:2012hb] treated as a small parameter of the same order as . Alarcón et al. [Alarcon:2020icz] used as an intermediate scale so that is the expansion parameter for the - corrections. This is referred to as the “-expansion scheme”. Additionally, Alarcón et al. included a phenomenological form factor, viz not derived within EFT, that suppresses the -dependence of observables at large . This clearly extends the range of applicability of the Alarcón et al. prediction [Alarcon:2020icz] to higher compared to that of Bernard et al. [Bernard:2012hb], albeit by introducing an extra scale factor.
An important outcome that can be drawn from the results shown in this section is that EFT, although successful in many instances, remains challenged by results from dedicated polarized experiments at low . This includes not only EG4, but also other experiments that used different detectors, methods, and spin observables [JeffersonLabE97-110:2019fsc, E97-110:2021mxm, JeffersonLabHallAg2p:2022qap]. To address this problem, further improvements on EFT would be required, although it is very difficult to upgrade the predictions to the next order. Since EFT is the leading QCD-rooted approach to calculating the effects of the strong force at large distances, this difficulty poses a problem in our pursuit of a complete description of nature. Alternatively or complementarily, it would be helpful to have predictions from other non-perturbative approaches to QCD, such as lattice QCD, the Dyson-Schwinger Equations, or the AdS/CFT methods. Recent progress in calculating doubly-virtual Compton scattering (VVCS) amplitudes in lattice QCD now makes moments involved in sum rules within reach of this method, see, e.g., Refs. [Lee:2023lnx, Wang:2023omf] for recent development. At present, only calculations for spin-independent VVCS amplitudes are available; these calculations are difficult because they involve computing 4-point correlation functions, in contrast to usual Lattice QCD calculations that require less computing-extensive 2- and 3-point correlations functions. Only recently, advances in methods and computer power have allowed such a computation. There is no difficulty in extending these calculations to computing the spin-flip VVCS amplitudes that enter our sum rules. Our results and their mixed level of agreement with EFT predictions are strong incentives to extend these calculations to spin-dependent VVCS amplitudes, allowing lattice QCD predictions to be extensively tested.
Apart from tests of fundamental predictions in the hadronic spin sector, our data will also provide input for calculations of two-photon effects in the hyperfine splitting of the hydrogen (or muonic hydrogen) ground state. The lack of proton spin data at low enough has made this contribution the dominant source of uncertainty in that calculation. The EG4 longitudinal spin data combined with the transverse spin data from the recent JLab Hall A experiment E08-027 [JeffersonLabHallAg2p:2022qap] will help to remedy this situation [Ruth:2024bsl].
VIII Summary and Conclusions
In this archival paper, we present the details of the Jefferson Lab EG4 experiment and our final results for proton and the deuteron spin structure function data for values as low as 0.012 and 0.017 GeV2, respectively. We include results on the proton and the deuteron , , and their moments, reported previously in Refs. [CLAS:2017ozc, CLAS:2021apd], along with details of the simulation and data analysis procedure that led to these results. Our results agree well with previous JLab experiments, but the significantly improved precision of the new data reveal that the world knowledge on low spin structure functions, mostly encapsulated in the parameterization [CLAS:2017qga], must be updated.
We also present new results on the neutron and various moments, extracted by combining the corresponding observable of the proton and deuteron and using a simple deuteron model to correct for Fermi smearing and the deuteron D-state component. Our results on the neutron moments agree with earlier results that used polarized 3He targets.
Extrapolating the lowest EG4 data to provides a check of the GDH sum rule validity, independent from exclusive photoproduction () method. This novel technique using quasi-real photons tests the sum rule with a competitive accuracy compared to real photon experiments. It provides results in excellent agreement with the GDH expectation for the proton, and reasonable agreement for the neutron and the deuteron. By combining the proton with neutron results, isovector or isoscalar moments can be formed, which provide complementary tests of theory predictions at low . We also compared the new results on moments to several phenomenological models, and found a mixed level of success.
Finally, our results test extensively EFT predictions using 8 distinct moments representing diverse -weightings and isospin components. The new data are of high precision and are at values well into domain of EFT. Despite these advantages and the recent improvements on the EFT calculation, there is mixed level of agreement between data and EFT depending on the observable, range, and type of calculations. Clearly, EFT remains challenged by the experimental data at low . We hope other methods, such as lattice QCD, Dyson-Schwinger Equations, and AdS/CFT, will soon provide predictions that can be compared with experimental results at low , advancing the theoretical non-perturbative approaches to the strong interaction.
Acknowledgements.
We would like to acknowledge the outstanding efforts of the staff of the Accelerator, the Target Group, and the Physics Divisions at Jefferson Lab that made this experiment possible. This work was supported in part by the U.S. Department of Energy, the U.S. National Science Foundation, the U.S. Jeffress Memorial Trust; the United Kingdom Science and Technology Facilities Council (STFC), the Italian Istituto Nazionale di Fisica Nucleare; the French Institut National de Physique Nucléaire et de Physique des Particules, the French Centre National de la Recherche Scientifique; and the National Research Foundation of Korea. J. Zhang and X. Zheng are supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract number DE–SC0014434. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177. We are grateful to J.M. Alarcón, A. Kotikov, R. Pasechnik, and D. Volkova for sending us the predictions from their models.Appendix A Table for Results on and
Results for and as shown in Figs. 10 and 12, respectively.
| W | Stat. | Syst. | Stat. | Syst. | ||||
| (GeV2) | (GeV) | |||||||
| 0.012 | 1.11 | 0.0330 | 0.115 | 0.090 | 0.013 | 0.145 | 0.120 | 0.013 |
| 0.012 | 1.13 | 0.0294 | -0.085 | 0.115 | 0.021 | -0.119 | 0.144 | 0.021 |
| 0.012 | 1.15 | 0.0264 | -0.108 | 0.120 | 0.022 | -0.146 | 0.144 | 0.022 |
| 0.012 | 1.17 | 0.0240 | -0.205 | 0.117 | 0.021 | -0.248 | 0.137 | 0.021 |
| 0.012 | 1.19 | 0.0219 | -0.745 | 0.136 | 0.038 | -0.854 | 0.155 | 0.038 |
| 0.012 | 1.21 | 0.0202 | -0.613 | 0.144 | 0.034 | -0.687 | 0.161 | 0.034 |
| 0.012 | 1.23 | 0.0186 | -0.912 | 0.140 | 0.037 | -1.007 | 0.154 | 0.037 |
| 0.012 | 1.25 | 0.0173 | -0.538 | 0.136 | 0.047 | -0.588 | 0.148 | 0.047 |
| 0.012 | 1.27 | 0.0161 | -0.200 | 0.134 | 0.028 | -0.219 | 0.144 | 0.028 |
| 0.012 | 1.29 | 0.0151 | -0.126 | 0.141 | 0.025 | -0.138 | 0.150 | 0.025 |
| 0.012 | 1.31 | 0.0142 | -0.316 | 0.133 | 0.038 | -0.339 | 0.141 | 0.038 |
| 0.012 | 1.33 | 0.0133 | 0.240 | 0.134 | 0.045 | 0.248 | 0.141 | 0.045 |
| 0.012 | 1.35 | 0.0126 | -0.121 | 0.151 | 0.053 | -0.132 | 0.158 | 0.053 |
| 0.012 | 1.37 | 0.0119 | 0.103 | 0.163 | 0.041 | 0.101 | 0.170 | 0.041 |
| 0.012 | 1.39 | 0.0113 | -0.177 | 0.186 | 0.037 | -0.190 | 0.193 | 0.037 |
| 0.012 | 1.41 | 0.0107 | 0.052 | 0.206 | 0.021 | 0.046 | 0.213 | 0.021 |
| 0.012 | 1.43 | 0.0102 | -0.319 | 0.227 | 0.019 | -0.337 | 0.234 | 0.019 |
| 0.012 | 1.45 | 0.0097 | -0.274 | 0.250 | 0.045 | -0.291 | 0.257 | 0.045 |
| 0.012 | 1.47 | 0.0093 | -0.115 | 0.467 | 0.151 | -0.127 | 0.479 | 0.151 |
| 0.0143 | 1.11 | 0.0391 | 0.109 | 0.084 | 0.022 | 0.139 | 0.115 | 0.022 |
| 0.0143 | 1.13 | 0.0348 | 0.056 | 0.083 | 0.021 | 0.057 | 0.107 | 0.021 |
| 0.0143 | 1.15 | 0.0313 | -0.069 | 0.098 | 0.020 | -0.099 | 0.121 | 0.020 |
| 0.0143 | 1.17 | 0.0284 | -0.134 | 0.108 | 0.017 | -0.173 | 0.129 | 0.017 |
| 0.0143 | 1.19 | 0.0260 | -0.651 | 0.120 | 0.028 | -0.761 | 0.140 | 0.028 |
| 0.0143 | 1.21 | 0.0239 | -0.546 | 0.117 | 0.015 | -0.623 | 0.134 | 0.015 |
| 0.0143 | 1.23 | 0.0221 | -0.645 | 0.119 | 0.037 | -0.723 | 0.133 | 0.037 |
| 0.0143 | 1.25 | 0.0205 | -0.524 | 0.115 | 0.032 | -0.581 | 0.127 | 0.032 |
| 0.0143 | 1.27 | 0.0191 | -0.358 | 0.114 | 0.028 | -0.394 | 0.125 | 0.028 |
| 0.0143 | 1.29 | 0.0179 | -0.284 | 0.118 | 0.032 | -0.312 | 0.127 | 0.032 |
| 0.0143 | 1.31 | 0.0168 | -0.041 | 0.102 | 0.033 | -0.049 | 0.109 | 0.033 |
| 0.0143 | 1.33 | 0.0158 | -0.110 | 0.118 | 0.039 | -0.122 | 0.125 | 0.039 |
| 0.0143 | 1.35 | 0.0150 | 0.021 | 0.127 | 0.051 | 0.015 | 0.134 | 0.051 |
| 0.0143 | 1.37 | 0.0141 | -0.086 | 0.130 | 0.059 | -0.096 | 0.136 | 0.059 |
| 0.0143 | 1.39 | 0.0134 | -0.187 | 0.145 | 0.092 | -0.203 | 0.151 | 0.092 |
| 0.0143 | 1.41 | 0.0127 | -0.122 | 0.158 | 0.029 | -0.135 | 0.164 | 0.029 |
| 0.0143 | 1.43 | 0.0121 | -0.059 | 0.191 | 0.026 | -0.071 | 0.198 | 0.026 |
| 0.0143 | 1.45 | 0.0116 | -0.322 | 0.235 | 0.040 | -0.342 | 0.243 | 0.040 |
| 0.0143 | 1.47 | 0.0110 | 0.422 | 0.427 | 0.243 | 0.425 | 0.441 | 0.243 |
| 0.0171 | 1.11 | 0.0463 | 0.126 | 0.072 | 0.054 | 0.172 | 0.105 | 0.054 |
| 0.0171 | 1.13 | 0.0413 | 0.133 | 0.077 | 0.011 | 0.159 | 0.105 | 0.011 |
| 0.0171 | 1.15 | 0.0372 | -0.046 | 0.077 | 0.019 | -0.076 | 0.098 | 0.019 |
| 0.0171 | 1.17 | 0.0338 | -0.209 | 0.075 | 0.033 | -0.270 | 0.092 | 0.033 |
| 0.0171 | 1.19 | 0.0309 | -0.407 | 0.077 | 0.032 | -0.491 | 0.092 | 0.032 |
| 0.0171 | 1.21 | 0.0284 | -0.510 | 0.077 | 0.027 | -0.597 | 0.090 | 0.027 |
| 0.0171 | 1.23 | 0.0263 | -0.590 | 0.076 | 0.028 | -0.675 | 0.087 | 0.028 |
| 0.0171 | 1.25 | 0.0244 | -0.393 | 0.067 | 0.024 | -0.445 | 0.075 | 0.024 |
| 0.0171 | 1.27 | 0.0228 | -0.281 | 0.065 | 0.031 | -0.318 | 0.072 | 0.031 |
| 0.0171 | 1.29 | 0.0213 | -0.057 | 0.065 | 0.029 | -0.068 | 0.071 | 0.029 |
| 0.0171 | 1.31 | 0.0200 | -0.102 | 0.066 | 0.036 | -0.116 | 0.072 | 0.036 |
| 0.0171 | 1.33 | 0.0189 | 0.050 | 0.065 | 0.046 | 0.046 | 0.070 | 0.046 |
| 0.0171 | 1.35 | 0.0178 | -0.084 | 0.072 | 0.053 | -0.096 | 0.077 | 0.053 |
| 0.0171 | 1.37 | 0.0169 | 0.002 | 0.074 | 0.050 | -0.006 | 0.078 | 0.050 |
| 0.0171 | 1.39 | 0.0160 | 0.032 | 0.080 | 0.046 | 0.025 | 0.084 | 0.046 |
| 0.0171 | 1.41 | 0.0152 | -0.213 | 0.086 | 0.023 | -0.234 | 0.091 | 0.023 |
| 0.0171 | 1.43 | 0.0145 | -0.198 | 0.096 | 0.021 | -0.218 | 0.100 | 0.021 |
| 0.0171 | 1.45 | 0.0138 | -0.416 | 0.109 | 0.051 | -0.445 | 0.113 | 0.051 |
| 0.0171 | 1.47 | 0.0132 | -0.129 | 0.130 | 0.044 | -0.146 | 0.135 | 0.044 |
| 0.0171 | 1.49 | 0.0126 | -0.232 | 0.146 | 0.063 | -0.252 | 0.150 | 0.063 |
| 0.0171 | 1.51 | 0.0121 | -0.229 | 0.141 | 0.077 | -0.246 | 0.145 | 0.077 |
| 0.0171 | 1.53 | 0.0116 | -0.109 | 0.152 | 0.051 | -0.120 | 0.156 | 0.051 |
| 0.0171 | 1.55 | 0.0111 | -0.153 | 0.160 | 0.031 | -0.165 | 0.164 | 0.031 |
| 0.0171 | 1.57 | 0.0107 | -0.094 | 0.199 | 0.060 | -0.102 | 0.204 | 0.060 |
| 0.0171 | 1.59 | 0.0103 | -0.339 | 0.198 | 0.065 | -0.351 | 0.203 | 0.065 |
| 0.0171 | 1.61 | 0.0099 | -0.327 | 0.224 | 0.081 | -0.337 | 0.229 | 0.081 |
| 0.0171 | 1.63 | 0.0095 | -0.164 | 0.297 | 0.239 | -0.170 | 0.303 | 0.239 |
| 0.0204 | 1.11 | 0.0549 | 0.093 | 0.047 | 0.021 | 0.129 | 0.071 | 0.021 |
| 0.0204 | 1.13 | 0.0490 | -0.080 | 0.047 | 0.025 | -0.133 | 0.066 | 0.025 |
| 0.0204 | 1.15 | 0.0442 | -0.062 | 0.052 | 0.024 | -0.102 | 0.070 | 0.024 |
| 0.0204 | 1.17 | 0.0401 | -0.247 | 0.055 | 0.034 | -0.326 | 0.070 | 0.034 |
| 0.0204 | 1.19 | 0.0367 | -0.433 | 0.057 | 0.032 | -0.538 | 0.069 | 0.032 |
| 0.0204 | 1.21 | 0.0338 | -0.591 | 0.058 | 0.022 | -0.707 | 0.069 | 0.022 |
| 0.0204 | 1.23 | 0.0313 | -0.552 | 0.058 | 0.036 | -0.646 | 0.068 | 0.036 |
| 0.0204 | 1.25 | 0.0291 | -0.337 | 0.058 | 0.036 | -0.390 | 0.067 | 0.036 |
| 0.0204 | 1.27 | 0.0271 | -0.282 | 0.059 | 0.032 | -0.325 | 0.067 | 0.032 |
| 0.0204 | 1.29 | 0.0254 | -0.127 | 0.060 | 0.027 | -0.149 | 0.067 | 0.027 |
| 0.0204 | 1.31 | 0.0239 | 0.006 | 0.060 | 0.034 | -0.001 | 0.066 | 0.034 |
| 0.0204 | 1.33 | 0.0225 | -0.026 | 0.063 | 0.060 | -0.036 | 0.068 | 0.060 |
| 0.0204 | 1.35 | 0.0212 | -0.060 | 0.067 | 0.061 | -0.074 | 0.072 | 0.061 |
| 0.0204 | 1.37 | 0.0201 | -0.137 | 0.071 | 0.051 | -0.156 | 0.076 | 0.051 |
| 0.0204 | 1.39 | 0.0191 | -0.001 | 0.077 | 0.026 | -0.011 | 0.082 | 0.026 |
| 0.0204 | 1.41 | 0.0181 | -0.106 | 0.079 | 0.033 | -0.123 | 0.084 | 0.033 |
| 0.0204 | 1.43 | 0.0172 | -0.237 | 0.088 | 0.014 | -0.262 | 0.093 | 0.014 |
| 0.0204 | 1.45 | 0.0164 | -0.194 | 0.093 | 0.027 | -0.217 | 0.097 | 0.027 |
| 0.0204 | 1.47 | 0.0157 | -0.123 | 0.108 | 0.053 | -0.143 | 0.113 | 0.053 |
| 0.0204 | 1.49 | 0.0150 | -0.205 | 0.120 | 0.072 | -0.228 | 0.125 | 0.072 |
| 0.0204 | 1.51 | 0.0144 | -0.313 | 0.128 | 0.082 | -0.335 | 0.133 | 0.082 |
| 0.0204 | 1.53 | 0.0138 | -0.254 | 0.141 | 0.064 | -0.273 | 0.146 | 0.064 |
| 0.0204 | 1.55 | 0.0132 | -0.314 | 0.150 | 0.025 | -0.332 | 0.154 | 0.025 |
| 0.0204 | 1.57 | 0.0127 | -0.086 | 0.161 | 0.053 | -0.096 | 0.165 | 0.053 |
| 0.0204 | 1.59 | 0.0122 | -0.099 | 0.171 | 0.059 | -0.107 | 0.175 | 0.059 |
| 0.0204 | 1.61 | 0.0118 | 0.152 | 0.210 | 0.057 | 0.152 | 0.215 | 0.057 |
| 0.0204 | 1.63 | 0.0114 | -0.523 | 0.276 | 0.210 | -0.538 | 0.282 | 0.210 |
| 0.0244 | 1.11 | 0.0648 | 0.005 | 0.034 | 0.009 | -0.011 | 0.054 | 0.009 |
| 0.0244 | 1.13 | 0.0579 | 0.059 | 0.037 | 0.014 | 0.063 | 0.054 | 0.014 |
| 0.0244 | 1.15 | 0.0522 | -0.094 | 0.044 | 0.023 | -0.155 | 0.062 | 0.023 |
| 0.0244 | 1.17 | 0.0475 | -0.212 | 0.048 | 0.033 | -0.294 | 0.063 | 0.033 |
| 0.0244 | 1.19 | 0.0435 | -0.404 | 0.051 | 0.026 | -0.519 | 0.065 | 0.026 |
| 0.0244 | 1.21 | 0.0401 | -0.565 | 0.052 | 0.018 | -0.694 | 0.064 | 0.018 |
| 0.0244 | 1.23 | 0.0371 | -0.595 | 0.052 | 0.034 | -0.716 | 0.063 | 0.034 |
| 0.0244 | 1.25 | 0.0345 | -0.451 | 0.054 | 0.044 | -0.534 | 0.063 | 0.044 |
| 0.0244 | 1.27 | 0.0322 | -0.318 | 0.054 | 0.030 | -0.373 | 0.062 | 0.030 |
| 0.0244 | 1.29 | 0.0301 | -0.145 | 0.055 | 0.034 | -0.172 | 0.062 | 0.034 |
| 0.0244 | 1.31 | 0.0283 | -0.105 | 0.055 | 0.033 | -0.126 | 0.062 | 0.033 |
| 0.0244 | 1.33 | 0.0267 | 0.010 | 0.058 | 0.042 | 0.001 | 0.064 | 0.042 |
| 0.0244 | 1.35 | 0.0252 | -0.095 | 0.059 | 0.053 | -0.114 | 0.064 | 0.053 |
| 0.0244 | 1.37 | 0.0239 | -0.029 | 0.057 | 0.051 | -0.043 | 0.062 | 0.051 |
| 0.0244 | 1.39 | 0.0226 | -0.024 | 0.060 | 0.042 | -0.039 | 0.064 | 0.042 |
| 0.0244 | 1.41 | 0.0215 | -0.195 | 0.067 | 0.026 | -0.222 | 0.072 | 0.026 |
| 0.0244 | 1.43 | 0.0205 | -0.191 | 0.073 | 0.033 | -0.219 | 0.077 | 0.033 |
| 0.0244 | 1.45 | 0.0195 | -0.216 | 0.080 | 0.017 | -0.245 | 0.085 | 0.017 |
| 0.0244 | 1.47 | 0.0187 | -0.102 | 0.103 | 0.035 | -0.126 | 0.108 | 0.035 |
| 0.0244 | 1.49 | 0.0179 | -0.274 | 0.112 | 0.069 | -0.303 | 0.117 | 0.069 |
| 0.0244 | 1.51 | 0.0171 | -0.120 | 0.122 | 0.056 | -0.139 | 0.127 | 0.056 |
| 0.0244 | 1.53 | 0.0164 | -0.309 | 0.133 | 0.039 | -0.333 | 0.139 | 0.039 |
| 0.0244 | 1.55 | 0.0158 | -0.284 | 0.139 | 0.027 | -0.305 | 0.144 | 0.027 |
| 0.0244 | 1.57 | 0.0151 | -0.089 | 0.141 | 0.077 | -0.101 | 0.146 | 0.077 |
| 0.0244 | 1.59 | 0.0146 | 0.018 | 0.151 | 0.053 | 0.012 | 0.155 | 0.053 |
| 0.0244 | 1.61 | 0.0140 | -0.193 | 0.173 | 0.044 | -0.204 | 0.178 | 0.044 |
| 0.0244 | 1.63 | 0.0135 | -0.243 | 0.236 | 0.203 | -0.253 | 0.242 | 0.203 |
| 0.029 | 1.11 | 0.0763 | 0.075 | 0.032 | 0.029 | 0.107 | 0.054 | 0.029 |
| 0.029 | 1.13 | 0.0682 | 0.042 | 0.036 | 0.010 | 0.040 | 0.056 | 0.010 |
| 0.029 | 1.15 | 0.0616 | -0.057 | 0.038 | 0.017 | -0.109 | 0.056 | 0.017 |
| 0.029 | 1.17 | 0.0561 | -0.260 | 0.042 | 0.035 | -0.378 | 0.058 | 0.035 |
| 0.029 | 1.19 | 0.0514 | -0.352 | 0.044 | 0.028 | -0.473 | 0.058 | 0.028 |
| 0.029 | 1.21 | 0.0474 | -0.454 | 0.045 | 0.020 | -0.579 | 0.057 | 0.020 |
| 0.029 | 1.23 | 0.0439 | -0.514 | 0.045 | 0.032 | -0.637 | 0.056 | 0.032 |
| 0.029 | 1.25 | 0.0408 | -0.385 | 0.046 | 0.033 | -0.471 | 0.056 | 0.033 |
| 0.029 | 1.27 | 0.0381 | -0.212 | 0.047 | 0.025 | -0.259 | 0.055 | 0.025 |
| 0.029 | 1.29 | 0.0357 | -0.094 | 0.045 | 0.028 | -0.119 | 0.053 | 0.028 |
| 0.029 | 1.31 | 0.0336 | -0.041 | 0.046 | 0.029 | -0.058 | 0.052 | 0.029 |
| 0.029 | 1.33 | 0.0317 | 0.012 | 0.045 | 0.039 | 0.001 | 0.051 | 0.039 |
| 0.029 | 1.35 | 0.0299 | -0.047 | 0.046 | 0.046 | -0.065 | 0.051 | 0.046 |
| 0.029 | 1.37 | 0.0283 | -0.057 | 0.049 | 0.045 | -0.076 | 0.053 | 0.045 |
| 0.029 | 1.39 | 0.0269 | -0.026 | 0.053 | 0.030 | -0.044 | 0.057 | 0.030 |
| 0.029 | 1.41 | 0.0255 | -0.064 | 0.056 | 0.028 | -0.086 | 0.061 | 0.028 |
| 0.029 | 1.43 | 0.0243 | -0.042 | 0.062 | 0.032 | -0.064 | 0.067 | 0.032 |
| 0.029 | 1.45 | 0.0232 | -0.151 | 0.068 | 0.023 | -0.182 | 0.073 | 0.023 |
| 0.029 | 1.47 | 0.0222 | -0.129 | 0.100 | 0.021 | -0.159 | 0.106 | 0.021 |
| 0.029 | 1.49 | 0.0212 | -0.107 | 0.108 | 0.042 | -0.134 | 0.114 | 0.042 |
| 0.029 | 1.51 | 0.0203 | -0.174 | 0.118 | 0.063 | -0.200 | 0.124 | 0.063 |
| 0.029 | 1.53 | 0.0195 | -0.097 | 0.117 | 0.018 | -0.115 | 0.122 | 0.018 |
| 0.029 | 1.55 | 0.0187 | -0.223 | 0.114 | 0.029 | -0.245 | 0.119 | 0.029 |
| 0.029 | 1.57 | 0.0180 | 0.104 | 0.126 | 0.042 | 0.097 | 0.131 | 0.042 |
| 0.029 | 1.59 | 0.0173 | 0.200 | 0.148 | 0.047 | 0.199 | 0.154 | 0.047 |
| 0.029 | 1.61 | 0.0167 | -0.082 | 0.163 | 0.034 | -0.092 | 0.168 | 0.034 |
| 0.029 | 1.63 | 0.0161 | -0.083 | 0.209 | 0.119 | -0.091 | 0.215 | 0.119 |
| 0.0347 | 1.11 | 0.0897 | 0.004 | 0.026 | 0.019 | -0.018 | 0.048 | 0.019 |
| 0.0347 | 1.13 | 0.0804 | 0.057 | 0.029 | 0.014 | 0.063 | 0.049 | 0.014 |
| 0.0347 | 1.15 | 0.0727 | -0.051 | 0.033 | 0.020 | -0.109 | 0.050 | 0.020 |
| 0.0347 | 1.17 | 0.0663 | -0.206 | 0.036 | 0.029 | -0.318 | 0.052 | 0.029 |
| 0.0347 | 1.19 | 0.0608 | -0.350 | 0.040 | 0.028 | -0.490 | 0.055 | 0.028 |
| 0.0347 | 1.21 | 0.0561 | -0.578 | 0.042 | 0.031 | -0.767 | 0.056 | 0.031 |
| 0.0347 | 1.23 | 0.0520 | -0.509 | 0.044 | 0.031 | -0.655 | 0.056 | 0.031 |
| 0.0347 | 1.25 | 0.0484 | -0.347 | 0.043 | 0.034 | -0.439 | 0.054 | 0.034 |
| 0.0347 | 1.27 | 0.0452 | -0.215 | 0.042 | 0.023 | -0.271 | 0.051 | 0.023 |
| 0.0347 | 1.29 | 0.0424 | -0.132 | 0.040 | 0.023 | -0.170 | 0.048 | 0.023 |
| 0.0347 | 1.31 | 0.0398 | -0.002 | 0.040 | 0.029 | -0.017 | 0.047 | 0.029 |
| 0.0347 | 1.33 | 0.0375 | -0.054 | 0.040 | 0.036 | -0.077 | 0.046 | 0.036 |
| 0.0347 | 1.35 | 0.0355 | -0.042 | 0.042 | 0.038 | -0.063 | 0.048 | 0.038 |
| 0.0347 | 1.37 | 0.0336 | 0.001 | 0.046 | 0.036 | -0.016 | 0.052 | 0.036 |
| 0.0347 | 1.39 | 0.0319 | -0.083 | 0.050 | 0.020 | -0.110 | 0.055 | 0.020 |
| 0.0347 | 1.41 | 0.0303 | -0.137 | 0.051 | 0.014 | -0.171 | 0.056 | 0.014 |
| 0.0347 | 1.43 | 0.0289 | -0.163 | 0.054 | 0.016 | -0.201 | 0.059 | 0.016 |
| 0.0347 | 1.45 | 0.0276 | -0.099 | 0.057 | 0.024 | -0.132 | 0.061 | 0.024 |
| 0.0347 | 1.47 | 0.0264 | -0.335 | 0.088 | 0.035 | -0.385 | 0.095 | 0.035 |
| 0.0347 | 1.49 | 0.0252 | -0.265 | 0.094 | 0.051 | -0.307 | 0.100 | 0.051 |
| 0.0347 | 1.51 | 0.0242 | -0.275 | 0.090 | 0.042 | -0.311 | 0.096 | 0.042 |
| 0.0347 | 1.53 | 0.0232 | -0.079 | 0.088 | 0.014 | -0.101 | 0.093 | 0.014 |
| 0.0347 | 1.55 | 0.0223 | -0.157 | 0.095 | 0.023 | -0.180 | 0.100 | 0.023 |
| 0.0347 | 1.57 | 0.0214 | -0.046 | 0.103 | 0.032 | -0.061 | 0.108 | 0.032 |
| 0.0347 | 1.59 | 0.0206 | -0.035 | 0.108 | 0.033 | -0.047 | 0.113 | 0.033 |
| 0.0347 | 1.61 | 0.0199 | -0.063 | 0.127 | 0.031 | -0.073 | 0.132 | 0.031 |
| 0.0347 | 1.63 | 0.0191 | -0.322 | 0.207 | 0.235 | -0.341 | 0.216 | 0.235 |
| 0.0414 | 1.11 | 0.1053 | 0.056 | 0.023 | 0.010 | 0.083 | 0.045 | 0.010 |
| 0.0414 | 1.13 | 0.0945 | 0.037 | 0.026 | 0.008 | 0.033 | 0.046 | 0.008 |
| 0.0414 | 1.15 | 0.0856 | -0.055 | 0.032 | 0.020 | -0.120 | 0.053 | 0.020 |
| 0.0414 | 1.17 | 0.0781 | -0.114 | 0.036 | 0.023 | -0.193 | 0.054 | 0.023 |
| 0.0414 | 1.19 | 0.0717 | -0.395 | 0.040 | 0.038 | -0.576 | 0.057 | 0.038 |
| 0.0414 | 1.21 | 0.0662 | -0.566 | 0.042 | 0.022 | -0.777 | 0.058 | 0.022 |
| 0.0414 | 1.23 | 0.0614 | -0.429 | 0.042 | 0.025 | -0.570 | 0.055 | 0.025 |
| 0.0414 | 1.25 | 0.0572 | -0.337 | 0.041 | 0.030 | -0.440 | 0.053 | 0.030 |
| 0.0414 | 1.27 | 0.0535 | -0.223 | 0.039 | 0.025 | -0.291 | 0.048 | 0.025 |
| 0.0414 | 1.29 | 0.0502 | -0.133 | 0.039 | 0.025 | -0.177 | 0.047 | 0.025 |
| 0.0414 | 1.31 | 0.0472 | -0.073 | 0.038 | 0.028 | -0.103 | 0.045 | 0.028 |
| 0.0414 | 1.33 | 0.0445 | -0.011 | 0.038 | 0.034 | -0.030 | 0.044 | 0.034 |
| 0.0414 | 1.35 | 0.0421 | -0.026 | 0.039 | 0.035 | -0.048 | 0.045 | 0.035 |
| 0.0414 | 1.37 | 0.0399 | -0.003 | 0.042 | 0.036 | -0.023 | 0.047 | 0.036 |
| 0.0414 | 1.39 | 0.0379 | -0.035 | 0.042 | 0.021 | -0.062 | 0.048 | 0.021 |
| 0.0414 | 1.41 | 0.0360 | -0.090 | 0.046 | 0.022 | -0.124 | 0.051 | 0.022 |
| 0.0414 | 1.43 | 0.0343 | -0.162 | 0.051 | 0.025 | -0.205 | 0.056 | 0.025 |
| 0.0414 | 1.45 | 0.0328 | -0.103 | 0.060 | 0.019 | -0.141 | 0.065 | 0.019 |
| 0.0414 | 1.47 | 0.0313 | -0.200 | 0.074 | 0.020 | -0.247 | 0.080 | 0.020 |
| 0.0414 | 1.49 | 0.0300 | -0.288 | 0.075 | 0.041 | -0.338 | 0.081 | 0.041 |
| 0.0414 | 1.51 | 0.0287 | -0.199 | 0.077 | 0.034 | -0.236 | 0.082 | 0.034 |
| 0.0414 | 1.53 | 0.0276 | -0.041 | 0.078 | 0.021 | -0.064 | 0.083 | 0.021 |
| 0.0414 | 1.55 | 0.0265 | -0.073 | 0.085 | 0.018 | -0.096 | 0.091 | 0.018 |
| 0.0414 | 1.57 | 0.0255 | -0.116 | 0.094 | 0.035 | -0.138 | 0.099 | 0.035 |
| 0.0414 | 1.59 | 0.0245 | -0.175 | 0.101 | 0.039 | -0.197 | 0.106 | 0.039 |
| 0.0414 | 1.61 | 0.0236 | 0.018 | 0.111 | 0.040 | 0.009 | 0.116 | 0.040 |
| 0.0414 | 1.63 | 0.0228 | 0.102 | 0.130 | 0.065 | 0.099 | 0.136 | 0.065 |
| 0.0414 | 1.65 | 0.0220 | -0.339 | 0.313 | 0.022 | -0.358 | 0.326 | 0.022 |
| 0.0414 | 1.67 | 0.0212 | -0.275 | 0.352 | 0.020 | -0.290 | 0.365 | 0.020 |
| 0.0414 | 1.69 | 0.0205 | -0.276 | 0.329 | 0.018 | -0.289 | 0.341 | 0.018 |
| 0.0414 | 1.71 | 0.0199 | -0.214 | 0.342 | 0.017 | -0.223 | 0.353 | 0.017 |
| 0.0414 | 1.73 | 0.0192 | -0.107 | 0.386 | 0.017 | -0.111 | 0.398 | 0.017 |
| 0.0414 | 1.75 | 0.0186 | -0.188 | 0.352 | 0.016 | -0.193 | 0.362 | 0.016 |
| 0.0414 | 1.77 | 0.0180 | -0.317 | 0.388 | 0.023 | -0.325 | 0.398 | 0.023 |
| 0.0414 | 1.79 | 0.0175 | 0.036 | 0.398 | 0.013 | 0.038 | 0.408 | 0.013 |
| 0.0414 | 1.81 | 0.0170 | 0.274 | 0.457 | 0.024 | 0.282 | 0.469 | 0.024 |
| 0.0414 | 1.83 | 0.0165 | -0.208 | 0.515 | 0.018 | -0.211 | 0.527 | 0.018 |
| 0.0414 | 1.85 | 0.0160 | 0.622 | 0.569 | 0.040 | 0.638 | 0.582 | 0.040 |
| 0.0414 | 1.87 | 0.0156 | 0.083 | 0.669 | 0.035 | 0.086 | 0.683 | 0.035 |
| 0.0414 | 1.89 | 0.0151 | 0.672 | 0.760 | 0.096 | 0.687 | 0.775 | 0.096 |
| 0.0414 | 1.91 | 0.0147 | -0.107 | 0.773 | 0.042 | -0.107 | 0.787 | 0.042 |
| 0.0414 | 1.93 | 0.0143 | 0.051 | 1.111 | 0.137 | 0.054 | 1.131 | 0.137 |
| 0.0494 | 1.11 | 0.1232 | 0.052 | 0.021 | 0.010 | 0.083 | 0.044 | 0.010 |
| 0.0494 | 1.13 | 0.1108 | 0.033 | 0.028 | 0.011 | 0.027 | 0.052 | 0.011 |
| 0.0494 | 1.15 | 0.1005 | -0.014 | 0.031 | 0.016 | -0.057 | 0.053 | 0.016 |
| 0.0494 | 1.17 | 0.0918 | -0.200 | 0.034 | 0.025 | -0.342 | 0.054 | 0.025 |
| 0.0494 | 1.19 | 0.0844 | -0.354 | 0.037 | 0.028 | -0.540 | 0.056 | 0.028 |
| 0.0494 | 1.21 | 0.0780 | -0.528 | 0.039 | 0.018 | -0.755 | 0.055 | 0.018 |
| 0.0494 | 1.23 | 0.0725 | -0.412 | 0.039 | 0.025 | -0.568 | 0.053 | 0.025 |
| 0.0494 | 1.25 | 0.0675 | -0.329 | 0.037 | 0.027 | -0.446 | 0.050 | 0.027 |
| 0.0494 | 1.27 | 0.0632 | -0.183 | 0.036 | 0.023 | -0.251 | 0.047 | 0.023 |
| 0.0494 | 1.29 | 0.0593 | -0.112 | 0.035 | 0.023 | -0.159 | 0.044 | 0.023 |
| 0.0494 | 1.31 | 0.0558 | -0.033 | 0.034 | 0.025 | -0.060 | 0.042 | 0.025 |
| 0.0494 | 1.33 | 0.0527 | -0.019 | 0.035 | 0.030 | -0.043 | 0.041 | 0.030 |
| 0.0494 | 1.35 | 0.0498 | -0.006 | 0.036 | 0.030 | -0.029 | 0.043 | 0.030 |
| 0.0494 | 1.37 | 0.0472 | -0.060 | 0.038 | 0.030 | -0.093 | 0.044 | 0.030 |
| 0.0494 | 1.39 | 0.0449 | 0.049 | 0.039 | 0.019 | 0.030 | 0.045 | 0.019 |
| 0.0494 | 1.41 | 0.0427 | 0.002 | 0.042 | 0.019 | -0.026 | 0.048 | 0.019 |
| 0.0494 | 1.43 | 0.0407 | -0.078 | 0.047 | 0.010 | -0.118 | 0.052 | 0.010 |
| 0.0494 | 1.45 | 0.0389 | -0.013 | 0.050 | 0.019 | -0.048 | 0.055 | 0.019 |
| 0.0494 | 1.47 | 0.0372 | -0.138 | 0.064 | 0.032 | -0.188 | 0.070 | 0.032 |
| 0.0494 | 1.49 | 0.0356 | -0.047 | 0.068 | 0.022 | -0.084 | 0.074 | 0.022 |
| 0.0494 | 1.51 | 0.0341 | -0.173 | 0.071 | 0.030 | -0.215 | 0.076 | 0.030 |
| 0.0494 | 1.53 | 0.0327 | -0.035 | 0.071 | 0.016 | -0.062 | 0.077 | 0.016 |
| 0.0494 | 1.55 | 0.0314 | -0.143 | 0.077 | 0.024 | -0.175 | 0.082 | 0.024 |
| 0.0494 | 1.57 | 0.0302 | -0.096 | 0.082 | 0.021 | -0.121 | 0.088 | 0.021 |
| 0.0494 | 1.59 | 0.0291 | -0.063 | 0.087 | 0.021 | -0.082 | 0.092 | 0.021 |
| 0.0494 | 1.61 | 0.0281 | -0.091 | 0.093 | 0.052 | -0.109 | 0.098 | 0.052 |
| 0.0494 | 1.63 | 0.0271 | -0.099 | 0.140 | 0.087 | -0.113 | 0.147 | 0.087 |
| 0.0494 | 1.65 | 0.0261 | 0.264 | 0.306 | 0.024 | 0.269 | 0.321 | 0.024 |
| 0.0494 | 1.67 | 0.0252 | -1.119 | 0.309 | 0.056 | -1.175 | 0.323 | 0.056 |
| 0.0494 | 1.69 | 0.0244 | -0.195 | 0.313 | 0.015 | -0.208 | 0.326 | 0.015 |
| 0.0494 | 1.71 | 0.0236 | -0.780 | 0.301 | 0.059 | -0.813 | 0.313 | 0.059 |
| 0.0494 | 1.73 | 0.0229 | 0.024 | 0.312 | 0.016 | 0.024 | 0.323 | 0.016 |
| 0.0494 | 1.75 | 0.0221 | 0.054 | 0.341 | 0.024 | 0.056 | 0.352 | 0.024 |
| 0.0494 | 1.77 | 0.0215 | -0.517 | 0.383 | 0.044 | -0.534 | 0.396 | 0.044 |
| 0.0494 | 1.79 | 0.0208 | -0.804 | 0.403 | 0.060 | -0.828 | 0.415 | 0.060 |
| 0.0494 | 1.81 | 0.0202 | -0.137 | 0.434 | 0.042 | -0.139 | 0.446 | 0.042 |
| 0.0494 | 1.83 | 0.0196 | -0.174 | 0.433 | 0.013 | -0.177 | 0.445 | 0.013 |
| 0.0494 | 1.85 | 0.0191 | -0.249 | 0.515 | 0.027 | -0.253 | 0.528 | 0.027 |
| 0.0494 | 1.87 | 0.0185 | 0.408 | 0.551 | 0.052 | 0.419 | 0.564 | 0.052 |
| 0.0494 | 1.89 | 0.0180 | 1.063 | 0.558 | 0.126 | 1.090 | 0.571 | 0.126 |
| 0.0494 | 1.91 | 0.0175 | -0.446 | 0.582 | 0.049 | -0.453 | 0.595 | 0.049 |
| 0.0494 | 1.93 | 0.0171 | 0.082 | 1.031 | 0.117 | 0.086 | 1.052 | 0.117 |
| 0.059 | 1.11 | 0.1437 | 0.039 | 0.020 | 0.007 | 0.060 | 0.045 | 0.007 |
| 0.059 | 1.13 | 0.1296 | 0.009 | 0.025 | 0.014 | -0.020 | 0.051 | 0.014 |
| 0.059 | 1.15 | 0.1178 | -0.035 | 0.030 | 0.016 | -0.099 | 0.055 | 0.016 |
| 0.059 | 1.17 | 0.1078 | -0.204 | 0.033 | 0.031 | -0.369 | 0.055 | 0.031 |
| 0.059 | 1.19 | 0.0992 | -0.355 | 0.034 | 0.029 | -0.571 | 0.055 | 0.029 |
| 0.059 | 1.21 | 0.0918 | -0.533 | 0.038 | 0.019 | -0.801 | 0.057 | 0.019 |
| 0.059 | 1.23 | 0.0853 | -0.432 | 0.037 | 0.030 | -0.624 | 0.054 | 0.030 |
| 0.059 | 1.25 | 0.0796 | -0.286 | 0.036 | 0.026 | -0.407 | 0.050 | 0.026 |
| 0.059 | 1.27 | 0.0746 | -0.184 | 0.034 | 0.027 | -0.264 | 0.046 | 0.027 |
| 0.059 | 1.29 | 0.0700 | -0.079 | 0.033 | 0.020 | -0.124 | 0.042 | 0.020 |
| 0.059 | 1.31 | 0.0660 | -0.068 | 0.032 | 0.023 | -0.108 | 0.040 | 0.023 |
| 0.059 | 1.33 | 0.0623 | -0.005 | 0.032 | 0.024 | -0.031 | 0.039 | 0.024 |
| 0.059 | 1.35 | 0.0590 | 0.003 | 0.034 | 0.025 | -0.022 | 0.042 | 0.025 |
| 0.059 | 1.37 | 0.0559 | -0.039 | 0.035 | 0.023 | -0.075 | 0.042 | 0.023 |
| 0.059 | 1.39 | 0.0531 | -0.043 | 0.036 | 0.018 | -0.080 | 0.042 | 0.018 |
| 0.059 | 1.41 | 0.0506 | 0.037 | 0.037 | 0.027 | 0.009 | 0.043 | 0.027 |
| 0.059 | 1.43 | 0.0482 | -0.072 | 0.040 | 0.011 | -0.119 | 0.046 | 0.011 |
| 0.059 | 1.45 | 0.0461 | -0.136 | 0.049 | 0.020 | -0.194 | 0.055 | 0.020 |
| 0.059 | 1.47 | 0.0441 | -0.091 | 0.060 | 0.018 | -0.143 | 0.067 | 0.018 |
| 0.059 | 1.49 | 0.0422 | -0.003 | 0.062 | 0.021 | -0.043 | 0.069 | 0.021 |
| 0.059 | 1.51 | 0.0405 | -0.076 | 0.065 | 0.024 | -0.115 | 0.071 | 0.024 |
| 0.059 | 1.53 | 0.0388 | -0.062 | 0.068 | 0.016 | -0.096 | 0.075 | 0.016 |
| 0.059 | 1.55 | 0.0373 | -0.166 | 0.070 | 0.019 | -0.206 | 0.075 | 0.019 |
| 0.059 | 1.57 | 0.0359 | -0.074 | 0.072 | 0.019 | -0.102 | 0.078 | 0.019 |
| 0.059 | 1.59 | 0.0346 | -0.014 | 0.077 | 0.050 | -0.033 | 0.082 | 0.050 |
| 0.059 | 1.61 | 0.0333 | -0.026 | 0.080 | 0.019 | -0.043 | 0.085 | 0.019 |
| 0.059 | 1.63 | 0.0322 | -0.042 | 0.106 | 0.038 | -0.057 | 0.113 | 0.038 |
| 0.059 | 1.65 | 0.0310 | -0.085 | 0.247 | 0.010 | -0.099 | 0.262 | 0.010 |
| 0.059 | 1.67 | 0.0300 | -0.228 | 0.266 | 0.011 | -0.247 | 0.281 | 0.011 |
| 0.059 | 1.69 | 0.0290 | -0.407 | 0.257 | 0.023 | -0.432 | 0.269 | 0.023 |
| 0.059 | 1.71 | 0.0281 | -0.081 | 0.295 | 0.013 | -0.088 | 0.309 | 0.013 |
| 0.059 | 1.73 | 0.0272 | -0.095 | 0.317 | 0.016 | -0.101 | 0.331 | 0.016 |
| 0.059 | 1.75 | 0.0263 | 0.150 | 0.292 | 0.021 | 0.156 | 0.304 | 0.021 |
| 0.059 | 1.77 | 0.0255 | -0.218 | 0.304 | 0.030 | -0.227 | 0.316 | 0.030 |
| 0.059 | 1.79 | 0.0248 | 0.096 | 0.310 | 0.027 | 0.101 | 0.321 | 0.027 |
| 0.059 | 1.81 | 0.0240 | 0.469 | 0.353 | 0.050 | 0.486 | 0.365 | 0.050 |
| 0.059 | 1.83 | 0.0234 | -0.302 | 0.362 | 0.019 | -0.310 | 0.374 | 0.019 |
| 0.059 | 1.85 | 0.0227 | 0.056 | 0.402 | 0.026 | 0.059 | 0.415 | 0.026 |
| 0.059 | 1.87 | 0.0221 | -0.558 | 0.435 | 0.035 | -0.572 | 0.447 | 0.035 |
| 0.059 | 1.89 | 0.0215 | -0.323 | 0.435 | 0.052 | -0.330 | 0.447 | 0.052 |
| 0.059 | 1.91 | 0.0209 | -0.288 | 0.468 | 0.036 | -0.293 | 0.481 | 0.036 |
| 0.059 | 1.93 | 0.0203 | -0.319 | 0.839 | 0.295 | -0.325 | 0.861 | 0.295 |
| 0.0705 | 1.11 | 0.1669 | 0.005 | 0.022 | 0.014 | -0.018 | 0.051 | 0.014 |
| 0.0705 | 1.13 | 0.1509 | -0.041 | 0.023 | 0.015 | -0.126 | 0.050 | 0.015 |
| 0.0705 | 1.15 | 0.1375 | -0.088 | 0.028 | 0.021 | -0.208 | 0.053 | 0.021 |
| 0.0705 | 1.17 | 0.1261 | -0.182 | 0.031 | 0.027 | -0.350 | 0.055 | 0.027 |
| 0.0705 | 1.19 | 0.1163 | -0.344 | 0.034 | 0.029 | -0.581 | 0.056 | 0.029 |
| 0.0705 | 1.21 | 0.1077 | -0.465 | 0.036 | 0.021 | -0.730 | 0.057 | 0.021 |
| 0.0705 | 1.23 | 0.1002 | -0.439 | 0.034 | 0.025 | -0.662 | 0.051 | 0.025 |
| 0.0705 | 1.25 | 0.0936 | -0.273 | 0.033 | 0.032 | -0.405 | 0.047 | 0.032 |
| 0.0705 | 1.27 | 0.0878 | -0.215 | 0.032 | 0.024 | -0.319 | 0.045 | 0.024 |
| 0.0705 | 1.29 | 0.0825 | -0.082 | 0.033 | 0.021 | -0.134 | 0.044 | 0.021 |
| 0.0705 | 1.31 | 0.0778 | -0.080 | 0.032 | 0.019 | -0.130 | 0.041 | 0.019 |
| 0.0705 | 1.33 | 0.0735 | -0.016 | 0.030 | 0.019 | -0.048 | 0.038 | 0.019 |
| 0.0705 | 1.35 | 0.0696 | 0.028 | 0.029 | 0.026 | 0.005 | 0.037 | 0.026 |
| 0.0705 | 1.37 | 0.0661 | -0.008 | 0.032 | 0.017 | -0.042 | 0.039 | 0.017 |
| 0.0705 | 1.39 | 0.0628 | -0.007 | 0.034 | 0.013 | -0.044 | 0.041 | 0.013 |
| 0.0705 | 1.41 | 0.0598 | -0.008 | 0.037 | 0.014 | -0.048 | 0.044 | 0.014 |
| 0.0705 | 1.43 | 0.0571 | -0.033 | 0.042 | 0.013 | -0.081 | 0.049 | 0.013 |
| 0.0705 | 1.45 | 0.0545 | -0.047 | 0.049 | 0.013 | -0.099 | 0.057 | 0.013 |
| 0.0705 | 1.47 | 0.0522 | 0.048 | 0.053 | 0.020 | 0.006 | 0.061 | 0.020 |
| 0.0705 | 1.49 | 0.0500 | 0.027 | 0.059 | 0.022 | -0.016 | 0.066 | 0.022 |
| 0.0705 | 1.51 | 0.0479 | -0.059 | 0.061 | 0.016 | -0.103 | 0.068 | 0.016 |
| 0.0705 | 1.53 | 0.0460 | -0.021 | 0.062 | 0.015 | -0.057 | 0.069 | 0.015 |
| 0.0705 | 1.55 | 0.0443 | 0.009 | 0.065 | 0.024 | -0.020 | 0.071 | 0.024 |
| 0.0705 | 1.57 | 0.0426 | 0.018 | 0.065 | 0.017 | -0.007 | 0.070 | 0.017 |
| 0.0705 | 1.59 | 0.0410 | 0.065 | 0.066 | 0.023 | 0.049 | 0.071 | 0.023 |
| 0.0705 | 1.61 | 0.0395 | -0.093 | 0.072 | 0.010 | -0.119 | 0.078 | 0.010 |
| 0.0705 | 1.63 | 0.0382 | -0.296 | 0.219 | 0.014 | -0.332 | 0.235 | 0.014 |
| 0.0705 | 1.65 | 0.0369 | -0.399 | 0.246 | 0.022 | -0.438 | 0.263 | 0.022 |
| 0.0705 | 1.67 | 0.0356 | -0.649 | 0.259 | 0.023 | -0.699 | 0.276 | 0.023 |
| 0.0705 | 1.69 | 0.0344 | -0.470 | 0.265 | 0.018 | -0.504 | 0.281 | 0.018 |
| 0.0705 | 1.71 | 0.0333 | -0.476 | 0.277 | 0.039 | -0.507 | 0.292 | 0.039 |
| 0.0705 | 1.73 | 0.0323 | -0.132 | 0.244 | 0.009 | -0.142 | 0.257 | 0.009 |
| 0.0705 | 1.75 | 0.0313 | -0.245 | 0.275 | 0.016 | -0.258 | 0.289 | 0.016 |
| 0.0705 | 1.77 | 0.0303 | -0.353 | 0.259 | 0.028 | -0.370 | 0.271 | 0.028 |
| 0.0705 | 1.79 | 0.0294 | -0.191 | 0.285 | 0.022 | -0.200 | 0.298 | 0.022 |
| 0.0705 | 1.81 | 0.0286 | 0.165 | 0.295 | 0.033 | 0.172 | 0.307 | 0.033 |
| 0.0705 | 1.83 | 0.0278 | 0.011 | 0.312 | 0.022 | 0.012 | 0.324 | 0.022 |
| 0.0705 | 1.85 | 0.0270 | -0.420 | 0.361 | 0.024 | -0.433 | 0.374 | 0.024 |
| 0.0705 | 1.87 | 0.0262 | -0.008 | 0.369 | 0.016 | -0.007 | 0.382 | 0.016 |
| 0.0705 | 1.89 | 0.0255 | -0.016 | 0.425 | 0.037 | -0.014 | 0.439 | 0.037 |
| 0.0705 | 1.91 | 0.0248 | 0.156 | 0.493 | 0.044 | 0.163 | 0.508 | 0.044 |
| 0.0705 | 1.93 | 0.0242 | 0.778 | 0.721 | 0.107 | 0.803 | 0.743 | 0.107 |
| 0.0841 | 1.11 | 0.1930 | -0.001 | 0.020 | 0.008 | -0.034 | 0.050 | 0.008 |
| 0.0841 | 1.13 | 0.1750 | 0.042 | 0.021 | 0.017 | 0.053 | 0.048 | 0.017 |
| 0.0841 | 1.15 | 0.1598 | -0.092 | 0.026 | 0.020 | -0.228 | 0.053 | 0.020 |
| 0.0841 | 1.17 | 0.1469 | -0.183 | 0.028 | 0.027 | -0.370 | 0.054 | 0.027 |
| 0.0841 | 1.19 | 0.1357 | -0.359 | 0.033 | 0.027 | -0.641 | 0.059 | 0.027 |
| 0.0841 | 1.21 | 0.1260 | -0.429 | 0.035 | 0.024 | -0.708 | 0.058 | 0.024 |
| 0.0841 | 1.23 | 0.1174 | -0.395 | 0.034 | 0.026 | -0.624 | 0.054 | 0.026 |
| 0.0841 | 1.25 | 0.1098 | -0.282 | 0.033 | 0.029 | -0.438 | 0.050 | 0.029 |
| 0.0841 | 1.27 | 0.1030 | -0.119 | 0.030 | 0.022 | -0.195 | 0.043 | 0.022 |
| 0.0841 | 1.29 | 0.0969 | -0.058 | 0.029 | 0.019 | -0.108 | 0.041 | 0.019 |
| 0.0841 | 1.31 | 0.0915 | -0.010 | 0.029 | 0.017 | -0.045 | 0.039 | 0.017 |
| 0.0841 | 1.33 | 0.0865 | -0.012 | 0.029 | 0.017 | -0.048 | 0.038 | 0.017 |
| 0.0841 | 1.35 | 0.0820 | 0.004 | 0.029 | 0.016 | -0.030 | 0.038 | 0.016 |
| 0.0841 | 1.37 | 0.0778 | 0.089 | 0.033 | 0.013 | 0.075 | 0.042 | 0.013 |
| 0.0841 | 1.39 | 0.0741 | 0.028 | 0.037 | 0.015 | -0.005 | 0.046 | 0.015 |
| 0.0841 | 1.41 | 0.0706 | -0.032 | 0.040 | 0.010 | -0.083 | 0.048 | 0.010 |
| 0.0841 | 1.43 | 0.0674 | -0.006 | 0.043 | 0.015 | -0.056 | 0.051 | 0.015 |
| 0.0841 | 1.45 | 0.0644 | 0.043 | 0.048 | 0.016 | -0.002 | 0.056 | 0.016 |
| 0.0841 | 1.47 | 0.0616 | -0.154 | 0.053 | 0.020 | -0.234 | 0.062 | 0.020 |
| 0.0841 | 1.49 | 0.0591 | -0.087 | 0.053 | 0.020 | -0.152 | 0.061 | 0.020 |
| 0.0841 | 1.51 | 0.0567 | -0.055 | 0.055 | 0.024 | -0.106 | 0.063 | 0.024 |
| 0.0841 | 1.53 | 0.0545 | 0.008 | 0.056 | 0.021 | -0.030 | 0.063 | 0.021 |
| 0.0841 | 1.55 | 0.0524 | 0.006 | 0.057 | 0.017 | -0.028 | 0.064 | 0.017 |
| 0.0841 | 1.57 | 0.0504 | 0.067 | 0.059 | 0.015 | 0.043 | 0.066 | 0.015 |
| 0.0841 | 1.59 | 0.0486 | 0.053 | 0.066 | 0.015 | 0.032 | 0.072 | 0.015 |
| 0.0841 | 1.61 | 0.0468 | -0.119 | 0.081 | 0.020 | -0.152 | 0.089 | 0.020 |
| 0.0841 | 1.63 | 0.0452 | -0.083 | 0.224 | 0.017 | -0.108 | 0.244 | 0.017 |
| 0.0841 | 1.65 | 0.0437 | -0.542 | 0.250 | 0.023 | -0.600 | 0.270 | 0.023 |
| 0.0841 | 1.67 | 0.0422 | -0.120 | 0.258 | 0.008 | -0.141 | 0.277 | 0.008 |
| 0.0841 | 1.69 | 0.0408 | -0.236 | 0.228 | 0.025 | -0.263 | 0.244 | 0.025 |
| 0.0841 | 1.71 | 0.0395 | -0.097 | 0.237 | 0.021 | -0.111 | 0.253 | 0.021 |
| 0.0841 | 1.73 | 0.0383 | -0.280 | 0.232 | 0.011 | -0.302 | 0.246 | 0.011 |
| 0.0841 | 1.75 | 0.0371 | 0.249 | 0.250 | 0.019 | 0.261 | 0.264 | 0.019 |
| 0.0841 | 1.77 | 0.0360 | -0.266 | 0.249 | 0.015 | -0.282 | 0.262 | 0.015 |
| 0.0841 | 1.79 | 0.0349 | -0.151 | 0.252 | 0.016 | -0.160 | 0.265 | 0.016 |
| 0.0841 | 1.81 | 0.0339 | -0.553 | 0.313 | 0.025 | -0.579 | 0.328 | 0.025 |
| 0.0841 | 1.83 | 0.0330 | -0.370 | 0.334 | 0.021 | -0.386 | 0.349 | 0.021 |
| 0.0841 | 1.85 | 0.0320 | -0.396 | 0.295 | 0.051 | -0.412 | 0.307 | 0.051 |
| 0.0841 | 1.87 | 0.0312 | -0.181 | 0.315 | 0.020 | -0.188 | 0.328 | 0.020 |
| 0.0841 | 1.89 | 0.0303 | -0.370 | 0.359 | 0.019 | -0.383 | 0.373 | 0.019 |
| 0.0841 | 1.91 | 0.0295 | -0.522 | 0.398 | 0.081 | -0.539 | 0.412 | 0.081 |
| 0.0841 | 1.93 | 0.0287 | -2.096 | 1.902 | 1.222 | -2.170 | 1.970 | 1.222 |
| 0.1005 | 1.11 | 0.2223 | 0.013 | 0.018 | 0.012 | 0.005 | 0.049 | 0.012 |
| 0.1005 | 1.13 | 0.2023 | 0.001 | 0.021 | 0.009 | -0.037 | 0.050 | 0.009 |
| 0.1005 | 1.15 | 0.1853 | -0.025 | 0.024 | 0.014 | -0.094 | 0.052 | 0.014 |
| 0.1005 | 1.17 | 0.1707 | -0.186 | 0.025 | 0.024 | -0.398 | 0.051 | 0.024 |
| 0.1005 | 1.19 | 0.1580 | -0.310 | 0.028 | 0.028 | -0.580 | 0.052 | 0.028 |
| 0.1005 | 1.21 | 0.1469 | -0.458 | 0.030 | 0.025 | -0.796 | 0.053 | 0.025 |
| 0.1005 | 1.23 | 0.1372 | -0.370 | 0.030 | 0.030 | -0.614 | 0.049 | 0.030 |
| 0.1005 | 1.25 | 0.1285 | -0.281 | 0.029 | 0.027 | -0.457 | 0.045 | 0.027 |
| 0.1005 | 1.27 | 0.1207 | -0.158 | 0.029 | 0.023 | -0.264 | 0.044 | 0.023 |
| 0.1005 | 1.29 | 0.1137 | -0.090 | 0.029 | 0.019 | -0.160 | 0.043 | 0.019 |
| 0.1005 | 1.31 | 0.1074 | -0.057 | 0.028 | 0.016 | -0.115 | 0.040 | 0.016 |
| 0.1005 | 1.33 | 0.1017 | 0.007 | 0.029 | 0.015 | -0.028 | 0.040 | 0.015 |
| 0.1005 | 1.35 | 0.0964 | 0.045 | 0.031 | 0.013 | 0.020 | 0.041 | 0.013 |
| 0.1005 | 1.37 | 0.0917 | 0.037 | 0.016 | 0.025 | 0.006 | 0.021 | 0.025 |
| 0.1005 | 1.39 | 0.0873 | -0.042 | 0.033 | 0.009 | -0.100 | 0.042 | 0.009 |
| 0.1005 | 1.41 | 0.0832 | 0.019 | 0.034 | 0.015 | -0.027 | 0.042 | 0.015 |
| 0.1005 | 1.43 | 0.0795 | -0.001 | 0.035 | 0.012 | -0.056 | 0.043 | 0.012 |
| 0.1005 | 1.45 | 0.0760 | -0.028 | 0.041 | 0.010 | -0.093 | 0.049 | 0.010 |
| 0.1005 | 1.47 | 0.0728 | -0.064 | 0.043 | 0.021 | -0.140 | 0.051 | 0.021 |
| 0.1005 | 1.49 | 0.0698 | -0.026 | 0.045 | 0.020 | -0.090 | 0.053 | 0.020 |
| 0.1005 | 1.51 | 0.0670 | -0.052 | 0.045 | 0.020 | -0.110 | 0.052 | 0.020 |
| 0.1005 | 1.53 | 0.0644 | 0.054 | 0.047 | 0.019 | 0.017 | 0.054 | 0.019 |
| 0.1005 | 1.55 | 0.0620 | -0.008 | 0.050 | 0.028 | -0.050 | 0.057 | 0.028 |
| 0.1005 | 1.57 | 0.0597 | -0.019 | 0.055 | 0.021 | -0.059 | 0.062 | 0.021 |
| 0.1005 | 1.59 | 0.0575 | 0.034 | 0.063 | 0.018 | 0.007 | 0.071 | 0.018 |
| 0.1005 | 1.61 | 0.0555 | -0.113 | 0.078 | 0.020 | -0.152 | 0.087 | 0.020 |
| 0.1005 | 1.63 | 0.0536 | -0.283 | 0.093 | 0.039 | -0.333 | 0.102 | 0.039 |
| 0.1005 | 1.65 | 0.0518 | 0.009 | 0.089 | 0.008 | -0.008 | 0.097 | 0.008 |
| 0.1005 | 1.67 | 0.0500 | -0.299 | 0.093 | 0.015 | -0.340 | 0.101 | 0.015 |
| 0.1005 | 1.69 | 0.0484 | -0.136 | 0.098 | 0.007 | -0.159 | 0.106 | 0.007 |
| 0.1005 | 1.71 | 0.0469 | -0.064 | 0.093 | 0.026 | -0.078 | 0.101 | 0.026 |
| 0.1005 | 1.73 | 0.0454 | -0.259 | 0.094 | 0.020 | -0.284 | 0.101 | 0.020 |
| 0.1005 | 1.75 | 0.0441 | -0.133 | 0.096 | 0.008 | -0.147 | 0.102 | 0.008 |
| 0.1005 | 1.77 | 0.0427 | -0.052 | 0.103 | 0.010 | -0.059 | 0.110 | 0.010 |
| 0.1005 | 1.79 | 0.0415 | -0.148 | 0.107 | 0.011 | -0.159 | 0.114 | 0.011 |
| 0.1005 | 1.81 | 0.0403 | -0.136 | 0.104 | 0.022 | -0.145 | 0.109 | 0.022 |
| 0.1005 | 1.83 | 0.0391 | -0.060 | 0.103 | 0.024 | -0.064 | 0.109 | 0.024 |
| 0.1005 | 1.85 | 0.0381 | 0.111 | 0.113 | 0.016 | 0.116 | 0.119 | 0.016 |
| 0.1005 | 1.87 | 0.0370 | -0.180 | 0.117 | 0.014 | -0.189 | 0.122 | 0.014 |
| 0.1005 | 1.89 | 0.0360 | -0.045 | 0.122 | 0.017 | -0.046 | 0.127 | 0.017 |
| 0.1005 | 1.91 | 0.0351 | -0.173 | 0.132 | 0.020 | -0.180 | 0.137 | 0.020 |
| 0.1005 | 1.93 | 0.0341 | -0.268 | 0.149 | 0.026 | -0.277 | 0.155 | 0.026 |
| 0.1005 | 1.95 | 0.0333 | -0.047 | 0.158 | 0.039 | -0.046 | 0.164 | 0.039 |
| 0.1005 | 1.97 | 0.0324 | 0.033 | 0.176 | 0.036 | 0.037 | 0.183 | 0.036 |
| 0.1005 | 1.99 | 0.0316 | -0.012 | 0.163 | 0.033 | -0.009 | 0.169 | 0.033 |
| 0.1005 | 2.01 | 0.0308 | -0.075 | 0.190 | 0.027 | -0.074 | 0.196 | 0.027 |
| 0.1005 | 2.03 | 0.0301 | -0.177 | 0.221 | 0.028 | -0.179 | 0.228 | 0.028 |
| 0.1005 | 2.05 | 0.0294 | 0.284 | 0.388 | 0.065 | 0.296 | 0.400 | 0.065 |
| 0.12 | 1.11 | 0.2544 | 0.003 | 0.020 | 0.010 | -0.021 | 0.056 | 0.010 |
| 0.12 | 1.13 | 0.2324 | -0.013 | 0.019 | 0.011 | -0.076 | 0.050 | 0.011 |
| 0.12 | 1.15 | 0.2135 | -0.048 | 0.022 | 0.017 | -0.149 | 0.051 | 0.017 |
| 0.12 | 1.17 | 0.1973 | -0.166 | 0.026 | 0.025 | -0.374 | 0.056 | 0.025 |
| 0.12 | 1.19 | 0.1831 | -0.268 | 0.030 | 0.029 | -0.531 | 0.060 | 0.029 |
| 0.12 | 1.21 | 0.1706 | -0.402 | 0.034 | 0.028 | -0.733 | 0.064 | 0.028 |
| 0.12 | 1.23 | 0.1595 | -0.336 | 0.035 | 0.027 | -0.586 | 0.061 | 0.027 |
| 0.12 | 1.25 | 0.1496 | -0.281 | 0.034 | 0.029 | -0.479 | 0.056 | 0.029 |
| 0.12 | 1.27 | 0.1408 | -0.118 | 0.026 | 0.028 | -0.213 | 0.041 | 0.028 |
| 0.12 | 1.29 | 0.1328 | -0.036 | 0.034 | 0.017 | -0.090 | 0.052 | 0.017 |
| 0.12 | 1.31 | 0.1256 | -0.013 | 0.029 | 0.017 | -0.057 | 0.042 | 0.017 |
| 0.12 | 1.33 | 0.1190 | -0.011 | 0.028 | 0.016 | -0.057 | 0.039 | 0.016 |
| 0.12 | 1.35 | 0.1130 | -0.055 | 0.028 | 0.011 | -0.120 | 0.038 | 0.011 |
| 0.12 | 1.37 | 0.1075 | 0.013 | 0.030 | 0.016 | -0.030 | 0.040 | 0.016 |
| 0.12 | 1.39 | 0.1024 | 0.013 | 0.032 | 0.014 | -0.035 | 0.041 | 0.014 |
| 0.12 | 1.41 | 0.0978 | -0.002 | 0.034 | 0.016 | -0.059 | 0.044 | 0.016 |
| 0.12 | 1.43 | 0.0935 | -0.016 | 0.035 | 0.009 | -0.081 | 0.044 | 0.009 |
| 0.12 | 1.45 | 0.0894 | -0.023 | 0.038 | 0.014 | -0.095 | 0.047 | 0.014 |
| 0.12 | 1.47 | 0.0857 | 0.004 | 0.044 | 0.020 | -0.065 | 0.053 | 0.020 |
| 0.12 | 1.49 | 0.0822 | 0.015 | 0.050 | 0.024 | -0.050 | 0.060 | 0.024 |
| 0.12 | 1.51 | 0.0790 | 0.108 | 0.057 | 0.045 | 0.070 | 0.068 | 0.045 |
| 0.12 | 1.53 | 0.0760 | 0.031 | 0.067 | 0.027 | -0.016 | 0.078 | 0.027 |
| 0.12 | 1.55 | 0.0731 | 0.031 | 0.070 | 0.086 | -0.012 | 0.081 | 0.086 |
| 0.12 | 1.57 | 0.0704 | 0.053 | 0.052 | 0.023 | 0.017 | 0.060 | 0.023 |
| 0.12 | 1.59 | 0.0679 | 0.019 | 0.071 | 0.032 | -0.016 | 0.081 | 0.032 |
| 0.12 | 1.61 | 0.0655 | -0.133 | 0.092 | 0.024 | -0.182 | 0.104 | 0.024 |
| 0.12 | 1.63 | 0.0633 | 0.002 | 0.094 | 0.012 | -0.025 | 0.105 | 0.012 |
| 0.12 | 1.65 | 0.0612 | -0.045 | 0.091 | 0.009 | -0.072 | 0.101 | 0.009 |
| 0.12 | 1.67 | 0.0592 | -0.012 | 0.094 | 0.016 | -0.032 | 0.103 | 0.016 |
| 0.12 | 1.69 | 0.0573 | -0.139 | 0.097 | 0.009 | -0.168 | 0.106 | 0.009 |
| 0.12 | 1.71 | 0.0555 | -0.033 | 0.094 | 0.008 | -0.049 | 0.102 | 0.008 |
| 0.12 | 1.73 | 0.0538 | -0.063 | 0.092 | 0.020 | -0.077 | 0.100 | 0.020 |
| 0.12 | 1.75 | 0.0521 | 0.050 | 0.093 | 0.016 | 0.047 | 0.101 | 0.016 |
| 0.12 | 1.77 | 0.0506 | 0.024 | 0.092 | 0.015 | 0.021 | 0.099 | 0.015 |
| 0.12 | 1.79 | 0.0491 | 0.057 | 0.094 | 0.012 | 0.057 | 0.100 | 0.012 |
| 0.12 | 1.81 | 0.0477 | -0.149 | 0.092 | 0.012 | -0.162 | 0.098 | 0.012 |
| 0.12 | 1.83 | 0.0464 | -0.102 | 0.103 | 0.013 | -0.111 | 0.110 | 0.013 |
| 0.12 | 1.85 | 0.0451 | -0.021 | 0.107 | 0.016 | -0.024 | 0.113 | 0.016 |
| 0.12 | 1.87 | 0.0439 | -0.262 | 0.108 | 0.041 | -0.278 | 0.114 | 0.041 |
| 0.12 | 1.89 | 0.0427 | 0.133 | 0.113 | 0.030 | 0.139 | 0.119 | 0.030 |
| 0.12 | 1.91 | 0.0416 | 0.059 | 0.123 | 0.019 | 0.062 | 0.129 | 0.019 |
| 0.12 | 1.93 | 0.0405 | -0.039 | 0.133 | 0.020 | -0.041 | 0.139 | 0.020 |
| 0.12 | 1.95 | 0.0395 | -0.017 | 0.140 | 0.025 | -0.017 | 0.146 | 0.025 |
| 0.12 | 1.97 | 0.0385 | -0.202 | 0.150 | 0.028 | -0.208 | 0.157 | 0.028 |
| 0.12 | 1.99 | 0.0375 | -0.164 | 0.144 | 0.025 | -0.168 | 0.149 | 0.025 |
| 0.12 | 2.01 | 0.0366 | -0.050 | 0.166 | 0.023 | -0.049 | 0.173 | 0.023 |
| 0.12 | 2.03 | 0.0357 | -0.072 | 0.181 | 0.024 | -0.071 | 0.188 | 0.024 |
| 0.143 | 1.11 | 0.2890 | 0.041 | 0.021 | 0.007 | 0.097 | 0.066 | 0.007 |
| 0.143 | 1.13 | 0.2650 | 0.001 | 0.028 | 0.010 | -0.036 | 0.076 | 0.010 |
| 0.143 | 1.15 | 0.2443 | -0.011 | 0.028 | 0.035 | -0.057 | 0.069 | 0.035 |
| 0.143 | 1.17 | 0.2264 | -0.123 | 0.025 | 0.050 | -0.293 | 0.056 | 0.050 |
| 0.143 | 1.19 | 0.2106 | -0.315 | 0.035 | 0.028 | -0.649 | 0.072 | 0.028 |
| 0.143 | 1.21 | 0.1967 | -0.348 | 0.034 | 0.036 | -0.657 | 0.067 | 0.036 |
| 0.143 | 1.23 | 0.1843 | -0.369 | 0.033 | 0.029 | -0.668 | 0.060 | 0.029 |
| 0.143 | 1.25 | 0.1733 | -0.182 | 0.030 | 0.025 | -0.328 | 0.052 | 0.025 |
| 0.143 | 1.27 | 0.1633 | -0.162 | 0.028 | 0.022 | -0.294 | 0.047 | 0.022 |
| 0.143 | 1.29 | 0.1543 | -0.121 | 0.027 | 0.017 | -0.227 | 0.043 | 0.017 |
| 0.143 | 1.31 | 0.1461 | -0.041 | 0.026 | 0.014 | -0.103 | 0.040 | 0.014 |
| 0.143 | 1.33 | 0.1386 | 0.022 | 0.027 | 0.010 | -0.013 | 0.039 | 0.010 |
| 0.143 | 1.35 | 0.1317 | 0.013 | 0.029 | 0.007 | -0.029 | 0.041 | 0.007 |
| 0.143 | 1.37 | 0.1255 | 0.020 | 0.031 | 0.022 | -0.024 | 0.043 | 0.022 |
| 0.143 | 1.39 | 0.1197 | 0.007 | 0.033 | 0.013 | -0.047 | 0.045 | 0.013 |
| 0.143 | 1.41 | 0.1143 | 0.046 | 0.037 | 0.006 | 0.000 | 0.049 | 0.006 |
| 0.143 | 1.43 | 0.1093 | 0.083 | 0.046 | 0.013 | 0.041 | 0.059 | 0.013 |
| 0.143 | 1.45 | 0.1047 | 0.018 | 0.054 | 0.016 | -0.049 | 0.069 | 0.016 |
| 0.143 | 1.47 | 0.1004 | -0.049 | 0.066 | 0.033 | -0.139 | 0.083 | 0.033 |
| 0.143 | 1.49 | 0.0964 | -0.020 | 0.057 | 0.086 | -0.099 | 0.071 | 0.086 |
| 0.143 | 1.51 | 0.0927 | 0.030 | 0.056 | 0.043 | -0.027 | 0.067 | 0.043 |
| 0.143 | 1.53 | 0.0892 | 0.104 | 0.063 | 0.026 | 0.067 | 0.075 | 0.026 |
| 0.143 | 1.55 | 0.0859 | 0.010 | 0.068 | 0.031 | -0.042 | 0.081 | 0.031 |
| 0.143 | 1.57 | 0.0828 | 0.047 | 0.070 | 0.030 | 0.006 | 0.082 | 0.030 |
| 0.143 | 1.59 | 0.0798 | -0.016 | 0.071 | 0.018 | -0.063 | 0.082 | 0.018 |
| 0.143 | 1.61 | 0.0771 | 0.028 | 0.087 | 0.010 | -0.006 | 0.100 | 0.010 |
| 0.143 | 1.63 | 0.0745 | -0.044 | 0.084 | 0.009 | -0.082 | 0.096 | 0.009 |
| 0.143 | 1.65 | 0.0720 | 0.030 | 0.091 | 0.011 | 0.006 | 0.102 | 0.011 |
| 0.143 | 1.67 | 0.0697 | -0.083 | 0.087 | 0.016 | -0.116 | 0.097 | 0.016 |
| 0.143 | 1.69 | 0.0675 | -0.037 | 0.090 | 0.010 | -0.061 | 0.100 | 0.010 |
| 0.143 | 1.71 | 0.0654 | -0.225 | 0.089 | 0.027 | -0.264 | 0.098 | 0.027 |
| 0.143 | 1.73 | 0.0634 | 0.047 | 0.089 | 0.020 | 0.039 | 0.098 | 0.020 |
| 0.143 | 1.75 | 0.0615 | -0.125 | 0.090 | 0.010 | -0.146 | 0.099 | 0.010 |
| 0.143 | 1.77 | 0.0597 | -0.049 | 0.094 | 0.026 | -0.061 | 0.102 | 0.026 |
| 0.143 | 1.79 | 0.0580 | -0.070 | 0.091 | 0.011 | -0.081 | 0.098 | 0.011 |
| 0.143 | 1.81 | 0.0563 | -0.087 | 0.095 | 0.013 | -0.098 | 0.103 | 0.013 |
| 0.143 | 1.83 | 0.0547 | -0.198 | 0.098 | 0.044 | -0.216 | 0.105 | 0.044 |
| 0.143 | 1.85 | 0.0532 | -0.073 | 0.102 | 0.015 | -0.081 | 0.109 | 0.015 |
| 0.143 | 1.87 | 0.0518 | -0.042 | 0.097 | 0.016 | -0.047 | 0.103 | 0.016 |
| 0.143 | 1.89 | 0.0504 | -0.007 | 0.115 | 0.039 | -0.010 | 0.122 | 0.039 |
| 0.143 | 1.91 | 0.0491 | -0.144 | 0.126 | 0.019 | -0.154 | 0.133 | 0.019 |
| 0.143 | 1.93 | 0.0479 | -0.105 | 0.129 | 0.024 | -0.112 | 0.136 | 0.024 |
| 0.143 | 1.95 | 0.0466 | -0.014 | 0.132 | 0.035 | -0.015 | 0.139 | 0.035 |
| 0.143 | 1.97 | 0.0455 | -0.049 | 0.144 | 0.025 | -0.050 | 0.151 | 0.025 |
| 0.143 | 1.99 | 0.0444 | 0.164 | 0.152 | 0.029 | 0.173 | 0.160 | 0.029 |
| 0.143 | 2.01 | 0.0433 | 0.105 | 0.152 | 0.031 | 0.113 | 0.158 | 0.031 |
| 0.143 | 2.03 | 0.0423 | 0.052 | 0.312 | 0.124 | 0.056 | 0.326 | 0.124 |
| 0.1708 | 1.11 | 0.3269 | 0.006 | 0.015 | 0.012 | -0.007 | 0.049 | 0.012 |
| 0.1708 | 1.13 | 0.3011 | 0.006 | 0.019 | 0.008 | -0.018 | 0.056 | 0.008 |
| 0.1708 | 1.15 | 0.2787 | -0.065 | 0.021 | 0.018 | -0.199 | 0.053 | 0.018 |
| 0.1708 | 1.17 | 0.2590 | -0.139 | 0.022 | 0.025 | -0.344 | 0.052 | 0.025 |
| 0.1708 | 1.19 | 0.2417 | -0.257 | 0.025 | 0.028 | -0.554 | 0.056 | 0.028 |
| 0.1708 | 1.21 | 0.2264 | -0.358 | 0.027 | 0.030 | -0.714 | 0.055 | 0.030 |
| 0.1708 | 1.23 | 0.2126 | -0.321 | 0.028 | 0.028 | -0.612 | 0.055 | 0.028 |
| 0.1708 | 1.25 | 0.2003 | -0.182 | 0.026 | 0.024 | -0.343 | 0.048 | 0.024 |
| 0.1708 | 1.27 | 0.1891 | -0.113 | 0.025 | 0.019 | -0.224 | 0.044 | 0.019 |
| 0.1708 | 1.29 | 0.1789 | -0.093 | 0.026 | 0.018 | -0.192 | 0.044 | 0.018 |
| 0.1708 | 1.31 | 0.1697 | 0.016 | 0.027 | 0.012 | -0.017 | 0.044 | 0.012 |
| 0.1708 | 1.33 | 0.1612 | -0.013 | 0.029 | 0.013 | -0.068 | 0.045 | 0.013 |
| 0.1708 | 1.35 | 0.1535 | 0.008 | 0.033 | 0.007 | -0.041 | 0.049 | 0.007 |
| 0.1708 | 1.37 | 0.1463 | 0.078 | 0.037 | 0.012 | 0.057 | 0.054 | 0.012 |
| 0.1708 | 1.39 | 0.1397 | 0.096 | 0.041 | 0.015 | 0.074 | 0.057 | 0.015 |
| 0.1708 | 1.41 | 0.1336 | -0.035 | 0.046 | 0.023 | -0.114 | 0.063 | 0.023 |
| 0.1708 | 1.43 | 0.1279 | 0.036 | 0.031 | 0.037 | -0.023 | 0.042 | 0.037 |
| 0.1708 | 1.45 | 0.1226 | 0.089 | 0.041 | 0.022 | 0.039 | 0.053 | 0.022 |
| 0.1708 | 1.47 | 0.1177 | 0.130 | 0.042 | 0.024 | 0.082 | 0.054 | 0.024 |
| 0.1708 | 1.49 | 0.1131 | 0.060 | 0.042 | 0.026 | -0.007 | 0.052 | 0.026 |
| 0.1708 | 1.51 | 0.1088 | 0.104 | 0.043 | 0.022 | 0.058 | 0.053 | 0.022 |
| 0.1708 | 1.53 | 0.1047 | 0.077 | 0.045 | 0.030 | 0.029 | 0.055 | 0.030 |
| 0.1708 | 1.55 | 0.1009 | 0.079 | 0.046 | 0.031 | 0.035 | 0.056 | 0.031 |
| 0.1708 | 1.57 | 0.0973 | 0.100 | 0.053 | 0.017 | 0.062 | 0.063 | 0.017 |
| 0.1708 | 1.59 | 0.0939 | 0.049 | 0.050 | 0.014 | 0.007 | 0.060 | 0.014 |
| 0.1708 | 1.61 | 0.0907 | 0.032 | 0.050 | 0.015 | -0.008 | 0.058 | 0.015 |
| 0.1708 | 1.63 | 0.0877 | 0.000 | 0.052 | 0.012 | -0.040 | 0.061 | 0.012 |
| 0.1708 | 1.65 | 0.0849 | -0.013 | 0.052 | 0.017 | -0.050 | 0.060 | 0.017 |
| 0.1708 | 1.67 | 0.0821 | 0.007 | 0.054 | 0.012 | -0.022 | 0.061 | 0.012 |
| 0.1708 | 1.69 | 0.0796 | -0.069 | 0.055 | 0.011 | -0.103 | 0.062 | 0.011 |
| 0.1708 | 1.71 | 0.0771 | 0.030 | 0.056 | 0.010 | 0.013 | 0.063 | 0.010 |
| 0.1708 | 1.73 | 0.0748 | -0.016 | 0.053 | 0.013 | -0.034 | 0.059 | 0.013 |
| 0.1708 | 1.75 | 0.0726 | 0.033 | 0.050 | 0.010 | 0.023 | 0.056 | 0.010 |
| 0.1708 | 1.77 | 0.0705 | -0.029 | 0.052 | 0.010 | -0.043 | 0.057 | 0.010 |
| 0.1708 | 1.79 | 0.0685 | -0.104 | 0.054 | 0.014 | -0.123 | 0.059 | 0.014 |
| 0.1708 | 1.81 | 0.0666 | -0.148 | 0.057 | 0.017 | -0.169 | 0.062 | 0.017 |
| 0.1708 | 1.83 | 0.0647 | -0.051 | 0.059 | 0.023 | -0.062 | 0.065 | 0.023 |
| 0.1708 | 1.85 | 0.0630 | -0.032 | 0.061 | 0.015 | -0.041 | 0.066 | 0.015 |
| 0.1708 | 1.87 | 0.0613 | 0.069 | 0.059 | 0.016 | 0.069 | 0.064 | 0.016 |
| 0.1708 | 1.89 | 0.0597 | -0.079 | 0.060 | 0.016 | -0.089 | 0.064 | 0.016 |
| 0.1708 | 1.91 | 0.0581 | -0.023 | 0.062 | 0.020 | -0.028 | 0.066 | 0.020 |
| 0.1708 | 1.93 | 0.0566 | 0.003 | 0.062 | 0.017 | 0.000 | 0.067 | 0.017 |
| 0.1708 | 1.95 | 0.0552 | 0.014 | 0.068 | 0.019 | 0.013 | 0.072 | 0.019 |
| 0.1708 | 1.97 | 0.0539 | 0.010 | 0.066 | 0.022 | 0.011 | 0.070 | 0.022 |
| 0.1708 | 1.99 | 0.0525 | -0.067 | 0.068 | 0.029 | -0.070 | 0.072 | 0.029 |
| 0.1708 | 2.01 | 0.0513 | -0.084 | 0.076 | 0.020 | -0.088 | 0.080 | 0.020 |
| 0.1708 | 2.03 | 0.0501 | -0.004 | 0.080 | 0.016 | -0.003 | 0.084 | 0.016 |
| 0.1708 | 2.05 | 0.0489 | -0.116 | 0.084 | 0.020 | -0.121 | 0.088 | 0.020 |
| 0.1708 | 2.07 | 0.0478 | 0.024 | 0.085 | 0.021 | 0.026 | 0.089 | 0.021 |
| 0.1708 | 2.09 | 0.0467 | -0.105 | 0.090 | 0.016 | -0.109 | 0.094 | 0.016 |
| 0.1708 | 2.11 | 0.0456 | -0.069 | 0.094 | 0.019 | -0.072 | 0.098 | 0.019 |
| 0.1708 | 2.13 | 0.0446 | 0.012 | 0.104 | 0.016 | 0.013 | 0.109 | 0.016 |
| 0.1708 | 2.15 | 0.0437 | 0.152 | 0.100 | 0.028 | 0.159 | 0.104 | 0.028 |
| 0.1708 | 2.17 | 0.0427 | 0.056 | 0.105 | 0.020 | 0.058 | 0.109 | 0.020 |
| 0.1708 | 2.19 | 0.0418 | 0.187 | 0.106 | 0.019 | 0.193 | 0.110 | 0.019 |
| 0.1708 | 2.21 | 0.0409 | -0.090 | 0.109 | 0.020 | -0.093 | 0.112 | 0.020 |
| 0.1708 | 2.23 | 0.0401 | 0.128 | 0.124 | 0.026 | 0.132 | 0.128 | 0.026 |
| 0.1708 | 2.25 | 0.0392 | -0.218 | 0.142 | 0.034 | -0.225 | 0.147 | 0.034 |
| 0.1708 | 2.27 | 0.0384 | -0.081 | 0.139 | 0.026 | -0.083 | 0.144 | 0.026 |
| 0.1708 | 2.29 | 0.0377 | 0.321 | 0.218 | 0.210 | 0.330 | 0.225 | 0.210 |
| 0.2042 | 1.11 | 0.3673 | 0.016 | 0.014 | 0.013 | 0.029 | 0.047 | 0.013 |
| 0.2042 | 1.13 | 0.3399 | 0.006 | 0.018 | 0.014 | -0.012 | 0.053 | 0.014 |
| 0.2042 | 1.15 | 0.3159 | -0.023 | 0.020 | 0.015 | -0.087 | 0.056 | 0.015 |
| 0.2042 | 1.17 | 0.2948 | -0.128 | 0.027 | 0.022 | -0.326 | 0.067 | 0.022 |
| 0.2042 | 1.19 | 0.2760 | -0.240 | 0.031 | 0.025 | -0.534 | 0.071 | 0.025 |
| 0.2042 | 1.21 | 0.2592 | -0.278 | 0.034 | 0.027 | -0.569 | 0.075 | 0.027 |
| 0.2042 | 1.23 | 0.2441 | -0.318 | 0.039 | 0.027 | -0.632 | 0.080 | 0.027 |
| 0.2042 | 1.25 | 0.2304 | -0.189 | 0.041 | 0.027 | -0.369 | 0.079 | 0.027 |
| 0.2042 | 1.27 | 0.2180 | -0.095 | 0.038 | 0.022 | -0.199 | 0.070 | 0.022 |
| 0.2042 | 1.29 | 0.2067 | 0.008 | 0.038 | 0.018 | -0.023 | 0.067 | 0.018 |
| 0.2042 | 1.31 | 0.1964 | -0.056 | 0.044 | 0.014 | -0.139 | 0.073 | 0.014 |
| 0.2042 | 1.33 | 0.1869 | 0.073 | 0.038 | 0.018 | 0.063 | 0.061 | 0.018 |
| 0.2042 | 1.35 | 0.1781 | 0.028 | 0.032 | 0.006 | -0.013 | 0.049 | 0.006 |
| 0.2042 | 1.37 | 0.1701 | 0.051 | 0.031 | 0.008 | 0.017 | 0.047 | 0.008 |
| 0.2042 | 1.39 | 0.1626 | -0.009 | 0.032 | 0.016 | -0.077 | 0.046 | 0.016 |
| 0.2042 | 1.41 | 0.1557 | 0.028 | 0.033 | 0.010 | -0.030 | 0.046 | 0.010 |
| 0.2042 | 1.43 | 0.1492 | 0.044 | 0.034 | 0.035 | -0.015 | 0.047 | 0.035 |
| 0.2042 | 1.45 | 0.1432 | 0.015 | 0.040 | 0.036 | -0.063 | 0.053 | 0.036 |
| 0.2042 | 1.47 | 0.1375 | 0.091 | 0.045 | 0.023 | 0.030 | 0.059 | 0.023 |
| 0.2042 | 1.49 | 0.1323 | 0.121 | 0.046 | 0.029 | 0.068 | 0.060 | 0.029 |
| 0.2042 | 1.51 | 0.1273 | 0.066 | 0.052 | 0.025 | 0.005 | 0.066 | 0.025 |
| 0.2042 | 1.53 | 0.1227 | 0.024 | 0.052 | 0.022 | -0.041 | 0.065 | 0.022 |
| 0.2042 | 1.55 | 0.1183 | 0.068 | 0.052 | 0.021 | 0.016 | 0.064 | 0.021 |
| 0.2042 | 1.57 | 0.1142 | 0.063 | 0.050 | 0.018 | 0.013 | 0.061 | 0.018 |
| 0.2042 | 1.59 | 0.1103 | 0.119 | 0.046 | 0.017 | 0.085 | 0.056 | 0.017 |
| 0.2042 | 1.61 | 0.1066 | 0.047 | 0.048 | 0.011 | 0.003 | 0.057 | 0.011 |
| 0.2042 | 1.63 | 0.1031 | 0.085 | 0.047 | 0.011 | 0.053 | 0.055 | 0.011 |
| 0.2042 | 1.65 | 0.0998 | -0.062 | 0.051 | 0.016 | -0.115 | 0.060 | 0.016 |
| 0.2042 | 1.67 | 0.0967 | -0.062 | 0.055 | 0.019 | -0.109 | 0.064 | 0.019 |
| 0.2042 | 1.69 | 0.0937 | 0.005 | 0.054 | 0.021 | -0.025 | 0.063 | 0.021 |
| 0.2042 | 1.71 | 0.0908 | -0.101 | 0.053 | 0.011 | -0.140 | 0.060 | 0.011 |
| 0.2042 | 1.73 | 0.0881 | 0.021 | 0.053 | 0.010 | 0.003 | 0.060 | 0.010 |
| 0.2042 | 1.75 | 0.0856 | 0.006 | 0.054 | 0.010 | -0.011 | 0.061 | 0.010 |
| 0.2042 | 1.77 | 0.0831 | 0.053 | 0.056 | 0.013 | 0.045 | 0.062 | 0.013 |
| 0.2042 | 1.79 | 0.0808 | -0.038 | 0.058 | 0.014 | -0.054 | 0.065 | 0.014 |
| 0.2042 | 1.81 | 0.0785 | -0.040 | 0.057 | 0.013 | -0.054 | 0.063 | 0.013 |
| 0.2042 | 1.83 | 0.0764 | -0.138 | 0.057 | 0.014 | -0.161 | 0.063 | 0.014 |
| 0.2042 | 1.85 | 0.0744 | 0.014 | 0.059 | 0.016 | 0.006 | 0.064 | 0.016 |
| 0.2042 | 1.87 | 0.0724 | -0.050 | 0.059 | 0.016 | -0.062 | 0.065 | 0.016 |
| 0.2042 | 1.89 | 0.0705 | -0.063 | 0.062 | 0.021 | -0.076 | 0.067 | 0.021 |
| 0.2042 | 1.91 | 0.0687 | -0.072 | 0.060 | 0.019 | -0.084 | 0.065 | 0.019 |
| 0.2042 | 1.93 | 0.0670 | -0.066 | 0.062 | 0.019 | -0.076 | 0.067 | 0.019 |
| 0.2042 | 1.95 | 0.0653 | 0.008 | 0.063 | 0.020 | 0.005 | 0.068 | 0.020 |
| 0.2042 | 1.97 | 0.0637 | -0.041 | 0.067 | 0.022 | -0.047 | 0.071 | 0.022 |
| 0.2042 | 1.99 | 0.0622 | 0.026 | 0.077 | 0.024 | 0.026 | 0.082 | 0.024 |
| 0.2042 | 2.01 | 0.0607 | -0.064 | 0.081 | 0.019 | -0.069 | 0.086 | 0.019 |
| 0.2042 | 2.03 | 0.0593 | -0.117 | 0.094 | 0.017 | -0.124 | 0.099 | 0.017 |
| 0.2042 | 2.05 | 0.0579 | -0.057 | 0.074 | 0.016 | -0.061 | 0.078 | 0.016 |
| 0.2042 | 2.07 | 0.0566 | 0.023 | 0.077 | 0.017 | 0.023 | 0.082 | 0.017 |
| 0.2042 | 2.09 | 0.0553 | -0.146 | 0.087 | 0.020 | -0.155 | 0.091 | 0.020 |
| 0.2042 | 2.11 | 0.0541 | 0.048 | 0.088 | 0.020 | 0.050 | 0.092 | 0.020 |
| 0.2042 | 2.13 | 0.0529 | -0.144 | 0.091 | 0.022 | -0.151 | 0.096 | 0.022 |
| 0.2042 | 2.15 | 0.0518 | -0.028 | 0.098 | 0.016 | -0.030 | 0.102 | 0.016 |
| 0.2042 | 2.17 | 0.0506 | 0.205 | 0.097 | 0.021 | 0.213 | 0.101 | 0.021 |
| 0.2042 | 2.19 | 0.0496 | -0.043 | 0.102 | 0.018 | -0.046 | 0.106 | 0.018 |
| 0.2042 | 2.21 | 0.0485 | -0.178 | 0.111 | 0.019 | -0.186 | 0.115 | 0.019 |
| 0.2042 | 2.23 | 0.0475 | -0.040 | 0.110 | 0.016 | -0.043 | 0.114 | 0.016 |
| 0.2042 | 2.25 | 0.0466 | 0.107 | 0.111 | 0.036 | 0.109 | 0.115 | 0.036 |
| 0.2042 | 2.27 | 0.0456 | 0.057 | 0.107 | 0.040 | 0.057 | 0.111 | 0.040 |
| 0.2436 | 1.11 | 0.4091 | -0.001 | 0.024 | 0.014 | -0.024 | 0.078 | 0.014 |
| 0.2436 | 1.13 | 0.3805 | -0.002 | 0.014 | 0.030 | -0.034 | 0.044 | 0.030 |
| 0.2436 | 1.15 | 0.3552 | -0.053 | 0.032 | 0.010 | -0.174 | 0.092 | 0.010 |
| 0.2436 | 1.17 | 0.3327 | -0.130 | 0.025 | 0.061 | -0.324 | 0.065 | 0.061 |
| 0.2436 | 1.19 | 0.3125 | -0.172 | 0.041 | 0.028 | -0.391 | 0.097 | 0.028 |
| 0.2436 | 1.21 | 0.2944 | -0.328 | 0.038 | 0.033 | -0.703 | 0.084 | 0.033 |
| 0.2436 | 1.23 | 0.2780 | -0.256 | 0.034 | 0.022 | -0.523 | 0.072 | 0.022 |
| 0.2436 | 1.25 | 0.2631 | -0.149 | 0.034 | 0.021 | -0.304 | 0.067 | 0.021 |
| 0.2436 | 1.27 | 0.2495 | -0.077 | 0.029 | 0.015 | -0.173 | 0.054 | 0.015 |
| 0.2436 | 1.29 | 0.2371 | -0.018 | 0.026 | 0.014 | -0.071 | 0.048 | 0.014 |
| 0.2436 | 1.31 | 0.2257 | 0.013 | 0.025 | 0.010 | -0.025 | 0.043 | 0.010 |
| 0.2436 | 1.33 | 0.2151 | 0.024 | 0.024 | 0.008 | -0.015 | 0.041 | 0.008 |
| 0.2436 | 1.35 | 0.2054 | 0.048 | 0.025 | 0.006 | 0.018 | 0.040 | 0.006 |
| 0.2436 | 1.37 | 0.1964 | 0.054 | 0.029 | 0.014 | 0.020 | 0.045 | 0.014 |
| 0.2436 | 1.39 | 0.1880 | 0.067 | 0.031 | 0.009 | 0.035 | 0.047 | 0.009 |
| 0.2436 | 1.41 | 0.1802 | 0.054 | 0.034 | 0.026 | 0.006 | 0.049 | 0.026 |
| 0.2436 | 1.43 | 0.1730 | 0.126 | 0.036 | 0.029 | 0.102 | 0.052 | 0.029 |
| 0.2436 | 1.45 | 0.1662 | 0.116 | 0.041 | 0.011 | 0.075 | 0.058 | 0.011 |
| 0.2436 | 1.47 | 0.1598 | 0.050 | 0.045 | 0.022 | -0.027 | 0.062 | 0.022 |
| 0.2436 | 1.49 | 0.1538 | 0.169 | 0.050 | 0.024 | 0.133 | 0.066 | 0.024 |
| 0.2436 | 1.51 | 0.1482 | 0.275 | 0.053 | 0.025 | 0.277 | 0.070 | 0.025 |
| 0.2436 | 1.53 | 0.1429 | 0.105 | 0.050 | 0.024 | 0.056 | 0.064 | 0.024 |
| 0.2436 | 1.55 | 0.1379 | 0.099 | 0.048 | 0.023 | 0.051 | 0.061 | 0.023 |
| 0.2436 | 1.57 | 0.1332 | 0.051 | 0.047 | 0.019 | -0.008 | 0.059 | 0.019 |
| 0.2436 | 1.59 | 0.1288 | 0.143 | 0.048 | 0.016 | 0.111 | 0.059 | 0.016 |
| 0.2436 | 1.61 | 0.1246 | 0.036 | 0.050 | 0.014 | -0.018 | 0.061 | 0.014 |
| 0.2436 | 1.63 | 0.1206 | -0.002 | 0.050 | 0.010 | -0.059 | 0.060 | 0.010 |
| 0.2436 | 1.65 | 0.1168 | -0.012 | 0.050 | 0.011 | -0.065 | 0.060 | 0.011 |
| 0.2436 | 1.67 | 0.1132 | 0.052 | 0.053 | 0.016 | 0.016 | 0.063 | 0.016 |
| 0.2436 | 1.69 | 0.1097 | 0.102 | 0.055 | 0.013 | 0.080 | 0.064 | 0.013 |
| 0.2436 | 1.71 | 0.1065 | 0.053 | 0.053 | 0.012 | 0.029 | 0.062 | 0.012 |
| 0.2436 | 1.73 | 0.1034 | 0.048 | 0.056 | 0.013 | 0.028 | 0.064 | 0.013 |
| 0.2436 | 1.75 | 0.1004 | 0.015 | 0.048 | 0.015 | -0.005 | 0.055 | 0.015 |
| 0.2436 | 1.77 | 0.0976 | 0.057 | 0.049 | 0.011 | 0.045 | 0.055 | 0.011 |
| 0.2436 | 1.79 | 0.0949 | -0.056 | 0.054 | 0.012 | -0.079 | 0.061 | 0.012 |
| 0.2436 | 1.81 | 0.0923 | -0.096 | 0.055 | 0.012 | -0.123 | 0.062 | 0.012 |
| 0.2436 | 1.83 | 0.0898 | 0.102 | 0.057 | 0.017 | 0.100 | 0.064 | 0.017 |
| 0.2436 | 1.85 | 0.0874 | -0.042 | 0.053 | 0.016 | -0.059 | 0.059 | 0.016 |
| 0.2436 | 1.87 | 0.0852 | -0.065 | 0.059 | 0.017 | -0.084 | 0.065 | 0.017 |
| 0.2436 | 1.89 | 0.0830 | -0.013 | 0.060 | 0.016 | -0.025 | 0.065 | 0.016 |
| 0.2436 | 1.91 | 0.0809 | 0.067 | 0.061 | 0.020 | 0.065 | 0.067 | 0.020 |
| 0.2436 | 1.93 | 0.0789 | -0.018 | 0.067 | 0.019 | -0.027 | 0.073 | 0.019 |
| 0.2436 | 1.95 | 0.0769 | 0.004 | 0.069 | 0.018 | -0.002 | 0.075 | 0.018 |
| 0.2436 | 1.97 | 0.0751 | -0.027 | 0.078 | 0.019 | -0.034 | 0.085 | 0.019 |
| 0.2436 | 1.99 | 0.0733 | -0.032 | 0.079 | 0.023 | -0.038 | 0.085 | 0.023 |
| 0.2436 | 2.01 | 0.0716 | 0.032 | 0.073 | 0.019 | 0.031 | 0.078 | 0.019 |
| 0.2436 | 2.03 | 0.0699 | -0.021 | 0.071 | 0.024 | -0.026 | 0.076 | 0.024 |
| 0.2436 | 2.05 | 0.0683 | 0.039 | 0.073 | 0.017 | 0.038 | 0.078 | 0.017 |
| 0.2436 | 2.07 | 0.0668 | -0.040 | 0.081 | 0.020 | -0.046 | 0.086 | 0.020 |
| 0.2436 | 2.09 | 0.0653 | -0.007 | 0.077 | 0.030 | -0.011 | 0.082 | 0.030 |
| 0.2436 | 2.11 | 0.0638 | -0.106 | 0.079 | 0.018 | -0.115 | 0.084 | 0.018 |
| 0.2436 | 2.13 | 0.0625 | 0.034 | 0.090 | 0.027 | 0.033 | 0.095 | 0.027 |
| 0.2436 | 2.15 | 0.0611 | -0.014 | 0.091 | 0.017 | -0.017 | 0.096 | 0.017 |
| 0.2436 | 2.17 | 0.0598 | 0.046 | 0.088 | 0.016 | 0.046 | 0.092 | 0.016 |
| 0.2436 | 2.19 | 0.0586 | 0.053 | 0.081 | 0.026 | 0.053 | 0.085 | 0.026 |
| 0.2436 | 2.21 | 0.0573 | 0.225 | 0.088 | 0.023 | 0.232 | 0.093 | 0.023 |
| 0.2436 | 2.23 | 0.0562 | -0.111 | 0.100 | 0.017 | -0.118 | 0.104 | 0.017 |
| 0.2436 | 2.25 | 0.0550 | -0.044 | 0.090 | 0.021 | -0.049 | 0.094 | 0.021 |
| 0.2436 | 2.27 | 0.0539 | 0.241 | 0.124 | 0.037 | 0.249 | 0.129 | 0.037 |
| 0.2904 | 1.11 | 0.4522 | 0.006 | 0.024 | 0.014 | 0.002 | 0.083 | 0.014 |
| 0.2904 | 1.13 | 0.4227 | 0.023 | 0.022 | 0.014 | 0.050 | 0.070 | 0.014 |
| 0.2904 | 1.15 | 0.3964 | -0.064 | 0.023 | 0.016 | -0.198 | 0.067 | 0.016 |
| 0.2904 | 1.17 | 0.3728 | -0.103 | 0.024 | 0.019 | -0.271 | 0.065 | 0.019 |
| 0.2904 | 1.19 | 0.3515 | -0.218 | 0.026 | 0.018 | -0.517 | 0.065 | 0.018 |
| 0.2904 | 1.21 | 0.3322 | -0.235 | 0.027 | 0.021 | -0.517 | 0.064 | 0.021 |
| 0.2904 | 1.23 | 0.3146 | -0.208 | 0.027 | 0.023 | -0.439 | 0.060 | 0.023 |
| 0.2904 | 1.25 | 0.2986 | -0.158 | 0.027 | 0.017 | -0.336 | 0.057 | 0.017 |
| 0.2904 | 1.27 | 0.2839 | -0.090 | 0.028 | 0.017 | -0.203 | 0.056 | 0.017 |
| 0.2904 | 1.29 | 0.2703 | 0.004 | 0.030 | 0.013 | -0.031 | 0.057 | 0.013 |
| 0.2904 | 1.31 | 0.2579 | -0.011 | 0.032 | 0.015 | -0.071 | 0.057 | 0.015 |
| 0.2904 | 1.33 | 0.2463 | 0.053 | 0.028 | 0.015 | 0.035 | 0.049 | 0.015 |
| 0.2904 | 1.35 | 0.2356 | 0.019 | 0.027 | 0.010 | -0.031 | 0.046 | 0.010 |
| 0.2904 | 1.37 | 0.2256 | 0.067 | 0.035 | 0.008 | 0.042 | 0.057 | 0.008 |
| 0.2904 | 1.39 | 0.2164 | 0.081 | 0.036 | 0.009 | 0.056 | 0.057 | 0.009 |
| 0.2904 | 1.41 | 0.2077 | 0.050 | 0.038 | 0.008 | 0.001 | 0.058 | 0.008 |
| 0.2904 | 1.43 | 0.1996 | 0.129 | 0.040 | 0.009 | 0.110 | 0.059 | 0.009 |
| 0.2904 | 1.45 | 0.1920 | 0.108 | 0.046 | 0.011 | 0.067 | 0.066 | 0.011 |
| 0.2904 | 1.47 | 0.1849 | 0.182 | 0.047 | 0.020 | 0.159 | 0.067 | 0.020 |
| 0.2904 | 1.49 | 0.1781 | 0.230 | 0.047 | 0.022 | 0.218 | 0.065 | 0.022 |
| 0.2904 | 1.51 | 0.1718 | 0.241 | 0.045 | 0.024 | 0.234 | 0.061 | 0.024 |
| 0.2904 | 1.53 | 0.1658 | 0.184 | 0.043 | 0.024 | 0.158 | 0.057 | 0.024 |
| 0.2904 | 1.55 | 0.1602 | 0.157 | 0.044 | 0.027 | 0.123 | 0.058 | 0.027 |
| 0.2904 | 1.57 | 0.1549 | 0.154 | 0.048 | 0.028 | 0.118 | 0.062 | 0.028 |
| 0.2904 | 1.59 | 0.1498 | 0.137 | 0.048 | 0.020 | 0.098 | 0.061 | 0.020 |
| 0.2904 | 1.61 | 0.1450 | 0.134 | 0.046 | 0.015 | 0.096 | 0.058 | 0.015 |
| 0.2904 | 1.63 | 0.1405 | 0.143 | 0.047 | 0.020 | 0.111 | 0.058 | 0.020 |
| 0.2904 | 1.65 | 0.1362 | 0.045 | 0.050 | 0.017 | -0.007 | 0.061 | 0.017 |
| 0.2904 | 1.67 | 0.1321 | 0.027 | 0.051 | 0.010 | -0.023 | 0.062 | 0.010 |
| 0.2904 | 1.69 | 0.1281 | 0.014 | 0.053 | 0.015 | -0.033 | 0.063 | 0.015 |
| 0.2904 | 1.71 | 0.1244 | 0.051 | 0.050 | 0.014 | 0.020 | 0.059 | 0.014 |
| 0.2904 | 1.73 | 0.1209 | 0.137 | 0.049 | 0.010 | 0.127 | 0.057 | 0.010 |
| 0.2904 | 1.75 | 0.1174 | 0.013 | 0.048 | 0.009 | -0.014 | 0.056 | 0.009 |
| 0.2904 | 1.77 | 0.1142 | 0.122 | 0.052 | 0.017 | 0.117 | 0.060 | 0.017 |
| 0.2904 | 1.79 | 0.1111 | 0.124 | 0.055 | 0.016 | 0.122 | 0.064 | 0.016 |
| 0.2904 | 1.81 | 0.1081 | 0.041 | 0.058 | 0.015 | 0.027 | 0.066 | 0.015 |
| 0.2904 | 1.83 | 0.1053 | -0.073 | 0.061 | 0.017 | -0.101 | 0.069 | 0.017 |
| 0.2904 | 1.85 | 0.1025 | 0.043 | 0.063 | 0.018 | 0.032 | 0.071 | 0.018 |
| 0.2904 | 1.87 | 0.0999 | -0.082 | 0.068 | 0.017 | -0.108 | 0.076 | 0.017 |
| 0.2904 | 1.89 | 0.0974 | -0.063 | 0.065 | 0.019 | -0.085 | 0.072 | 0.019 |
| 0.2904 | 1.91 | 0.0950 | -0.003 | 0.059 | 0.018 | -0.016 | 0.065 | 0.018 |
| 0.2904 | 1.93 | 0.0926 | 0.052 | 0.071 | 0.019 | 0.046 | 0.079 | 0.019 |
| 0.2904 | 1.95 | 0.0904 | -0.081 | 0.063 | 0.020 | -0.098 | 0.069 | 0.020 |
| 0.2904 | 1.97 | 0.0882 | 0.072 | 0.067 | 0.022 | 0.071 | 0.073 | 0.022 |
| 0.2904 | 1.99 | 0.0862 | -0.033 | 0.066 | 0.018 | -0.043 | 0.072 | 0.018 |
| 0.2904 | 2.01 | 0.0842 | 0.100 | 0.062 | 0.019 | 0.103 | 0.067 | 0.019 |
| 0.2904 | 2.03 | 0.0822 | 0.031 | 0.072 | 0.016 | 0.028 | 0.078 | 0.016 |
| 0.2904 | 2.05 | 0.0804 | 0.028 | 0.072 | 0.016 | 0.024 | 0.078 | 0.016 |
| 0.2904 | 2.07 | 0.0786 | 0.056 | 0.073 | 0.019 | 0.055 | 0.079 | 0.019 |
| 0.2904 | 2.09 | 0.0769 | -0.135 | 0.073 | 0.024 | -0.150 | 0.078 | 0.024 |
| 0.2904 | 2.11 | 0.0752 | -0.013 | 0.075 | 0.015 | -0.020 | 0.080 | 0.015 |
| 0.2904 | 2.13 | 0.0736 | -0.001 | 0.077 | 0.018 | -0.007 | 0.082 | 0.018 |
| 0.2904 | 2.15 | 0.0720 | -0.026 | 0.081 | 0.019 | -0.032 | 0.087 | 0.019 |
| 0.2904 | 2.17 | 0.0705 | 0.046 | 0.084 | 0.021 | 0.044 | 0.089 | 0.021 |
| 0.2904 | 2.19 | 0.0690 | -0.040 | 0.076 | 0.033 | -0.048 | 0.081 | 0.033 |
| 0.2904 | 2.21 | 0.0676 | 0.020 | 0.087 | 0.040 | 0.016 | 0.092 | 0.040 |
| 0.2904 | 2.23 | 0.0663 | -0.121 | 0.109 | 0.022 | -0.132 | 0.114 | 0.022 |
| 0.2904 | 2.25 | 0.0649 | -0.009 | 0.119 | 0.041 | -0.014 | 0.126 | 0.041 |
| 0.2904 | 2.27 | 0.0636 | 0.056 | 0.168 | 0.046 | 0.054 | 0.176 | 0.046 |
| 0.3466 | 1.11 | 0.4963 | 0.007 | 0.017 | 0.016 | 0.012 | 0.059 | 0.016 |
| 0.3466 | 1.13 | 0.4664 | 0.007 | 0.022 | 0.013 | 0.005 | 0.072 | 0.013 |
| 0.3466 | 1.15 | 0.4394 | -0.021 | 0.025 | 0.012 | -0.074 | 0.074 | 0.012 |
| 0.3466 | 1.17 | 0.4150 | -0.079 | 0.027 | 0.014 | -0.210 | 0.074 | 0.014 |
| 0.3466 | 1.19 | 0.3928 | -0.152 | 0.030 | 0.017 | -0.358 | 0.076 | 0.017 |
| 0.3466 | 1.21 | 0.3726 | -0.251 | 0.030 | 0.016 | -0.572 | 0.073 | 0.016 |
| 0.3466 | 1.23 | 0.3540 | -0.210 | 0.030 | 0.016 | -0.459 | 0.068 | 0.016 |
| 0.3466 | 1.25 | 0.3369 | -0.089 | 0.028 | 0.024 | -0.196 | 0.061 | 0.024 |
| 0.3466 | 1.27 | 0.3212 | -0.053 | 0.027 | 0.052 | -0.134 | 0.056 | 0.052 |
| 0.3466 | 1.29 | 0.3066 | -0.009 | 0.028 | 0.009 | -0.054 | 0.054 | 0.009 |
| 0.3466 | 1.31 | 0.2932 | -0.015 | 0.029 | 0.015 | -0.079 | 0.055 | 0.015 |
| 0.3466 | 1.33 | 0.2806 | 0.097 | 0.033 | 0.011 | 0.116 | 0.059 | 0.011 |
| 0.3466 | 1.35 | 0.2690 | 0.076 | 0.031 | 0.011 | 0.067 | 0.055 | 0.011 |
| 0.3466 | 1.37 | 0.2581 | 0.102 | 0.034 | 0.010 | 0.104 | 0.058 | 0.010 |
| 0.3466 | 1.39 | 0.2479 | 0.059 | 0.036 | 0.008 | 0.023 | 0.059 | 0.008 |
| 0.3466 | 1.41 | 0.2383 | 0.054 | 0.036 | 0.012 | 0.007 | 0.057 | 0.012 |
| 0.3466 | 1.43 | 0.2294 | 0.041 | 0.036 | 0.009 | -0.022 | 0.055 | 0.009 |
| 0.3466 | 1.45 | 0.2210 | 0.090 | 0.038 | 0.011 | 0.043 | 0.058 | 0.011 |
| 0.3466 | 1.47 | 0.2130 | 0.189 | 0.041 | 0.019 | 0.176 | 0.061 | 0.019 |
| 0.3466 | 1.49 | 0.2055 | 0.204 | 0.044 | 0.022 | 0.188 | 0.063 | 0.022 |
| 0.3466 | 1.51 | 0.1985 | 0.209 | 0.045 | 0.024 | 0.192 | 0.063 | 0.024 |
| 0.3466 | 1.53 | 0.1918 | 0.260 | 0.042 | 0.028 | 0.261 | 0.057 | 0.028 |
| 0.3466 | 1.55 | 0.1855 | 0.155 | 0.041 | 0.026 | 0.118 | 0.055 | 0.026 |
| 0.3466 | 1.57 | 0.1795 | 0.099 | 0.044 | 0.021 | 0.041 | 0.058 | 0.021 |
| 0.3466 | 1.59 | 0.1738 | 0.053 | 0.039 | 0.037 | -0.018 | 0.052 | 0.037 |
| 0.3466 | 1.61 | 0.1684 | 0.105 | 0.045 | 0.021 | 0.052 | 0.058 | 0.021 |
| 0.3466 | 1.63 | 0.1633 | 0.052 | 0.044 | 0.015 | -0.014 | 0.056 | 0.015 |
| 0.3466 | 1.65 | 0.1584 | 0.053 | 0.046 | 0.020 | -0.008 | 0.058 | 0.020 |
| 0.3466 | 1.67 | 0.1537 | 0.058 | 0.048 | 0.016 | 0.003 | 0.060 | 0.016 |
| 0.3466 | 1.69 | 0.1493 | 0.099 | 0.052 | 0.019 | 0.061 | 0.064 | 0.019 |
| 0.3466 | 1.71 | 0.1450 | 0.166 | 0.049 | 0.017 | 0.151 | 0.060 | 0.017 |
| 0.3466 | 1.73 | 0.1410 | 0.133 | 0.052 | 0.012 | 0.118 | 0.062 | 0.012 |
| 0.3466 | 1.75 | 0.1371 | 0.114 | 0.059 | 0.018 | 0.101 | 0.070 | 0.018 |
| 0.3466 | 1.77 | 0.1334 | 0.094 | 0.058 | 0.016 | 0.081 | 0.069 | 0.016 |
| 0.3466 | 1.79 | 0.1298 | 0.097 | 0.056 | 0.023 | 0.088 | 0.066 | 0.023 |
| 0.3466 | 1.81 | 0.1264 | -0.067 | 0.059 | 0.021 | -0.102 | 0.069 | 0.021 |
| 0.3466 | 1.83 | 0.1231 | 0.065 | 0.056 | 0.017 | 0.052 | 0.065 | 0.017 |
| 0.3466 | 1.85 | 0.1200 | 0.052 | 0.060 | 0.021 | 0.038 | 0.069 | 0.021 |
| 0.3466 | 1.87 | 0.1170 | 0.033 | 0.058 | 0.018 | 0.017 | 0.066 | 0.018 |
| 0.3466 | 1.89 | 0.1141 | 0.092 | 0.057 | 0.018 | 0.086 | 0.065 | 0.018 |
| 0.3466 | 1.91 | 0.1113 | 0.084 | 0.066 | 0.020 | 0.078 | 0.074 | 0.020 |
| 0.3466 | 1.93 | 0.1086 | 0.099 | 0.065 | 0.020 | 0.096 | 0.073 | 0.020 |
| 0.3466 | 1.95 | 0.1060 | 0.015 | 0.062 | 0.018 | 0.003 | 0.070 | 0.018 |
| 0.3466 | 1.97 | 0.1036 | 0.028 | 0.071 | 0.032 | 0.019 | 0.079 | 0.032 |
| 0.3466 | 1.99 | 0.1012 | 0.048 | 0.068 | 0.030 | 0.043 | 0.075 | 0.030 |
| 0.3466 | 2.01 | 0.0989 | -0.007 | 0.060 | 0.030 | -0.017 | 0.066 | 0.030 |
| 0.3466 | 2.03 | 0.0966 | -0.038 | 0.066 | 0.023 | -0.051 | 0.073 | 0.023 |
| 0.3466 | 2.05 | 0.0945 | -0.031 | 0.068 | 0.042 | -0.043 | 0.075 | 0.042 |
| 0.3466 | 2.07 | 0.0924 | 0.064 | 0.060 | 0.026 | 0.061 | 0.066 | 0.026 |
| 0.3466 | 2.09 | 0.0904 | 0.034 | 0.069 | 0.021 | 0.028 | 0.075 | 0.021 |
| 0.3466 | 2.11 | 0.0885 | 0.100 | 0.075 | 0.019 | 0.100 | 0.081 | 0.019 |
| 0.3466 | 2.13 | 0.0866 | -0.012 | 0.074 | 0.025 | -0.021 | 0.080 | 0.025 |
| 0.3466 | 2.15 | 0.0848 | 0.097 | 0.073 | 0.017 | 0.096 | 0.078 | 0.017 |
| 0.3466 | 2.17 | 0.0830 | -0.070 | 0.096 | 0.023 | -0.083 | 0.103 | 0.023 |
| 0.3466 | 2.19 | 0.0813 | 0.181 | 0.098 | 0.029 | 0.186 | 0.105 | 0.029 |
| 0.3466 | 2.21 | 0.0797 | 0.003 | 0.102 | 0.037 | -0.005 | 0.108 | 0.037 |
| 0.3466 | 2.23 | 0.0781 | 0.264 | 0.141 | 0.066 | 0.273 | 0.150 | 0.066 |
| 0.3466 | 2.25 | 0.0765 | 0.618 | 0.438 | 0.374 | 0.649 | 0.464 | 0.374 |
| 0.4139 | 1.11 | 0.5406 | 0.002 | 0.014 | 0.020 | -0.003 | 0.048 | 0.020 |
| 0.4139 | 1.13 | 0.5107 | -0.002 | 0.017 | 0.008 | -0.022 | 0.055 | 0.008 |
| 0.4139 | 1.15 | 0.4835 | -0.061 | 0.022 | 0.025 | -0.191 | 0.067 | 0.025 |
| 0.4139 | 1.17 | 0.4586 | -0.111 | 0.026 | 0.011 | -0.302 | 0.072 | 0.011 |
| 0.4139 | 1.19 | 0.4358 | -0.168 | 0.031 | 0.008 | -0.407 | 0.080 | 0.008 |
| 0.4139 | 1.21 | 0.4149 | -0.185 | 0.033 | 0.015 | -0.416 | 0.080 | 0.015 |
| 0.4139 | 1.23 | 0.3955 | -0.178 | 0.035 | 0.009 | -0.395 | 0.082 | 0.009 |
| 0.4139 | 1.25 | 0.3776 | -0.096 | 0.034 | 0.011 | -0.214 | 0.075 | 0.011 |
| 0.4139 | 1.27 | 0.3610 | -0.061 | 0.034 | 0.010 | -0.154 | 0.072 | 0.010 |
| 0.4139 | 1.29 | 0.3456 | -0.052 | 0.034 | 0.012 | -0.144 | 0.069 | 0.012 |
| 0.4139 | 1.31 | 0.3312 | 0.017 | 0.031 | 0.011 | -0.016 | 0.060 | 0.011 |
| 0.4139 | 1.33 | 0.3178 | 0.031 | 0.031 | 0.011 | -0.003 | 0.057 | 0.011 |
| 0.4139 | 1.35 | 0.3052 | 0.107 | 0.031 | 0.014 | 0.124 | 0.056 | 0.014 |
| 0.4139 | 1.37 | 0.2935 | 0.071 | 0.035 | 0.011 | 0.052 | 0.061 | 0.011 |
| 0.4139 | 1.39 | 0.2824 | 0.086 | 0.032 | 0.010 | 0.068 | 0.053 | 0.010 |
| 0.4139 | 1.41 | 0.2720 | 0.058 | 0.031 | 0.013 | 0.013 | 0.050 | 0.013 |
| 0.4139 | 1.43 | 0.2622 | 0.149 | 0.035 | 0.011 | 0.150 | 0.056 | 0.011 |
| 0.4139 | 1.45 | 0.2530 | 0.133 | 0.037 | 0.012 | 0.110 | 0.057 | 0.012 |
| 0.4139 | 1.47 | 0.2443 | 0.207 | 0.041 | 0.022 | 0.205 | 0.061 | 0.022 |
| 0.4139 | 1.49 | 0.2360 | 0.256 | 0.044 | 0.026 | 0.264 | 0.064 | 0.026 |
| 0.4139 | 1.51 | 0.2282 | 0.228 | 0.043 | 0.025 | 0.218 | 0.063 | 0.025 |
| 0.4139 | 1.53 | 0.2208 | 0.227 | 0.046 | 0.028 | 0.216 | 0.065 | 0.028 |
| 0.4139 | 1.55 | 0.2138 | 0.135 | 0.042 | 0.023 | 0.086 | 0.059 | 0.023 |
| 0.4139 | 1.57 | 0.2071 | 0.164 | 0.042 | 0.023 | 0.125 | 0.058 | 0.023 |
| 0.4139 | 1.59 | 0.2008 | 0.141 | 0.037 | 0.025 | 0.092 | 0.049 | 0.025 |
| 0.4139 | 1.61 | 0.1947 | 0.165 | 0.042 | 0.024 | 0.124 | 0.055 | 0.024 |
| 0.4139 | 1.63 | 0.1890 | 0.077 | 0.047 | 0.033 | 0.009 | 0.061 | 0.033 |
| 0.4139 | 1.65 | 0.1835 | 0.093 | 0.054 | 0.019 | 0.032 | 0.069 | 0.019 |
| 0.4139 | 1.67 | 0.1782 | 0.046 | 0.062 | 0.027 | -0.024 | 0.078 | 0.027 |
| 0.4139 | 1.69 | 0.1732 | 0.020 | 0.060 | 0.025 | -0.046 | 0.075 | 0.025 |
| 0.4139 | 1.71 | 0.1684 | 0.101 | 0.057 | 0.028 | 0.065 | 0.071 | 0.028 |
| 0.4139 | 1.73 | 0.1638 | 0.051 | 0.054 | 0.009 | 0.013 | 0.066 | 0.009 |
| 0.4139 | 1.75 | 0.1594 | 0.075 | 0.060 | 0.018 | 0.049 | 0.073 | 0.018 |
| 0.4139 | 1.77 | 0.1552 | 0.032 | 0.054 | 0.010 | 0.003 | 0.065 | 0.010 |
| 0.4139 | 1.79 | 0.1512 | 0.039 | 0.052 | 0.016 | 0.016 | 0.062 | 0.016 |
| 0.4139 | 1.81 | 0.1473 | 0.085 | 0.055 | 0.021 | 0.073 | 0.065 | 0.021 |
| 0.4139 | 1.83 | 0.1436 | 0.092 | 0.051 | 0.015 | 0.082 | 0.060 | 0.015 |
| 0.4139 | 1.85 | 0.1400 | 0.022 | 0.059 | 0.021 | 0.001 | 0.068 | 0.021 |
| 0.4139 | 1.87 | 0.1366 | 0.008 | 0.056 | 0.031 | -0.015 | 0.065 | 0.031 |
| 0.4139 | 1.89 | 0.1333 | 0.053 | 0.052 | 0.017 | 0.039 | 0.059 | 0.017 |
| 0.4139 | 1.91 | 0.1301 | -0.036 | 0.064 | 0.016 | -0.061 | 0.073 | 0.016 |
| 0.4139 | 1.93 | 0.1270 | 0.004 | 0.061 | 0.022 | -0.014 | 0.069 | 0.022 |
| 0.4139 | 1.95 | 0.1241 | 0.044 | 0.070 | 0.016 | 0.033 | 0.079 | 0.016 |
| 0.4139 | 1.97 | 0.1212 | 0.041 | 0.069 | 0.017 | 0.031 | 0.077 | 0.017 |
| 0.4139 | 1.99 | 0.1185 | 0.010 | 0.065 | 0.025 | -0.002 | 0.072 | 0.025 |
| 0.4139 | 2.01 | 0.1158 | -0.042 | 0.061 | 0.019 | -0.059 | 0.068 | 0.019 |
| 0.4139 | 2.03 | 0.1133 | 0.112 | 0.067 | 0.020 | 0.111 | 0.075 | 0.020 |
| 0.4139 | 2.05 | 0.1108 | 0.022 | 0.063 | 0.017 | 0.012 | 0.069 | 0.017 |
| 0.4139 | 2.07 | 0.1084 | 0.142 | 0.069 | 0.023 | 0.145 | 0.075 | 0.023 |
| 0.4139 | 2.09 | 0.1061 | 0.043 | 0.064 | 0.022 | 0.036 | 0.070 | 0.022 |
| 0.4139 | 2.11 | 0.1039 | 0.070 | 0.071 | 0.020 | 0.065 | 0.078 | 0.020 |
| 0.4139 | 2.13 | 0.1017 | -0.001 | 0.094 | 0.034 | -0.013 | 0.102 | 0.034 |
| 0.4139 | 2.15 | 0.0996 | 0.183 | 0.168 | 0.038 | 0.188 | 0.182 | 0.038 |
| 0.4139 | 2.17 | 0.0976 | -0.560 | 0.406 | 0.415 | -0.616 | 0.439 | 0.415 |
| 0.4139 | 2.19 | 0.0956 | -0.041 | 0.251 | 0.109 | -0.054 | 0.270 | 0.109 |
| 0.4139 | 2.21 | 0.0937 | -0.158 | 0.147 | 0.057 | -0.180 | 0.158 | 0.057 |
| 0.4139 | 2.23 | 0.0918 | 0.080 | 0.169 | 0.025 | 0.075 | 0.182 | 0.025 |
| 0.494 | 1.11 | 0.5841 | 0.005 | 0.016 | 0.003 | 0.009 | 0.055 | 0.003 |
| 0.494 | 1.13 | 0.5547 | 0.006 | 0.029 | 0.005 | 0.011 | 0.092 | 0.005 |
| 0.494 | 1.15 | 0.5277 | -0.019 | 0.025 | 0.005 | -0.061 | 0.075 | 0.005 |
| 0.494 | 1.17 | 0.5028 | -0.008 | 0.029 | 0.011 | -0.012 | 0.080 | 0.011 |
| 0.494 | 1.19 | 0.4798 | -0.166 | 0.033 | 0.008 | -0.411 | 0.089 | 0.008 |
| 0.494 | 1.21 | 0.4584 | -0.146 | 0.038 | 0.010 | -0.328 | 0.096 | 0.010 |
| 0.494 | 1.23 | 0.4385 | -0.086 | 0.033 | 0.006 | -0.186 | 0.079 | 0.006 |
| 0.494 | 1.25 | 0.4200 | -0.085 | 0.030 | 0.008 | -0.197 | 0.069 | 0.008 |
| 0.494 | 1.27 | 0.4028 | -0.030 | 0.031 | 0.008 | -0.087 | 0.067 | 0.008 |
| 0.494 | 1.29 | 0.3866 | 0.041 | 0.031 | 0.008 | 0.048 | 0.064 | 0.008 |
| 0.494 | 1.31 | 0.3715 | 0.036 | 0.031 | 0.015 | 0.023 | 0.061 | 0.015 |
| 0.494 | 1.33 | 0.3573 | -0.002 | 0.028 | 0.032 | -0.067 | 0.054 | 0.032 |
| 0.494 | 1.35 | 0.3440 | 0.087 | 0.028 | 0.014 | 0.087 | 0.051 | 0.014 |
| 0.494 | 1.37 | 0.3314 | 0.050 | 0.029 | 0.014 | 0.012 | 0.052 | 0.014 |
| 0.494 | 1.39 | 0.3196 | 0.089 | 0.032 | 0.015 | 0.074 | 0.055 | 0.015 |
| 0.494 | 1.41 | 0.3084 | 0.088 | 0.029 | 0.015 | 0.062 | 0.049 | 0.015 |
| 0.494 | 1.43 | 0.2979 | 0.160 | 0.033 | 0.017 | 0.170 | 0.053 | 0.017 |
| 0.494 | 1.45 | 0.2879 | 0.168 | 0.036 | 0.020 | 0.167 | 0.058 | 0.020 |
| 0.494 | 1.47 | 0.2784 | 0.183 | 0.044 | 0.024 | 0.170 | 0.068 | 0.024 |
| 0.494 | 1.49 | 0.2694 | 0.215 | 0.043 | 0.029 | 0.203 | 0.066 | 0.029 |
| 0.494 | 1.51 | 0.2609 | 0.252 | 0.046 | 0.030 | 0.250 | 0.068 | 0.030 |
| 0.494 | 1.53 | 0.2528 | 0.215 | 0.048 | 0.026 | 0.195 | 0.069 | 0.026 |
| 0.494 | 1.55 | 0.2450 | 0.138 | 0.046 | 0.026 | 0.084 | 0.066 | 0.026 |
| 0.494 | 1.57 | 0.2377 | 0.090 | 0.048 | 0.025 | 0.015 | 0.067 | 0.025 |
| 0.494 | 1.59 | 0.2307 | 0.196 | 0.039 | 0.034 | 0.161 | 0.054 | 0.034 |
| 0.494 | 1.61 | 0.2240 | 0.080 | 0.047 | 0.027 | -0.001 | 0.063 | 0.027 |
| 0.494 | 1.63 | 0.2176 | 0.113 | 0.049 | 0.041 | 0.043 | 0.066 | 0.041 |
| 0.494 | 1.65 | 0.2115 | 0.181 | 0.049 | 0.018 | 0.134 | 0.064 | 0.018 |
| 0.494 | 1.67 | 0.2056 | 0.046 | 0.050 | 0.022 | -0.038 | 0.065 | 0.022 |
| 0.494 | 1.69 | 0.2000 | 0.145 | 0.042 | 0.024 | 0.099 | 0.054 | 0.024 |
| 0.494 | 1.71 | 0.1947 | 0.146 | 0.047 | 0.017 | 0.112 | 0.060 | 0.017 |
| 0.494 | 1.73 | 0.1895 | 0.076 | 0.046 | 0.007 | 0.035 | 0.058 | 0.007 |
| 0.494 | 1.75 | 0.1846 | 0.059 | 0.048 | 0.025 | 0.023 | 0.059 | 0.025 |
| 0.494 | 1.77 | 0.1799 | 0.119 | 0.045 | 0.011 | 0.104 | 0.056 | 0.011 |
| 0.494 | 1.79 | 0.1753 | 0.187 | 0.046 | 0.015 | 0.192 | 0.056 | 0.015 |
| 0.494 | 1.81 | 0.1710 | 0.056 | 0.048 | 0.015 | 0.036 | 0.058 | 0.015 |
| 0.494 | 1.83 | 0.1668 | 0.125 | 0.053 | 0.019 | 0.119 | 0.064 | 0.019 |
| 0.494 | 1.85 | 0.1627 | 0.094 | 0.060 | 0.019 | 0.083 | 0.072 | 0.019 |
| 0.494 | 1.87 | 0.1588 | 0.048 | 0.051 | 0.016 | 0.028 | 0.061 | 0.016 |
| 0.494 | 1.89 | 0.1551 | 0.045 | 0.054 | 0.016 | 0.025 | 0.063 | 0.016 |
| 0.494 | 1.91 | 0.1515 | -0.049 | 0.057 | 0.024 | -0.082 | 0.066 | 0.024 |
| 0.494 | 1.93 | 0.1480 | 0.041 | 0.057 | 0.019 | 0.024 | 0.066 | 0.019 |
| 0.494 | 1.95 | 0.1446 | -0.069 | 0.068 | 0.032 | -0.100 | 0.079 | 0.032 |
| 0.494 | 1.97 | 0.1414 | -0.026 | 0.067 | 0.032 | -0.049 | 0.076 | 0.032 |
| 0.494 | 1.99 | 0.1382 | 0.119 | 0.066 | 0.021 | 0.117 | 0.075 | 0.021 |
| 0.494 | 2.01 | 0.1352 | 0.068 | 0.067 | 0.020 | 0.060 | 0.076 | 0.020 |
| 0.494 | 2.03 | 0.1323 | 0.073 | 0.090 | 0.018 | 0.066 | 0.102 | 0.018 |
| 0.494 | 2.05 | 0.1295 | 0.156 | 0.104 | 0.019 | 0.158 | 0.117 | 0.019 |
| 0.494 | 2.07 | 0.1267 | -0.051 | 0.187 | 0.089 | -0.073 | 0.209 | 0.089 |
| 0.494 | 2.09 | 0.1241 | 0.247 | 0.195 | 0.179 | 0.258 | 0.216 | 0.179 |
| 0.494 | 2.11 | 0.1215 | 0.107 | 0.183 | 0.032 | 0.103 | 0.201 | 0.032 |
| 0.494 | 2.13 | 0.1190 | 0.072 | 0.128 | 0.037 | 0.064 | 0.141 | 0.037 |
| 0.494 | 2.15 | 0.1166 | 0.106 | 0.074 | 0.035 | 0.102 | 0.081 | 0.035 |
| 0.494 | 2.17 | 0.1143 | -0.020 | 0.081 | 0.037 | -0.036 | 0.089 | 0.037 |
| 0.494 | 2.19 | 0.1120 | 0.162 | 0.143 | 0.075 | 0.162 | 0.156 | 0.075 |
| 0.494 | 2.21 | 0.1098 | 0.038 | 0.213 | 0.017 | 0.028 | 0.231 | 0.017 |
| 0.5902 | 1.11 | 0.6266 | 0.002 | 0.018 | 0.003 | 0.002 | 0.060 | 0.003 |
| 0.5902 | 1.13 | 0.5981 | 0.001 | 0.015 | 0.015 | -0.005 | 0.046 | 0.015 |
| 0.5902 | 1.15 | 0.5717 | -0.050 | 0.026 | 0.013 | -0.150 | 0.078 | 0.013 |
| 0.5902 | 1.17 | 0.5471 | -0.020 | 0.030 | 0.006 | -0.045 | 0.085 | 0.006 |
| 0.5902 | 1.19 | 0.5242 | -0.056 | 0.032 | 0.006 | -0.121 | 0.084 | 0.006 |
| 0.5902 | 1.21 | 0.5027 | -0.169 | 0.036 | 0.011 | -0.395 | 0.091 | 0.011 |
| 0.5902 | 1.23 | 0.4827 | -0.135 | 0.032 | 0.010 | -0.308 | 0.077 | 0.010 |
| 0.5902 | 1.25 | 0.4639 | -0.021 | 0.028 | 0.010 | -0.053 | 0.064 | 0.010 |
| 0.5902 | 1.27 | 0.4462 | -0.040 | 0.026 | 0.010 | -0.110 | 0.058 | 0.010 |
| 0.5902 | 1.29 | 0.4296 | 0.026 | 0.026 | 0.010 | 0.021 | 0.055 | 0.010 |
| 0.5902 | 1.31 | 0.4139 | 0.002 | 0.026 | 0.013 | -0.046 | 0.052 | 0.013 |
| 0.5902 | 1.33 | 0.3991 | 0.048 | 0.025 | 0.016 | 0.028 | 0.049 | 0.016 |
| 0.5902 | 1.35 | 0.3852 | 0.092 | 0.025 | 0.021 | 0.097 | 0.047 | 0.021 |
| 0.5902 | 1.37 | 0.3719 | 0.143 | 0.035 | 0.016 | 0.179 | 0.065 | 0.016 |
| 0.5902 | 1.39 | 0.3594 | 0.094 | 0.032 | 0.017 | 0.079 | 0.056 | 0.017 |
| 0.5902 | 1.41 | 0.3476 | 0.120 | 0.035 | 0.019 | 0.115 | 0.060 | 0.019 |
| 0.5902 | 1.43 | 0.3363 | 0.136 | 0.033 | 0.019 | 0.129 | 0.056 | 0.019 |
| 0.5902 | 1.45 | 0.3256 | 0.157 | 0.034 | 0.022 | 0.149 | 0.056 | 0.022 |
| 0.5902 | 1.47 | 0.3155 | 0.166 | 0.035 | 0.029 | 0.140 | 0.056 | 0.029 |
| 0.5902 | 1.49 | 0.3058 | 0.165 | 0.044 | 0.025 | 0.120 | 0.069 | 0.025 |
| 0.5902 | 1.51 | 0.2966 | 0.240 | 0.042 | 0.033 | 0.225 | 0.064 | 0.033 |
| 0.5902 | 1.53 | 0.2878 | 0.241 | 0.042 | 0.030 | 0.227 | 0.062 | 0.030 |
| 0.5902 | 1.55 | 0.2794 | 0.155 | 0.042 | 0.032 | 0.098 | 0.062 | 0.032 |
| 0.5902 | 1.57 | 0.2714 | 0.192 | 0.037 | 0.031 | 0.151 | 0.054 | 0.031 |
| 0.5902 | 1.59 | 0.2637 | 0.161 | 0.037 | 0.029 | 0.102 | 0.053 | 0.029 |
| 0.5902 | 1.61 | 0.2564 | 0.191 | 0.042 | 0.032 | 0.139 | 0.059 | 0.032 |
| 0.5902 | 1.63 | 0.2494 | 0.099 | 0.042 | 0.031 | 0.010 | 0.058 | 0.031 |
| 0.5902 | 1.65 | 0.2426 | 0.077 | 0.044 | 0.031 | -0.021 | 0.059 | 0.031 |
| 0.5902 | 1.67 | 0.2362 | 0.133 | 0.049 | 0.030 | 0.058 | 0.065 | 0.030 |
| 0.5902 | 1.69 | 0.2300 | 0.136 | 0.053 | 0.024 | 0.072 | 0.070 | 0.024 |
| 0.5902 | 1.71 | 0.2241 | 0.204 | 0.044 | 0.019 | 0.177 | 0.057 | 0.019 |
| 0.5902 | 1.73 | 0.2184 | 0.109 | 0.044 | 0.013 | 0.067 | 0.056 | 0.013 |
| 0.5902 | 1.75 | 0.2129 | 0.120 | 0.044 | 0.011 | 0.094 | 0.056 | 0.011 |
| 0.5902 | 1.77 | 0.2076 | 0.157 | 0.043 | 0.013 | 0.148 | 0.054 | 0.013 |
| 0.5902 | 1.79 | 0.2025 | 0.070 | 0.045 | 0.016 | 0.048 | 0.056 | 0.016 |
| 0.5902 | 1.81 | 0.1977 | 0.107 | 0.049 | 0.031 | 0.098 | 0.060 | 0.031 |
| 0.5902 | 1.83 | 0.1930 | 0.124 | 0.048 | 0.023 | 0.118 | 0.059 | 0.023 |
| 0.5902 | 1.85 | 0.1884 | -0.060 | 0.060 | 0.026 | -0.106 | 0.073 | 0.026 |
| 0.5902 | 1.87 | 0.1840 | 0.143 | 0.076 | 0.016 | 0.140 | 0.091 | 0.016 |
| 0.5902 | 1.89 | 0.1798 | 0.061 | 0.062 | 0.013 | 0.043 | 0.074 | 0.013 |
| 0.5902 | 1.91 | 0.1758 | -0.052 | 0.091 | 0.018 | -0.090 | 0.108 | 0.018 |
| 0.5902 | 1.93 | 0.1718 | -0.078 | 0.110 | 0.048 | -0.119 | 0.130 | 0.048 |
| 0.5902 | 1.95 | 0.1680 | 0.052 | 0.128 | 0.044 | 0.036 | 0.150 | 0.044 |
| 0.5902 | 1.97 | 0.1644 | 0.480 | 0.245 | 0.158 | 0.536 | 0.285 | 0.158 |
| 0.5902 | 1.99 | 0.1608 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 |
| 0.5902 | 2.01 | 0.1574 | 0.096 | 0.145 | 0.068 | 0.089 | 0.166 | 0.068 |
| 0.5902 | 2.03 | 0.1541 | 0.025 | 0.128 | 0.107 | 0.009 | 0.146 | 0.107 |
| 0.5902 | 2.05 | 0.1509 | 0.115 | 0.092 | 0.016 | 0.110 | 0.104 | 0.016 |
| 0.5902 | 2.07 | 0.1477 | 0.084 | 0.084 | 0.027 | 0.076 | 0.095 | 0.027 |
| 0.5902 | 2.09 | 0.1447 | -0.054 | 0.096 | 0.018 | -0.080 | 0.108 | 0.018 |
| 0.5902 | 2.11 | 0.1418 | 0.123 | 0.089 | 0.023 | 0.119 | 0.100 | 0.023 |
| 0.5902 | 2.13 | 0.1390 | 0.095 | 0.079 | 0.019 | 0.088 | 0.088 | 0.019 |
| 0.5902 | 2.15 | 0.1362 | 0.156 | 0.121 | 0.038 | 0.156 | 0.134 | 0.038 |
| 0.5902 | 2.17 | 0.1336 | 0.001 | 0.215 | 0.290 | -0.017 | 0.238 | 0.290 |
| 0.7047 | 1.11 | 0.6671 | 0.005 | 0.033 | 0.008 | 0.014 | 0.106 | 0.008 |
| 0.7047 | 1.13 | 0.6399 | 0.006 | 0.026 | 0.009 | 0.014 | 0.080 | 0.009 |
| 0.7047 | 1.15 | 0.6145 | -0.057 | 0.040 | 0.008 | -0.170 | 0.118 | 0.008 |
| 0.7047 | 1.17 | 0.5906 | -0.057 | 0.033 | 0.008 | -0.144 | 0.091 | 0.008 |
| 0.7047 | 1.19 | 0.5681 | -0.053 | 0.040 | 0.012 | -0.116 | 0.103 | 0.012 |
| 0.7047 | 1.21 | 0.5470 | -0.057 | 0.030 | 0.005 | -0.119 | 0.074 | 0.005 |
| 0.7047 | 1.23 | 0.5270 | -0.080 | 0.031 | 0.007 | -0.179 | 0.075 | 0.007 |
| 0.7047 | 1.25 | 0.5081 | -0.041 | 0.029 | 0.010 | -0.101 | 0.067 | 0.010 |
| 0.7047 | 1.27 | 0.4903 | -0.071 | 0.028 | 0.011 | -0.177 | 0.062 | 0.011 |
| 0.7047 | 1.29 | 0.4735 | 0.041 | 0.025 | 0.009 | 0.053 | 0.053 | 0.009 |
| 0.7047 | 1.31 | 0.4575 | 0.024 | 0.022 | 0.012 | 0.001 | 0.045 | 0.012 |
| 0.7047 | 1.33 | 0.4423 | 0.039 | 0.024 | 0.022 | 0.013 | 0.047 | 0.022 |
| 0.7047 | 1.35 | 0.4279 | 0.088 | 0.021 | 0.022 | 0.091 | 0.041 | 0.022 |
| 0.7047 | 1.37 | 0.4142 | 0.097 | 0.026 | 0.019 | 0.094 | 0.048 | 0.019 |
| 0.7047 | 1.39 | 0.4012 | 0.091 | 0.028 | 0.018 | 0.072 | 0.051 | 0.018 |
| 0.7047 | 1.41 | 0.3888 | 0.087 | 0.029 | 0.020 | 0.058 | 0.051 | 0.020 |
| 0.7047 | 1.43 | 0.3770 | 0.098 | 0.026 | 0.022 | 0.065 | 0.044 | 0.022 |
| 0.7047 | 1.45 | 0.3657 | 0.114 | 0.027 | 0.026 | 0.076 | 0.046 | 0.026 |
| 0.7047 | 1.47 | 0.3550 | 0.180 | 0.032 | 0.030 | 0.163 | 0.053 | 0.030 |
| 0.7047 | 1.49 | 0.3447 | 0.202 | 0.034 | 0.041 | 0.172 | 0.055 | 0.041 |
| 0.7047 | 1.51 | 0.3349 | 0.215 | 0.039 | 0.038 | 0.178 | 0.061 | 0.038 |
| 0.7047 | 1.53 | 0.3255 | 0.226 | 0.035 | 0.040 | 0.199 | 0.053 | 0.040 |
| 0.7047 | 1.55 | 0.3165 | 0.180 | 0.034 | 0.041 | 0.129 | 0.052 | 0.041 |
| 0.7047 | 1.57 | 0.3078 | 0.193 | 0.035 | 0.036 | 0.146 | 0.052 | 0.036 |
| 0.7047 | 1.59 | 0.2996 | 0.161 | 0.038 | 0.039 | 0.095 | 0.055 | 0.039 |
| 0.7047 | 1.61 | 0.2916 | 0.174 | 0.040 | 0.034 | 0.109 | 0.057 | 0.034 |
| 0.7047 | 1.63 | 0.2840 | 0.148 | 0.040 | 0.042 | 0.067 | 0.055 | 0.042 |
| 0.7047 | 1.65 | 0.2767 | 0.185 | 0.046 | 0.039 | 0.115 | 0.063 | 0.039 |
| 0.7047 | 1.67 | 0.2697 | 0.194 | 0.049 | 0.037 | 0.130 | 0.067 | 0.037 |
| 0.7047 | 1.69 | 0.2629 | 0.232 | 0.051 | 0.037 | 0.191 | 0.069 | 0.037 |
| 0.7047 | 1.71 | 0.2564 | 0.141 | 0.054 | 0.033 | 0.086 | 0.072 | 0.033 |
| 0.7047 | 1.73 | 0.2502 | 0.191 | 0.060 | 0.014 | 0.170 | 0.079 | 0.014 |
| 0.7047 | 1.75 | 0.2441 | 0.122 | 0.068 | 0.028 | 0.091 | 0.088 | 0.028 |
| 0.7047 | 1.77 | 0.2383 | 0.081 | 0.078 | 0.020 | 0.050 | 0.100 | 0.020 |
| 0.7047 | 1.79 | 0.2327 | 0.093 | 0.117 | 0.013 | 0.077 | 0.149 | 0.013 |
| 0.7047 | 1.81 | 0.2273 | 0.166 | 0.166 | 0.063 | 0.174 | 0.209 | 0.063 |
| 0.7047 | 1.83 | 0.2221 | 0.637 | 0.506 | 0.449 | 0.763 | 0.633 | 0.449 |
| 0.7047 | 1.85 | 0.2171 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 |
| 0.7047 | 1.87 | 0.2122 | -0.082 | 0.566 | 0.090 | -0.132 | 0.691 | 0.090 |
| 0.7047 | 1.89 | 0.2075 | -0.204 | 0.184 | 0.091 | -0.278 | 0.223 | 0.091 |
| 0.7047 | 1.91 | 0.2030 | -0.090 | 0.082 | 0.018 | -0.138 | 0.099 | 0.018 |
| 0.7047 | 1.93 | 0.1986 | 0.081 | 0.055 | 0.020 | 0.069 | 0.066 | 0.020 |
| 0.7047 | 1.95 | 0.1943 | 0.043 | 0.072 | 0.049 | 0.024 | 0.086 | 0.049 |
| 0.7047 | 1.97 | 0.1902 | 0.094 | 0.068 | 0.032 | 0.085 | 0.080 | 0.032 |
| 0.7047 | 1.99 | 0.1862 | 0.122 | 0.053 | 0.024 | 0.119 | 0.063 | 0.024 |
| 0.7047 | 2.01 | 0.1824 | 0.151 | 0.065 | 0.019 | 0.152 | 0.076 | 0.019 |
| 0.7047 | 2.03 | 0.1786 | 0.183 | 0.071 | 0.030 | 0.188 | 0.082 | 0.030 |
| 0.7047 | 2.05 | 0.1750 | 0.135 | 0.090 | 0.026 | 0.133 | 0.104 | 0.026 |
| 0.7047 | 2.07 | 0.1715 | 0.186 | 0.143 | 0.055 | 0.190 | 0.164 | 0.055 |
| 0.7047 | 2.09 | 0.1681 | 0.157 | 0.332 | 0.090 | 0.157 | 0.379 | 0.090 |
| 0.7047 | 2.11 | 0.1648 | -0.226 | 0.356 | 0.200 | -0.279 | 0.404 | 0.200 |
| 0.8412 | 1.11 | 0.7052 | 0.005 | 0.008 | 0.007 | 0.014 | 0.025 | 0.007 |
| 0.8412 | 1.13 | 0.6796 | -0.003 | 0.009 | 0.005 | -0.011 | 0.025 | 0.005 |
| 0.8412 | 1.15 | 0.6555 | -0.020 | 0.012 | 0.006 | -0.057 | 0.033 | 0.006 |
| 0.8412 | 1.17 | 0.6326 | -0.041 | 0.015 | 0.007 | -0.104 | 0.042 | 0.007 |
| 0.8412 | 1.19 | 0.6109 | -0.064 | 0.023 | 0.005 | -0.150 | 0.060 | 0.005 |
| 0.8412 | 1.21 | 0.5903 | -0.074 | 0.023 | 0.005 | -0.163 | 0.056 | 0.005 |
| 0.8412 | 1.23 | 0.5708 | -0.040 | 0.024 | 0.007 | -0.085 | 0.057 | 0.007 |
| 0.8412 | 1.25 | 0.5522 | -0.057 | 0.025 | 0.008 | -0.136 | 0.057 | 0.008 |
| 0.8412 | 1.27 | 0.5345 | -0.018 | 0.023 | 0.010 | -0.060 | 0.050 | 0.010 |
| 0.8412 | 1.29 | 0.5177 | 0.024 | 0.021 | 0.010 | 0.021 | 0.044 | 0.010 |
| 0.8412 | 1.31 | 0.5016 | 0.014 | 0.026 | 0.013 | -0.015 | 0.054 | 0.013 |
| 0.8412 | 1.33 | 0.4863 | 0.060 | 0.019 | 0.016 | 0.058 | 0.037 | 0.016 |
| 0.8412 | 1.35 | 0.4717 | 0.073 | 0.024 | 0.018 | 0.070 | 0.047 | 0.018 |
| 0.8412 | 1.37 | 0.4577 | 0.097 | 0.022 | 0.023 | 0.100 | 0.041 | 0.023 |
| 0.8412 | 1.39 | 0.4444 | 0.108 | 0.022 | 0.023 | 0.113 | 0.041 | 0.023 |
| 0.8412 | 1.41 | 0.4316 | 0.098 | 0.026 | 0.023 | 0.085 | 0.047 | 0.023 |
| 0.8412 | 1.43 | 0.4194 | 0.086 | 0.029 | 0.022 | 0.054 | 0.051 | 0.022 |
| 0.8412 | 1.45 | 0.4077 | 0.137 | 0.025 | 0.030 | 0.127 | 0.042 | 0.030 |
| 0.8412 | 1.47 | 0.3965 | 0.140 | 0.040 | 0.030 | 0.111 | 0.066 | 0.030 |
| 0.8412 | 1.49 | 0.3857 | 0.237 | 0.043 | 0.039 | 0.244 | 0.070 | 0.039 |
| 0.8412 | 1.51 | 0.3754 | 0.187 | 0.044 | 0.045 | 0.152 | 0.069 | 0.045 |
| 0.8412 | 1.53 | 0.3655 | 0.240 | 0.048 | 0.047 | 0.236 | 0.076 | 0.047 |
| 0.8412 | 1.55 | 0.3559 | 0.156 | 0.062 | 0.046 | 0.108 | 0.095 | 0.046 |
| 0.8412 | 1.57 | 0.3468 | 0.182 | 0.062 | 0.040 | 0.145 | 0.093 | 0.040 |
| 0.8412 | 1.59 | 0.3380 | 0.192 | 0.065 | 0.044 | 0.153 | 0.097 | 0.044 |
| 0.8412 | 1.61 | 0.3295 | 0.212 | 0.061 | 0.061 | 0.174 | 0.088 | 0.061 |
| 0.8412 | 1.63 | 0.3214 | 0.171 | 0.065 | 0.077 | 0.110 | 0.093 | 0.077 |
| 0.8412 | 1.65 | 0.3135 | 0.379 | 0.174 | 0.175 | 0.400 | 0.246 | 0.175 |
| 0.8412 | 1.67 | 0.3059 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 |
| 0.8412 | 1.69 | 0.2986 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 | 0.000 |
| 0.8412 | 1.71 | 0.2916 | 0.271 | 0.252 | 0.185 | 0.270 | 0.341 | 0.185 |
| 0.8412 | 1.73 | 0.2848 | 0.090 | 0.109 | 0.048 | 0.042 | 0.146 | 0.048 |
| 0.8412 | 1.75 | 0.2782 | 0.000 | 0.082 | 0.034 | -0.064 | 0.109 | 0.034 |
| 0.8412 | 1.77 | 0.2719 | 0.117 | 0.091 | 0.041 | 0.103 | 0.118 | 0.041 |
| 0.8412 | 1.79 | 0.2658 | 0.036 | 0.057 | 0.031 | 0.007 | 0.074 | 0.031 |
| 0.8412 | 1.81 | 0.2599 | 0.103 | 0.052 | 0.023 | 0.098 | 0.067 | 0.023 |
| 0.8412 | 1.83 | 0.2542 | 0.096 | 0.040 | 0.016 | 0.089 | 0.051 | 0.016 |
| 0.8412 | 1.85 | 0.2486 | 0.113 | 0.043 | 0.020 | 0.109 | 0.055 | 0.020 |
| 0.8412 | 1.87 | 0.2433 | 0.111 | 0.058 | 0.020 | 0.106 | 0.072 | 0.020 |
| 0.8412 | 1.89 | 0.2381 | 0.066 | 0.060 | 0.022 | 0.051 | 0.074 | 0.022 |
| 0.8412 | 1.91 | 0.2331 | 0.000 | 0.060 | 0.012 | -0.030 | 0.073 | 0.012 |
| 0.8412 | 1.93 | 0.2282 | 0.070 | 0.057 | 0.021 | 0.057 | 0.069 | 0.021 |
| 0.8412 | 1.95 | 0.2235 | 0.069 | 0.095 | 0.016 | 0.056 | 0.114 | 0.016 |
| 0.8412 | 1.97 | 0.2190 | 0.136 | 0.083 | 0.016 | 0.137 | 0.100 | 0.016 |
| 0.8412 | 1.99 | 0.2145 | 0.059 | 0.133 | 0.047 | 0.045 | 0.158 | 0.047 |
| 0.8412 | 2.01 | 0.2103 | 0.205 | 0.126 | 0.111 | 0.218 | 0.149 | 0.111 |
| 0.8412 | 2.03 | 0.2061 | 0.219 | 0.137 | 0.190 | 0.234 | 0.162 | 0.190 |
Appendix B Table for Results on and
Results for and as shown in Figs. 11 and 13, respectively.
| W | Stat. | Syst. | Stat. | Syst. | ||||
| (GeV2) | (GeV) | |||||||
| 0.0171 | 1.13 | 0.0413 | 0.353 | 0.413 | 0.060 | 0.436 | 0.558 | 0.077 |
| 0.0171 | 1.15 | 0.0372 | 0.760 | 0.362 | 0.102 | 0.936 | 0.464 | 0.129 |
| 0.0171 | 1.17 | 0.0338 | -0.536 | 0.348 | 0.170 | -0.692 | 0.430 | 0.214 |
| 0.0171 | 1.19 | 0.0309 | -0.400 | 0.353 | 0.160 | -0.496 | 0.422 | 0.208 |
| 0.0171 | 1.21 | 0.0285 | -0.900 | 0.369 | 0.174 | -1.060 | 0.431 | 0.221 |
| 0.0171 | 1.23 | 0.0263 | -1.209 | 0.349 | 0.183 | -1.391 | 0.398 | 0.220 |
| 0.0171 | 1.25 | 0.0245 | -0.691 | 0.337 | 0.133 | -0.790 | 0.378 | 0.155 |
| 0.0171 | 1.27 | 0.0228 | -0.580 | 0.319 | 0.109 | -0.658 | 0.353 | 0.122 |
| 0.0171 | 1.29 | 0.0214 | -0.610 | 0.334 | 0.098 | -0.683 | 0.365 | 0.106 |
| 0.0171 | 1.31 | 0.0201 | -0.431 | 0.325 | 0.061 | -0.484 | 0.351 | 0.064 |
| 0.0171 | 1.33 | 0.0189 | -0.516 | 0.321 | 0.057 | -0.570 | 0.344 | 0.061 |
| 0.0171 | 1.35 | 0.0178 | 0.033 | 0.335 | 0.027 | 0.018 | 0.357 | 0.028 |
| 0.0171 | 1.37 | 0.0169 | -0.305 | 0.341 | 0.033 | -0.340 | 0.360 | 0.035 |
| 0.0171 | 1.39 | 0.0160 | 0.290 | 0.395 | 0.052 | 0.287 | 0.416 | 0.057 |
| 0.0171 | 1.41 | 0.0152 | -0.257 | 0.385 | 0.028 | -0.288 | 0.403 | 0.032 |
| 0.0171 | 1.43 | 0.0145 | 0.126 | 0.416 | 0.047 | 0.112 | 0.434 | 0.051 |
| 0.0171 | 1.45 | 0.0138 | -0.174 | 0.433 | 0.033 | -0.200 | 0.450 | 0.032 |
| 0.0171 | 1.47 | 0.0132 | -0.620 | 0.473 | 0.039 | -0.661 | 0.490 | 0.041 |
| 0.0171 | 1.49 | 0.0126 | -0.641 | 0.441 | 0.047 | -0.680 | 0.455 | 0.047 |
| 0.0171 | 1.51 | 0.0121 | -0.333 | 0.490 | 0.032 | -0.359 | 0.505 | 0.034 |
| 0.0171 | 1.53 | 0.0116 | -0.837 | 0.468 | 0.058 | -0.875 | 0.481 | 0.060 |
| 0.0171 | 1.55 | 0.0111 | 0.400 | 0.555 | 0.081 | 0.395 | 0.569 | 0.084 |
| 0.0171 | 1.57 | 0.0107 | -0.554 | 0.582 | 0.061 | -0.581 | 0.596 | 0.063 |
| 0.0171 | 1.59 | 0.0103 | 0.098 | 0.579 | 0.100 | 0.088 | 0.591 | 0.102 |
| 0.0171 | 1.61 | 0.0099 | -0.990 | 0.629 | 0.102 | -1.021 | 0.641 | 0.104 |
| 0.0171 | 1.75 | 0.0078 | -1.629 | 0.753 | 0.144 | -1.651 | 0.763 | 0.146 |
| 0.0171 | 1.77 | 0.0075 | 0.293 | 0.537 | 0.088 | 0.294 | 0.543 | 0.090 |
| 0.0171 | 1.79 | 0.0073 | -0.879 | 0.400 | 0.088 | -0.891 | 0.405 | 0.089 |
| 0.0171 | 1.81 | 0.0071 | -0.507 | 0.390 | 0.044 | -0.514 | 0.394 | 0.044 |
| 0.0171 | 1.83 | 0.0069 | -0.746 | 0.431 | 0.043 | -0.755 | 0.435 | 0.043 |
| 0.0171 | 1.85 | 0.0067 | -0.443 | 0.401 | 0.031 | -0.448 | 0.405 | 0.031 |
| 0.0171 | 1.87 | 0.0065 | 0.019 | 0.425 | 0.046 | 0.019 | 0.429 | 0.047 |
| 0.0171 | 1.89 | 0.0063 | -0.533 | 0.463 | 0.020 | -0.538 | 0.466 | 0.020 |
| 0.0171 | 1.91 | 0.0061 | -1.100 | 0.491 | 0.077 | -1.108 | 0.494 | 0.078 |
| 0.0204 | 1.13 | 0.0489 | 0.267 | 0.289 | 0.054 | 0.331 | 0.407 | 0.076 |
| 0.0204 | 1.15 | 0.0441 | -0.018 | 0.274 | 0.075 | -0.069 | 0.365 | 0.098 |
| 0.0204 | 1.17 | 0.0401 | -0.716 | 0.294 | 0.152 | -0.946 | 0.374 | 0.199 |
| 0.0204 | 1.19 | 0.0367 | -0.554 | 0.324 | 0.164 | -0.700 | 0.398 | 0.220 |
| 0.0204 | 1.21 | 0.0338 | -1.557 | 0.325 | 0.215 | -1.869 | 0.387 | 0.274 |
| 0.0204 | 1.23 | 0.0312 | -1.420 | 0.313 | 0.195 | -1.668 | 0.365 | 0.238 |
| 0.0204 | 1.25 | 0.0290 | -0.569 | 0.309 | 0.145 | -0.666 | 0.353 | 0.172 |
| 0.0204 | 1.27 | 0.0271 | -0.194 | 0.306 | 0.104 | -0.237 | 0.344 | 0.120 |
| 0.0204 | 1.29 | 0.0254 | -0.242 | 0.278 | 0.080 | -0.288 | 0.308 | 0.088 |
| 0.0204 | 1.31 | 0.0238 | 0.179 | 0.267 | 0.055 | 0.177 | 0.293 | 0.056 |
| 0.0204 | 1.33 | 0.0224 | -0.298 | 0.277 | 0.046 | -0.343 | 0.301 | 0.049 |
| 0.0204 | 1.35 | 0.0212 | -0.190 | 0.292 | 0.033 | -0.224 | 0.314 | 0.035 |
| 0.0204 | 1.37 | 0.0201 | -0.140 | 0.323 | 0.023 | -0.169 | 0.345 | 0.025 |
| 0.0204 | 1.39 | 0.0190 | -0.150 | 0.366 | 0.029 | -0.180 | 0.389 | 0.034 |
| 0.0204 | 1.41 | 0.0181 | -0.443 | 0.360 | 0.045 | -0.490 | 0.380 | 0.049 |
| 0.0204 | 1.43 | 0.0172 | -0.142 | 0.384 | 0.034 | -0.172 | 0.404 | 0.036 |
| 0.0204 | 1.45 | 0.0164 | -0.506 | 0.407 | 0.047 | -0.551 | 0.425 | 0.051 |
| 0.0204 | 1.47 | 0.0157 | -0.370 | 0.437 | 0.028 | -0.408 | 0.455 | 0.032 |
| 0.0204 | 1.49 | 0.0150 | -1.031 | 0.391 | 0.078 | -1.091 | 0.406 | 0.081 |
| 0.0204 | 1.51 | 0.0144 | -0.417 | 0.428 | 0.020 | -0.451 | 0.443 | 0.025 |
| 0.0204 | 1.53 | 0.0138 | -0.514 | 0.425 | 0.032 | -0.549 | 0.439 | 0.034 |
| 0.0204 | 1.55 | 0.0132 | -0.200 | 0.525 | 0.050 | -0.224 | 0.540 | 0.054 |
| 0.0204 | 1.57 | 0.0127 | -0.942 | 0.471 | 0.084 | -0.985 | 0.484 | 0.087 |
| 0.0204 | 1.59 | 0.0122 | 0.056 | 0.505 | 0.065 | 0.043 | 0.517 | 0.067 |
| 0.0204 | 1.65 | 0.0110 | 0.626 | 0.744 | 0.081 | 0.631 | 0.759 | 0.083 |
| 0.0204 | 1.67 | 0.0106 | -0.766 | 0.499 | 0.054 | -0.786 | 0.509 | 0.054 |
| 0.0204 | 1.69 | 0.0102 | -0.566 | 0.299 | 0.032 | -0.580 | 0.304 | 0.033 |
| 0.0204 | 1.71 | 0.0099 | -0.327 | 0.266 | 0.023 | -0.335 | 0.271 | 0.023 |
| 0.0204 | 1.73 | 0.0096 | -1.021 | 0.314 | 0.086 | -1.039 | 0.318 | 0.087 |
| 0.0204 | 1.75 | 0.0093 | -0.606 | 0.312 | 0.068 | -0.617 | 0.317 | 0.069 |
| 0.0204 | 1.77 | 0.0090 | -0.249 | 0.326 | 0.060 | -0.255 | 0.330 | 0.061 |
| 0.0204 | 1.79 | 0.0087 | -0.798 | 0.313 | 0.085 | -0.811 | 0.317 | 0.086 |
| 0.0204 | 1.81 | 0.0084 | -0.283 | 0.350 | 0.036 | -0.289 | 0.354 | 0.036 |
| 0.0204 | 1.83 | 0.0082 | -0.738 | 0.384 | 0.056 | -0.748 | 0.388 | 0.057 |
| 0.0204 | 1.85 | 0.0080 | -0.885 | 0.432 | 0.057 | -0.896 | 0.436 | 0.058 |
| 0.0204 | 1.87 | 0.0077 | -0.740 | 0.429 | 0.056 | -0.748 | 0.433 | 0.057 |
| 0.0204 | 1.89 | 0.0075 | -0.314 | 0.480 | 0.020 | -0.317 | 0.485 | 0.022 |
| 0.0204 | 1.91 | 0.0073 | -0.350 | 0.519 | 0.016 | -0.353 | 0.524 | 0.016 |
| 0.0244 | 1.11 | 0.0649 | -0.206 | 0.228 | 0.030 | -0.368 | 0.365 | 0.053 |
| 0.0244 | 1.13 | 0.0580 | -0.606 | 0.294 | 0.077 | -0.947 | 0.435 | 0.117 |
| 0.0244 | 1.15 | 0.0523 | -0.379 | 0.258 | 0.088 | -0.578 | 0.359 | 0.121 |
| 0.0244 | 1.17 | 0.0476 | -0.307 | 0.290 | 0.133 | -0.445 | 0.383 | 0.182 |
| 0.0244 | 1.19 | 0.0436 | -1.022 | 0.283 | 0.170 | -1.320 | 0.360 | 0.235 |
| 0.0244 | 1.21 | 0.0401 | -0.920 | 0.290 | 0.163 | -1.144 | 0.356 | 0.225 |
| 0.0244 | 1.23 | 0.0371 | -0.709 | 0.288 | 0.147 | -0.862 | 0.345 | 0.194 |
| 0.0244 | 1.25 | 0.0345 | -0.578 | 0.302 | 0.146 | -0.695 | 0.353 | 0.178 |
| 0.0244 | 1.27 | 0.0322 | -0.135 | 0.271 | 0.094 | -0.177 | 0.311 | 0.111 |
| 0.0244 | 1.29 | 0.0302 | -0.438 | 0.265 | 0.094 | -0.519 | 0.300 | 0.106 |
| 0.0244 | 1.31 | 0.0284 | -0.060 | 0.251 | 0.060 | -0.091 | 0.280 | 0.063 |
| 0.0244 | 1.33 | 0.0267 | 0.352 | 0.265 | 0.047 | 0.364 | 0.292 | 0.048 |
| 0.0244 | 1.35 | 0.0252 | -0.607 | 0.259 | 0.068 | -0.686 | 0.282 | 0.073 |
| 0.0244 | 1.37 | 0.0239 | -0.318 | 0.267 | 0.039 | -0.369 | 0.289 | 0.041 |
| 0.0244 | 1.39 | 0.0227 | -0.215 | 0.275 | 0.024 | -0.256 | 0.295 | 0.025 |
| 0.0244 | 1.41 | 0.0216 | -0.018 | 0.327 | 0.026 | -0.045 | 0.348 | 0.027 |
| 0.0244 | 1.43 | 0.0205 | -0.876 | 0.287 | 0.081 | -0.956 | 0.304 | 0.084 |
| 0.0244 | 1.45 | 0.0196 | -0.950 | 0.341 | 0.076 | -1.029 | 0.360 | 0.080 |
| 0.0244 | 1.47 | 0.0187 | -0.369 | 0.360 | 0.024 | -0.414 | 0.378 | 0.026 |
| 0.0244 | 1.49 | 0.0179 | -0.119 | 0.361 | 0.016 | -0.150 | 0.377 | 0.021 |
| 0.0244 | 1.51 | 0.0171 | -1.586 | 0.406 | 0.130 | -1.675 | 0.423 | 0.136 |
| 0.0244 | 1.53 | 0.0164 | -0.769 | 0.394 | 0.066 | -0.821 | 0.409 | 0.068 |
| 0.0244 | 1.55 | 0.0158 | -0.180 | 0.327 | 0.038 | -0.207 | 0.339 | 0.038 |
| 0.0244 | 1.57 | 0.0152 | -0.453 | 0.282 | 0.040 | -0.488 | 0.291 | 0.041 |
| 0.0244 | 1.59 | 0.0146 | -0.325 | 0.233 | 0.034 | -0.352 | 0.240 | 0.034 |
| 0.0244 | 1.61 | 0.0141 | -0.369 | 0.237 | 0.030 | -0.394 | 0.244 | 0.031 |
| 0.0244 | 1.63 | 0.0135 | 0.033 | 0.244 | 0.026 | 0.022 | 0.250 | 0.026 |
| 0.0244 | 1.65 | 0.0131 | -0.260 | 0.240 | 0.013 | -0.275 | 0.246 | 0.014 |
| 0.0244 | 1.67 | 0.0126 | -0.656 | 0.221 | 0.044 | -0.678 | 0.226 | 0.045 |
| 0.0244 | 1.69 | 0.0122 | -0.198 | 0.206 | 0.011 | -0.207 | 0.211 | 0.011 |
| 0.0244 | 1.71 | 0.0118 | -0.070 | 0.242 | 0.028 | -0.074 | 0.247 | 0.027 |
| 0.0244 | 1.73 | 0.0114 | -0.721 | 0.259 | 0.061 | -0.737 | 0.263 | 0.062 |
| 0.0244 | 1.75 | 0.0111 | -0.305 | 0.281 | 0.059 | -0.313 | 0.286 | 0.060 |
| 0.0244 | 1.77 | 0.0107 | -0.730 | 0.302 | 0.074 | -0.745 | 0.307 | 0.076 |
| 0.0244 | 1.79 | 0.0104 | -0.206 | 0.304 | 0.059 | -0.212 | 0.309 | 0.060 |
| 0.0244 | 1.81 | 0.0101 | -0.238 | 0.328 | 0.051 | -0.245 | 0.333 | 0.051 |
| 0.0244 | 1.83 | 0.0098 | -0.323 | 0.387 | 0.031 | -0.331 | 0.392 | 0.032 |
| 0.0244 | 1.85 | 0.0095 | -0.770 | 0.396 | 0.052 | -0.782 | 0.401 | 0.053 |
| 0.0244 | 1.87 | 0.0092 | 0.616 | 0.430 | 0.097 | 0.622 | 0.436 | 0.098 |
| 0.0244 | 1.89 | 0.0090 | -0.141 | 0.425 | 0.032 | -0.142 | 0.430 | 0.034 |
| 0.0244 | 1.91 | 0.0087 | -0.626 | 0.464 | 0.037 | -0.632 | 0.469 | 0.039 |
| 0.029 | 1.15 | 0.0616 | -0.073 | 0.212 | 0.066 | -0.164 | 0.310 | 0.096 |
| 0.029 | 1.17 | 0.0560 | -0.416 | 0.250 | 0.129 | -0.620 | 0.345 | 0.186 |
| 0.029 | 1.19 | 0.0514 | -0.830 | 0.254 | 0.152 | -1.121 | 0.335 | 0.225 |
| 0.029 | 1.21 | 0.0473 | -1.323 | 0.271 | 0.170 | -1.698 | 0.345 | 0.242 |
| 0.029 | 1.23 | 0.0438 | -1.043 | 0.259 | 0.154 | -1.304 | 0.320 | 0.208 |
| 0.029 | 1.25 | 0.0408 | -1.263 | 0.253 | 0.166 | -1.541 | 0.304 | 0.207 |
| 0.029 | 1.27 | 0.0381 | -0.829 | 0.237 | 0.127 | -1.002 | 0.278 | 0.153 |
| 0.029 | 1.29 | 0.0357 | -0.489 | 0.228 | 0.088 | -0.594 | 0.263 | 0.101 |
| 0.029 | 1.31 | 0.0335 | -0.190 | 0.232 | 0.062 | -0.245 | 0.264 | 0.066 |
| 0.029 | 1.33 | 0.0316 | -0.130 | 0.219 | 0.043 | -0.174 | 0.245 | 0.042 |
| 0.029 | 1.35 | 0.0299 | -0.030 | 0.233 | 0.031 | -0.063 | 0.258 | 0.033 |
| 0.029 | 1.37 | 0.0283 | -0.122 | 0.228 | 0.022 | -0.164 | 0.251 | 0.023 |
| 0.029 | 1.39 | 0.0268 | -0.151 | 0.246 | 0.024 | -0.195 | 0.267 | 0.025 |
| 0.029 | 1.41 | 0.0255 | -0.236 | 0.213 | 0.026 | -0.287 | 0.230 | 0.028 |
| 0.029 | 1.43 | 0.0243 | -0.112 | 0.214 | 0.055 | -0.152 | 0.229 | 0.058 |
| 0.029 | 1.45 | 0.0232 | -0.601 | 0.203 | 0.062 | -0.674 | 0.216 | 0.066 |
| 0.029 | 1.47 | 0.0221 | -0.377 | 0.209 | 0.029 | -0.432 | 0.221 | 0.030 |
| 0.029 | 1.49 | 0.0212 | -0.819 | 0.205 | 0.065 | -0.895 | 0.216 | 0.068 |
| 0.029 | 1.51 | 0.0203 | -0.614 | 0.192 | 0.050 | -0.672 | 0.201 | 0.052 |
| 0.029 | 1.53 | 0.0195 | -0.185 | 0.177 | 0.021 | -0.220 | 0.186 | 0.021 |
| 0.029 | 1.55 | 0.0187 | -0.264 | 0.178 | 0.029 | -0.301 | 0.185 | 0.029 |
| 0.029 | 1.57 | 0.0180 | -0.294 | 0.179 | 0.036 | -0.329 | 0.186 | 0.036 |
| 0.029 | 1.59 | 0.0173 | -0.511 | 0.175 | 0.051 | -0.550 | 0.181 | 0.052 |
| 0.029 | 1.61 | 0.0167 | -0.460 | 0.221 | 0.036 | -0.494 | 0.228 | 0.038 |
| 0.029 | 1.63 | 0.0161 | -0.293 | 0.214 | 0.019 | -0.316 | 0.220 | 0.019 |
| 0.029 | 1.65 | 0.0155 | -0.067 | 0.207 | 0.012 | -0.080 | 0.213 | 0.012 |
| 0.029 | 1.67 | 0.0150 | -0.801 | 0.196 | 0.060 | -0.832 | 0.202 | 0.061 |
| 0.029 | 1.69 | 0.0145 | -0.845 | 0.230 | 0.063 | -0.872 | 0.236 | 0.064 |
| 0.029 | 1.71 | 0.0140 | -0.521 | 0.218 | 0.039 | -0.537 | 0.223 | 0.040 |
| 0.029 | 1.73 | 0.0135 | -0.549 | 0.235 | 0.052 | -0.563 | 0.240 | 0.053 |
| 0.029 | 1.75 | 0.0131 | -0.103 | 0.249 | 0.061 | -0.108 | 0.254 | 0.062 |
| 0.029 | 1.77 | 0.0127 | -0.237 | 0.273 | 0.059 | -0.245 | 0.278 | 0.060 |
| 0.029 | 1.79 | 0.0123 | -0.383 | 0.282 | 0.057 | -0.394 | 0.287 | 0.058 |
| 0.029 | 1.81 | 0.0120 | 0.328 | 0.320 | 0.074 | 0.330 | 0.325 | 0.076 |
| 0.029 | 1.83 | 0.0116 | -0.845 | 0.326 | 0.065 | -0.862 | 0.332 | 0.067 |
| 0.029 | 1.85 | 0.0113 | -0.585 | 0.310 | 0.036 | -0.596 | 0.315 | 0.036 |
| 0.029 | 1.87 | 0.0110 | -0.363 | 0.350 | 0.016 | -0.370 | 0.355 | 0.017 |
| 0.029 | 1.89 | 0.0107 | -0.511 | 0.353 | 0.019 | -0.519 | 0.358 | 0.020 |
| 0.029 | 1.91 | 0.0104 | -0.544 | 0.377 | 0.022 | -0.551 | 0.382 | 0.025 |
| 0.0347 | 1.15 | 0.0728 | -0.155 | 0.211 | 0.071 | -0.304 | 0.325 | 0.109 |
| 0.0347 | 1.17 | 0.0663 | -0.404 | 0.219 | 0.125 | -0.635 | 0.318 | 0.193 |
| 0.0347 | 1.19 | 0.0608 | -0.901 | 0.214 | 0.170 | -1.270 | 0.295 | 0.231 |
| 0.0347 | 1.21 | 0.0561 | -0.986 | 0.220 | 0.156 | -1.321 | 0.291 | 0.199 |
| 0.0347 | 1.23 | 0.0520 | -1.041 | 0.214 | 0.134 | -1.349 | 0.274 | 0.165 |
| 0.0347 | 1.25 | 0.0484 | -0.812 | 0.192 | 0.106 | -1.034 | 0.238 | 0.127 |
| 0.0347 | 1.27 | 0.0452 | -0.517 | 0.173 | 0.080 | -0.658 | 0.210 | 0.093 |
| 0.0347 | 1.29 | 0.0424 | -0.179 | 0.154 | 0.048 | -0.246 | 0.182 | 0.053 |
| 0.0347 | 1.31 | 0.0399 | -0.045 | 0.141 | 0.037 | -0.087 | 0.164 | 0.042 |
| 0.0347 | 1.33 | 0.0376 | 0.011 | 0.133 | 0.030 | -0.022 | 0.153 | 0.033 |
| 0.0347 | 1.35 | 0.0355 | 0.057 | 0.128 | 0.019 | 0.029 | 0.145 | 0.021 |
| 0.0347 | 1.37 | 0.0337 | -0.148 | 0.122 | 0.025 | -0.202 | 0.137 | 0.028 |
| 0.0347 | 1.39 | 0.0319 | -0.017 | 0.129 | 0.014 | -0.056 | 0.143 | 0.016 |
| 0.0347 | 1.41 | 0.0304 | -0.036 | 0.136 | 0.018 | -0.078 | 0.149 | 0.019 |
| 0.0347 | 1.43 | 0.0289 | -0.440 | 0.145 | 0.041 | -0.518 | 0.157 | 0.044 |
| 0.0347 | 1.45 | 0.0276 | -0.486 | 0.149 | 0.044 | -0.564 | 0.161 | 0.047 |
| 0.0347 | 1.47 | 0.0264 | -0.139 | 0.159 | 0.035 | -0.188 | 0.171 | 0.037 |
| 0.0347 | 1.49 | 0.0252 | -0.428 | 0.157 | 0.044 | -0.493 | 0.168 | 0.047 |
| 0.0347 | 1.51 | 0.0242 | -0.135 | 0.161 | 0.020 | -0.177 | 0.171 | 0.020 |
| 0.0347 | 1.53 | 0.0232 | -0.106 | 0.161 | 0.031 | -0.144 | 0.170 | 0.031 |
| 0.0347 | 1.55 | 0.0223 | -0.432 | 0.166 | 0.037 | -0.484 | 0.174 | 0.038 |
| 0.0347 | 1.57 | 0.0214 | -0.063 | 0.165 | 0.033 | -0.094 | 0.172 | 0.033 |
| 0.0347 | 1.59 | 0.0206 | -0.394 | 0.170 | 0.036 | -0.437 | 0.177 | 0.036 |
| 0.0347 | 1.61 | 0.0199 | 0.015 | 0.194 | 0.018 | -0.006 | 0.202 | 0.019 |
| 0.0347 | 1.63 | 0.0192 | -0.381 | 0.194 | 0.026 | -0.412 | 0.202 | 0.027 |
| 0.0347 | 1.65 | 0.0185 | -0.261 | 0.182 | 0.015 | -0.284 | 0.188 | 0.015 |
| 0.0347 | 1.67 | 0.0179 | -0.373 | 0.202 | 0.023 | -0.395 | 0.209 | 0.024 |
| 0.0347 | 1.69 | 0.0173 | -0.647 | 0.205 | 0.047 | -0.674 | 0.211 | 0.048 |
| 0.0347 | 1.71 | 0.0167 | -0.260 | 0.198 | 0.028 | -0.272 | 0.203 | 0.027 |
| 0.0347 | 1.73 | 0.0162 | -0.188 | 0.226 | 0.043 | -0.196 | 0.232 | 0.045 |
| 0.0347 | 1.75 | 0.0157 | -0.055 | 0.220 | 0.059 | -0.060 | 0.225 | 0.061 |
| 0.0347 | 1.77 | 0.0152 | 0.044 | 0.222 | 0.064 | 0.041 | 0.227 | 0.065 |
| 0.0347 | 1.79 | 0.0147 | -0.042 | 0.276 | 0.050 | -0.048 | 0.282 | 0.051 |
| 0.0347 | 1.81 | 0.0143 | -0.596 | 0.270 | 0.057 | -0.613 | 0.275 | 0.059 |
| 0.0347 | 1.83 | 0.0139 | -0.256 | 0.255 | 0.032 | -0.265 | 0.260 | 0.033 |
| 0.0347 | 1.85 | 0.0135 | -0.352 | 0.271 | 0.024 | -0.361 | 0.276 | 0.025 |
| 0.0347 | 1.87 | 0.0131 | -0.061 | 0.307 | 0.028 | -0.064 | 0.312 | 0.030 |
| 0.0347 | 1.89 | 0.0127 | -0.805 | 0.328 | 0.055 | -0.819 | 0.334 | 0.057 |
| 0.0347 | 1.91 | 0.0124 | -0.266 | 0.348 | 0.030 | -0.269 | 0.353 | 0.034 |
| 0.0414 | 1.11 | 0.1053 | 0.088 | 0.098 | 0.020 | 0.120 | 0.191 | 0.042 |
| 0.0414 | 1.13 | 0.0945 | 0.132 | 0.106 | 0.036 | 0.163 | 0.186 | 0.064 |
| 0.0414 | 1.15 | 0.0856 | -0.215 | 0.108 | 0.063 | -0.418 | 0.174 | 0.101 |
| 0.0414 | 1.17 | 0.0781 | -0.406 | 0.141 | 0.119 | -0.669 | 0.213 | 0.179 |
| 0.0414 | 1.19 | 0.0717 | -0.699 | 0.149 | 0.131 | -1.032 | 0.214 | 0.185 |
| 0.0414 | 1.21 | 0.0662 | -1.277 | 0.156 | 0.128 | -1.767 | 0.214 | 0.170 |
| 0.0414 | 1.23 | 0.0614 | -0.776 | 0.159 | 0.102 | -1.043 | 0.209 | 0.128 |
| 0.0414 | 1.25 | 0.0572 | -0.531 | 0.145 | 0.092 | -0.708 | 0.185 | 0.114 |
| 0.0414 | 1.27 | 0.0535 | -0.168 | 0.132 | 0.065 | -0.246 | 0.164 | 0.079 |
| 0.0414 | 1.29 | 0.0502 | -0.207 | 0.121 | 0.050 | -0.291 | 0.147 | 0.060 |
| 0.0414 | 1.31 | 0.0472 | -0.041 | 0.120 | 0.037 | -0.090 | 0.142 | 0.042 |
| 0.0414 | 1.33 | 0.0445 | -0.120 | 0.114 | 0.032 | -0.182 | 0.133 | 0.037 |
| 0.0414 | 1.35 | 0.0421 | 0.009 | 0.112 | 0.017 | -0.032 | 0.129 | 0.019 |
| 0.0414 | 1.37 | 0.0399 | -0.135 | 0.113 | 0.022 | -0.196 | 0.128 | 0.024 |
| 0.0414 | 1.39 | 0.0379 | -0.410 | 0.122 | 0.044 | -0.504 | 0.137 | 0.048 |
| 0.0414 | 1.41 | 0.0360 | -0.159 | 0.125 | 0.028 | -0.222 | 0.139 | 0.030 |
| 0.0414 | 1.43 | 0.0343 | -0.115 | 0.130 | 0.016 | -0.173 | 0.143 | 0.016 |
| 0.0414 | 1.45 | 0.0328 | -0.098 | 0.140 | 0.022 | -0.153 | 0.152 | 0.023 |
| 0.0414 | 1.47 | 0.0313 | -0.458 | 0.142 | 0.042 | -0.541 | 0.154 | 0.045 |
| 0.0414 | 1.49 | 0.0300 | -0.290 | 0.150 | 0.030 | -0.355 | 0.161 | 0.031 |
| 0.0414 | 1.51 | 0.0287 | -0.501 | 0.155 | 0.050 | -0.576 | 0.165 | 0.052 |
| 0.0414 | 1.53 | 0.0276 | -0.213 | 0.156 | 0.029 | -0.264 | 0.166 | 0.030 |
| 0.0414 | 1.55 | 0.0265 | -0.641 | 0.153 | 0.055 | -0.714 | 0.162 | 0.057 |
| 0.0414 | 1.57 | 0.0255 | -0.176 | 0.155 | 0.023 | -0.218 | 0.164 | 0.023 |
| 0.0414 | 1.59 | 0.0245 | -0.177 | 0.161 | 0.034 | -0.216 | 0.169 | 0.035 |
| 0.0414 | 1.61 | 0.0236 | -0.123 | 0.179 | 0.009 | -0.154 | 0.188 | 0.010 |
| 0.0414 | 1.63 | 0.0228 | -0.439 | 0.175 | 0.032 | -0.479 | 0.183 | 0.033 |
| 0.0414 | 1.65 | 0.0220 | -0.208 | 0.171 | 0.011 | -0.232 | 0.178 | 0.011 |
| 0.0414 | 1.67 | 0.0212 | -0.902 | 0.192 | 0.073 | -0.949 | 0.199 | 0.076 |
| 0.0414 | 1.69 | 0.0205 | -0.478 | 0.175 | 0.034 | -0.503 | 0.181 | 0.035 |
| 0.0414 | 1.71 | 0.0199 | -0.171 | 0.186 | 0.028 | -0.183 | 0.192 | 0.030 |
| 0.0414 | 1.73 | 0.0192 | -0.642 | 0.198 | 0.065 | -0.666 | 0.204 | 0.067 |
| 0.0414 | 1.75 | 0.0186 | -0.115 | 0.200 | 0.052 | -0.122 | 0.205 | 0.054 |
| 0.0414 | 1.77 | 0.0180 | -0.486 | 0.232 | 0.059 | -0.504 | 0.238 | 0.061 |
| 0.0414 | 1.79 | 0.0175 | -0.365 | 0.238 | 0.048 | -0.380 | 0.244 | 0.049 |
| 0.0414 | 1.81 | 0.0170 | -0.246 | 0.219 | 0.045 | -0.257 | 0.224 | 0.046 |
| 0.0414 | 1.83 | 0.0165 | -0.442 | 0.237 | 0.040 | -0.457 | 0.243 | 0.042 |
| 0.0414 | 1.85 | 0.0160 | -0.150 | 0.257 | 0.023 | -0.157 | 0.263 | 0.024 |
| 0.0414 | 1.87 | 0.0156 | -0.232 | 0.278 | 0.019 | -0.239 | 0.284 | 0.021 |
| 0.0414 | 1.89 | 0.0151 | -0.657 | 0.307 | 0.038 | -0.670 | 0.313 | 0.040 |
| 0.0414 | 1.91 | 0.0147 | -0.150 | 0.309 | 0.027 | -0.152 | 0.315 | 0.030 |
| 0.0494 | 1.11 | 0.1232 | -0.015 | 0.165 | 0.032 | -0.085 | 0.343 | 0.073 |
| 0.0494 | 1.13 | 0.1108 | 0.041 | 0.089 | 0.038 | 0.002 | 0.167 | 0.070 |
| 0.0494 | 1.15 | 0.1005 | -0.104 | 0.098 | 0.060 | -0.255 | 0.169 | 0.103 |
| 0.0494 | 1.17 | 0.0918 | -0.540 | 0.120 | 0.116 | -0.920 | 0.192 | 0.184 |
| 0.0494 | 1.19 | 0.0844 | -0.774 | 0.142 | 0.126 | -1.196 | 0.214 | 0.187 |
| 0.0494 | 1.21 | 0.0780 | -0.717 | 0.145 | 0.085 | -1.044 | 0.207 | 0.115 |
| 0.0494 | 1.23 | 0.0724 | -1.270 | 0.142 | 0.126 | -1.765 | 0.195 | 0.168 |
| 0.0494 | 1.25 | 0.0675 | -0.602 | 0.130 | 0.092 | -0.830 | 0.172 | 0.118 |
| 0.0494 | 1.27 | 0.0632 | -0.565 | 0.122 | 0.085 | -0.767 | 0.156 | 0.108 |
| 0.0494 | 1.29 | 0.0593 | -0.355 | 0.115 | 0.059 | -0.490 | 0.143 | 0.072 |
| 0.0494 | 1.31 | 0.0558 | 0.010 | 0.111 | 0.037 | -0.036 | 0.136 | 0.044 |
| 0.0494 | 1.33 | 0.0527 | -0.143 | 0.107 | 0.030 | -0.221 | 0.128 | 0.035 |
| 0.0494 | 1.35 | 0.0498 | 0.174 | 0.104 | 0.025 | 0.154 | 0.122 | 0.028 |
| 0.0494 | 1.37 | 0.0472 | 0.045 | 0.101 | 0.013 | 0.000 | 0.117 | 0.014 |
| 0.0494 | 1.39 | 0.0449 | -0.143 | 0.110 | 0.025 | -0.217 | 0.125 | 0.028 |
| 0.0494 | 1.41 | 0.0427 | -0.279 | 0.112 | 0.030 | -0.370 | 0.126 | 0.034 |
| 0.0494 | 1.43 | 0.0407 | -0.122 | 0.122 | 0.019 | -0.191 | 0.136 | 0.020 |
| 0.0494 | 1.45 | 0.0389 | -0.477 | 0.125 | 0.047 | -0.583 | 0.138 | 0.051 |
| 0.0494 | 1.47 | 0.0371 | -0.150 | 0.132 | 0.023 | -0.220 | 0.145 | 0.024 |
| 0.0494 | 1.49 | 0.0356 | -0.474 | 0.135 | 0.044 | -0.568 | 0.148 | 0.046 |
| 0.0494 | 1.51 | 0.0341 | -0.247 | 0.138 | 0.027 | -0.314 | 0.149 | 0.028 |
| 0.0494 | 1.53 | 0.0327 | -0.139 | 0.141 | 0.027 | -0.193 | 0.151 | 0.027 |
| 0.0494 | 1.55 | 0.0314 | -0.268 | 0.146 | 0.034 | -0.329 | 0.156 | 0.035 |
| 0.0494 | 1.57 | 0.0302 | -0.159 | 0.145 | 0.017 | -0.209 | 0.154 | 0.018 |
| 0.0494 | 1.59 | 0.0291 | 0.038 | 0.148 | 0.020 | 0.005 | 0.157 | 0.021 |
| 0.0494 | 1.61 | 0.0281 | -0.403 | 0.162 | 0.031 | -0.456 | 0.171 | 0.033 |
| 0.0494 | 1.63 | 0.0271 | -0.020 | 0.160 | 0.008 | -0.046 | 0.168 | 0.009 |
| 0.0494 | 1.65 | 0.0261 | -0.062 | 0.159 | 0.009 | -0.084 | 0.167 | 0.011 |
| 0.0494 | 1.67 | 0.0252 | -0.118 | 0.170 | 0.013 | -0.138 | 0.178 | 0.011 |
| 0.0494 | 1.69 | 0.0244 | -0.434 | 0.172 | 0.033 | -0.463 | 0.179 | 0.033 |
| 0.0494 | 1.71 | 0.0236 | -0.422 | 0.177 | 0.037 | -0.446 | 0.184 | 0.037 |
| 0.0494 | 1.73 | 0.0229 | -0.182 | 0.184 | 0.048 | -0.194 | 0.190 | 0.050 |
| 0.0494 | 1.75 | 0.0221 | -0.536 | 0.193 | 0.060 | -0.559 | 0.200 | 0.062 |
| 0.0494 | 1.77 | 0.0215 | -0.031 | 0.194 | 0.062 | -0.037 | 0.200 | 0.064 |
| 0.0494 | 1.79 | 0.0208 | -0.282 | 0.195 | 0.049 | -0.297 | 0.201 | 0.050 |
| 0.0494 | 1.81 | 0.0202 | -0.294 | 0.220 | 0.030 | -0.308 | 0.226 | 0.030 |
| 0.0494 | 1.83 | 0.0196 | -0.085 | 0.230 | 0.029 | -0.093 | 0.236 | 0.030 |
| 0.0494 | 1.85 | 0.0191 | -0.287 | 0.223 | 0.034 | -0.299 | 0.229 | 0.036 |
| 0.0494 | 1.87 | 0.0185 | -0.648 | 0.270 | 0.041 | -0.667 | 0.276 | 0.043 |
| 0.0494 | 1.89 | 0.0180 | -0.564 | 0.278 | 0.030 | -0.578 | 0.284 | 0.032 |
| 0.0494 | 1.91 | 0.0175 | -0.201 | 0.310 | 0.019 | -0.204 | 0.317 | 0.023 |
| 0.059 | 1.11 | 0.1437 | 0.085 | 0.102 | 0.042 | 0.131 | 0.227 | 0.091 |
| 0.059 | 1.13 | 0.1295 | -0.082 | 0.069 | 0.026 | -0.244 | 0.138 | 0.049 |
| 0.059 | 1.15 | 0.1177 | -0.179 | 0.099 | 0.063 | -0.408 | 0.182 | 0.113 |
| 0.059 | 1.17 | 0.1078 | -0.652 | 0.118 | 0.124 | -1.165 | 0.199 | 0.209 |
| 0.059 | 1.19 | 0.0992 | -0.844 | 0.136 | 0.137 | -1.372 | 0.217 | 0.214 |
| 0.059 | 1.21 | 0.0918 | -0.740 | 0.139 | 0.086 | -1.129 | 0.209 | 0.122 |
| 0.059 | 1.23 | 0.0853 | -0.803 | 0.130 | 0.085 | -1.176 | 0.186 | 0.116 |
| 0.059 | 1.25 | 0.0796 | -0.730 | 0.127 | 0.098 | -1.044 | 0.174 | 0.132 |
| 0.059 | 1.27 | 0.0745 | -0.488 | 0.118 | 0.079 | -0.698 | 0.158 | 0.102 |
| 0.059 | 1.29 | 0.0700 | -0.133 | 0.108 | 0.048 | -0.226 | 0.140 | 0.061 |
| 0.059 | 1.31 | 0.0659 | -0.087 | 0.099 | 0.031 | -0.167 | 0.125 | 0.038 |
| 0.059 | 1.33 | 0.0623 | -0.041 | 0.097 | 0.025 | -0.109 | 0.120 | 0.031 |
| 0.059 | 1.35 | 0.0589 | -0.145 | 0.101 | 0.025 | -0.236 | 0.122 | 0.029 |
| 0.059 | 1.37 | 0.0559 | 0.034 | 0.107 | 0.019 | -0.022 | 0.127 | 0.021 |
| 0.059 | 1.39 | 0.0531 | -0.091 | 0.100 | 0.015 | -0.170 | 0.116 | 0.017 |
| 0.059 | 1.41 | 0.0506 | -0.004 | 0.108 | 0.017 | -0.069 | 0.124 | 0.017 |
| 0.059 | 1.43 | 0.0482 | -0.276 | 0.110 | 0.030 | -0.380 | 0.125 | 0.033 |
| 0.059 | 1.45 | 0.0461 | -0.168 | 0.113 | 0.031 | -0.255 | 0.127 | 0.033 |
| 0.059 | 1.47 | 0.0440 | -0.090 | 0.123 | 0.021 | -0.165 | 0.137 | 0.022 |
| 0.059 | 1.49 | 0.0422 | -0.146 | 0.131 | 0.031 | -0.222 | 0.145 | 0.031 |
| 0.059 | 1.51 | 0.0404 | -0.369 | 0.133 | 0.035 | -0.461 | 0.146 | 0.037 |
| 0.059 | 1.53 | 0.0388 | -0.112 | 0.130 | 0.026 | -0.174 | 0.142 | 0.027 |
| 0.059 | 1.55 | 0.0373 | 0.167 | 0.130 | 0.032 | 0.132 | 0.140 | 0.033 |
| 0.059 | 1.57 | 0.0359 | 0.001 | 0.130 | 0.016 | -0.045 | 0.140 | 0.017 |
| 0.059 | 1.59 | 0.0346 | -0.075 | 0.130 | 0.019 | -0.122 | 0.139 | 0.020 |
| 0.059 | 1.61 | 0.0333 | -0.157 | 0.153 | 0.013 | -0.203 | 0.163 | 0.013 |
| 0.059 | 1.63 | 0.0321 | -0.204 | 0.151 | 0.019 | -0.246 | 0.160 | 0.020 |
| 0.059 | 1.65 | 0.0310 | -0.238 | 0.148 | 0.018 | -0.275 | 0.157 | 0.019 |
| 0.059 | 1.67 | 0.0300 | -0.164 | 0.158 | 0.013 | -0.191 | 0.167 | 0.015 |
| 0.059 | 1.69 | 0.0290 | -0.282 | 0.163 | 0.026 | -0.309 | 0.171 | 0.027 |
| 0.059 | 1.71 | 0.0281 | -0.596 | 0.179 | 0.050 | -0.633 | 0.187 | 0.051 |
| 0.059 | 1.73 | 0.0272 | -0.358 | 0.183 | 0.044 | -0.379 | 0.191 | 0.045 |
| 0.059 | 1.75 | 0.0263 | -0.338 | 0.172 | 0.054 | -0.357 | 0.179 | 0.056 |
| 0.059 | 1.77 | 0.0255 | -0.098 | 0.188 | 0.053 | -0.109 | 0.195 | 0.055 |
| 0.059 | 1.79 | 0.0248 | -0.392 | 0.191 | 0.054 | -0.413 | 0.198 | 0.055 |
| 0.059 | 1.81 | 0.0240 | -0.511 | 0.191 | 0.050 | -0.535 | 0.197 | 0.052 |
| 0.059 | 1.83 | 0.0233 | -0.123 | 0.207 | 0.023 | -0.134 | 0.214 | 0.023 |
| 0.059 | 1.85 | 0.0227 | -0.130 | 0.202 | 0.023 | -0.139 | 0.208 | 0.023 |
| 0.059 | 1.87 | 0.0221 | -0.262 | 0.230 | 0.016 | -0.273 | 0.237 | 0.017 |
| 0.059 | 1.89 | 0.0214 | -0.664 | 0.235 | 0.042 | -0.684 | 0.242 | 0.043 |
| 0.059 | 1.91 | 0.0209 | -0.271 | 0.274 | 0.010 | -0.278 | 0.281 | 0.008 |
| 0.0705 | 1.11 | 0.1670 | -0.007 | 0.036 | 0.018 | -0.080 | 0.086 | 0.034 |
| 0.0705 | 1.13 | 0.1510 | -0.061 | 0.095 | 0.027 | -0.214 | 0.202 | 0.058 |
| 0.0705 | 1.15 | 0.1375 | -0.075 | 0.091 | 0.061 | -0.230 | 0.177 | 0.117 |
| 0.0705 | 1.17 | 0.1261 | -0.410 | 0.110 | 0.106 | -0.796 | 0.198 | 0.188 |
| 0.0705 | 1.19 | 0.1163 | -0.997 | 0.127 | 0.139 | -1.698 | 0.213 | 0.228 |
| 0.0705 | 1.21 | 0.1078 | -0.968 | 0.134 | 0.104 | -1.542 | 0.212 | 0.156 |
| 0.0705 | 1.23 | 0.1003 | -0.881 | 0.125 | 0.090 | -1.345 | 0.188 | 0.129 |
| 0.0705 | 1.25 | 0.0937 | -0.742 | 0.120 | 0.098 | -1.106 | 0.173 | 0.137 |
| 0.0705 | 1.27 | 0.0878 | -0.234 | 0.111 | 0.065 | -0.378 | 0.154 | 0.087 |
| 0.0705 | 1.29 | 0.0825 | -0.087 | 0.106 | 0.053 | -0.178 | 0.141 | 0.069 |
| 0.0705 | 1.31 | 0.0778 | -0.050 | 0.094 | 0.035 | -0.131 | 0.122 | 0.044 |
| 0.0705 | 1.33 | 0.0735 | -0.113 | 0.092 | 0.026 | -0.211 | 0.117 | 0.031 |
| 0.0705 | 1.35 | 0.0696 | -0.127 | 0.091 | 0.021 | -0.227 | 0.113 | 0.025 |
| 0.0705 | 1.37 | 0.0661 | -0.098 | 0.088 | 0.020 | -0.192 | 0.107 | 0.022 |
| 0.0705 | 1.39 | 0.0628 | -0.073 | 0.095 | 0.017 | -0.162 | 0.113 | 0.018 |
| 0.0705 | 1.41 | 0.0598 | -0.126 | 0.090 | 0.022 | -0.224 | 0.107 | 0.024 |
| 0.0705 | 1.43 | 0.0571 | 0.009 | 0.103 | 0.014 | -0.067 | 0.119 | 0.014 |
| 0.0705 | 1.45 | 0.0545 | -0.132 | 0.109 | 0.022 | -0.229 | 0.125 | 0.023 |
| 0.0705 | 1.47 | 0.0522 | -0.276 | 0.117 | 0.033 | -0.389 | 0.133 | 0.036 |
| 0.0705 | 1.49 | 0.0500 | -0.356 | 0.122 | 0.038 | -0.471 | 0.137 | 0.040 |
| 0.0705 | 1.51 | 0.0480 | 0.101 | 0.125 | 0.029 | 0.048 | 0.140 | 0.030 |
| 0.0705 | 1.53 | 0.0460 | -0.158 | 0.126 | 0.019 | -0.235 | 0.139 | 0.020 |
| 0.0705 | 1.55 | 0.0443 | -0.164 | 0.114 | 0.026 | -0.237 | 0.125 | 0.028 |
| 0.0705 | 1.57 | 0.0426 | -0.089 | 0.122 | 0.015 | -0.151 | 0.133 | 0.016 |
| 0.0705 | 1.59 | 0.0410 | -0.198 | 0.122 | 0.019 | -0.263 | 0.132 | 0.020 |
| 0.0705 | 1.61 | 0.0396 | -0.142 | 0.132 | 0.016 | -0.196 | 0.142 | 0.018 |
| 0.0705 | 1.63 | 0.0382 | 0.132 | 0.140 | 0.024 | 0.107 | 0.150 | 0.025 |
| 0.0705 | 1.65 | 0.0369 | -0.148 | 0.146 | 0.017 | -0.186 | 0.156 | 0.017 |
| 0.0705 | 1.67 | 0.0356 | -0.433 | 0.156 | 0.035 | -0.482 | 0.166 | 0.036 |
| 0.0705 | 1.69 | 0.0345 | -0.286 | 0.161 | 0.025 | -0.319 | 0.171 | 0.024 |
| 0.0705 | 1.71 | 0.0333 | -0.450 | 0.155 | 0.044 | -0.485 | 0.163 | 0.047 |
| 0.0705 | 1.73 | 0.0323 | -0.481 | 0.156 | 0.052 | -0.513 | 0.164 | 0.054 |
| 0.0705 | 1.75 | 0.0313 | -0.457 | 0.162 | 0.054 | -0.486 | 0.170 | 0.056 |
| 0.0705 | 1.77 | 0.0303 | -0.298 | 0.166 | 0.050 | -0.319 | 0.174 | 0.052 |
| 0.0705 | 1.79 | 0.0294 | -0.072 | 0.174 | 0.044 | -0.084 | 0.182 | 0.045 |
| 0.0705 | 1.81 | 0.0286 | -0.574 | 0.180 | 0.054 | -0.605 | 0.187 | 0.056 |
| 0.0705 | 1.83 | 0.0278 | -0.464 | 0.194 | 0.038 | -0.489 | 0.202 | 0.038 |
| 0.0705 | 1.85 | 0.0270 | -0.414 | 0.182 | 0.034 | -0.435 | 0.189 | 0.034 |
| 0.0705 | 1.87 | 0.0262 | -0.331 | 0.213 | 0.016 | -0.347 | 0.220 | 0.016 |
| 0.0705 | 1.89 | 0.0255 | -0.343 | 0.216 | 0.015 | -0.356 | 0.223 | 0.014 |
| 0.0705 | 1.91 | 0.0248 | -0.415 | 0.246 | 0.019 | -0.428 | 0.254 | 0.020 |
| 0.0841 | 1.11 | 0.1930 | 0.009 | 0.078 | 0.025 | -0.042 | 0.199 | 0.062 |
| 0.0841 | 1.13 | 0.1750 | -0.008 | 0.090 | 0.031 | -0.104 | 0.206 | 0.072 |
| 0.0841 | 1.15 | 0.1598 | -0.142 | 0.089 | 0.063 | -0.381 | 0.184 | 0.129 |
| 0.0841 | 1.17 | 0.1469 | -0.443 | 0.103 | 0.099 | -0.904 | 0.195 | 0.186 |
| 0.0841 | 1.19 | 0.1357 | -0.802 | 0.118 | 0.119 | -1.447 | 0.208 | 0.205 |
| 0.0841 | 1.21 | 0.1259 | -0.729 | 0.122 | 0.081 | -1.226 | 0.203 | 0.126 |
| 0.0841 | 1.23 | 0.1174 | -0.768 | 0.117 | 0.079 | -1.233 | 0.185 | 0.118 |
| 0.0841 | 1.25 | 0.1098 | -0.476 | 0.118 | 0.082 | -0.759 | 0.178 | 0.118 |
| 0.0841 | 1.27 | 0.1030 | -0.365 | 0.105 | 0.065 | -0.586 | 0.151 | 0.090 |
| 0.0841 | 1.29 | 0.0969 | -0.199 | 0.096 | 0.048 | -0.347 | 0.133 | 0.065 |
| 0.0841 | 1.31 | 0.0914 | 0.000 | 0.091 | 0.033 | -0.075 | 0.123 | 0.043 |
| 0.0841 | 1.33 | 0.0865 | 0.002 | 0.087 | 0.028 | -0.076 | 0.114 | 0.034 |
| 0.0841 | 1.35 | 0.0820 | 0.154 | 0.087 | 0.025 | 0.115 | 0.111 | 0.030 |
| 0.0841 | 1.37 | 0.0778 | -0.079 | 0.082 | 0.026 | -0.184 | 0.103 | 0.030 |
| 0.0841 | 1.39 | 0.0740 | -0.067 | 0.087 | 0.018 | -0.170 | 0.107 | 0.019 |
| 0.0841 | 1.41 | 0.0706 | -0.055 | 0.090 | 0.017 | -0.156 | 0.109 | 0.018 |
| 0.0841 | 1.43 | 0.0674 | -0.206 | 0.098 | 0.025 | -0.336 | 0.117 | 0.028 |
| 0.0841 | 1.45 | 0.0644 | -0.015 | 0.101 | 0.019 | -0.109 | 0.118 | 0.020 |
| 0.0841 | 1.47 | 0.0616 | -0.255 | 0.112 | 0.032 | -0.384 | 0.130 | 0.035 |
| 0.0841 | 1.49 | 0.0591 | -0.094 | 0.115 | 0.025 | -0.191 | 0.132 | 0.026 |
| 0.0841 | 1.51 | 0.0567 | -0.270 | 0.118 | 0.033 | -0.382 | 0.134 | 0.036 |
| 0.0841 | 1.53 | 0.0544 | -0.174 | 0.118 | 0.030 | -0.265 | 0.132 | 0.032 |
| 0.0841 | 1.55 | 0.0524 | -0.200 | 0.112 | 0.022 | -0.290 | 0.125 | 0.024 |
| 0.0841 | 1.57 | 0.0504 | -0.079 | 0.112 | 0.017 | -0.150 | 0.124 | 0.018 |
| 0.0841 | 1.59 | 0.0486 | -0.076 | 0.128 | 0.014 | -0.141 | 0.141 | 0.016 |
| 0.0841 | 1.61 | 0.0468 | -0.024 | 0.125 | 0.013 | -0.076 | 0.136 | 0.015 |
| 0.0841 | 1.63 | 0.0452 | 0.114 | 0.138 | 0.024 | 0.082 | 0.150 | 0.026 |
| 0.0841 | 1.65 | 0.0437 | -0.313 | 0.135 | 0.028 | -0.372 | 0.146 | 0.031 |
| 0.0841 | 1.67 | 0.0422 | -0.278 | 0.140 | 0.022 | -0.325 | 0.150 | 0.022 |
| 0.0841 | 1.69 | 0.0408 | -0.147 | 0.139 | 0.018 | -0.177 | 0.149 | 0.015 |
| 0.0841 | 1.71 | 0.0395 | -0.282 | 0.142 | 0.032 | -0.314 | 0.151 | 0.033 |
| 0.0841 | 1.73 | 0.0383 | -0.559 | 0.140 | 0.055 | -0.602 | 0.149 | 0.058 |
| 0.0841 | 1.75 | 0.0371 | -0.199 | 0.148 | 0.050 | -0.219 | 0.157 | 0.050 |
| 0.0841 | 1.77 | 0.0360 | -0.333 | 0.153 | 0.050 | -0.360 | 0.161 | 0.053 |
| 0.0841 | 1.79 | 0.0349 | -0.389 | 0.156 | 0.048 | -0.419 | 0.164 | 0.050 |
| 0.0841 | 1.81 | 0.0339 | -0.461 | 0.164 | 0.045 | -0.493 | 0.171 | 0.047 |
| 0.0841 | 1.83 | 0.0329 | -0.227 | 0.166 | 0.030 | -0.246 | 0.174 | 0.031 |
| 0.0841 | 1.85 | 0.0320 | -0.189 | 0.180 | 0.018 | -0.204 | 0.188 | 0.020 |
| 0.0841 | 1.87 | 0.0311 | -0.748 | 0.188 | 0.055 | -0.783 | 0.195 | 0.057 |
| 0.0841 | 1.89 | 0.0303 | -0.285 | 0.206 | 0.011 | -0.299 | 0.214 | 0.010 |
| 0.0841 | 1.91 | 0.0295 | -0.819 | 0.227 | 0.063 | -0.849 | 0.235 | 0.065 |
| 0.1005 | 1.11 | 0.2223 | -0.056 | 0.129 | 0.040 | -0.219 | 0.352 | 0.109 |
| 0.1005 | 1.13 | 0.2022 | -0.012 | 0.054 | 0.023 | -0.115 | 0.131 | 0.062 |
| 0.1005 | 1.15 | 0.1852 | -0.102 | 0.087 | 0.058 | -0.308 | 0.190 | 0.125 |
| 0.1005 | 1.17 | 0.1706 | -0.358 | 0.104 | 0.097 | -0.780 | 0.210 | 0.192 |
| 0.1005 | 1.19 | 0.1580 | -0.545 | 0.115 | 0.106 | -1.044 | 0.215 | 0.192 |
| 0.1005 | 1.21 | 0.1469 | -0.796 | 0.111 | 0.083 | -1.403 | 0.194 | 0.135 |
| 0.1005 | 1.23 | 0.1371 | -0.839 | 0.114 | 0.086 | -1.409 | 0.189 | 0.132 |
| 0.1005 | 1.25 | 0.1284 | -0.567 | 0.107 | 0.078 | -0.936 | 0.168 | 0.116 |
| 0.1005 | 1.27 | 0.1206 | -0.264 | 0.101 | 0.062 | -0.462 | 0.151 | 0.089 |
| 0.1005 | 1.29 | 0.1137 | -0.358 | 0.091 | 0.053 | -0.594 | 0.131 | 0.073 |
| 0.1005 | 1.31 | 0.1073 | -0.088 | 0.086 | 0.035 | -0.207 | 0.121 | 0.045 |
| 0.1005 | 1.33 | 0.1016 | -0.011 | 0.081 | 0.026 | -0.104 | 0.110 | 0.033 |
| 0.1005 | 1.35 | 0.0964 | -0.095 | 0.081 | 0.026 | -0.218 | 0.107 | 0.031 |
| 0.1005 | 1.37 | 0.0916 | -0.097 | 0.081 | 0.022 | -0.221 | 0.105 | 0.025 |
| 0.1005 | 1.39 | 0.0872 | -0.019 | 0.082 | 0.018 | -0.123 | 0.104 | 0.019 |
| 0.1005 | 1.41 | 0.0832 | 0.048 | 0.081 | 0.016 | -0.043 | 0.100 | 0.016 |
| 0.1005 | 1.43 | 0.0794 | 0.041 | 0.091 | 0.014 | -0.054 | 0.111 | 0.015 |
| 0.1005 | 1.45 | 0.0760 | -0.104 | 0.095 | 0.017 | -0.229 | 0.114 | 0.018 |
| 0.1005 | 1.47 | 0.0728 | -0.131 | 0.104 | 0.023 | -0.257 | 0.123 | 0.025 |
| 0.1005 | 1.49 | 0.0698 | -0.160 | 0.109 | 0.024 | -0.282 | 0.128 | 0.025 |
| 0.1005 | 1.51 | 0.0670 | -0.002 | 0.106 | 0.024 | -0.089 | 0.122 | 0.024 |
| 0.1005 | 1.53 | 0.0644 | -0.156 | 0.109 | 0.026 | -0.260 | 0.125 | 0.028 |
| 0.1005 | 1.55 | 0.0619 | -0.119 | 0.111 | 0.027 | -0.211 | 0.126 | 0.030 |
| 0.1005 | 1.57 | 0.0596 | -0.030 | 0.116 | 0.020 | -0.106 | 0.130 | 0.021 |
| 0.1005 | 1.59 | 0.0575 | -0.003 | 0.127 | 0.023 | -0.069 | 0.142 | 0.022 |
| 0.1005 | 1.61 | 0.0555 | 0.006 | 0.129 | 0.019 | -0.052 | 0.143 | 0.017 |
| 0.1005 | 1.63 | 0.0535 | -0.135 | 0.120 | 0.025 | -0.198 | 0.131 | 0.024 |
| 0.1005 | 1.65 | 0.0517 | -0.170 | 0.126 | 0.015 | -0.225 | 0.138 | 0.014 |
| 0.1005 | 1.67 | 0.0500 | -0.167 | 0.120 | 0.017 | -0.213 | 0.131 | 0.019 |
| 0.1005 | 1.69 | 0.0484 | -0.250 | 0.140 | 0.020 | -0.295 | 0.151 | 0.021 |
| 0.1005 | 1.71 | 0.0469 | -0.050 | 0.136 | 0.024 | -0.071 | 0.147 | 0.025 |
| 0.1005 | 1.73 | 0.0454 | 0.058 | 0.146 | 0.033 | 0.050 | 0.156 | 0.034 |
| 0.1005 | 1.75 | 0.0440 | -0.013 | 0.143 | 0.037 | -0.025 | 0.153 | 0.039 |
| 0.1005 | 1.77 | 0.0427 | -0.192 | 0.144 | 0.045 | -0.216 | 0.154 | 0.047 |
| 0.1005 | 1.79 | 0.0415 | -0.282 | 0.157 | 0.035 | -0.311 | 0.166 | 0.036 |
| 0.1005 | 1.81 | 0.0403 | -0.212 | 0.146 | 0.038 | -0.236 | 0.154 | 0.040 |
| 0.1005 | 1.83 | 0.0391 | -0.297 | 0.156 | 0.032 | -0.323 | 0.165 | 0.034 |
| 0.1005 | 1.85 | 0.0380 | -0.096 | 0.160 | 0.020 | -0.109 | 0.168 | 0.019 |
| 0.1005 | 1.87 | 0.0370 | 0.097 | 0.175 | 0.032 | 0.095 | 0.183 | 0.033 |
| 0.1005 | 1.89 | 0.0360 | -0.603 | 0.189 | 0.041 | -0.634 | 0.198 | 0.043 |
| 0.1005 | 1.91 | 0.0350 | -0.626 | 0.190 | 0.045 | -0.654 | 0.198 | 0.047 |
| 0.12 | 1.11 | 0.2544 | 0.057 | 0.034 | 0.022 | 0.099 | 0.098 | 0.058 |
| 0.12 | 1.13 | 0.2323 | 0.040 | 0.116 | 0.038 | 0.017 | 0.300 | 0.102 |
| 0.12 | 1.15 | 0.2135 | -0.283 | 0.087 | 0.064 | -0.744 | 0.203 | 0.146 |
| 0.12 | 1.17 | 0.1972 | -0.475 | 0.096 | 0.091 | -1.069 | 0.206 | 0.192 |
| 0.12 | 1.19 | 0.1830 | -0.734 | 0.112 | 0.106 | -1.470 | 0.222 | 0.205 |
| 0.12 | 1.21 | 0.1705 | -1.041 | 0.112 | 0.103 | -1.931 | 0.208 | 0.182 |
| 0.12 | 1.23 | 0.1595 | -0.983 | 0.107 | 0.095 | -1.731 | 0.187 | 0.157 |
| 0.12 | 1.25 | 0.1496 | -0.582 | 0.103 | 0.076 | -1.005 | 0.170 | 0.118 |
| 0.12 | 1.27 | 0.1408 | -0.344 | 0.094 | 0.065 | -0.608 | 0.149 | 0.097 |
| 0.12 | 1.29 | 0.1328 | -0.251 | 0.085 | 0.048 | -0.462 | 0.129 | 0.068 |
| 0.12 | 1.31 | 0.1256 | -0.005 | 0.082 | 0.031 | -0.098 | 0.121 | 0.041 |
| 0.12 | 1.33 | 0.1190 | -0.032 | 0.082 | 0.024 | -0.144 | 0.116 | 0.031 |
| 0.12 | 1.35 | 0.1130 | 0.093 | 0.076 | 0.021 | 0.024 | 0.104 | 0.025 |
| 0.12 | 1.37 | 0.1075 | 0.128 | 0.082 | 0.019 | 0.062 | 0.110 | 0.021 |
| 0.12 | 1.39 | 0.1024 | -0.033 | 0.078 | 0.023 | -0.157 | 0.101 | 0.025 |
| 0.12 | 1.41 | 0.0977 | -0.100 | 0.090 | 0.017 | -0.245 | 0.115 | 0.019 |
| 0.12 | 1.43 | 0.0934 | -0.004 | 0.087 | 0.020 | -0.125 | 0.110 | 0.021 |
| 0.12 | 1.45 | 0.0894 | -0.055 | 0.090 | 0.015 | -0.187 | 0.110 | 0.017 |
| 0.12 | 1.47 | 0.0857 | 0.111 | 0.089 | 0.025 | 0.018 | 0.109 | 0.027 |
| 0.12 | 1.49 | 0.0822 | 0.085 | 0.106 | 0.022 | -0.007 | 0.127 | 0.024 |
| 0.12 | 1.51 | 0.0790 | 0.004 | 0.106 | 0.024 | -0.094 | 0.125 | 0.025 |
| 0.12 | 1.53 | 0.0759 | -0.153 | 0.114 | 0.026 | -0.271 | 0.133 | 0.028 |
| 0.12 | 1.55 | 0.0731 | 0.209 | 0.126 | 0.035 | 0.154 | 0.146 | 0.038 |
| 0.12 | 1.57 | 0.0704 | 0.014 | 0.116 | 0.030 | -0.067 | 0.132 | 0.028 |
| 0.12 | 1.59 | 0.0679 | 0.158 | 0.110 | 0.031 | 0.103 | 0.125 | 0.031 |
| 0.12 | 1.61 | 0.0655 | -0.083 | 0.124 | 0.027 | -0.161 | 0.139 | 0.028 |
| 0.12 | 1.63 | 0.0633 | -0.110 | 0.122 | 0.027 | -0.181 | 0.136 | 0.029 |
| 0.12 | 1.65 | 0.0612 | -0.253 | 0.118 | 0.026 | -0.328 | 0.131 | 0.028 |
| 0.12 | 1.67 | 0.0592 | -0.182 | 0.125 | 0.019 | -0.239 | 0.137 | 0.022 |
| 0.12 | 1.69 | 0.0573 | -0.096 | 0.134 | 0.012 | -0.134 | 0.147 | 0.014 |
| 0.12 | 1.71 | 0.0555 | -0.287 | 0.123 | 0.028 | -0.335 | 0.134 | 0.031 |
| 0.12 | 1.73 | 0.0538 | -0.142 | 0.144 | 0.025 | -0.170 | 0.157 | 0.026 |
| 0.12 | 1.75 | 0.0521 | -0.016 | 0.133 | 0.042 | -0.031 | 0.144 | 0.045 |
| 0.12 | 1.77 | 0.0506 | 0.031 | 0.139 | 0.037 | 0.019 | 0.150 | 0.040 |
| 0.12 | 1.79 | 0.0491 | -0.141 | 0.146 | 0.029 | -0.165 | 0.157 | 0.031 |
| 0.12 | 1.81 | 0.0477 | -0.083 | 0.145 | 0.026 | -0.102 | 0.155 | 0.027 |
| 0.12 | 1.83 | 0.0464 | -0.131 | 0.144 | 0.023 | -0.152 | 0.153 | 0.024 |
| 0.12 | 1.85 | 0.0451 | -0.267 | 0.161 | 0.016 | -0.293 | 0.171 | 0.016 |
| 0.12 | 1.87 | 0.0439 | -0.274 | 0.168 | 0.018 | -0.297 | 0.178 | 0.020 |
| 0.12 | 1.89 | 0.0427 | -0.258 | 0.174 | 0.013 | -0.277 | 0.183 | 0.011 |
| 0.12 | 1.91 | 0.0416 | -0.407 | 0.207 | 0.023 | -0.430 | 0.218 | 0.026 |
| 0.143 | 1.15 | 0.2444 | -0.233 | 0.078 | 0.055 | -0.648 | 0.192 | 0.133 |
| 0.143 | 1.17 | 0.2264 | -0.392 | 0.086 | 0.078 | -0.925 | 0.192 | 0.172 |
| 0.143 | 1.19 | 0.2107 | -0.564 | 0.099 | 0.084 | -1.181 | 0.206 | 0.167 |
| 0.143 | 1.21 | 0.1968 | -0.555 | 0.108 | 0.065 | -1.071 | 0.210 | 0.114 |
| 0.143 | 1.23 | 0.1844 | -0.650 | 0.105 | 0.067 | -1.196 | 0.191 | 0.110 |
| 0.143 | 1.25 | 0.1733 | -0.720 | 0.101 | 0.086 | -1.282 | 0.175 | 0.140 |
| 0.143 | 1.27 | 0.1633 | -0.295 | 0.090 | 0.053 | -0.550 | 0.148 | 0.080 |
| 0.143 | 1.29 | 0.1543 | -0.174 | 0.087 | 0.049 | -0.357 | 0.138 | 0.071 |
| 0.143 | 1.31 | 0.1461 | -0.134 | 0.082 | 0.035 | -0.300 | 0.125 | 0.048 |
| 0.143 | 1.33 | 0.1386 | 0.118 | 0.077 | 0.024 | 0.066 | 0.113 | 0.031 |
| 0.143 | 1.35 | 0.1318 | 0.008 | 0.074 | 0.020 | -0.102 | 0.105 | 0.022 |
| 0.143 | 1.37 | 0.1255 | 0.098 | 0.075 | 0.021 | 0.016 | 0.103 | 0.025 |
| 0.143 | 1.39 | 0.1197 | -0.035 | 0.075 | 0.017 | -0.172 | 0.101 | 0.019 |
| 0.143 | 1.41 | 0.1143 | -0.026 | 0.081 | 0.017 | -0.164 | 0.107 | 0.019 |
| 0.143 | 1.43 | 0.1094 | 0.046 | 0.085 | 0.014 | -0.073 | 0.110 | 0.015 |
| 0.143 | 1.45 | 0.1048 | -0.117 | 0.096 | 0.018 | -0.280 | 0.121 | 0.020 |
| 0.143 | 1.47 | 0.1005 | -0.006 | 0.108 | 0.019 | -0.137 | 0.135 | 0.020 |
| 0.143 | 1.49 | 0.0964 | 0.006 | 0.110 | 0.021 | -0.113 | 0.134 | 0.022 |
| 0.143 | 1.51 | 0.0927 | 0.066 | 0.103 | 0.025 | -0.030 | 0.124 | 0.026 |
| 0.143 | 1.53 | 0.0892 | 0.104 | 0.106 | 0.036 | 0.022 | 0.126 | 0.040 |
| 0.143 | 1.55 | 0.0859 | 0.065 | 0.106 | 0.029 | -0.021 | 0.124 | 0.030 |
| 0.143 | 1.57 | 0.0828 | 0.010 | 0.114 | 0.036 | -0.080 | 0.132 | 0.037 |
| 0.143 | 1.59 | 0.0799 | 0.111 | 0.110 | 0.034 | 0.042 | 0.126 | 0.036 |
| 0.143 | 1.61 | 0.0771 | 0.100 | 0.106 | 0.034 | 0.038 | 0.121 | 0.029 |
| 0.143 | 1.63 | 0.0745 | -0.191 | 0.109 | 0.030 | -0.282 | 0.124 | 0.031 |
| 0.143 | 1.65 | 0.0720 | 0.151 | 0.131 | 0.023 | 0.116 | 0.148 | 0.032 |
| 0.143 | 1.67 | 0.0697 | -0.175 | 0.130 | 0.019 | -0.240 | 0.146 | 0.022 |
| 0.143 | 1.69 | 0.0675 | -0.051 | 0.123 | 0.026 | -0.092 | 0.137 | 0.026 |
| 0.143 | 1.71 | 0.0654 | -0.210 | 0.130 | 0.025 | -0.259 | 0.143 | 0.027 |
| 0.143 | 1.73 | 0.0634 | -0.062 | 0.124 | 0.031 | -0.088 | 0.136 | 0.034 |
| 0.143 | 1.75 | 0.0615 | -0.191 | 0.129 | 0.033 | -0.225 | 0.141 | 0.036 |
| 0.143 | 1.77 | 0.0597 | 0.116 | 0.121 | 0.044 | 0.108 | 0.131 | 0.048 |
| 0.143 | 1.79 | 0.0580 | -0.055 | 0.128 | 0.032 | -0.077 | 0.138 | 0.034 |
| 0.143 | 1.81 | 0.0563 | -0.317 | 0.135 | 0.034 | -0.357 | 0.145 | 0.037 |
| 0.143 | 1.83 | 0.0548 | -0.238 | 0.143 | 0.024 | -0.270 | 0.153 | 0.026 |
| 0.143 | 1.85 | 0.0533 | -0.336 | 0.153 | 0.024 | -0.372 | 0.164 | 0.026 |
| 0.143 | 1.87 | 0.0518 | -0.178 | 0.150 | 0.018 | -0.199 | 0.159 | 0.020 |
| 0.143 | 1.89 | 0.0504 | 0.095 | 0.190 | 0.033 | 0.094 | 0.202 | 0.035 |
| 0.143 | 1.91 | 0.0491 | 0.004 | 0.239 | 0.033 | 0.000 | 0.253 | 0.036 |
| 0.1708 | 1.11 | 0.3269 | 0.105 | 0.069 | 0.036 | 0.276 | 0.221 | 0.114 |
| 0.1708 | 1.13 | 0.3011 | -0.024 | 0.056 | 0.022 | -0.143 | 0.162 | 0.079 |
| 0.1708 | 1.15 | 0.2787 | -0.089 | 0.076 | 0.049 | -0.296 | 0.197 | 0.124 |
| 0.1708 | 1.17 | 0.2591 | -0.368 | 0.081 | 0.070 | -0.910 | 0.192 | 0.164 |
| 0.1708 | 1.19 | 0.2418 | -0.487 | 0.096 | 0.076 | -1.069 | 0.211 | 0.162 |
| 0.1708 | 1.21 | 0.2264 | -0.651 | 0.096 | 0.066 | -1.319 | 0.196 | 0.125 |
| 0.1708 | 1.23 | 0.2126 | -0.766 | 0.096 | 0.074 | -1.477 | 0.186 | 0.132 |
| 0.1708 | 1.25 | 0.2003 | -0.395 | 0.090 | 0.056 | -0.751 | 0.164 | 0.092 |
| 0.1708 | 1.27 | 0.1891 | -0.280 | 0.086 | 0.054 | -0.547 | 0.149 | 0.087 |
| 0.1708 | 1.29 | 0.1789 | -0.240 | 0.081 | 0.044 | -0.482 | 0.134 | 0.067 |
| 0.1708 | 1.31 | 0.1697 | -0.099 | 0.074 | 0.033 | -0.257 | 0.118 | 0.048 |
| 0.1708 | 1.33 | 0.1612 | 0.034 | 0.077 | 0.026 | -0.062 | 0.118 | 0.036 |
| 0.1708 | 1.35 | 0.1535 | 0.052 | 0.074 | 0.020 | -0.045 | 0.110 | 0.025 |
| 0.1708 | 1.37 | 0.1463 | -0.017 | 0.080 | 0.023 | -0.155 | 0.116 | 0.028 |
| 0.1708 | 1.39 | 0.1397 | -0.051 | 0.076 | 0.021 | -0.209 | 0.107 | 0.024 |
| 0.1708 | 1.41 | 0.1336 | 0.130 | 0.085 | 0.014 | 0.035 | 0.116 | 0.018 |
| 0.1708 | 1.43 | 0.1279 | 0.057 | 0.083 | 0.016 | -0.070 | 0.111 | 0.017 |
| 0.1708 | 1.45 | 0.1226 | -0.035 | 0.086 | 0.016 | -0.193 | 0.113 | 0.017 |
| 0.1708 | 1.47 | 0.1177 | 0.148 | 0.084 | 0.023 | 0.046 | 0.108 | 0.025 |
| 0.1708 | 1.49 | 0.1131 | 0.062 | 0.089 | 0.022 | -0.057 | 0.113 | 0.023 |
| 0.1708 | 1.51 | 0.1088 | 0.209 | 0.095 | 0.034 | 0.137 | 0.119 | 0.037 |
| 0.1708 | 1.53 | 0.1047 | 0.045 | 0.093 | 0.030 | -0.059 | 0.113 | 0.032 |
| 0.1708 | 1.55 | 0.1009 | 0.031 | 0.115 | 0.032 | -0.072 | 0.139 | 0.049 |
| 0.1708 | 1.57 | 0.0973 | 0.215 | 0.113 | 0.043 | 0.153 | 0.135 | 0.047 |
| 0.1708 | 1.59 | 0.0939 | 0.122 | 0.109 | 0.035 | 0.047 | 0.129 | 0.043 |
| 0.1708 | 1.61 | 0.0907 | 0.167 | 0.109 | 0.038 | 0.108 | 0.127 | 0.038 |
| 0.1708 | 1.63 | 0.0877 | -0.204 | 0.113 | 0.032 | -0.313 | 0.131 | 0.037 |
| 0.1708 | 1.65 | 0.0849 | -0.001 | 0.109 | 0.030 | -0.065 | 0.125 | 0.025 |
| 0.1708 | 1.67 | 0.0821 | -0.115 | 0.125 | 0.016 | -0.184 | 0.142 | 0.018 |
| 0.1708 | 1.69 | 0.0796 | -0.099 | 0.117 | 0.017 | -0.156 | 0.132 | 0.017 |
| 0.1708 | 1.71 | 0.0771 | -0.115 | 0.134 | 0.023 | -0.164 | 0.150 | 0.023 |
| 0.1708 | 1.73 | 0.0748 | -0.071 | 0.110 | 0.023 | -0.107 | 0.122 | 0.026 |
| 0.1708 | 1.75 | 0.0726 | 0.035 | 0.114 | 0.030 | 0.015 | 0.126 | 0.033 |
| 0.1708 | 1.77 | 0.0705 | -0.234 | 0.115 | 0.034 | -0.279 | 0.127 | 0.038 |
| 0.1708 | 1.79 | 0.0685 | -0.037 | 0.125 | 0.029 | -0.061 | 0.137 | 0.032 |
| 0.1708 | 1.81 | 0.0665 | 0.011 | 0.131 | 0.029 | -0.007 | 0.143 | 0.031 |
| 0.1708 | 1.83 | 0.0647 | -0.265 | 0.133 | 0.029 | -0.305 | 0.145 | 0.031 |
| 0.1708 | 1.85 | 0.0630 | -0.466 | 0.167 | 0.034 | -0.519 | 0.180 | 0.037 |
| 0.1708 | 1.87 | 0.0613 | -0.284 | 0.173 | 0.014 | -0.318 | 0.186 | 0.017 |
| 0.1708 | 1.89 | 0.0597 | -0.478 | 0.247 | 0.034 | -0.521 | 0.265 | 0.036 |
| 0.2042 | 1.11 | 0.3673 | 0.008 | 0.039 | 0.013 | -0.030 | 0.131 | 0.043 |
| 0.2042 | 1.13 | 0.3399 | -0.143 | 0.068 | 0.027 | -0.493 | 0.202 | 0.080 |
| 0.2042 | 1.15 | 0.3160 | -0.077 | 0.058 | 0.039 | -0.263 | 0.157 | 0.105 |
| 0.2042 | 1.17 | 0.2948 | -0.194 | 0.077 | 0.056 | -0.503 | 0.193 | 0.135 |
| 0.2042 | 1.19 | 0.2760 | -0.569 | 0.094 | 0.066 | -1.295 | 0.216 | 0.145 |
| 0.2042 | 1.21 | 0.2592 | -0.526 | 0.097 | 0.059 | -1.103 | 0.208 | 0.114 |
| 0.2042 | 1.23 | 0.2441 | -0.671 | 0.104 | 0.083 | -1.343 | 0.210 | 0.155 |
| 0.2042 | 1.25 | 0.2304 | -0.360 | 0.106 | 0.080 | -0.711 | 0.201 | 0.142 |
| 0.2042 | 1.27 | 0.2180 | -0.208 | 0.099 | 0.060 | -0.434 | 0.179 | 0.099 |
| 0.2042 | 1.29 | 0.2067 | -0.101 | 0.088 | 0.035 | -0.256 | 0.152 | 0.051 |
| 0.2042 | 1.31 | 0.1964 | -0.147 | 0.080 | 0.029 | -0.345 | 0.132 | 0.043 |
| 0.2042 | 1.33 | 0.1869 | 0.030 | 0.069 | 0.017 | -0.069 | 0.110 | 0.023 |
| 0.2042 | 1.35 | 0.1781 | -0.011 | 0.067 | 0.017 | -0.146 | 0.103 | 0.022 |
| 0.2042 | 1.37 | 0.1701 | 0.084 | 0.066 | 0.019 | -0.012 | 0.098 | 0.021 |
| 0.2042 | 1.39 | 0.1626 | 0.045 | 0.073 | 0.016 | -0.080 | 0.106 | 0.018 |
| 0.2042 | 1.41 | 0.1557 | 0.021 | 0.073 | 0.019 | -0.122 | 0.103 | 0.021 |
| 0.2042 | 1.43 | 0.1492 | 0.073 | 0.075 | 0.017 | -0.056 | 0.103 | 0.020 |
| 0.2042 | 1.45 | 0.1432 | 0.021 | 0.077 | 0.013 | -0.128 | 0.104 | 0.013 |
| 0.2042 | 1.47 | 0.1375 | 0.261 | 0.087 | 0.030 | 0.191 | 0.114 | 0.036 |
| 0.2042 | 1.49 | 0.1323 | 0.162 | 0.092 | 0.028 | 0.065 | 0.119 | 0.031 |
| 0.2042 | 1.51 | 0.1273 | 0.106 | 0.113 | 0.035 | 0.002 | 0.144 | 0.052 |
| 0.2042 | 1.53 | 0.1227 | 0.216 | 0.103 | 0.043 | 0.146 | 0.130 | 0.056 |
| 0.2042 | 1.55 | 0.1183 | 0.135 | 0.096 | 0.050 | 0.048 | 0.119 | 0.046 |
| 0.2042 | 1.57 | 0.1142 | 0.101 | 0.106 | 0.040 | 0.009 | 0.129 | 0.048 |
| 0.2042 | 1.59 | 0.1103 | 0.167 | 0.100 | 0.044 | 0.095 | 0.121 | 0.047 |
| 0.2042 | 1.61 | 0.1066 | -0.142 | 0.114 | 0.033 | -0.267 | 0.136 | 0.038 |
| 0.2042 | 1.63 | 0.1031 | 0.006 | 0.101 | 0.030 | -0.079 | 0.120 | 0.035 |
| 0.2042 | 1.65 | 0.0998 | 0.133 | 0.100 | 0.034 | 0.083 | 0.116 | 0.033 |
| 0.2042 | 1.67 | 0.0967 | -0.183 | 0.115 | 0.022 | -0.273 | 0.133 | 0.034 |
| 0.2042 | 1.69 | 0.0937 | -0.024 | 0.120 | 0.013 | -0.080 | 0.138 | 0.034 |
| 0.2042 | 1.71 | 0.0908 | -0.163 | 0.125 | 0.014 | -0.227 | 0.142 | 0.015 |
| 0.2042 | 1.73 | 0.0881 | -0.010 | 0.107 | 0.019 | -0.045 | 0.121 | 0.028 |
| 0.2042 | 1.75 | 0.0856 | -0.036 | 0.110 | 0.025 | -0.069 | 0.124 | 0.028 |
| 0.2042 | 1.77 | 0.0831 | -0.247 | 0.118 | 0.029 | -0.302 | 0.132 | 0.033 |
| 0.2042 | 1.79 | 0.0808 | -0.143 | 0.129 | 0.027 | -0.183 | 0.144 | 0.031 |
| 0.2042 | 1.81 | 0.0785 | -0.196 | 0.136 | 0.023 | -0.239 | 0.150 | 0.025 |
| 0.2042 | 1.83 | 0.0764 | -0.291 | 0.157 | 0.028 | -0.340 | 0.173 | 0.031 |
| 0.2042 | 1.85 | 0.0744 | 0.100 | 0.271 | 0.027 | 0.091 | 0.296 | 0.029 |
| 0.2436 | 1.11 | 0.4092 | -0.009 | 0.027 | 0.011 | -0.079 | 0.092 | 0.043 |
| 0.2436 | 1.13 | 0.3806 | 0.013 | 0.048 | 0.023 | -0.012 | 0.147 | 0.066 |
| 0.2436 | 1.15 | 0.3553 | -0.147 | 0.071 | 0.036 | -0.455 | 0.200 | 0.100 |
| 0.2436 | 1.17 | 0.3327 | -0.203 | 0.083 | 0.056 | -0.535 | 0.214 | 0.141 |
| 0.2436 | 1.19 | 0.3126 | -0.428 | 0.089 | 0.070 | -1.006 | 0.213 | 0.163 |
| 0.2436 | 1.21 | 0.2945 | -0.589 | 0.088 | 0.067 | -1.288 | 0.198 | 0.144 |
| 0.2436 | 1.23 | 0.2781 | -0.506 | 0.087 | 0.052 | -1.050 | 0.183 | 0.099 |
| 0.2436 | 1.25 | 0.2632 | -0.514 | 0.082 | 0.054 | -1.040 | 0.164 | 0.100 |
| 0.2436 | 1.27 | 0.2496 | -0.374 | 0.073 | 0.044 | -0.757 | 0.137 | 0.078 |
| 0.2436 | 1.29 | 0.2371 | -0.281 | 0.071 | 0.040 | -0.584 | 0.127 | 0.068 |
| 0.2436 | 1.31 | 0.2257 | -0.093 | 0.064 | 0.022 | -0.261 | 0.110 | 0.034 |
| 0.2436 | 1.33 | 0.2152 | -0.116 | 0.064 | 0.024 | -0.310 | 0.107 | 0.037 |
| 0.2436 | 1.35 | 0.2054 | 0.038 | 0.063 | 0.016 | -0.072 | 0.101 | 0.023 |
| 0.2436 | 1.37 | 0.1964 | 0.068 | 0.060 | 0.018 | -0.037 | 0.093 | 0.021 |
| 0.2436 | 1.39 | 0.1881 | 0.088 | 0.062 | 0.020 | -0.019 | 0.094 | 0.022 |
| 0.2436 | 1.41 | 0.1803 | 0.108 | 0.072 | 0.017 | -0.002 | 0.106 | 0.021 |
| 0.2436 | 1.43 | 0.1730 | 0.094 | 0.076 | 0.017 | -0.030 | 0.109 | 0.020 |
| 0.2436 | 1.45 | 0.1662 | 0.213 | 0.085 | 0.035 | 0.132 | 0.119 | 0.045 |
| 0.2436 | 1.47 | 0.1598 | 0.061 | 0.088 | 0.045 | -0.079 | 0.120 | 0.033 |
| 0.2436 | 1.49 | 0.1539 | 0.158 | 0.093 | 0.038 | 0.058 | 0.125 | 0.051 |
| 0.2436 | 1.51 | 0.1482 | 0.137 | 0.103 | 0.052 | 0.037 | 0.136 | 0.059 |
| 0.2436 | 1.53 | 0.1429 | 0.149 | 0.089 | 0.056 | 0.058 | 0.116 | 0.043 |
| 0.2436 | 1.55 | 0.1380 | 0.102 | 0.088 | 0.062 | 0.000 | 0.112 | 0.053 |
| 0.2436 | 1.57 | 0.1333 | 0.269 | 0.092 | 0.049 | 0.213 | 0.115 | 0.054 |
| 0.2436 | 1.59 | 0.1288 | 0.184 | 0.091 | 0.057 | 0.109 | 0.112 | 0.053 |
| 0.2436 | 1.61 | 0.1246 | 0.051 | 0.104 | 0.038 | -0.047 | 0.127 | 0.046 |
| 0.2436 | 1.63 | 0.1206 | -0.043 | 0.095 | 0.042 | -0.149 | 0.114 | 0.039 |
| 0.2436 | 1.65 | 0.1168 | 0.108 | 0.107 | 0.030 | 0.045 | 0.127 | 0.034 |
| 0.2436 | 1.67 | 0.1132 | -0.223 | 0.089 | 0.047 | -0.336 | 0.105 | 0.039 |
| 0.2436 | 1.69 | 0.1098 | -0.163 | 0.102 | 0.034 | -0.253 | 0.120 | 0.029 |
| 0.2436 | 1.71 | 0.1065 | 0.050 | 0.101 | 0.034 | 0.008 | 0.118 | 0.033 |
| 0.2436 | 1.73 | 0.1034 | -0.020 | 0.108 | 0.036 | -0.065 | 0.125 | 0.039 |
| 0.2436 | 1.75 | 0.1004 | 0.107 | 0.125 | 0.032 | 0.086 | 0.143 | 0.035 |
| 0.2436 | 1.77 | 0.0976 | -0.032 | 0.140 | 0.025 | -0.069 | 0.159 | 0.030 |
| 0.2436 | 1.79 | 0.0949 | -0.051 | 0.231 | 0.028 | -0.087 | 0.261 | 0.031 |
| 0.2436 | 1.81 | 0.0923 | -0.380 | 0.248 | 0.047 | -0.452 | 0.278 | 0.054 |
| 0.2436 | 1.83 | 0.0898 | 0.012 | 0.188 | 0.042 | -0.010 | 0.209 | 0.048 |
| 0.2904 | 1.11 | 0.4523 | 0.025 | 0.036 | 0.017 | 0.049 | 0.125 | 0.061 |
| 0.2904 | 1.13 | 0.4228 | -0.127 | 0.045 | 0.023 | -0.445 | 0.141 | 0.071 |
| 0.2904 | 1.15 | 0.3965 | -0.203 | 0.074 | 0.041 | -0.616 | 0.217 | 0.116 |
| 0.2904 | 1.17 | 0.3728 | -0.189 | 0.064 | 0.042 | -0.504 | 0.173 | 0.111 |
| 0.2904 | 1.19 | 0.3515 | -0.459 | 0.073 | 0.050 | -1.114 | 0.182 | 0.122 |
| 0.2904 | 1.21 | 0.3322 | -0.479 | 0.077 | 0.043 | -1.076 | 0.180 | 0.092 |
| 0.2904 | 1.23 | 0.3147 | -0.488 | 0.083 | 0.051 | -1.049 | 0.182 | 0.105 |
| 0.2904 | 1.25 | 0.2986 | -0.298 | 0.082 | 0.045 | -0.627 | 0.170 | 0.084 |
| 0.2904 | 1.27 | 0.2839 | -0.309 | 0.078 | 0.047 | -0.653 | 0.154 | 0.086 |
| 0.2904 | 1.29 | 0.2704 | -0.102 | 0.069 | 0.029 | -0.264 | 0.129 | 0.051 |
| 0.2904 | 1.31 | 0.2579 | -0.051 | 0.069 | 0.027 | -0.188 | 0.123 | 0.046 |
| 0.2904 | 1.33 | 0.2463 | 0.004 | 0.065 | 0.016 | -0.109 | 0.112 | 0.027 |
| 0.2904 | 1.35 | 0.2356 | 0.024 | 0.060 | 0.011 | -0.092 | 0.101 | 0.017 |
| 0.2904 | 1.37 | 0.2257 | 0.127 | 0.069 | 0.016 | 0.061 | 0.112 | 0.022 |
| 0.2904 | 1.39 | 0.2164 | 0.025 | 0.068 | 0.023 | -0.116 | 0.107 | 0.031 |
| 0.2904 | 1.41 | 0.2077 | 0.080 | 0.068 | 0.041 | -0.041 | 0.103 | 0.038 |
| 0.2904 | 1.43 | 0.1996 | 0.042 | 0.071 | 0.045 | -0.107 | 0.105 | 0.040 |
| 0.2904 | 1.45 | 0.1920 | 0.099 | 0.085 | 0.036 | -0.028 | 0.122 | 0.054 |
| 0.2904 | 1.47 | 0.1849 | 0.146 | 0.080 | 0.040 | 0.038 | 0.113 | 0.053 |
| 0.2904 | 1.49 | 0.1781 | 0.156 | 0.081 | 0.043 | 0.054 | 0.112 | 0.045 |
| 0.2904 | 1.51 | 0.1718 | 0.136 | 0.081 | 0.051 | 0.033 | 0.110 | 0.042 |
| 0.2904 | 1.53 | 0.1659 | 0.322 | 0.079 | 0.065 | 0.285 | 0.105 | 0.065 |
| 0.2904 | 1.55 | 0.1602 | 0.193 | 0.088 | 0.057 | 0.115 | 0.115 | 0.052 |
| 0.2904 | 1.57 | 0.1549 | 0.152 | 0.086 | 0.046 | 0.062 | 0.111 | 0.051 |
| 0.2904 | 1.59 | 0.1498 | 0.177 | 0.093 | 0.046 | 0.097 | 0.119 | 0.059 |
| 0.2904 | 1.61 | 0.1450 | 0.122 | 0.090 | 0.042 | 0.033 | 0.113 | 0.047 |
| 0.2904 | 1.63 | 0.1405 | 0.034 | 0.096 | 0.047 | -0.067 | 0.119 | 0.043 |
| 0.2904 | 1.65 | 0.1362 | -0.008 | 0.139 | 0.032 | -0.106 | 0.170 | 0.061 |
| 0.2904 | 1.67 | 0.1321 | 0.106 | 0.121 | 0.036 | 0.044 | 0.146 | 0.033 |
| 0.2904 | 1.69 | 0.1282 | 0.093 | 0.173 | 0.046 | 0.038 | 0.207 | 0.048 |
| 0.2904 | 1.71 | 0.1244 | -0.018 | 0.164 | 0.045 | -0.083 | 0.195 | 0.036 |
| 0.2904 | 1.73 | 0.1209 | 0.086 | 0.174 | 0.032 | 0.050 | 0.205 | 0.036 |
| 0.2904 | 1.75 | 0.1175 | -0.042 | 0.157 | 0.025 | -0.093 | 0.183 | 0.027 |
| 0.2904 | 1.77 | 0.1142 | -0.172 | 0.135 | 0.032 | -0.238 | 0.157 | 0.036 |
| 0.2904 | 1.79 | 0.1111 | -0.214 | 0.135 | 0.034 | -0.281 | 0.155 | 0.037 |
| 0.3466 | 1.11 | 0.4964 | -0.101 | 0.067 | 0.010 | -0.385 | 0.235 | 0.039 |
| 0.3466 | 1.13 | 0.4664 | 0.030 | 0.051 | 0.027 | 0.070 | 0.165 | 0.087 |
| 0.3466 | 1.15 | 0.4395 | -0.071 | 0.053 | 0.027 | -0.223 | 0.158 | 0.080 |
| 0.3466 | 1.17 | 0.4150 | -0.209 | 0.069 | 0.036 | -0.560 | 0.190 | 0.097 |
| 0.3466 | 1.19 | 0.3928 | -0.235 | 0.078 | 0.032 | -0.565 | 0.200 | 0.076 |
| 0.3466 | 1.21 | 0.3726 | -0.295 | 0.077 | 0.030 | -0.667 | 0.187 | 0.064 |
| 0.3466 | 1.23 | 0.3540 | -0.444 | 0.084 | 0.043 | -0.980 | 0.191 | 0.088 |
| 0.3466 | 1.25 | 0.3369 | -0.248 | 0.078 | 0.036 | -0.538 | 0.167 | 0.068 |
| 0.3466 | 1.27 | 0.3212 | -0.191 | 0.071 | 0.034 | -0.427 | 0.146 | 0.065 |
| 0.3466 | 1.29 | 0.3066 | -0.069 | 0.070 | 0.030 | -0.200 | 0.137 | 0.055 |
| 0.3466 | 1.31 | 0.2932 | -0.056 | 0.065 | 0.021 | -0.194 | 0.122 | 0.038 |
| 0.3466 | 1.33 | 0.2806 | -0.067 | 0.065 | 0.022 | -0.233 | 0.116 | 0.029 |
| 0.3466 | 1.35 | 0.2690 | 0.001 | 0.062 | 0.026 | -0.128 | 0.108 | 0.020 |
| 0.3466 | 1.37 | 0.2581 | 0.042 | 0.066 | 0.027 | -0.073 | 0.110 | 0.024 |
| 0.3466 | 1.39 | 0.2479 | 0.050 | 0.062 | 0.031 | -0.073 | 0.101 | 0.029 |
| 0.3466 | 1.41 | 0.2383 | 0.031 | 0.066 | 0.029 | -0.114 | 0.104 | 0.033 |
| 0.3466 | 1.43 | 0.2294 | 0.068 | 0.074 | 0.037 | -0.065 | 0.114 | 0.031 |
| 0.3466 | 1.45 | 0.2209 | 0.160 | 0.068 | 0.046 | 0.066 | 0.102 | 0.038 |
| 0.3466 | 1.47 | 0.2130 | 0.307 | 0.066 | 0.059 | 0.277 | 0.096 | 0.046 |
| 0.3466 | 1.49 | 0.2055 | 0.390 | 0.082 | 0.050 | 0.392 | 0.118 | 0.074 |
| 0.3466 | 1.51 | 0.1985 | 0.213 | 0.076 | 0.050 | 0.141 | 0.106 | 0.052 |
| 0.3466 | 1.53 | 0.1918 | 0.206 | 0.089 | 0.058 | 0.133 | 0.122 | 0.054 |
| 0.3466 | 1.55 | 0.1855 | 0.203 | 0.092 | 0.079 | 0.129 | 0.124 | 0.057 |
| 0.3466 | 1.57 | 0.1795 | 0.195 | 0.115 | 0.075 | 0.117 | 0.153 | 0.063 |
| 0.3466 | 1.59 | 0.1738 | 0.142 | 0.123 | 0.059 | 0.048 | 0.161 | 0.039 |
| 0.3466 | 1.61 | 0.1684 | 0.216 | 0.129 | 0.093 | 0.148 | 0.167 | 0.070 |
| 0.3466 | 1.63 | 0.1633 | 0.000 | 0.164 | 0.066 | -0.120 | 0.209 | 0.083 |
| 0.3466 | 1.65 | 0.1584 | 0.027 | 0.105 | 0.038 | -0.076 | 0.132 | 0.043 |
| 0.3466 | 1.67 | 0.1537 | 0.328 | 0.120 | 0.042 | 0.310 | 0.149 | 0.060 |
| 0.3466 | 1.69 | 0.1492 | 0.230 | 0.096 | 0.032 | 0.197 | 0.118 | 0.044 |
| 0.3466 | 1.71 | 0.1450 | 0.044 | 0.103 | 0.020 | -0.019 | 0.125 | 0.029 |
| 0.3466 | 1.73 | 0.1409 | -0.002 | 0.128 | 0.033 | -0.064 | 0.154 | 0.025 |
| 0.3466 | 1.75 | 0.1371 | 0.041 | 0.162 | 0.057 | -0.004 | 0.193 | 0.062 |
| 0.3466 | 1.77 | 0.1334 | -0.363 | 0.267 | 0.080 | -0.475 | 0.316 | 0.094 |
| 0.4139 | 1.11 | 0.5406 | -0.040 | 0.040 | 0.014 | -0.160 | 0.139 | 0.046 |
| 0.4139 | 1.13 | 0.5107 | -0.135 | 0.059 | 0.020 | -0.452 | 0.189 | 0.072 |
| 0.4139 | 1.15 | 0.4835 | -0.007 | 0.057 | 0.026 | -0.025 | 0.171 | 0.075 |
| 0.4139 | 1.17 | 0.4587 | -0.216 | 0.069 | 0.031 | -0.582 | 0.192 | 0.084 |
| 0.4139 | 1.19 | 0.4359 | -0.219 | 0.079 | 0.028 | -0.531 | 0.207 | 0.070 |
| 0.4139 | 1.21 | 0.4149 | -0.215 | 0.082 | 0.040 | -0.486 | 0.202 | 0.092 |
| 0.4139 | 1.23 | 0.3955 | -0.084 | 0.083 | 0.021 | -0.165 | 0.194 | 0.088 |
| 0.4139 | 1.25 | 0.3776 | -0.268 | 0.077 | 0.033 | -0.591 | 0.170 | 0.090 |
| 0.4139 | 1.27 | 0.3610 | -0.194 | 0.075 | 0.035 | -0.438 | 0.158 | 0.078 |
| 0.4139 | 1.29 | 0.3456 | -0.093 | 0.066 | 0.022 | -0.245 | 0.132 | 0.044 |
| 0.4139 | 1.31 | 0.3312 | -0.017 | 0.062 | 0.016 | -0.117 | 0.119 | 0.031 |
| 0.4139 | 1.33 | 0.3178 | 0.050 | 0.060 | 0.018 | -0.013 | 0.112 | 0.024 |
| 0.4139 | 1.35 | 0.3052 | -0.080 | 0.066 | 0.025 | -0.268 | 0.119 | 0.032 |
| 0.4139 | 1.37 | 0.2935 | 0.139 | 0.069 | 0.031 | 0.100 | 0.119 | 0.022 |
| 0.4139 | 1.39 | 0.2824 | 0.099 | 0.059 | 0.031 | 0.015 | 0.099 | 0.032 |
| 0.4139 | 1.41 | 0.2720 | 0.048 | 0.082 | 0.034 | -0.083 | 0.133 | 0.043 |
| 0.4139 | 1.43 | 0.2622 | 0.091 | 0.082 | 0.054 | -0.023 | 0.129 | 0.041 |
| 0.4139 | 1.45 | 0.2530 | 0.059 | 0.106 | 0.049 | -0.080 | 0.164 | 0.043 |
| 0.4139 | 1.47 | 0.2443 | 0.204 | 0.082 | 0.067 | 0.134 | 0.124 | 0.050 |
| 0.4139 | 1.49 | 0.2360 | 0.146 | 0.124 | 0.055 | 0.047 | 0.183 | 0.083 |
| 0.4139 | 1.51 | 0.2282 | 0.306 | 0.115 | 0.077 | 0.278 | 0.166 | 0.092 |
| 0.4139 | 1.53 | 0.2208 | 0.201 | 0.092 | 0.063 | 0.127 | 0.130 | 0.067 |
| 0.4139 | 1.55 | 0.2138 | 0.259 | 0.088 | 0.056 | 0.207 | 0.122 | 0.071 |
| 0.4139 | 1.57 | 0.2071 | 0.280 | 0.086 | 0.061 | 0.232 | 0.117 | 0.075 |
| 0.4139 | 1.59 | 0.2008 | 0.288 | 0.080 | 0.065 | 0.240 | 0.108 | 0.066 |
| 0.4139 | 1.61 | 0.1947 | 0.122 | 0.074 | 0.065 | 0.021 | 0.098 | 0.042 |
| 0.4139 | 1.63 | 0.1890 | 0.022 | 0.093 | 0.052 | -0.103 | 0.121 | 0.036 |
| 0.4139 | 1.65 | 0.1835 | -0.024 | 0.076 | 0.062 | -0.150 | 0.098 | 0.041 |
| 0.4139 | 1.67 | 0.1782 | 0.264 | 0.081 | 0.057 | 0.226 | 0.102 | 0.049 |
| 0.4139 | 1.69 | 0.1732 | 0.383 | 0.142 | 0.065 | 0.384 | 0.177 | 0.061 |
| 0.4139 | 1.71 | 0.1684 | 0.077 | 0.132 | 0.086 | 0.012 | 0.163 | 0.085 |
| 0.494 | 1.11 | 0.5841 | -0.088 | 0.031 | 0.012 | -0.315 | 0.107 | 0.040 |
| 0.494 | 1.13 | 0.5547 | -0.043 | 0.096 | 0.030 | -0.147 | 0.308 | 0.094 |
| 0.494 | 1.15 | 0.5277 | -0.141 | 0.066 | 0.024 | -0.416 | 0.196 | 0.082 |
| 0.494 | 1.17 | 0.5028 | -0.264 | 0.073 | 0.029 | -0.717 | 0.205 | 0.085 |
| 0.494 | 1.19 | 0.4798 | -0.164 | 0.096 | 0.022 | -0.395 | 0.253 | 0.078 |
| 0.494 | 1.21 | 0.4584 | -0.287 | 0.108 | 0.032 | -0.675 | 0.270 | 0.093 |
| 0.494 | 1.23 | 0.4385 | -0.231 | 0.103 | 0.033 | -0.517 | 0.244 | 0.097 |
| 0.494 | 1.25 | 0.4200 | -0.420 | 0.105 | 0.043 | -0.942 | 0.236 | 0.096 |
| 0.494 | 1.27 | 0.4028 | -0.056 | 0.104 | 0.023 | -0.145 | 0.224 | 0.056 |
| 0.494 | 1.29 | 0.3866 | -0.018 | 0.096 | 0.017 | -0.089 | 0.196 | 0.034 |
| 0.494 | 1.31 | 0.3715 | -0.039 | 0.093 | 0.012 | -0.154 | 0.183 | 0.022 |
| 0.494 | 1.33 | 0.3573 | 0.089 | 0.078 | 0.018 | 0.070 | 0.148 | 0.019 |
| 0.494 | 1.35 | 0.3440 | 0.104 | 0.063 | 0.026 | 0.073 | 0.115 | 0.023 |
| 0.494 | 1.37 | 0.3314 | 0.124 | 0.064 | 0.029 | 0.085 | 0.114 | 0.035 |
| 0.494 | 1.39 | 0.3196 | 0.067 | 0.054 | 0.036 | -0.030 | 0.092 | 0.026 |
| 0.494 | 1.41 | 0.3084 | 0.034 | 0.056 | 0.034 | -0.099 | 0.093 | 0.034 |
| 0.494 | 1.43 | 0.2979 | 0.095 | 0.072 | 0.031 | -0.008 | 0.117 | 0.056 |
| 0.494 | 1.45 | 0.2879 | 0.229 | 0.052 | 0.062 | 0.194 | 0.082 | 0.052 |
| 0.494 | 1.47 | 0.2784 | 0.138 | 0.059 | 0.059 | 0.042 | 0.091 | 0.051 |
| 0.494 | 1.49 | 0.2694 | 0.306 | 0.062 | 0.050 | 0.292 | 0.094 | 0.066 |
| 0.494 | 1.51 | 0.2609 | 0.223 | 0.076 | 0.082 | 0.164 | 0.113 | 0.060 |
| 0.494 | 1.53 | 0.2528 | 0.262 | 0.083 | 0.073 | 0.218 | 0.121 | 0.072 |
| 0.494 | 1.55 | 0.2450 | 0.098 | 0.139 | 0.041 | -0.019 | 0.198 | 0.101 |
| 0.494 | 1.57 | 0.2377 | 0.386 | 0.100 | 0.080 | 0.383 | 0.139 | 0.080 |
| 0.494 | 1.59 | 0.2307 | 0.212 | 0.115 | 0.056 | 0.137 | 0.158 | 0.056 |
| 0.494 | 1.61 | 0.2240 | 0.023 | 0.150 | 0.078 | -0.119 | 0.203 | 0.080 |
| 0.494 | 1.63 | 0.2176 | 0.255 | 0.125 | 0.093 | 0.198 | 0.167 | 0.056 |
| 0.494 | 1.65 | 0.2115 | 0.054 | 0.453 | 0.305 | -0.062 | 0.597 | 0.406 |
| 0.5902 | 1.11 | 0.6266 | -0.026 | 0.014 | 0.005 | -0.092 | 0.048 | 0.017 |
| 0.5902 | 1.13 | 0.5982 | -0.098 | 0.063 | 0.014 | -0.309 | 0.199 | 0.045 |
| 0.5902 | 1.15 | 0.5717 | -0.117 | 0.079 | 0.013 | -0.336 | 0.232 | 0.055 |
| 0.5902 | 1.17 | 0.5471 | -0.155 | 0.064 | 0.021 | -0.410 | 0.178 | 0.063 |
| 0.5902 | 1.19 | 0.5242 | -0.197 | 0.070 | 0.025 | -0.486 | 0.185 | 0.068 |
| 0.5902 | 1.21 | 0.5028 | -0.312 | 0.080 | 0.039 | -0.747 | 0.202 | 0.091 |
| 0.5902 | 1.23 | 0.4827 | -0.171 | 0.071 | 0.023 | -0.384 | 0.168 | 0.055 |
| 0.5902 | 1.25 | 0.4639 | -0.162 | 0.065 | 0.021 | -0.367 | 0.149 | 0.044 |
| 0.5902 | 1.27 | 0.4462 | -0.028 | 0.061 | 0.014 | -0.085 | 0.134 | 0.029 |
| 0.5902 | 1.29 | 0.4296 | -0.104 | 0.056 | 0.015 | -0.264 | 0.118 | 0.032 |
| 0.5902 | 1.31 | 0.4139 | 0.008 | 0.051 | 0.011 | -0.055 | 0.103 | 0.017 |
| 0.5902 | 1.33 | 0.3991 | 0.069 | 0.048 | 0.018 | 0.041 | 0.093 | 0.018 |
| 0.5902 | 1.35 | 0.3852 | 0.072 | 0.073 | 0.021 | 0.025 | 0.136 | 0.029 |
| 0.5902 | 1.37 | 0.3720 | 0.101 | 0.052 | 0.035 | 0.058 | 0.094 | 0.020 |
| 0.5902 | 1.39 | 0.3595 | 0.082 | 0.066 | 0.028 | 0.005 | 0.117 | 0.033 |
| 0.5902 | 1.41 | 0.3476 | 0.030 | 0.081 | 0.027 | -0.099 | 0.139 | 0.073 |
| 0.5902 | 1.43 | 0.3364 | 0.088 | 0.064 | 0.043 | -0.013 | 0.107 | 0.071 |
| 0.5902 | 1.45 | 0.3257 | 0.224 | 0.077 | 0.058 | 0.199 | 0.125 | 0.050 |
| 0.5902 | 1.47 | 0.3155 | 0.256 | 0.069 | 0.066 | 0.236 | 0.110 | 0.040 |
| 0.5902 | 1.49 | 0.3058 | 0.273 | 0.098 | 0.081 | 0.251 | 0.153 | 0.089 |
| 0.5902 | 1.51 | 0.2966 | 0.297 | 0.119 | 0.052 | 0.279 | 0.182 | 0.075 |
| 0.5902 | 1.53 | 0.2878 | 0.322 | 0.151 | 0.059 | 0.311 | 0.226 | 0.099 |
| 0.7047 | 1.11 | 0.6671 | 0.009 | 0.921 | 0.294 | 0.027 | 2.976 | 0.952 |
| 0.7047 | 1.13 | 0.6399 | 0.277 | 0.358 | 0.119 | 0.847 | 1.092 | 0.358 |
| 0.7047 | 1.15 | 0.6145 | -0.071 | 0.044 | 0.015 | -0.196 | 0.128 | 0.050 |
| 0.7047 | 1.17 | 0.5906 | -0.027 | 0.059 | 0.018 | -0.055 | 0.162 | 0.044 |
| 0.7047 | 1.19 | 0.5681 | -0.102 | 0.060 | 0.013 | -0.238 | 0.157 | 0.042 |
| 0.7047 | 1.21 | 0.5470 | -0.125 | 0.076 | 0.021 | -0.286 | 0.189 | 0.039 |
| 0.7047 | 1.23 | 0.5270 | -0.084 | 0.062 | 0.010 | -0.182 | 0.148 | 0.032 |
| 0.7047 | 1.25 | 0.5082 | -0.079 | 0.073 | 0.021 | -0.182 | 0.167 | 0.045 |
| 0.7047 | 1.27 | 0.4903 | -0.033 | 0.061 | 0.012 | -0.094 | 0.133 | 0.025 |
| 0.7047 | 1.29 | 0.4735 | 0.058 | 0.071 | 0.013 | 0.079 | 0.149 | 0.027 |
| 0.7047 | 1.31 | 0.4575 | 0.172 | 0.075 | 0.020 | 0.286 | 0.153 | 0.040 |
| 0.7047 | 1.33 | 0.4423 | -0.018 | 0.055 | 0.017 | -0.120 | 0.108 | 0.019 |
| 0.7047 | 1.35 | 0.4279 | -0.019 | 0.054 | 0.025 | -0.139 | 0.104 | 0.020 |
| 0.7047 | 1.37 | 0.4142 | 0.087 | 0.058 | 0.032 | 0.043 | 0.107 | 0.024 |
| 0.7047 | 1.39 | 0.4012 | 0.071 | 0.090 | 0.036 | -0.004 | 0.162 | 0.035 |
| 0.7047 | 1.41 | 0.3888 | 0.026 | 0.137 | 0.038 | -0.097 | 0.240 | 0.087 |
Appendix C Table for Results on
Results for as shown in Fig. 14.
| W | Stat. | Syst. | |||
| (GeV2) | (GeV) | ||||
| 0.0171 | 1.15 | 0.0372 | 0.8620 | 0.3999 | 0.1226 |
| 0.0171 | 1.17 | 0.0338 | -0.3898 | 0.3905 | 0.1905 |
| 0.0171 | 1.19 | 0.0309 | -0.0851 | 0.4073 | 0.2234 |
| 0.0171 | 1.21 | 0.0285 | -0.5346 | 0.4303 | 0.2233 |
| 0.0171 | 1.23 | 0.0263 | -0.8129 | 0.3972 | 0.2110 |
| 0.0171 | 1.25 | 0.0245 | -0.3160 | 0.3701 | 0.1938 |
| 0.0171 | 1.27 | 0.0228 | -0.3262 | 0.3447 | 0.1645 |
| 0.0171 | 1.29 | 0.0214 | -0.5077 | 0.3622 | 0.1237 |
| 0.0171 | 1.31 | 0.0201 | -0.3829 | 0.3546 | 0.0797 |
| 0.0171 | 1.33 | 0.0189 | -0.5266 | 0.3524 | 0.1183 |
| 0.0171 | 1.35 | 0.0178 | 0.0724 | 0.3694 | 0.0515 |
| 0.0171 | 1.37 | 0.0169 | -0.3035 | 0.3771 | 0.0568 |
| 0.0171 | 1.39 | 0.0160 | 0.3570 | 0.4348 | 0.0713 |
| 0.0171 | 1.41 | 0.0152 | -0.1365 | 0.4247 | 0.0560 |
| 0.0171 | 1.43 | 0.0145 | 0.3522 | 0.4587 | 0.0769 |
| 0.0171 | 1.45 | 0.0138 | 0.0758 | 0.4816 | 0.0931 |
| 0.0171 | 1.47 | 0.0132 | -0.4618 | 0.5426 | 0.1103 |
| 0.0171 | 1.49 | 0.0126 | -0.5344 | 0.5338 | 0.1001 |
| 0.0171 | 1.51 | 0.0121 | -0.1916 | 0.5820 | 0.0789 |
| 0.0171 | 1.53 | 0.0116 | -0.7417 | 0.5324 | 0.0851 |
| 0.0171 | 1.55 | 0.0111 | 0.5890 | 0.6109 | 0.1275 |
| 0.0171 | 1.57 | 0.0107 | -0.3961 | 0.6446 | 0.1287 |
| 0.0171 | 1.59 | 0.0103 | 0.3460 | 0.6539 | 0.1812 |
| 0.0171 | 1.61 | 0.0099 | -0.8135 | 0.7181 | 0.1712 |
| 0.0204 | 1.13 | 0.0489 | 0.3355 | 0.3162 | 0.0945 |
| 0.0204 | 1.15 | 0.0441 | 0.0850 | 0.3007 | 0.1002 |
| 0.0204 | 1.17 | 0.0401 | -0.5514 | 0.3279 | 0.1712 |
| 0.0204 | 1.19 | 0.0367 | -0.2209 | 0.3712 | 0.2239 |
| 0.0204 | 1.21 | 0.0338 | -1.2408 | 0.3773 | 0.2645 |
| 0.0204 | 1.23 | 0.0312 | -1.0576 | 0.3547 | 0.2243 |
| 0.0204 | 1.25 | 0.0290 | -0.2166 | 0.3385 | 0.2024 |
| 0.0204 | 1.27 | 0.0271 | 0.0781 | 0.3298 | 0.1603 |
| 0.0204 | 1.29 | 0.0254 | -0.0939 | 0.3007 | 0.1154 |
| 0.0204 | 1.31 | 0.0238 | 0.2498 | 0.2925 | 0.0751 |
| 0.0204 | 1.33 | 0.0224 | -0.2802 | 0.3050 | 0.0670 |
| 0.0204 | 1.35 | 0.0212 | -0.1394 | 0.3229 | 0.0648 |
| 0.0204 | 1.37 | 0.0201 | -0.0654 | 0.3571 | 0.0499 |
| 0.0204 | 1.39 | 0.0190 | -0.0893 | 0.4030 | 0.0513 |
| 0.0204 | 1.41 | 0.0181 | -0.3663 | 0.3968 | 0.0644 |
| 0.0204 | 1.43 | 0.0172 | 0.0173 | 0.4227 | 0.0653 |
| 0.0204 | 1.45 | 0.0164 | -0.3680 | 0.4506 | 0.0967 |
| 0.0204 | 1.47 | 0.0157 | -0.2411 | 0.4990 | 0.1037 |
| 0.0204 | 1.49 | 0.0150 | -0.9938 | 0.4725 | 0.1243 |
| 0.0204 | 1.51 | 0.0144 | -0.2335 | 0.5095 | 0.0754 |
| 0.0204 | 1.53 | 0.0138 | -0.2975 | 0.4835 | 0.0681 |
| 0.0204 | 1.55 | 0.0132 | 0.0290 | 0.5753 | 0.1038 |
| 0.0204 | 1.57 | 0.0127 | -0.8249 | 0.5252 | 0.1727 |
| 0.0204 | 1.59 | 0.0122 | 0.1724 | 0.5709 | 0.1444 |
| 0.0244 | 1.11 | 0.0649 | -0.2512 | 0.2538 | 0.1420 |
| 0.0244 | 1.13 | 0.0580 | -0.6642 | 0.3195 | 0.1070 |
| 0.0244 | 1.15 | 0.0523 | -0.3208 | 0.2819 | 0.1097 |
| 0.0244 | 1.17 | 0.0476 | -0.1248 | 0.3226 | 0.1516 |
| 0.0244 | 1.19 | 0.0436 | -0.7771 | 0.3246 | 0.2257 |
| 0.0244 | 1.21 | 0.0401 | -0.5262 | 0.3371 | 0.2080 |
| 0.0244 | 1.23 | 0.0371 | -0.2295 | 0.3265 | 0.1726 |
| 0.0244 | 1.25 | 0.0345 | -0.1545 | 0.3303 | 0.2013 |
| 0.0244 | 1.27 | 0.0322 | 0.1925 | 0.2924 | 0.1524 |
| 0.0244 | 1.29 | 0.0302 | -0.2594 | 0.2860 | 0.1268 |
| 0.0244 | 1.31 | 0.0284 | 0.0522 | 0.2733 | 0.0821 |
| 0.0244 | 1.33 | 0.0267 | 0.4397 | 0.2910 | 0.0642 |
| 0.0244 | 1.35 | 0.0252 | -0.5946 | 0.2859 | 0.0883 |
| 0.0244 | 1.37 | 0.0239 | -0.2896 | 0.2950 | 0.0811 |
| 0.0244 | 1.39 | 0.0227 | -0.1598 | 0.3038 | 0.0502 |
| 0.0244 | 1.41 | 0.0216 | 0.1198 | 0.3589 | 0.0514 |
| 0.0244 | 1.43 | 0.0205 | -0.7648 | 0.3178 | 0.1014 |
| 0.0244 | 1.45 | 0.0196 | -0.8454 | 0.3783 | 0.1142 |
| 0.0244 | 1.47 | 0.0187 | -0.2395 | 0.4139 | 0.0982 |
| 0.0244 | 1.49 | 0.0179 | 0.0497 | 0.4374 | 0.0797 |
| 0.0244 | 1.51 | 0.0171 | -1.6200 | 0.4836 | 0.1618 |
| 0.0244 | 1.53 | 0.0164 | -0.6038 | 0.4487 | 0.0873 |
| 0.0244 | 1.55 | 0.0158 | 0.0364 | 0.3728 | 0.0936 |
| 0.0244 | 1.57 | 0.0152 | -0.3169 | 0.3325 | 0.1313 |
| 0.0244 | 1.59 | 0.0146 | -0.2261 | 0.2986 | 0.1290 |
| 0.0244 | 1.61 | 0.0141 | -0.2609 | 0.3176 | 0.1058 |
| 0.0244 | 1.63 | 0.0135 | 0.1788 | 0.3508 | 0.0738 |
| 0.0290 | 1.15 | 0.0616 | -0.0032 | 0.2319 | 0.0881 |
| 0.0290 | 1.17 | 0.0560 | -0.2364 | 0.2781 | 0.1468 |
| 0.0290 | 1.19 | 0.0514 | -0.5998 | 0.2913 | 0.2050 |
| 0.0290 | 1.21 | 0.0473 | -1.0787 | 0.3148 | 0.2132 |
| 0.0290 | 1.23 | 0.0438 | -0.6907 | 0.2937 | 0.1795 |
| 0.0290 | 1.25 | 0.0408 | -0.9716 | 0.2771 | 0.2160 |
| 0.0290 | 1.27 | 0.0381 | -0.6237 | 0.2555 | 0.1750 |
| 0.0290 | 1.29 | 0.0357 | -0.3775 | 0.2457 | 0.1207 |
| 0.0290 | 1.31 | 0.0335 | -0.1352 | 0.2512 | 0.0827 |
| 0.0290 | 1.33 | 0.0316 | -0.1068 | 0.2391 | 0.0602 |
| 0.0290 | 1.35 | 0.0299 | 0.0062 | 0.2550 | 0.0498 |
| 0.0290 | 1.37 | 0.0283 | -0.0837 | 0.2514 | 0.0434 |
| 0.0290 | 1.39 | 0.0268 | -0.1164 | 0.2710 | 0.0487 |
| 0.0290 | 1.41 | 0.0255 | -0.1977 | 0.2369 | 0.0510 |
| 0.0290 | 1.43 | 0.0243 | -0.0477 | 0.2391 | 0.0764 |
| 0.0290 | 1.45 | 0.0232 | -0.5433 | 0.2324 | 0.1016 |
| 0.0290 | 1.47 | 0.0221 | -0.3036 | 0.2525 | 0.0960 |
| 0.0290 | 1.49 | 0.0212 | -0.8379 | 0.2662 | 0.1036 |
| 0.0290 | 1.51 | 0.0203 | -0.5813 | 0.2532 | 0.0805 |
| 0.0290 | 1.53 | 0.0195 | -0.0728 | 0.2276 | 0.0466 |
| 0.0290 | 1.55 | 0.0187 | -0.1727 | 0.2245 | 0.0833 |
| 0.0290 | 1.57 | 0.0180 | -0.2876 | 0.2307 | 0.1068 |
| 0.0290 | 1.59 | 0.0173 | -0.5716 | 0.2378 | 0.1269 |
| 0.0290 | 1.61 | 0.0167 | -0.4824 | 0.2947 | 0.0945 |
| 0.0290 | 1.63 | 0.0161 | -0.2598 | 0.3110 | 0.0554 |
| 0.0347 | 1.15 | 0.0728 | -0.1039 | 0.2294 | 0.0906 |
| 0.0347 | 1.17 | 0.0663 | -0.2533 | 0.2435 | 0.1420 |
| 0.0347 | 1.19 | 0.0608 | -0.6814 | 0.2474 | 0.2280 |
| 0.0347 | 1.21 | 0.0561 | -0.6353 | 0.2588 | 0.2036 |
| 0.0347 | 1.23 | 0.0520 | -0.6680 | 0.2441 | 0.1582 |
| 0.0347 | 1.25 | 0.0484 | -0.4984 | 0.2120 | 0.1632 |
| 0.0347 | 1.27 | 0.0452 | -0.2943 | 0.1882 | 0.1382 |
| 0.0347 | 1.29 | 0.0424 | -0.0391 | 0.1681 | 0.0912 |
| 0.0347 | 1.31 | 0.0399 | 0.0239 | 0.1556 | 0.0630 |
| 0.0347 | 1.33 | 0.0376 | 0.0628 | 0.1484 | 0.0485 |
| 0.0347 | 1.35 | 0.0355 | 0.1050 | 0.1444 | 0.0386 |
| 0.0347 | 1.37 | 0.0337 | -0.1162 | 0.1395 | 0.0411 |
| 0.0347 | 1.39 | 0.0319 | 0.0584 | 0.1475 | 0.0337 |
| 0.0347 | 1.41 | 0.0304 | 0.0806 | 0.1559 | 0.0359 |
| 0.0347 | 1.43 | 0.0289 | -0.3288 | 0.1661 | 0.0829 |
| 0.0347 | 1.45 | 0.0276 | -0.3554 | 0.1741 | 0.0883 |
| 0.0347 | 1.47 | 0.0264 | 0.0778 | 0.1962 | 0.0970 |
| 0.0347 | 1.49 | 0.0252 | -0.2461 | 0.2081 | 0.0863 |
| 0.0347 | 1.51 | 0.0242 | 0.0615 | 0.2102 | 0.0577 |
| 0.0347 | 1.53 | 0.0232 | 0.0496 | 0.1996 | 0.0488 |
| 0.0347 | 1.55 | 0.0223 | -0.3200 | 0.2006 | 0.0986 |
| 0.0347 | 1.57 | 0.0214 | 0.0375 | 0.2036 | 0.0910 |
| 0.0347 | 1.59 | 0.0206 | -0.3357 | 0.2154 | 0.1168 |
| 0.0347 | 1.61 | 0.0199 | 0.1159 | 0.2558 | 0.0910 |
| 0.0347 | 1.63 | 0.0192 | -0.2931 | 0.2850 | 0.0654 |
| 0.0414 | 1.11 | 0.1053 | 0.0359 | 0.1109 | 0.1500 |
| 0.0414 | 1.13 | 0.0945 | 0.1266 | 0.1168 | 0.0757 |
| 0.0414 | 1.15 | 0.0856 | -0.1762 | 0.1203 | 0.0840 |
| 0.0414 | 1.17 | 0.0781 | -0.2960 | 0.1593 | 0.1353 |
| 0.0414 | 1.19 | 0.0717 | -0.4545 | 0.1751 | 0.1864 |
| 0.0414 | 1.21 | 0.0662 | -0.9961 | 0.1866 | 0.1722 |
| 0.0414 | 1.23 | 0.0614 | -0.4219 | 0.1846 | 0.1248 |
| 0.0414 | 1.25 | 0.0572 | -0.2187 | 0.1623 | 0.1489 |
| 0.0414 | 1.27 | 0.0535 | 0.0742 | 0.1455 | 0.1264 |
| 0.0414 | 1.29 | 0.0502 | -0.0561 | 0.1339 | 0.0918 |
| 0.0414 | 1.31 | 0.0472 | 0.0507 | 0.1334 | 0.0631 |
| 0.0414 | 1.33 | 0.0445 | -0.0807 | 0.1281 | 0.0500 |
| 0.0414 | 1.35 | 0.0421 | 0.0394 | 0.1269 | 0.0369 |
| 0.0414 | 1.37 | 0.0399 | -0.1174 | 0.1287 | 0.0387 |
| 0.0414 | 1.39 | 0.0379 | -0.3916 | 0.1384 | 0.0558 |
| 0.0414 | 1.41 | 0.0360 | -0.0841 | 0.1424 | 0.0442 |
| 0.0414 | 1.43 | 0.0343 | -0.0008 | 0.1495 | 0.0434 |
| 0.0414 | 1.45 | 0.0328 | 0.0347 | 0.1635 | 0.0733 |
| 0.0414 | 1.47 | 0.0313 | -0.3306 | 0.1748 | 0.0988 |
| 0.0414 | 1.49 | 0.0300 | -0.1303 | 0.1941 | 0.0721 |
| 0.0414 | 1.51 | 0.0287 | -0.4081 | 0.1976 | 0.0796 |
| 0.0414 | 1.53 | 0.0276 | -0.1110 | 0.1894 | 0.0536 |
| 0.0414 | 1.55 | 0.0265 | -0.5651 | 0.1839 | 0.1005 |
| 0.0414 | 1.57 | 0.0255 | -0.0669 | 0.1899 | 0.0809 |
| 0.0414 | 1.59 | 0.0245 | -0.0796 | 0.2036 | 0.1074 |
| 0.0414 | 1.61 | 0.0236 | -0.0631 | 0.2351 | 0.0732 |
| 0.0414 | 1.63 | 0.0228 | -0.3965 | 0.2650 | 0.0521 |
| 0.0414 | 1.65 | 0.0220 | -0.0610 | 0.3214 | 0.0390 |
| 0.0414 | 1.67 | 0.0212 | -0.7602 | 0.3819 | 0.0925 |
| 0.0414 | 1.69 | 0.0205 | -0.2819 | 0.3973 | 0.0555 |
| 0.0414 | 1.71 | 0.0199 | 0.0296 | 0.4192 | 0.1109 |
| 0.0414 | 1.73 | 0.0192 | -0.5173 | 0.4387 | 0.2111 |
| 0.0414 | 1.75 | 0.0186 | 0.0410 | 0.4515 | 0.1715 |
| 0.0414 | 1.77 | 0.0180 | -0.3964 | 0.4853 | 0.1045 |
| 0.0414 | 1.79 | 0.0175 | -0.3588 | 0.5117 | 0.0617 |
| 0.0414 | 1.81 | 0.0170 | -0.3242 | 0.5417 | 0.0590 |
| 0.0414 | 1.83 | 0.0165 | -0.6053 | 0.6056 | 0.0703 |
| 0.0414 | 1.85 | 0.0160 | -0.3747 | 0.6730 | 0.0850 |
| 0.0414 | 1.87 | 0.0156 | -0.5100 | 0.7596 | 0.0954 |
| 0.0414 | 1.89 | 0.0151 | -0.9606 | 0.8390 | 0.1213 |
| 0.0414 | 1.91 | 0.0147 | -0.3194 | 0.9403 | 0.1118 |
| 0.0494 | 1.11 | 0.1232 | -0.0537 | 0.1914 | 0.0963 |
| 0.0494 | 1.13 | 0.1108 | 0.0309 | 0.1008 | 0.0750 |
| 0.0494 | 1.15 | 0.1005 | -0.0636 | 0.1096 | 0.0785 |
| 0.0494 | 1.17 | 0.0918 | -0.4158 | 0.1366 | 0.1326 |
| 0.0494 | 1.19 | 0.0844 | -0.5494 | 0.1670 | 0.1768 |
| 0.0494 | 1.21 | 0.0780 | -0.3791 | 0.1738 | 0.1270 |
| 0.0494 | 1.23 | 0.0724 | -1.0036 | 0.1655 | 0.1491 |
| 0.0494 | 1.25 | 0.0675 | -0.3109 | 0.1457 | 0.1565 |
| 0.0494 | 1.27 | 0.0632 | -0.3724 | 0.1345 | 0.1362 |
| 0.0494 | 1.29 | 0.0593 | -0.2382 | 0.1267 | 0.0964 |
| 0.0494 | 1.31 | 0.0558 | 0.0808 | 0.1232 | 0.0624 |
| 0.0494 | 1.33 | 0.0527 | -0.1160 | 0.1199 | 0.0475 |
| 0.0494 | 1.35 | 0.0498 | 0.2128 | 0.1176 | 0.0399 |
| 0.0494 | 1.37 | 0.0472 | 0.0725 | 0.1153 | 0.0308 |
| 0.0494 | 1.39 | 0.0449 | -0.1481 | 0.1249 | 0.0372 |
| 0.0494 | 1.41 | 0.0427 | -0.2838 | 0.1278 | 0.0422 |
| 0.0494 | 1.43 | 0.0407 | -0.0925 | 0.1391 | 0.0454 |
| 0.0494 | 1.45 | 0.0389 | -0.4597 | 0.1451 | 0.0953 |
| 0.0494 | 1.47 | 0.0371 | -0.0926 | 0.1607 | 0.0882 |
| 0.0494 | 1.49 | 0.0356 | -0.4756 | 0.1747 | 0.1020 |
| 0.0494 | 1.51 | 0.0341 | -0.1950 | 0.1768 | 0.0654 |
| 0.0494 | 1.53 | 0.0327 | -0.0537 | 0.1714 | 0.0400 |
| 0.0494 | 1.55 | 0.0314 | -0.1674 | 0.1731 | 0.0635 |
| 0.0494 | 1.57 | 0.0302 | -0.0551 | 0.1752 | 0.0715 |
| 0.0494 | 1.59 | 0.0291 | 0.1432 | 0.1853 | 0.0973 |
| 0.0494 | 1.61 | 0.0281 | -0.3408 | 0.2137 | 0.0757 |
| 0.0494 | 1.63 | 0.0271 | 0.0802 | 0.2477 | 0.0446 |
| 0.0494 | 1.65 | 0.0261 | 0.1287 | 0.3024 | 0.0479 |
| 0.0494 | 1.67 | 0.0252 | 0.2622 | 0.3487 | 0.0442 |
| 0.0494 | 1.69 | 0.0244 | -0.0304 | 0.3695 | 0.0519 |
| 0.0494 | 1.71 | 0.0236 | -0.0444 | 0.3812 | 0.0992 |
| 0.0494 | 1.73 | 0.0229 | 0.0857 | 0.3958 | 0.2099 |
| 0.0494 | 1.75 | 0.0221 | -0.3381 | 0.4210 | 0.1853 |
| 0.0494 | 1.77 | 0.0215 | 0.3185 | 0.4461 | 0.1139 |
| 0.0494 | 1.79 | 0.0208 | 0.0775 | 0.4707 | 0.0694 |
| 0.0494 | 1.81 | 0.0202 | -0.0098 | 0.5093 | 0.0520 |
| 0.0494 | 1.83 | 0.0196 | 0.0816 | 0.5479 | 0.0670 |
| 0.0494 | 1.85 | 0.0191 | -0.2904 | 0.5798 | 0.0877 |
| 0.0494 | 1.87 | 0.0185 | -0.8560 | 0.6514 | 0.1023 |
| 0.0494 | 1.89 | 0.0180 | -0.8523 | 0.6964 | 0.1186 |
| 0.0494 | 1.91 | 0.0175 | -0.3603 | 0.8074 | 0.1103 |
| 0.0590 | 1.11 | 0.1437 | 0.0710 | 0.1266 | 0.0884 |
| 0.0590 | 1.13 | 0.1295 | -0.0791 | 0.0800 | 0.0656 |
| 0.0590 | 1.15 | 0.1177 | -0.1230 | 0.1099 | 0.0808 |
| 0.0590 | 1.17 | 0.1078 | -0.5296 | 0.1340 | 0.1408 |
| 0.0590 | 1.19 | 0.0992 | -0.6355 | 0.1609 | 0.1915 |
| 0.0590 | 1.21 | 0.0918 | -0.4015 | 0.1670 | 0.1248 |
| 0.0590 | 1.23 | 0.0853 | -0.4754 | 0.1524 | 0.1055 |
| 0.0590 | 1.25 | 0.0796 | -0.4645 | 0.1424 | 0.1563 |
| 0.0590 | 1.27 | 0.0745 | -0.2958 | 0.1292 | 0.1435 |
| 0.0590 | 1.29 | 0.0700 | -0.0143 | 0.1190 | 0.0880 |
| 0.0590 | 1.31 | 0.0659 | -0.0179 | 0.1101 | 0.0578 |
| 0.0590 | 1.33 | 0.0623 | -0.0079 | 0.1088 | 0.0431 |
| 0.0590 | 1.35 | 0.0589 | -0.1333 | 0.1137 | 0.0380 |
| 0.0590 | 1.37 | 0.0559 | 0.0660 | 0.1204 | 0.0323 |
| 0.0590 | 1.39 | 0.0531 | -0.0682 | 0.1136 | 0.0290 |
| 0.0590 | 1.41 | 0.0506 | 0.0223 | 0.1222 | 0.0313 |
| 0.0590 | 1.43 | 0.0482 | -0.2391 | 0.1253 | 0.0492 |
| 0.0590 | 1.45 | 0.0461 | -0.0980 | 0.1316 | 0.0793 |
| 0.0590 | 1.47 | 0.0440 | -0.0293 | 0.1501 | 0.0922 |
| 0.0590 | 1.49 | 0.0422 | -0.1299 | 0.1696 | 0.0732 |
| 0.0590 | 1.51 | 0.0404 | -0.3801 | 0.1692 | 0.0586 |
| 0.0590 | 1.53 | 0.0388 | -0.0406 | 0.1582 | 0.0345 |
| 0.0590 | 1.55 | 0.0373 | 0.2838 | 0.1548 | 0.0557 |
| 0.0590 | 1.57 | 0.0359 | 0.0963 | 0.1566 | 0.0640 |
| 0.0590 | 1.59 | 0.0346 | -0.0140 | 0.1622 | 0.0897 |
| 0.0590 | 1.61 | 0.0333 | -0.1131 | 0.1946 | 0.0585 |
| 0.0590 | 1.63 | 0.0321 | -0.1525 | 0.2183 | 0.0321 |
| 0.0590 | 1.65 | 0.0310 | -0.1453 | 0.2607 | 0.0307 |
| 0.0590 | 1.67 | 0.0300 | -0.0035 | 0.3039 | 0.0241 |
| 0.0590 | 1.69 | 0.0290 | -0.0973 | 0.3300 | 0.0343 |
| 0.0590 | 1.71 | 0.0281 | -0.4828 | 0.3563 | 0.0859 |
| 0.0590 | 1.73 | 0.0272 | -0.3051 | 0.3699 | 0.2122 |
| 0.0590 | 1.75 | 0.0263 | -0.3357 | 0.3701 | 0.1928 |
| 0.0590 | 1.77 | 0.0255 | -0.0976 | 0.3858 | 0.1103 |
| 0.0590 | 1.79 | 0.0248 | -0.4655 | 0.3990 | 0.0708 |
| 0.0590 | 1.81 | 0.0240 | -0.6150 | 0.4191 | 0.0643 |
| 0.0590 | 1.83 | 0.0233 | -0.1045 | 0.4535 | 0.0582 |
| 0.0590 | 1.85 | 0.0227 | -0.0061 | 0.4748 | 0.0767 |
| 0.0590 | 1.87 | 0.0221 | -0.0591 | 0.5256 | 0.0878 |
| 0.0590 | 1.89 | 0.0214 | -0.4645 | 0.5617 | 0.1200 |
| 0.0590 | 1.91 | 0.0209 | -0.0511 | 0.6611 | 0.1112 |
| 0.0705 | 1.11 | 0.1670 | -0.0051 | 0.0712 | 0.1750 |
| 0.0705 | 1.13 | 0.1510 | -0.0155 | 0.1044 | 0.0786 |
| 0.0705 | 1.15 | 0.1375 | 0.0225 | 0.1007 | 0.0769 |
| 0.0705 | 1.17 | 0.1261 | -0.2623 | 0.1249 | 0.1215 |
| 0.0705 | 1.19 | 0.1163 | -0.8330 | 0.1508 | 0.1881 |
| 0.0705 | 1.21 | 0.1078 | -0.7215 | 0.1627 | 0.1446 |
| 0.0705 | 1.23 | 0.1003 | -0.5827 | 0.1465 | 0.1099 |
| 0.0705 | 1.25 | 0.0937 | -0.4814 | 0.1344 | 0.1507 |
| 0.0705 | 1.27 | 0.0878 | -0.0191 | 0.1215 | 0.1303 |
| 0.0705 | 1.29 | 0.0825 | 0.0511 | 0.1156 | 0.1000 |
| 0.0705 | 1.31 | 0.0778 | 0.0295 | 0.1048 | 0.0592 |
| 0.0705 | 1.33 | 0.0735 | -0.0820 | 0.1031 | 0.0431 |
| 0.0705 | 1.35 | 0.0696 | -0.1274 | 0.1025 | 0.0339 |
| 0.0705 | 1.37 | 0.0661 | -0.0983 | 0.1001 | 0.0316 |
| 0.0705 | 1.39 | 0.0628 | -0.0661 | 0.1078 | 0.0272 |
| 0.0705 | 1.41 | 0.0598 | -0.1165 | 0.1039 | 0.0309 |
| 0.0705 | 1.43 | 0.0571 | 0.0337 | 0.1184 | 0.0350 |
| 0.0705 | 1.45 | 0.0545 | -0.1264 | 0.1273 | 0.0711 |
| 0.0705 | 1.47 | 0.0522 | -0.3201 | 0.1428 | 0.0911 |
| 0.0705 | 1.49 | 0.0500 | -0.4456 | 0.1584 | 0.0722 |
| 0.0705 | 1.51 | 0.0480 | 0.1109 | 0.1609 | 0.0563 |
| 0.0705 | 1.53 | 0.0460 | -0.1575 | 0.1528 | 0.0270 |
| 0.0705 | 1.55 | 0.0443 | -0.1552 | 0.1374 | 0.0439 |
| 0.0705 | 1.57 | 0.0426 | -0.0783 | 0.1468 | 0.0526 |
| 0.0705 | 1.59 | 0.0410 | -0.1701 | 0.1551 | 0.0792 |
| 0.0705 | 1.61 | 0.0396 | -0.0334 | 0.1860 | 0.0465 |
| 0.0705 | 1.63 | 0.0382 | 0.3842 | 0.2353 | 0.0305 |
| 0.0705 | 1.65 | 0.0369 | 0.2041 | 0.2776 | 0.0260 |
| 0.0705 | 1.67 | 0.0356 | -0.0222 | 0.3090 | 0.0467 |
| 0.0705 | 1.69 | 0.0345 | 0.1476 | 0.3251 | 0.0473 |
| 0.0705 | 1.71 | 0.0333 | -0.0850 | 0.3247 | 0.0677 |
| 0.0705 | 1.73 | 0.0323 | -0.2170 | 0.3255 | 0.2217 |
| 0.0705 | 1.75 | 0.0313 | -0.2375 | 0.3328 | 0.2061 |
| 0.0705 | 1.77 | 0.0303 | -0.0948 | 0.3408 | 0.1124 |
| 0.0705 | 1.79 | 0.0294 | 0.0835 | 0.3546 | 0.0609 |
| 0.0705 | 1.81 | 0.0286 | -0.5333 | 0.3739 | 0.0656 |
| 0.0705 | 1.83 | 0.0278 | -0.4282 | 0.4035 | 0.0623 |
| 0.0705 | 1.85 | 0.0270 | -0.3502 | 0.4231 | 0.0745 |
| 0.0705 | 1.87 | 0.0262 | -0.2970 | 0.4778 | 0.0760 |
| 0.0705 | 1.89 | 0.0255 | -0.3819 | 0.5200 | 0.0931 |
| 0.0705 | 1.91 | 0.0248 | -0.5360 | 0.6073 | 0.0925 |
| 0.0841 | 1.11 | 0.1930 | 0.0483 | 0.0951 | 0.0739 |
| 0.0841 | 1.13 | 0.1750 | 0.0139 | 0.0978 | 0.0614 |
| 0.0841 | 1.15 | 0.1598 | -0.0642 | 0.0977 | 0.0749 |
| 0.0841 | 1.17 | 0.1469 | -0.3013 | 0.1170 | 0.1149 |
| 0.0841 | 1.19 | 0.1357 | -0.6114 | 0.1407 | 0.1640 |
| 0.0841 | 1.21 | 0.1259 | -0.4651 | 0.1492 | 0.1177 |
| 0.0841 | 1.23 | 0.1174 | -0.4957 | 0.1388 | 0.0994 |
| 0.0841 | 1.25 | 0.1098 | -0.2233 | 0.1323 | 0.1328 |
| 0.0841 | 1.27 | 0.1030 | -0.2048 | 0.1149 | 0.1240 |
| 0.0841 | 1.29 | 0.0969 | -0.1115 | 0.1047 | 0.0930 |
| 0.0841 | 1.31 | 0.0914 | 0.0467 | 0.1000 | 0.0629 |
| 0.0841 | 1.33 | 0.0865 | 0.0185 | 0.0974 | 0.0430 |
| 0.0841 | 1.35 | 0.0820 | 0.1566 | 0.0982 | 0.0354 |
| 0.0841 | 1.37 | 0.0778 | -0.1158 | 0.0943 | 0.0352 |
| 0.0841 | 1.39 | 0.0740 | -0.0916 | 0.1005 | 0.0271 |
| 0.0841 | 1.41 | 0.0706 | -0.0543 | 0.1046 | 0.0255 |
| 0.0841 | 1.43 | 0.0674 | -0.2090 | 0.1137 | 0.0387 |
| 0.0841 | 1.45 | 0.0644 | 0.0068 | 0.1191 | 0.0657 |
| 0.0841 | 1.47 | 0.0616 | -0.2285 | 0.1370 | 0.0861 |
| 0.0841 | 1.49 | 0.0591 | -0.0523 | 0.1493 | 0.0577 |
| 0.0841 | 1.51 | 0.0567 | -0.2858 | 0.1513 | 0.0568 |
| 0.0841 | 1.53 | 0.0544 | -0.1703 | 0.1424 | 0.0375 |
| 0.0841 | 1.55 | 0.0524 | -0.2007 | 0.1328 | 0.0363 |
| 0.0841 | 1.57 | 0.0504 | -0.0759 | 0.1358 | 0.0496 |
| 0.0841 | 1.59 | 0.0486 | -0.0610 | 0.1602 | 0.0675 |
| 0.0841 | 1.61 | 0.0468 | 0.0569 | 0.1822 | 0.0365 |
| 0.0841 | 1.63 | 0.0452 | 0.2831 | 0.2357 | 0.0308 |
| 0.0841 | 1.65 | 0.0437 | -0.1007 | 0.2704 | 0.0355 |
| 0.0841 | 1.67 | 0.0422 | -0.0719 | 0.2924 | 0.0467 |
| 0.0841 | 1.69 | 0.0408 | 0.0467 | 0.2947 | 0.0583 |
| 0.0841 | 1.71 | 0.0395 | -0.1300 | 0.2943 | 0.0466 |
| 0.0841 | 1.73 | 0.0383 | -0.4862 | 0.2954 | 0.2296 |
| 0.0841 | 1.75 | 0.0371 | -0.1174 | 0.3062 | 0.2153 |
| 0.0841 | 1.77 | 0.0360 | -0.2129 | 0.3184 | 0.1162 |
| 0.0841 | 1.79 | 0.0349 | -0.1865 | 0.3317 | 0.0629 |
| 0.0841 | 1.81 | 0.0339 | -0.1856 | 0.3558 | 0.0564 |
| 0.0841 | 1.83 | 0.0329 | 0.1167 | 0.3893 | 0.0576 |
| 0.0841 | 1.85 | 0.0320 | 0.1530 | 0.4162 | 0.0766 |
| 0.0841 | 1.87 | 0.0311 | -0.4480 | 0.4895 | 0.1297 |
| 0.0841 | 1.89 | 0.0303 | 0.0862 | 0.6053 | 0.1908 |
| 0.0841 | 1.91 | 0.0295 | -0.4252 | 0.9198 | 0.2458 |
| 0.1005 | 1.11 | 0.2223 | -0.1279 | 0.1618 | 0.4271 |
| 0.1005 | 1.13 | 0.2022 | -0.0157 | 0.0722 | 0.1764 |
| 0.1005 | 1.15 | 0.1852 | -0.0470 | 0.0954 | 0.0931 |
| 0.1005 | 1.17 | 0.1706 | -0.2292 | 0.1172 | 0.1122 |
| 0.1005 | 1.19 | 0.1580 | -0.3457 | 0.1375 | 0.1517 |
| 0.1005 | 1.21 | 0.1469 | -0.5533 | 0.1359 | 0.1148 |
| 0.1005 | 1.23 | 0.1371 | -0.5860 | 0.1347 | 0.1067 |
| 0.1005 | 1.25 | 0.1284 | -0.3193 | 0.1199 | 0.1310 |
| 0.1005 | 1.27 | 0.1206 | -0.0771 | 0.1102 | 0.1144 |
| 0.1005 | 1.29 | 0.1137 | -0.2484 | 0.0993 | 0.0905 |
| 0.1005 | 1.31 | 0.1073 | -0.0224 | 0.0948 | 0.0613 |
| 0.1005 | 1.33 | 0.1016 | 0.0109 | 0.0906 | 0.0428 |
| 0.1005 | 1.35 | 0.0964 | -0.1086 | 0.0909 | 0.0371 |
| 0.1005 | 1.37 | 0.0916 | -0.1143 | 0.0907 | 0.0322 |
| 0.1005 | 1.39 | 0.0872 | -0.0139 | 0.0929 | 0.0275 |
| 0.1005 | 1.41 | 0.0832 | 0.0595 | 0.0931 | 0.0234 |
| 0.1005 | 1.43 | 0.0794 | 0.0558 | 0.1040 | 0.0285 |
| 0.1005 | 1.45 | 0.0760 | -0.0869 | 0.1100 | 0.0594 |
| 0.1005 | 1.47 | 0.0728 | -0.1159 | 0.1252 | 0.0757 |
| 0.1005 | 1.49 | 0.0698 | -0.1732 | 0.1391 | 0.0481 |
| 0.1005 | 1.51 | 0.0670 | -0.0012 | 0.1342 | 0.0455 |
| 0.1005 | 1.53 | 0.0644 | -0.1708 | 0.1294 | 0.0346 |
| 0.1005 | 1.55 | 0.0619 | -0.1081 | 0.1278 | 0.0350 |
| 0.1005 | 1.57 | 0.0596 | -0.0014 | 0.1344 | 0.0403 |
| 0.1005 | 1.59 | 0.0575 | 0.0510 | 0.1494 | 0.0647 |
| 0.1005 | 1.61 | 0.0555 | 0.1055 | 0.1575 | 0.0352 |
| 0.1005 | 1.63 | 0.0535 | -0.0107 | 0.1561 | 0.0336 |
| 0.1005 | 1.65 | 0.0517 | -0.0425 | 0.1656 | 0.0316 |
| 0.1005 | 1.67 | 0.0500 | -0.0178 | 0.1638 | 0.0610 |
| 0.1005 | 1.69 | 0.0484 | -0.1089 | 0.1828 | 0.0747 |
| 0.1005 | 1.71 | 0.0469 | 0.1036 | 0.1786 | 0.0335 |
| 0.1005 | 1.73 | 0.0454 | 0.2079 | 0.1880 | 0.2324 |
| 0.1005 | 1.75 | 0.0440 | 0.1147 | 0.1865 | 0.2210 |
| 0.1005 | 1.77 | 0.0427 | -0.0895 | 0.1895 | 0.1136 |
| 0.1005 | 1.79 | 0.0415 | -0.1949 | 0.2033 | 0.0499 |
| 0.1005 | 1.81 | 0.0403 | -0.1346 | 0.1946 | 0.0477 |
| 0.1005 | 1.83 | 0.0391 | -0.2509 | 0.2052 | 0.0525 |
| 0.1005 | 1.85 | 0.0380 | -0.0479 | 0.2116 | 0.0576 |
| 0.1005 | 1.87 | 0.0370 | 0.1798 | 0.2288 | 0.0689 |
| 0.1005 | 1.89 | 0.0360 | -0.5636 | 0.2475 | 0.0860 |
| 0.1005 | 1.91 | 0.0350 | -0.5694 | 0.2540 | 0.0915 |
| 0.120 | 1.11 | 0.2544 | 0.0782 | 0.0451 | 0.0469 |
| 0.120 | 1.13 | 0.2323 | 0.0758 | 0.1197 | 0.0544 |
| 0.120 | 1.15 | 0.2135 | -0.2224 | 0.0937 | 0.0702 |
| 0.120 | 1.17 | 0.1972 | -0.3711 | 0.1089 | 0.1076 |
| 0.120 | 1.19 | 0.1830 | -0.6033 | 0.1350 | 0.1461 |
| 0.120 | 1.21 | 0.1705 | -0.9125 | 0.1401 | 0.1357 |
| 0.120 | 1.23 | 0.1595 | -0.7915 | 0.1287 | 0.1170 |
| 0.120 | 1.25 | 0.1496 | -0.3618 | 0.1169 | 0.1228 |
| 0.120 | 1.27 | 0.1408 | -0.1865 | 0.1029 | 0.1175 |
| 0.120 | 1.29 | 0.1328 | -0.1738 | 0.0938 | 0.0817 |
| 0.120 | 1.31 | 0.1256 | 0.0407 | 0.0909 | 0.0540 |
| 0.120 | 1.33 | 0.1190 | -0.0009 | 0.0914 | 0.0388 |
| 0.120 | 1.35 | 0.1130 | 0.1280 | 0.0860 | 0.0311 |
| 0.120 | 1.37 | 0.1075 | 0.1515 | 0.0927 | 0.0288 |
| 0.120 | 1.39 | 0.1024 | -0.0299 | 0.0896 | 0.0306 |
| 0.120 | 1.41 | 0.0977 | -0.0982 | 0.1021 | 0.0237 |
| 0.120 | 1.43 | 0.0934 | 0.0100 | 0.0998 | 0.0307 |
| 0.120 | 1.45 | 0.0894 | -0.0494 | 0.1049 | 0.0543 |
| 0.120 | 1.47 | 0.0857 | 0.1143 | 0.1106 | 0.0688 |
| 0.120 | 1.49 | 0.0822 | 0.0550 | 0.1384 | 0.0433 |
| 0.120 | 1.51 | 0.0790 | -0.0677 | 0.1390 | 0.0459 |
| 0.120 | 1.53 | 0.0759 | -0.2170 | 0.1405 | 0.0458 |
| 0.120 | 1.55 | 0.0731 | 0.1938 | 0.1476 | 0.0513 |
| 0.120 | 1.57 | 0.0704 | 0.0055 | 0.1385 | 0.0474 |
| 0.120 | 1.59 | 0.0679 | 0.1834 | 0.1377 | 0.0571 |
| 0.120 | 1.61 | 0.0655 | -0.0507 | 0.1567 | 0.0366 |
| 0.120 | 1.63 | 0.0633 | -0.0783 | 0.1599 | 0.0336 |
| 0.120 | 1.65 | 0.0612 | -0.2344 | 0.1598 | 0.0493 |
| 0.120 | 1.67 | 0.0592 | -0.1464 | 0.1686 | 0.0852 |
| 0.120 | 1.69 | 0.0573 | -0.0349 | 0.1775 | 0.0828 |
| 0.120 | 1.71 | 0.0555 | -0.2503 | 0.1665 | 0.0340 |
| 0.120 | 1.73 | 0.0538 | -0.1338 | 0.1848 | 0.2342 |
| 0.120 | 1.75 | 0.0521 | -0.0220 | 0.1747 | 0.2244 |
| 0.120 | 1.77 | 0.0506 | 0.0308 | 0.1806 | 0.1075 |
| 0.120 | 1.79 | 0.0491 | -0.1348 | 0.1881 | 0.0420 |
| 0.120 | 1.81 | 0.0477 | -0.0394 | 0.1886 | 0.0364 |
| 0.120 | 1.83 | 0.0464 | -0.0682 | 0.1901 | 0.0453 |
| 0.120 | 1.85 | 0.0451 | -0.2178 | 0.2083 | 0.0522 |
| 0.120 | 1.87 | 0.0439 | -0.2373 | 0.2176 | 0.0585 |
| 0.120 | 1.89 | 0.0427 | -0.2513 | 0.2281 | 0.0698 |
| 0.120 | 1.91 | 0.0416 | -0.4381 | 0.2632 | 0.0766 |
| 0.143 | 1.15 | 0.2444 | -0.1950 | 0.0852 | 0.0614 |
| 0.143 | 1.17 | 0.2264 | -0.3025 | 0.0990 | 0.0974 |
| 0.143 | 1.19 | 0.2107 | -0.4154 | 0.1225 | 0.1252 |
| 0.143 | 1.21 | 0.1968 | -0.3442 | 0.1365 | 0.0910 |
| 0.143 | 1.23 | 0.1844 | -0.4355 | 0.1277 | 0.0906 |
| 0.143 | 1.25 | 0.1733 | -0.5362 | 0.1145 | 0.1241 |
| 0.143 | 1.27 | 0.1633 | -0.1255 | 0.0984 | 0.0982 |
| 0.143 | 1.29 | 0.1543 | -0.0507 | 0.0938 | 0.0820 |
| 0.143 | 1.31 | 0.1461 | -0.0722 | 0.0899 | 0.0504 |
| 0.143 | 1.33 | 0.1386 | 0.1459 | 0.0859 | 0.0341 |
| 0.143 | 1.35 | 0.1318 | 0.0089 | 0.0840 | 0.0276 |
| 0.143 | 1.37 | 0.1255 | 0.0968 | 0.0861 | 0.0283 |
| 0.143 | 1.39 | 0.1197 | -0.0520 | 0.0873 | 0.0243 |
| 0.143 | 1.41 | 0.1143 | -0.0546 | 0.0949 | 0.0232 |
| 0.143 | 1.43 | 0.1094 | 0.0145 | 0.1017 | 0.0258 |
| 0.143 | 1.45 | 0.1048 | -0.1436 | 0.1167 | 0.0529 |
| 0.143 | 1.47 | 0.1005 | -0.0076 | 0.1365 | 0.0626 |
| 0.143 | 1.49 | 0.0964 | -0.0102 | 0.1475 | 0.0474 |
| 0.143 | 1.51 | 0.0927 | 0.0311 | 0.1380 | 0.0517 |
| 0.143 | 1.53 | 0.0892 | 0.0664 | 0.1335 | 0.0513 |
| 0.143 | 1.55 | 0.0859 | 0.0404 | 0.1303 | 0.0406 |
| 0.143 | 1.57 | 0.0828 | -0.0020 | 0.1388 | 0.0439 |
| 0.143 | 1.59 | 0.0799 | 0.1167 | 0.1386 | 0.0434 |
| 0.143 | 1.61 | 0.0771 | 0.1118 | 0.1396 | 0.0403 |
| 0.143 | 1.63 | 0.0745 | -0.1935 | 0.1461 | 0.0342 |
| 0.143 | 1.65 | 0.0720 | 0.1904 | 0.1689 | 0.0581 |
| 0.143 | 1.67 | 0.0697 | -0.1423 | 0.1708 | 0.0962 |
| 0.143 | 1.69 | 0.0675 | 0.0243 | 0.1649 | 0.0968 |
| 0.143 | 1.71 | 0.0654 | -0.1378 | 0.1702 | 0.0334 |
| 0.143 | 1.73 | 0.0634 | -0.0056 | 0.1644 | 0.2334 |
| 0.143 | 1.75 | 0.0615 | -0.1484 | 0.1697 | 0.2167 |
| 0.143 | 1.77 | 0.0597 | 0.2006 | 0.1637 | 0.1035 |
| 0.143 | 1.79 | 0.0580 | 0.0223 | 0.1711 | 0.0429 |
| 0.143 | 1.81 | 0.0563 | -0.2492 | 0.1787 | 0.0447 |
| 0.143 | 1.83 | 0.0548 | -0.1566 | 0.1876 | 0.0469 |
| 0.143 | 1.85 | 0.0533 | -0.2761 | 0.1988 | 0.0535 |
| 0.143 | 1.87 | 0.0518 | -0.1215 | 0.1988 | 0.0521 |
| 0.143 | 1.89 | 0.0504 | 0.1713 | 0.2419 | 0.0711 |
| 0.143 | 1.91 | 0.0491 | 0.0748 | 0.2933 | 0.0769 |
| 0.171 | 1.11 | 0.3269 | 0.1059 | 0.0665 | 0.0707 |
| 0.171 | 1.13 | 0.3011 | 0.0044 | 0.0581 | 0.0378 |
| 0.171 | 1.15 | 0.2787 | -0.0225 | 0.0807 | 0.0538 |
| 0.171 | 1.17 | 0.2591 | -0.2662 | 0.0919 | 0.0881 |
| 0.171 | 1.19 | 0.2418 | -0.3473 | 0.1175 | 0.1091 |
| 0.171 | 1.21 | 0.2264 | -0.4856 | 0.1212 | 0.0869 |
| 0.171 | 1.23 | 0.2126 | -0.6009 | 0.1166 | 0.0987 |
| 0.171 | 1.25 | 0.2003 | -0.2076 | 0.1023 | 0.1009 |
| 0.171 | 1.27 | 0.1891 | -0.1403 | 0.0935 | 0.0892 |
| 0.171 | 1.29 | 0.1789 | -0.1501 | 0.0873 | 0.0691 |
| 0.171 | 1.31 | 0.1697 | -0.0539 | 0.0811 | 0.0475 |
| 0.171 | 1.33 | 0.1612 | 0.0494 | 0.0863 | 0.0323 |
| 0.171 | 1.35 | 0.1535 | 0.0445 | 0.0852 | 0.0252 |
| 0.171 | 1.37 | 0.1463 | -0.0549 | 0.0928 | 0.0282 |
| 0.171 | 1.39 | 0.1397 | -0.0955 | 0.0902 | 0.0273 |
| 0.171 | 1.41 | 0.1336 | 0.1115 | 0.0992 | 0.0236 |
| 0.171 | 1.43 | 0.1279 | 0.0248 | 0.0966 | 0.0321 |
| 0.171 | 1.45 | 0.1226 | -0.1038 | 0.1012 | 0.0496 |
| 0.171 | 1.47 | 0.1177 | 0.0737 | 0.1044 | 0.0532 |
| 0.171 | 1.49 | 0.1131 | -0.0359 | 0.1164 | 0.0348 |
| 0.171 | 1.51 | 0.1088 | 0.1313 | 0.1212 | 0.0470 |
| 0.171 | 1.53 | 0.1047 | -0.0421 | 0.1123 | 0.0392 |
| 0.171 | 1.55 | 0.1009 | -0.0407 | 0.1301 | 0.0437 |
| 0.171 | 1.57 | 0.0973 | 0.1640 | 0.1282 | 0.0500 |
| 0.171 | 1.59 | 0.0939 | 0.0849 | 0.1257 | 0.0397 |
| 0.171 | 1.61 | 0.0907 | 0.1488 | 0.1272 | 0.0494 |
| 0.171 | 1.63 | 0.0877 | -0.2310 | 0.1327 | 0.0363 |
| 0.171 | 1.65 | 0.0849 | 0.0010 | 0.1307 | 0.0699 |
| 0.171 | 1.67 | 0.0821 | -0.1142 | 0.1491 | 0.0988 |
| 0.171 | 1.69 | 0.0796 | -0.0854 | 0.1411 | 0.0907 |
| 0.171 | 1.71 | 0.0771 | -0.1065 | 0.1567 | 0.0313 |
| 0.171 | 1.73 | 0.0748 | -0.0842 | 0.1320 | 0.2450 |
| 0.171 | 1.75 | 0.0726 | 0.0403 | 0.1358 | 0.2297 |
| 0.171 | 1.77 | 0.0705 | -0.2224 | 0.1371 | 0.1066 |
| 0.171 | 1.79 | 0.0685 | 0.0233 | 0.1479 | 0.0362 |
| 0.171 | 1.81 | 0.0665 | 0.0863 | 0.1549 | 0.0396 |
| 0.171 | 1.83 | 0.0647 | -0.2280 | 0.1577 | 0.0469 |
| 0.171 | 1.85 | 0.0630 | -0.4711 | 0.1931 | 0.0549 |
| 0.171 | 1.87 | 0.0613 | -0.2894 | 0.1996 | 0.0438 |
| 0.171 | 1.89 | 0.0597 | -0.5063 | 0.2802 | 0.0649 |
| 0.204 | 1.11 | 0.3673 | 0.0185 | 0.0377 | 0.0912 |
| 0.204 | 1.13 | 0.3399 | -0.1158 | 0.0671 | 0.0493 |
| 0.204 | 1.15 | 0.3160 | -0.0262 | 0.0625 | 0.0516 |
| 0.204 | 1.17 | 0.2948 | -0.0988 | 0.0889 | 0.0782 |
| 0.204 | 1.19 | 0.2760 | -0.4801 | 0.1177 | 0.0939 |
| 0.204 | 1.21 | 0.2592 | -0.3904 | 0.1278 | 0.0780 |
| 0.204 | 1.23 | 0.2441 | -0.5227 | 0.1300 | 0.1102 |
| 0.204 | 1.25 | 0.2304 | -0.1860 | 0.1235 | 0.1144 |
| 0.204 | 1.27 | 0.2180 | -0.0850 | 0.1101 | 0.0915 |
| 0.204 | 1.29 | 0.2067 | -0.0364 | 0.0983 | 0.0554 |
| 0.204 | 1.31 | 0.1964 | -0.1202 | 0.0912 | 0.0396 |
| 0.204 | 1.33 | 0.1869 | 0.0284 | 0.0809 | 0.0236 |
| 0.204 | 1.35 | 0.1781 | -0.0359 | 0.0792 | 0.0216 |
| 0.204 | 1.37 | 0.1701 | 0.0660 | 0.0778 | 0.0233 |
| 0.204 | 1.39 | 0.1626 | 0.0317 | 0.0846 | 0.0214 |
| 0.204 | 1.41 | 0.1557 | 0.0047 | 0.0848 | 0.0261 |
| 0.204 | 1.43 | 0.1492 | 0.0515 | 0.0876 | 0.0316 |
| 0.204 | 1.45 | 0.1432 | -0.0170 | 0.0921 | 0.0493 |
| 0.204 | 1.47 | 0.1375 | 0.2261 | 0.1087 | 0.0559 |
| 0.204 | 1.49 | 0.1323 | 0.0939 | 0.1217 | 0.0431 |
| 0.204 | 1.51 | 0.1273 | 0.0287 | 0.1432 | 0.0506 |
| 0.204 | 1.53 | 0.1227 | 0.1683 | 0.1248 | 0.0510 |
| 0.204 | 1.55 | 0.1183 | 0.0850 | 0.1135 | 0.0600 |
| 0.204 | 1.57 | 0.1142 | 0.0467 | 0.1219 | 0.0499 |
| 0.204 | 1.59 | 0.1103 | 0.1159 | 0.1166 | 0.0505 |
| 0.204 | 1.61 | 0.1066 | -0.2039 | 0.1312 | 0.0513 |
| 0.204 | 1.63 | 0.1031 | -0.0232 | 0.1200 | 0.0347 |
| 0.204 | 1.65 | 0.0998 | 0.1468 | 0.1214 | 0.0784 |
| 0.204 | 1.67 | 0.0967 | -0.1756 | 0.1392 | 0.0989 |
| 0.204 | 1.69 | 0.0937 | 0.0113 | 0.1442 | 0.0817 |
| 0.204 | 1.71 | 0.0908 | -0.1441 | 0.1474 | 0.0304 |
| 0.204 | 1.73 | 0.0881 | -0.0089 | 0.1291 | 0.2363 |
| 0.204 | 1.75 | 0.0856 | -0.0464 | 0.1324 | 0.2124 |
| 0.204 | 1.77 | 0.0831 | -0.2643 | 0.1409 | 0.0907 |
| 0.204 | 1.79 | 0.0808 | -0.1352 | 0.1525 | 0.0353 |
| 0.204 | 1.81 | 0.0785 | -0.1765 | 0.1600 | 0.0340 |
| 0.204 | 1.83 | 0.0764 | -0.2669 | 0.1822 | 0.0448 |
| 0.204 | 1.85 | 0.0744 | 0.1580 | 0.3028 | 0.0471 |
| 0.244 | 1.11 | 0.4092 | 0.0102 | 0.0295 | 0.0247 |
| 0.244 | 1.13 | 0.3806 | 0.0417 | 0.0486 | 0.0313 |
| 0.244 | 1.15 | 0.3553 | -0.0863 | 0.0754 | 0.0517 |
| 0.244 | 1.17 | 0.3327 | -0.1114 | 0.0967 | 0.0825 |
| 0.244 | 1.19 | 0.3126 | -0.3386 | 0.1156 | 0.1004 |
| 0.244 | 1.21 | 0.2945 | -0.4920 | 0.1192 | 0.0898 |
| 0.244 | 1.23 | 0.2781 | -0.3692 | 0.1121 | 0.0853 |
| 0.244 | 1.25 | 0.2632 | -0.3888 | 0.0967 | 0.0902 |
| 0.244 | 1.27 | 0.2496 | -0.2762 | 0.0815 | 0.0736 |
| 0.244 | 1.29 | 0.2371 | -0.2306 | 0.0774 | 0.0526 |
| 0.244 | 1.31 | 0.2257 | -0.0769 | 0.0702 | 0.0288 |
| 0.244 | 1.33 | 0.2152 | -0.1269 | 0.0707 | 0.0270 |
| 0.244 | 1.35 | 0.2054 | 0.0126 | 0.0708 | 0.0195 |
| 0.244 | 1.37 | 0.1964 | 0.0277 | 0.0698 | 0.0218 |
| 0.244 | 1.39 | 0.1881 | 0.0384 | 0.0730 | 0.0248 |
| 0.244 | 1.41 | 0.1803 | 0.0492 | 0.0837 | 0.0240 |
| 0.244 | 1.43 | 0.1730 | 0.0170 | 0.0888 | 0.0292 |
| 0.244 | 1.45 | 0.1662 | 0.1331 | 0.1004 | 0.0557 |
| 0.244 | 1.47 | 0.1598 | -0.0492 | 0.1105 | 0.0624 |
| 0.244 | 1.49 | 0.1539 | 0.0211 | 0.1247 | 0.0538 |
| 0.244 | 1.51 | 0.1482 | -0.0323 | 0.1327 | 0.0694 |
| 0.244 | 1.53 | 0.1429 | 0.0106 | 0.1107 | 0.0644 |
| 0.244 | 1.55 | 0.1380 | -0.0014 | 0.1052 | 0.0711 |
| 0.244 | 1.57 | 0.1333 | 0.1969 | 0.1081 | 0.0600 |
| 0.244 | 1.59 | 0.1288 | 0.1221 | 0.1081 | 0.0663 |
| 0.244 | 1.61 | 0.1246 | 0.0002 | 0.1217 | 0.0590 |
| 0.244 | 1.63 | 0.1206 | -0.0773 | 0.1145 | 0.0483 |
| 0.244 | 1.65 | 0.1168 | 0.0915 | 0.1286 | 0.0803 |
| 0.244 | 1.67 | 0.1132 | -0.2860 | 0.1138 | 0.1042 |
| 0.244 | 1.69 | 0.1098 | -0.2236 | 0.1265 | 0.0758 |
| 0.244 | 1.71 | 0.1065 | 0.0105 | 0.1236 | 0.0519 |
| 0.244 | 1.73 | 0.1034 | -0.0793 | 0.1294 | 0.2311 |
| 0.244 | 1.75 | 0.1004 | 0.0751 | 0.1459 | 0.1950 |
| 0.244 | 1.77 | 0.0976 | -0.0497 | 0.1614 | 0.0750 |
| 0.244 | 1.79 | 0.0949 | -0.0510 | 0.2574 | 0.0339 |
| 0.244 | 1.81 | 0.0923 | -0.4026 | 0.2762 | 0.0568 |
| 0.244 | 1.83 | 0.0898 | 0.0215 | 0.2132 | 0.0571 |
| 0.290 | 1.11 | 0.4523 | 0.0327 | 0.0327 | 0.0244 |
| 0.290 | 1.13 | 0.4228 | -0.0893 | 0.0450 | 0.0339 |
| 0.290 | 1.15 | 0.3965 | -0.1466 | 0.0764 | 0.0577 |
| 0.290 | 1.17 | 0.3728 | -0.1034 | 0.0751 | 0.0698 |
| 0.290 | 1.19 | 0.3515 | -0.3952 | 0.0948 | 0.0729 |
| 0.290 | 1.21 | 0.3322 | -0.3977 | 0.1036 | 0.0602 |
| 0.290 | 1.23 | 0.3147 | -0.3899 | 0.1058 | 0.0871 |
| 0.290 | 1.25 | 0.2986 | -0.1689 | 0.0953 | 0.0806 |
| 0.290 | 1.27 | 0.2839 | -0.2146 | 0.0857 | 0.0702 |
| 0.290 | 1.29 | 0.2704 | -0.0464 | 0.0749 | 0.0429 |
| 0.290 | 1.31 | 0.2579 | -0.0338 | 0.0757 | 0.0309 |
| 0.290 | 1.33 | 0.2463 | -0.0036 | 0.0727 | 0.0203 |
| 0.290 | 1.35 | 0.2356 | -0.0011 | 0.0693 | 0.0172 |
| 0.290 | 1.37 | 0.2257 | 0.0890 | 0.0795 | 0.0202 |
| 0.290 | 1.39 | 0.2164 | -0.0320 | 0.0804 | 0.0263 |
| 0.290 | 1.41 | 0.2077 | 0.0127 | 0.0817 | 0.0445 |
| 0.290 | 1.43 | 0.1996 | -0.0493 | 0.0858 | 0.0505 |
| 0.290 | 1.45 | 0.1920 | -0.0164 | 0.1017 | 0.0519 |
| 0.290 | 1.47 | 0.1849 | -0.0028 | 0.1030 | 0.0542 |
| 0.290 | 1.49 | 0.1781 | -0.0325 | 0.1110 | 0.0591 |
| 0.290 | 1.51 | 0.1718 | -0.0707 | 0.1087 | 0.0713 |
| 0.290 | 1.53 | 0.1659 | 0.1553 | 0.0994 | 0.0753 |
| 0.290 | 1.55 | 0.1602 | 0.0368 | 0.1038 | 0.0656 |
| 0.290 | 1.57 | 0.1549 | 0.0117 | 0.1018 | 0.0578 |
| 0.290 | 1.59 | 0.1498 | 0.0522 | 0.1091 | 0.0592 |
| 0.290 | 1.61 | 0.1450 | 0.0066 | 0.1073 | 0.0644 |
| 0.290 | 1.63 | 0.1405 | -0.0672 | 0.1144 | 0.0570 |
| 0.290 | 1.65 | 0.1362 | -0.0857 | 0.1604 | 0.0840 |
| 0.290 | 1.67 | 0.1321 | 0.0636 | 0.1457 | 0.0912 |
| 0.290 | 1.69 | 0.1282 | 0.0591 | 0.2000 | 0.0709 |
| 0.290 | 1.71 | 0.1244 | -0.0724 | 0.1865 | 0.0698 |
| 0.290 | 1.73 | 0.1209 | 0.0098 | 0.1954 | 0.2241 |
| 0.290 | 1.75 | 0.1175 | -0.1287 | 0.1779 | 0.1754 |
| 0.290 | 1.77 | 0.1142 | -0.2598 | 0.1563 | 0.0653 |
| 0.290 | 1.79 | 0.1111 | -0.2977 | 0.1575 | 0.0397 |
| 0.347 | 1.11 | 0.4964 | -0.0629 | 0.0497 | 0.0267 |
| 0.347 | 1.13 | 0.4664 | 0.0415 | 0.0476 | 0.0436 |
| 0.347 | 1.15 | 0.4395 | -0.0342 | 0.0564 | 0.0569 |
| 0.347 | 1.17 | 0.4150 | -0.1560 | 0.0810 | 0.0706 |
| 0.347 | 1.19 | 0.3928 | -0.1558 | 0.1028 | 0.0517 |
| 0.347 | 1.21 | 0.3726 | -0.2020 | 0.1079 | 0.0517 |
| 0.347 | 1.23 | 0.3540 | -0.3668 | 0.1091 | 0.0843 |
| 0.347 | 1.25 | 0.3369 | -0.1483 | 0.0927 | 0.0787 |
| 0.347 | 1.27 | 0.3212 | -0.1182 | 0.0788 | 0.0594 |
| 0.347 | 1.29 | 0.3066 | -0.0285 | 0.0754 | 0.0409 |
| 0.347 | 1.31 | 0.2932 | -0.0484 | 0.0707 | 0.0257 |
| 0.347 | 1.33 | 0.2806 | -0.0989 | 0.0727 | 0.0270 |
| 0.347 | 1.35 | 0.2690 | -0.0536 | 0.0714 | 0.0315 |
| 0.347 | 1.37 | 0.2581 | -0.0191 | 0.0766 | 0.0318 |
| 0.347 | 1.39 | 0.2479 | -0.0092 | 0.0742 | 0.0348 |
| 0.347 | 1.41 | 0.2383 | -0.0262 | 0.0786 | 0.0328 |
| 0.347 | 1.43 | 0.2294 | 0.0047 | 0.0870 | 0.0420 |
| 0.347 | 1.45 | 0.2209 | 0.0736 | 0.0834 | 0.0579 |
| 0.347 | 1.47 | 0.2130 | 0.1979 | 0.0871 | 0.0719 |
| 0.347 | 1.49 | 0.2055 | 0.2668 | 0.1103 | 0.0665 |
| 0.347 | 1.51 | 0.1985 | 0.0297 | 0.1022 | 0.0706 |
| 0.347 | 1.53 | 0.1918 | 0.0285 | 0.1082 | 0.0696 |
| 0.347 | 1.55 | 0.1855 | 0.0539 | 0.1064 | 0.0864 |
| 0.347 | 1.57 | 0.1795 | 0.0820 | 0.1276 | 0.0840 |
| 0.347 | 1.59 | 0.1738 | 0.0555 | 0.1368 | 0.0720 |
| 0.347 | 1.61 | 0.1684 | 0.1474 | 0.1441 | 0.1093 |
| 0.347 | 1.63 | 0.1633 | -0.0706 | 0.1819 | 0.0791 |
| 0.347 | 1.65 | 0.1584 | -0.0379 | 0.1251 | 0.0874 |
| 0.347 | 1.67 | 0.1537 | 0.2916 | 0.1448 | 0.0855 |
| 0.347 | 1.69 | 0.1492 | 0.1670 | 0.1201 | 0.0486 |
| 0.347 | 1.71 | 0.1450 | -0.0572 | 0.1246 | 0.0691 |
| 0.347 | 1.73 | 0.1409 | -0.1314 | 0.1488 | 0.2184 |
| 0.347 | 1.75 | 0.1371 | -0.0698 | 0.1840 | 0.1670 |
| 0.347 | 1.77 | 0.1334 | -0.4829 | 0.2950 | 0.0967 |
| 0.414 | 1.11 | 0.5406 | -0.0029 | 0.0311 | 0.1114 |
| 0.414 | 1.13 | 0.5107 | -0.0807 | 0.0517 | 0.0776 |
| 0.414 | 1.15 | 0.4835 | 0.0490 | 0.0590 | 0.0800 |
| 0.414 | 1.17 | 0.4587 | -0.1467 | 0.0813 | 0.0769 |
| 0.414 | 1.19 | 0.4359 | -0.1367 | 0.1060 | 0.0467 |
| 0.414 | 1.21 | 0.4149 | -0.1193 | 0.1175 | 0.0679 |
| 0.414 | 1.23 | 0.3955 | 0.0540 | 0.1112 | 0.0732 |
| 0.414 | 1.25 | 0.3776 | -0.1810 | 0.0944 | 0.0741 |
| 0.414 | 1.27 | 0.3610 | -0.1174 | 0.0849 | 0.0571 |
| 0.414 | 1.29 | 0.3456 | -0.0439 | 0.0732 | 0.0281 |
| 0.414 | 1.31 | 0.3312 | -0.0005 | 0.0687 | 0.0189 |
| 0.414 | 1.33 | 0.3178 | 0.0341 | 0.0671 | 0.0257 |
| 0.414 | 1.35 | 0.3052 | -0.1279 | 0.0745 | 0.0334 |
| 0.414 | 1.37 | 0.2935 | 0.0810 | 0.0786 | 0.0375 |
| 0.414 | 1.39 | 0.2824 | 0.0307 | 0.0702 | 0.0358 |
| 0.414 | 1.41 | 0.2720 | -0.0357 | 0.0925 | 0.0380 |
| 0.414 | 1.43 | 0.2622 | -0.0138 | 0.0936 | 0.0594 |
| 0.414 | 1.45 | 0.2530 | -0.0796 | 0.1205 | 0.0591 |
| 0.414 | 1.47 | 0.2443 | 0.0448 | 0.1027 | 0.0802 |
| 0.414 | 1.49 | 0.2360 | -0.0560 | 0.1571 | 0.0716 |
| 0.414 | 1.51 | 0.2282 | 0.1217 | 0.1442 | 0.0983 |
| 0.414 | 1.53 | 0.2208 | 0.0143 | 0.1116 | 0.0763 |
| 0.414 | 1.55 | 0.2138 | 0.0972 | 0.1030 | 0.0624 |
| 0.414 | 1.57 | 0.2071 | 0.1384 | 0.0997 | 0.0684 |
| 0.414 | 1.59 | 0.2008 | 0.1619 | 0.0942 | 0.0774 |
| 0.414 | 1.61 | 0.1947 | 0.0031 | 0.0899 | 0.0824 |
| 0.414 | 1.63 | 0.1890 | -0.0798 | 0.1111 | 0.0714 |
| 0.414 | 1.65 | 0.1835 | -0.1115 | 0.0998 | 0.1034 |
| 0.414 | 1.67 | 0.1782 | 0.2295 | 0.1097 | 0.0897 |
| 0.414 | 1.69 | 0.1732 | 0.3735 | 0.1711 | 0.0774 |
| 0.414 | 1.71 | 0.1684 | 0.0257 | 0.1559 | 0.1219 |
| 0.494 | 1.11 | 0.5841 | -0.0464 | 0.0224 | 0.0329 |
| 0.494 | 1.13 | 0.5547 | -0.0227 | 0.0783 | 0.0574 |
| 0.494 | 1.15 | 0.5277 | -0.1125 | 0.0667 | 0.0685 |
| 0.494 | 1.17 | 0.5028 | -0.2476 | 0.0871 | 0.0692 |
| 0.494 | 1.19 | 0.4798 | -0.1231 | 0.1296 | 0.0353 |
| 0.494 | 1.21 | 0.4584 | -0.2816 | 0.1548 | 0.0621 |
| 0.494 | 1.23 | 0.4385 | -0.1909 | 0.1376 | 0.0837 |
| 0.494 | 1.25 | 0.4200 | -0.4055 | 0.1251 | 0.0807 |
| 0.494 | 1.27 | 0.4028 | -0.0155 | 0.1127 | 0.0462 |
| 0.494 | 1.29 | 0.3866 | -0.0058 | 0.0998 | 0.0223 |
| 0.494 | 1.31 | 0.3715 | -0.0438 | 0.0953 | 0.0194 |
| 0.494 | 1.33 | 0.3573 | 0.0685 | 0.0816 | 0.0308 |
| 0.494 | 1.35 | 0.3440 | 0.0675 | 0.0687 | 0.0403 |
| 0.494 | 1.37 | 0.3314 | 0.0708 | 0.0721 | 0.0390 |
| 0.494 | 1.39 | 0.3196 | -0.0039 | 0.0639 | 0.0425 |
| 0.494 | 1.41 | 0.3084 | -0.0588 | 0.0667 | 0.0393 |
| 0.494 | 1.43 | 0.2979 | -0.0202 | 0.0832 | 0.0375 |
| 0.494 | 1.45 | 0.2879 | 0.0997 | 0.0677 | 0.0718 |
| 0.494 | 1.47 | 0.2784 | -0.0264 | 0.0802 | 0.0722 |
| 0.494 | 1.49 | 0.2694 | 0.1482 | 0.0895 | 0.0654 |
| 0.494 | 1.51 | 0.2609 | 0.0362 | 0.1028 | 0.1032 |
| 0.494 | 1.53 | 0.2528 | 0.0906 | 0.1036 | 0.0883 |
| 0.494 | 1.55 | 0.2450 | -0.0629 | 0.1545 | 0.0498 |
| 0.494 | 1.57 | 0.2377 | 0.2604 | 0.1142 | 0.0868 |
| 0.494 | 1.59 | 0.2307 | 0.0884 | 0.1288 | 0.0680 |
| 0.494 | 1.61 | 0.2240 | -0.1018 | 0.1647 | 0.0921 |
| 0.494 | 1.63 | 0.2176 | 0.1548 | 0.1418 | 0.1130 |
| 0.494 | 1.65 | 0.2115 | -0.0587 | 0.5014 | 0.3446 |
| 0.590 | 1.11 | 0.6266 | -0.0073 | 0.0137 | 0.0356 |
| 0.590 | 1.13 | 0.5982 | -0.0606 | 0.0504 | 0.0604 |
| 0.590 | 1.15 | 0.5717 | -0.0859 | 0.0765 | 0.0721 |
| 0.590 | 1.17 | 0.5471 | -0.1338 | 0.0776 | 0.0671 |
| 0.590 | 1.19 | 0.5242 | -0.1889 | 0.0992 | 0.0373 |
| 0.590 | 1.21 | 0.5028 | -0.3412 | 0.1204 | 0.0749 |
| 0.590 | 1.23 | 0.4827 | -0.1255 | 0.1003 | 0.0827 |
| 0.590 | 1.25 | 0.4639 | -0.1203 | 0.0818 | 0.0683 |
| 0.590 | 1.27 | 0.4462 | 0.0080 | 0.0696 | 0.0384 |
| 0.590 | 1.29 | 0.4296 | -0.0889 | 0.0607 | 0.0188 |
| 0.590 | 1.31 | 0.4139 | -0.0013 | 0.0552 | 0.0219 |
| 0.590 | 1.33 | 0.3991 | 0.0358 | 0.0528 | 0.0356 |
| 0.590 | 1.35 | 0.3852 | 0.0140 | 0.0765 | 0.0424 |
| 0.590 | 1.37 | 0.3720 | 0.0232 | 0.0595 | 0.0494 |
| 0.590 | 1.39 | 0.3595 | -0.0120 | 0.0748 | 0.0382 |
| 0.590 | 1.41 | 0.3476 | -0.0771 | 0.0902 | 0.0345 |
| 0.590 | 1.43 | 0.3364 | -0.0301 | 0.0753 | 0.0495 |
| 0.590 | 1.45 | 0.3257 | 0.1002 | 0.0906 | 0.0676 |
| 0.590 | 1.47 | 0.3155 | 0.1240 | 0.0889 | 0.0797 |
| 0.590 | 1.49 | 0.3058 | 0.1267 | 0.1266 | 0.1017 |
| 0.590 | 1.51 | 0.2966 | 0.1324 | 0.1475 | 0.0730 |
| 0.590 | 1.53 | 0.2878 | 0.1506 | 0.1725 | 0.0771 |
| 0.705 | 1.11 | 0.6671 | 0.0107 | 0.4872 | 0.1701 |
| 0.705 | 1.13 | 0.6399 | 0.2159 | 0.2646 | 0.1173 |
| 0.705 | 1.15 | 0.6145 | -0.0428 | 0.0493 | 0.0824 |
| 0.705 | 1.17 | 0.5906 | 0.0051 | 0.0760 | 0.0736 |
| 0.705 | 1.19 | 0.5681 | -0.0874 | 0.0919 | 0.0286 |
| 0.705 | 1.21 | 0.5470 | -0.1202 | 0.1196 | 0.0667 |
| 0.705 | 1.23 | 0.5270 | -0.0511 | 0.0922 | 0.0820 |
| 0.705 | 1.25 | 0.5082 | -0.0437 | 0.0925 | 0.0676 |
| 0.705 | 1.27 | 0.4903 | -0.0049 | 0.0701 | 0.0346 |
| 0.705 | 1.29 | 0.4735 | 0.0652 | 0.0739 | 0.0161 |
| 0.705 | 1.31 | 0.4575 | 0.1525 | 0.0755 | 0.0310 |
| 0.705 | 1.33 | 0.4423 | -0.0455 | 0.0575 | 0.0402 |
| 0.705 | 1.35 | 0.4279 | -0.0658 | 0.0573 | 0.0494 |
| 0.705 | 1.37 | 0.4142 | 0.0246 | 0.0621 | 0.0515 |
| 0.705 | 1.39 | 0.4012 | -0.0045 | 0.0951 | 0.0473 |
| 0.705 | 1.41 | 0.3888 | -0.0593 | 0.1435 | 0.0461 |
Appendix D Table for Results on Moments
| (GeV2) | Stat. | Syst. | Stat. | Syst. | Stat. | Syst. | ||||||
| 0.0120 | -0.0061 | -0.0062 | 0.0007 | 0.0006 | -1.0058 | -1.0291 | 0.1088 | 0.0978 | -2.4741 | -2.2482 | 0.2926 | 0.1917 |
| 0.0143 | -0.0062 | -0.0063 | 0.0007 | 0.0006 | -0.8755 | -0.8991 | 0.0950 | 0.0863 | -1.9606 | -1.7464 | 0.2603 | 0.1710 |
| 0.0171 | -0.0068 | -0.0070 | 0.0006 | 0.0002 | -0.8247 | -0.8453 | 0.0663 | 0.0357 | -1.7245 | -1.5237 | 0.1782 | 0.1429 |
| 0.0204 | -0.0084 | -0.0085 | 0.0005 | 0.0002 | -0.8692 | -0.8875 | 0.0532 | 0.0375 | -1.8753 | -1.6853 | 0.1323 | 0.1497 |
| 0.0244 | -0.0103 | -0.0104 | 0.0006 | 0.0003 | -0.9246 | -0.9405 | 0.0479 | 0.0410 | -1.9733 | -1.7951 | 0.1176 | 0.1481 |
| 0.0290 | -0.0094 | -0.0094 | 0.0006 | 0.0002 | -0.7466 | -0.7600 | 0.0424 | 0.0442 | -1.6696 | -1.5045 | 0.1026 | 0.1522 |
| 0.0347 | -0.0122 | -0.0122 | 0.0006 | 0.0003 | -0.8319 | -0.8429 | 0.0379 | 0.0511 | -1.7124 | -1.5622 | 0.0904 | 0.1532 |
| 0.0414 | -0.0133 | -0.0122 | 0.0012 | 0.0003 | -0.7820 | -0.7452 | 0.0568 | 0.0600 | -1.5215 | -1.3801 | 0.0888 | 0.1592 |
| 0.0494 | -0.0158 | -0.0145 | 0.0013 | 0.0003 | -0.8082 | -0.7723 | 0.0513 | 0.0689 | -1.5316 | -1.4098 | 0.0828 | 0.1684 |
| 0.0590 | -0.0173 | -0.0159 | 0.0013 | 0.0003 | -0.7926 | -0.7582 | 0.0457 | 0.0808 | -1.5135 | -1.4131 | 0.0790 | 0.1824 |
| 0.0705 | -0.0201 | -0.0183 | 0.0014 | 0.0004 | -0.8031 | -0.7697 | 0.0431 | 0.0913 | -1.4813 | -1.4032 | 0.0722 | 0.1933 |
| 0.0841 | -0.0215 | -0.0195 | 0.0015 | 0.0005 | -0.7590 | -0.7268 | 0.0407 | 0.0999 | -1.3284 | -1.2728 | 0.0661 | 0.1977 |
| 0.1005 | -0.0223 | -0.0198 | 0.0011 | 0.0008 | -0.7055 | -0.6736 | 0.0299 | 0.1038 | -1.1878 | -1.1547 | 0.0555 | 0.1927 |
| 0.1200 | -0.0206 | -0.0176 | 0.0014 | 0.0010 | -0.6134 | -0.5818 | 0.0315 | 0.0989 | -1.0376 | -1.0244 | 0.0545 | 0.1732 |
| 0.1430 | -0.0209 | -0.0173 | 0.0016 | 0.0012 | -0.5522 | -0.5201 | 0.0326 | 0.0942 | -0.8477 | -0.8500 | 0.0546 | 0.1577 |
| 0.1708 | -0.0187 | -0.0149 | 0.0014 | 0.0015 | -0.4837 | -0.4577 | 0.0252 | 0.0899 | -0.7521 | -0.7653 | 0.0375 | 0.1362 |
| 0.2042 | -0.0175 | -0.0127 | 0.0019 | 0.0015 | -0.4016 | -0.3730 | 0.0298 | 0.0830 | -0.5467 | -0.5650 | 0.0403 | 0.1175 |
| 0.2436 | -0.0099 | -0.0034 | 0.0021 | 0.0015 | -0.2852 | -0.2498 | 0.0300 | 0.0547 | -0.4062 | -0.4244 | 0.0412 | 0.0739 |
| 0.2904 | -0.0048 | 0.0039 | 0.0020 | 0.0023 | -0.2168 | -0.1723 | 0.0224 | 0.0413 | -0.3054 | -0.3180 | 0.0242 | 0.0535 |
| 0.3466 | 0.0027 | 0.0141 | 0.0022 | 0.0023 | -0.1295 | -0.0803 | 0.0218 | 0.0192 | -0.1844 | -0.1975 | 0.0209 | 0.0243 |
| 0.4139 | 0.0037 | 0.0194 | 0.0024 | 0.0030 | -0.1208 | -0.0599 | 0.0206 | 0.0186 | -0.1530 | -0.1596 | 0.0167 | 0.0149 |
| 0.4940 | 0.0158 | 0.0366 | 0.0026 | 0.0037 | -0.0245 | 0.0453 | 0.0187 | 0.0248 | -0.0574 | -0.0606 | 0.0130 | 0.0087 |
| 0.5902 | 0.0216 | 0.0488 | 0.0026 | 0.0052 | 0.0067 | 0.0823 | 0.0154 | 0.0314 | -0.0228 | -0.0265 | 0.0086 | 0.0096 |
| 0.7047 | 0.0247 | 0.0626 | 0.0029 | 0.0068 | 0.0138 | 0.1060 | 0.0158 | 0.0311 | -0.0128 | -0.0119 | 0.0084 | 0.0079 |
| 0.8412 | 0.0201 | 0.0776 | 0.0023 | 0.0037 | 0.0157 | 0.1313 | 0.0087 | 0.0208 | -0.0056 | -0.0017 | 0.0029 | 0.0045 |
| (GeV2) | Stat. | Syst. | Stat. | Syst. | Stat. | Syst. | ||||||
| 0.0204 | -0.0194 | -0.0205 | 0.0025 | 0.0011 | -1.9738 | -2.0776 | 0.2489 | 0.1252 | -3.8999 | -3.2665 | 0.6048 | 0.3394 |
| 0.0244 | -0.0199 | -0.0201 | 0.0026 | 0.0012 | -1.7253 | -1.7485 | 0.2295 | 0.1151 | -3.0963 | -2.5206 | 0.5697 | 0.3085 |
| 0.0290 | -0.0286 | -0.0290 | 0.0027 | 0.0014 | -2.1487 | -2.1853 | 0.2021 | 0.1171 | -3.9462 | -3.4522 | 0.5043 | 0.3041 |
| 0.0347 | -0.0255 | -0.0263 | 0.0025 | 0.0015 | -1.7120 | -1.7651 | 0.1670 | 0.1035 | -3.3601 | -2.9548 | 0.4277 | 0.2799 |
| 0.0414 | -0.0289 | -0.0300 | 0.0021 | 0.0015 | -1.6658 | -1.7296 | 0.1212 | 0.0872 | -3.0780 | -2.7518 | 0.2878 | 0.2322 |
| 0.0494 | -0.0340 | -0.0355 | 0.0023 | 0.0017 | -1.7351 | -1.8125 | 0.1119 | 0.0849 | -3.2859 | -3.0435 | 0.2565 | 0.2191 |
| 0.0590 | -0.0367 | -0.0389 | 0.0025 | 0.0019 | -1.6804 | -1.7729 | 0.1087 | 0.0880 | -3.2418 | -3.0891 | 0.2441 | 0.2266 |
| 0.0705 | -0.0432 | -0.0460 | 0.0027 | 0.0022 | -1.7187 | -1.8226 | 0.1038 | 0.0859 | -3.1097 | -3.0412 | 0.2247 | 0.2072 |
| 0.0841 | -0.0422 | -0.0459 | 0.0029 | 0.0022 | -1.5006 | -1.6168 | 0.0977 | 0.0759 | -2.5957 | -2.6031 | 0.2016 | 0.1765 |
| 0.1005 | -0.0432 | -0.0479 | 0.0032 | 0.0024 | -1.3782 | -1.5049 | 0.0944 | 0.0726 | -2.2651 | -2.3351 | 0.1879 | 0.1598 |
| 0.1200 | -0.0525 | -0.0581 | 0.0035 | 0.0027 | -1.5293 | -1.6627 | 0.0903 | 0.0741 | -2.5625 | -2.6817 | 0.1679 | 0.1504 |
| 0.1430 | -0.0435 | -0.0500 | 0.0038 | 0.0025 | -1.1512 | -1.2839 | 0.0828 | 0.0574 | -1.7705 | -1.9131 | 0.1402 | 0.1089 |
| 0.1708 | -0.0447 | -0.0523 | 0.0039 | 0.0025 | -1.0652 | -1.1990 | 0.0758 | 0.0519 | -1.5226 | -1.6809 | 0.1183 | 0.0901 |
| 0.2042 | -0.0375 | -0.0461 | 0.0044 | 0.0027 | -0.8382 | -0.9677 | 0.0739 | 0.0454 | -1.1072 | -1.2590 | 0.1033 | 0.0681 |
| 0.2436 | -0.0401 | -0.0493 | 0.0044 | 0.0027 | -0.7906 | -0.9098 | 0.0646 | 0.0409 | -0.9348 | -1.0696 | 0.0838 | 0.0568 |
| 0.2904 | -0.0302 | -0.0390 | 0.0044 | 0.0024 | -0.5937 | -0.6940 | 0.0538 | 0.0294 | -0.6596 | -0.7657 | 0.0584 | 0.0350 |
| 0.3466 | -0.0158 | -0.0231 | 0.0048 | 0.0025 | -0.3757 | -0.4564 | 0.0504 | 0.0214 | -0.3849 | -0.4703 | 0.0483 | 0.0210 |
| 0.4139 | -0.0046 | -0.0083 | 0.0051 | 0.0027 | -0.2282 | -0.2805 | 0.0455 | 0.0198 | -0.2206 | -0.2760 | 0.0375 | 0.0158 |
| 0.4940 | -0.0062 | -0.0034 | 0.0066 | 0.0065 | -0.2115 | -0.2329 | 0.0492 | 0.0328 | -0.1806 | -0.2137 | 0.0336 | 0.0123 |
| (GeV2) | Stat. | Syst. | Stat. | Syst. | Stat. | Syst. | |||
| 0.0204 | -0.0139 | 0.0028 | 0.0013 | -1.3806 | 0.2769 | 0.1409 | -1.8808 | 0.6734 | 0.3974 |
| 0.0244 | -0.0115 | 0.0029 | 0.0013 | -0.9684 | 0.2551 | 0.1313 | -0.9566 | 0.6329 | 0.3657 |
| 0.0290 | -0.0223 | 0.0030 | 0.0015 | -1.6257 | 0.2247 | 0.1346 | -2.2643 | 0.5600 | 0.3633 |
| 0.0347 | -0.0165 | 0.0028 | 0.0017 | -1.0840 | 0.1862 | 0.1233 | -1.6636 | 0.4756 | 0.3399 |
| 0.0414 | -0.0205 | 0.0026 | 0.0016 | -1.1431 | 0.1439 | 0.1115 | -1.6240 | 0.3265 | 0.2967 |
| 0.0494 | -0.0243 | 0.0028 | 0.0018 | -1.2063 | 0.1325 | 0.1148 | -1.9128 | 0.2920 | 0.2908 |
| 0.0590 | -0.0266 | 0.0030 | 0.0021 | -1.1772 | 0.1272 | 0.1249 | -1.9592 | 0.2779 | 0.3057 |
| 0.0705 | -0.0319 | 0.0033 | 0.0024 | -1.2200 | 0.1213 | 0.1303 | -1.9169 | 0.2558 | 0.2960 |
| 0.0841 | -0.0307 | 0.0036 | 0.0024 | -1.0382 | 0.1141 | 0.1292 | -1.5690 | 0.2298 | 0.2748 |
| 0.1005 | -0.0325 | 0.0037 | 0.0026 | -0.9693 | 0.1073 | 0.1302 | -1.3946 | 0.2125 | 0.2591 |
| 0.1200 | -0.0459 | 0.0041 | 0.0030 | -1.2334 | 0.1035 | 0.1273 | -1.9032 | 0.1912 | 0.2377 |
| 0.1430 | -0.0372 | 0.0044 | 0.0028 | -0.8816 | 0.0961 | 0.1118 | -1.2385 | 0.1625 | 0.1938 |
| 0.1708 | -0.0421 | 0.0045 | 0.0028 | -0.8512 | 0.0865 | 0.1058 | -1.0698 | 0.1344 | 0.1671 |
| 0.2042 | -0.0376 | 0.0052 | 0.0030 | -0.6835 | 0.0860 | 0.0962 | -0.8095 | 0.1198 | 0.1383 |
| 0.2436 | -0.0504 | 0.0053 | 0.0031 | -0.7435 | 0.0766 | 0.0698 | -0.7433 | 0.1003 | 0.0955 |
| 0.2904 | -0.0465 | 0.0052 | 0.0027 | -0.5854 | 0.0629 | 0.0518 | -0.5180 | 0.0682 | 0.0652 |
| 0.3466 | -0.0394 | 0.0057 | 0.0028 | -0.4179 | 0.0592 | 0.0289 | -0.3160 | 0.0568 | 0.0324 |
| 0.4139 | -0.0284 | 0.0060 | 0.0030 | -0.2463 | 0.0538 | 0.0275 | -0.1417 | 0.0442 | 0.0216 |
| 0.4940 | -0.0404 | 0.0077 | 0.0071 | -0.2996 | 0.0569 | 0.0433 | -0.1727 | 0.0389 | 0.0158 |
| (GeV2) | (meas.) | (full) | Stat. | Syst. |
| 0.021 | 0.0042 | 0.0052 | ||
| 0.024 | 0.0008 | 0.0008 | ||
| 0.029 | 0.0122 | 0.0126 | ||
| 0.035 | 0.0031 | 0.0040 | ||
| 0.042 | 0.0046 | 0.0080 | ||
| 0.050 | 0.0051 | 0.0095 | ||
| 0.059 | 0.0050 | 0.0103 | ||
| 0.071 | 0.0065 | 0.0131 | ||
| 0.084 | 0.0026 | 0.0107 | ||
| 0.101 | 0.0022 | 0.0122 | ||
| 0.120 | 0.0156 | 0.0276 | ||
| 0.144 | 0.0051 | 0.0194 | ||
| 0.173 | 0.0109 | 0.0266 | ||
| 0.205 | 0.0055 | 0.0244 | ||
| 0.244 | 0.0236 | 0.0464 | ||
| 0.292 | 0.0230 | 0.0500 | ||
| 0.348 | 0.0226 | 0.0533 | ||
| 0.416 | 0.0123 | 0.0477 | ||
| 0.496 | 0.0384 | 0.0770 |
| (GeV2) | Stat. | Syst. | Stat. | Syst. | ||
| 0.020 | 0.161 | -3.531 | ||||
| 0.024 | -0.865 | -2.725 | ||||
| 0.029 | 0.723 | -3.732 | ||||
| 0.035 | 0.070 | -3.194 | ||||
| 0.042 | 0.215 | -2.975 | ||||
| 0.050 | 0.471 | -3.290 | ||||
| 0.059 | 0.513 | -3.340 | ||||
| 0.071 | 0.481 | -3.288 | ||||
| 0.084 | 0.269 | -2.814 | ||||
| 0.101 | 0.215 | -2.524 | ||||
| 0.120 | 0.850 | -2.899 | ||||
| 0.144 | 0.368 | -2.068 | ||||
| 0.171 | 0.287 | -1.817 | ||||
| 0.205 | 0.231 | -1.361 | ||||
| 0.244 | 0.308 | -1.156 | ||||
| 0.292 | 0.192 | -0.828 | ||||
| 0.348 | 0.113 | -0.508 | ||||
| 0.416 | -0.021 | -0.298 | ||||
| 0.496 | 0.110 | -0.231 |